### v1

# A Tutte-type canonical decomposition of 4-connected graphs

Abstract.
We provide a unique decomposition of every 4-connected graph into parts that are either quasi-5-connected, cycles of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices, generalised double-wheels, or thickened $K_{4,m}$’s. The decomposition can be described in terms of a tree-decomposition but with edges allowed in the adhesion-sets. Our construction is explicit, canonical, and exhibits a defining property of the Tutte-decomposition.
As a corollary, we obtain a new Tutte-type canonical decomposition of 3-connected graphs into parts that are either quasi-4-connected, generalised wheels or thickened $K_{3,m}$’s. This decomposition is similar yet different from the tri-separation decomposition [Tridecomp].
As an application of the decomposition for 4-connectivity, we obtain a new theorem characterising all 4-connected vertex-transitive finite graphs as quasi-5-connected, the $K_{4}$-expansion of a quasi-5-connected graph, or on a short explicit list of graphs.
Key words and phrases:
4-connected, decomposition, Tutte, canonical, double wheel, 4-separation, 4-separator, tetra-separation, tetra-separator
2020 Mathematics Subject Classification:
05C40, 05C75, 05C83, 05E18
[Uncaptioned image] Funded by DFG, project number 546892829.
1. Introduction
Problem 1.1.
Is there for every $k\in\mathbb{N}$ an explicit canonical way to decompose every $k$-connected graph along $k$-separators into smaller pieces that are $(k+1)$-connected or ‘basic’?
[…]
1.1 has a trivial solution for $k=0$, and for $k=1$ a positive answer is provided by the block-cutvertex-decomposition. Already for $k=2$, the solution is a classical theorem of Tutte [TutteCon] from 1966. It was not known how to extend Tutte’s approach to $k=3$ until recently, when Carmesin and Kurkofka discovered a new notion of 3-separators which allowed them to find a Tutte-type canonical decomposition of 3-connected graphs [Tridecomp]. However, examples show that the obvious extension of their solution to $k=4$ fails. In particular, it was not clear whether there even exists a notion of 4-separators that allows for such an extension.
In this paper, we introduce a new type of 4-separator and, as our main result, use it to prove a Tutte-type canonical decomposition theorem for 4-connectivity. As a corollary, we obtain a new Tutte-type canonical decomposition for 3-connectivity that provides an alternative perspective to [Tridecomp]. We also obtain a new theorem characterising vertex-transitive 4-connected finite graphs.
Background. To solve 1.1 for $k=0$ and $k=1$, one simply cuts along all $k$-separators of the graph. But already for $k=2$, this approach fails. Indeed, different 2-separators can cross in that they cut through each other, as happens in a cycle. Then if we cut at one of the 2-separators, we loose the other. Tutte realised that the solution is to cut the graph at all 2-separators that are totally-nested in that they are not crossed by any 2-separators, and indeed no conflicts arise when cutting in this way. The resulting pieces of Tutte’s decomposition are 3-connected, cycles or $K_{2}$’s. All three decompositions for $k\leqslant 3$ can be described in terms of tree-decompositions. For instance, Tutte’s selection of 2-separators gives a recipe for a tree-decomposition with all adhesion-sets of size $2$ and all whose torsos [^1] A torso of a tree-decomposition is obtained from a bag by making each of its adhesion-sets into a clique. are 3-connected, cycles or $K_{2}$’s.
Tutte’s approach, when extended verbatim, fails for $k\geqslant 3$. There are at least two main challenges that must be overcome.
Challenge 1.
There are $k$-connected graphs which just slightly fail to be $(k+1)$-connected, but all whose $k$-separators are crossed. For example, if $G$ is $k$-connected and the neighbourhoods of its vertices are precisely the $k$-separators of $G$, then the $k$-separators around neighbouring vertices cross, but $G$ fails to be $(k+1)$-connected. Such $G$ exist for all $k\geqslant 3$: take a circular saw as in Figure 1.1, for instance.
Figure 1.1. A circular saw has vertex-set $\mathbb{Z}_{n}\times\mathbb{Z}_{2}$ with $n\gg k$ and every vertex $(v,0)$ sends edges to the $k$ vertices $(v,1),\ldots,(v+k-1,1)$. It is $k$-connected, and when $k\geqslant 3$ its neighbourhoods of vertices are precisely its $k$-separators (Lemma A.2).
Challenge 2.
We consider the $k$-connected graph $G$ depicted in Figure 1.2 if $k$ is odd and in Figure 1.3 if $k$ is even. Every $k$-separator of $G$ is formed from a red set, a non-adjacent blue set, and if $k$ is even also the apex-vertex in the centre. Hence every $k$-separator of $G$ is crossed by another $k$-separator. However, $G$ is far from $(k+1)$-connected, and in no way ‘basic’: every $42k$-clique can be replaced with an arbitrarily large $(k+1)$-connected graph, so $G$ represents a whole zoo of examples.
Figure 1.2. $k$ is odd. Red: $\lfloor k/2\rfloor$-cliques. Blue: $\lceil k/2\rceil$-cliques. Grey indicates all possible edges.
Figure 1.3. $k$ is even. Red: $(\tfrac{k}{2}-1)$-cliques. Blue: $\tfrac{k}{2}$-cliques. Grey indicates all possible edges. The central vertex is an apex.
Decomposing 4-connected and 3-connected graphs. We found a way to overcome the two challenges for $k=4$ and $k=3$ with a twofold approach, which we first explain for $k=4$.
Ad 1. We relax the notion of $5$-connected to quasi-5-connected, an idea that we learned about from Grohe’s work [grohe2016quasi]. A graph $G$ is quasi-5-connected if it is 4-connected and every 4-separation $(A,B)$ of $G$ satisfies $|A\smallsetminus B|=1$ or $|B\smallsetminus A|=1$.
Ad 2. We introduce the new notion of a tetra-separation, which we use instead of 4-separators. There are two key differences. On the one hand, tetra-separations may use edges in addition to vertices to cut up the graph, an idea that we learned about from the work of Carmesin and Kurkofka [Tridecomp] (we will discuss the relation later). On the other hand, every vertex used by a tetra-separation to cut up the graph into two sides must have at least two neighbours in each side, not counting neighbours that are also used by the tetra-separation. This degree-condition is new.
Recall that a mixed-separation of a graph $G$ is a pair $(A,B)$ with $A\cup B=V(G)$ and $A\smallsetminus B\neq\emptyset\neq B\smallsetminus A$. The separator of $(A,B)$ is the set of all vertices in $A\cap B$ and all edges in $G$ with one end in $A\smallsetminus B$ and the other in $B\smallsetminus A$. A tetra-separation of $G$ is a mixed-separation $(A,B)$ of $G$ whose separator has size four and which satisfies the following two conditions:
  1. degree-condition:
    every vertex in the separator has $\geqslant 2$ neighbours in both $A\smallsetminus B$ and $B\smallsetminus A$;
  2. matching-condition:
    the edges in the separator form a matching.
Two tetra-separations $(A,B)$ and $(C,D)$ are nested if, after possibly exchanging the name $A$ with $B$ or the name $C$ with $D$, we have $A\subseteq C$ and $B\supseteq D$. A tetra-separation of $G$ is totally-nested if it is nested with every tetra-separation of $G$. Figure 1.4 shows the totally-nested tetra-separations of an example-graph similar to the one shown in Figure 1.3.
Figure 1.4. The totally-nested tetra-separations of a graph. Their separators are coloured.
While the decompositions for $k$-connectivity for $k\leqslant 2$ such as the Tutte-decomposition can be described as tree-decompositions, the decomposition of 4-connected graphs via tetra-separations requires a notion that sits in between the notions of tree-decomposition (which allows only vertices in adhesion-sets) and Wollan’s notion of tree-cut decomposition (which allows only edges in adhesion-sets) [WollanExcludedImmersion]. To this end, the notion of mixed-tree-decomposition has been introduced which allows both vertices and edges in adhesion-sets [Tridecomp]. The set of totally-nested tetra-separations of a 4-connected graph $G$ canonically corresponds to a unique mixed-tree-decomposition of $G$. See Section 2.2 for details.
In order to state our decomposition-result for 4-connected graphs, we need the following definitions to describe the ‘basic’ parts of the decomposition – those that are not quasi-5-connected. Recall that graph-decompositions generalise tree-decompositions in that the decomposition-tree may take the form of an arbitrary graph, such as a cycle, in which case we speak of a cycle-decomposition [canonicalGraphDec]. Let $X$ be a graph. We call $X$ a generalised double-wheel with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that $X-u-v$ has a cycle-decomposition of adhesion 1 into $K_{2}$’s and triangles (Figure 1.5). We call $X$ a cycle of $X_{1}$-torsos, …, $X_{k}$-torsos where the $X_{i}$ are graphs, if $X$ is 4-connected and has a cycle-decomposition with all adhesion-sets of size 2 such that every torso is isomorphic to some $X_{i}$ (Figure 1.6). Recall that the only 3-connected graphs on $\leqslant 5$ vertices are $K_{4}$, the $4$-wheel, $K_{5}$ minus an edge, and $K_{5}$. A thickened $K_{4,m}$ is obtained from a $K_{4,m}$ by turning the left side into a $K_{4}$. A sprinkled $K_{4,m}$ is obtained from a $K_{4,m}$ by adding any number of edges to the left side.
Figure 1.5. A generalised double-wheel
Figure 1.6. A cycle of triangle-torsos, $K_{4}$-torsos and $K_{5}$-torsos
Theorem 1.
Let $G$ be a 4-connected graph. Let $\mathcal{T}(G)$ denote the mixed-tree-decomposition of $G$ that is uniquely determined by the set $N(G)$ of all totally-nested tetra-separations of $G$. Every torso $\tau$ of $\mathcal{T}(G)$ satisfies one of the following:
  1. (i)
    $\tau$ is quasi-5-connected;
  2. (ii)
    $\tau$ is a cycle of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices;
  3. (iii)
    $\tau$ is a generalised double-wheel;
  4. (iv)
    $\tau$ is a thickened $K_{4,m}$ with $m\geqslant 0$, or the entire graph $G$ is a sprinkled $K_{4,m}$ with $m\geqslant 4$.
Remark 1.2.
Both $N(G)$ and $\mathcal{T}(G)$ in Theorem 1 are canonical: if $\varphi\colon G\to G^{\prime}$ is a graph-isomorphism, then $N(\varphi(G))=\varphi(N(G))$ and $\mathcal{T}(\varphi(G))=\varphi(\mathcal{T}(G))$.
Remark 1.3.
Theorem 1 includes infinite graphs $G$ with no additional restrictions whatsoever. More generally, all graphs in this paper are allowed to be infinite, unless stated otherwise.
Theorem 1 implies a new Tutte-type canonical decomposition for 3-connectivity. A strict tri-separation is defined like a tetra-separation, except that the separator is required to have size three instead of four. A strict tri-separation of a graph $G$ is totally-nested if it is nested with every strict tri-separation of $G$. A generalised wheel with centre $v$ is a 3-connected graph $X$ with a vertex $v$ such that $X-v$ has a cycle-decomposition with all adhesion-sets of size 1 and into $K_{2}$’s and triangles.
Corollary 2.
Let $G$ be a 3-connected graph. Let $M(G)$ denote the set of all strict tri-separations of $G$ that are totally-nested (with regard to the strict tri-separations). Let $\mathcal{T}(G)$ denote the mixed-tree-decomposition of $G$ that is uniquely determined by $M(G)$. Every torso $\tau$ of $\mathcal{T}(G)$ satisfies one of the following:
  1. (i)
    $\tau$ is quasi-4-connected;
  2. (ii)
    $\tau$ is a generalised wheel;
  3. (iii)
    $\tau$ is a thickened $K_{3,m}$ with $m\geqslant 0$, or the entire graph $G$ is a sprinkled $K_{3,m}$ with $m\geqslant 4$.
Both Remarks 1.2 and 1.3 also apply to 2.
Vertex-transitive graphs. It is well known that all vertex-transitive finite connected graphs are either 3-connected, cycles, or $K_{2}$’s [infiniteSPQR]. This fact has then been refined in [Tridecomp] by analysing the 3-connected outcome further; see [Tridecomp, Corollary 2]. However, [Tridecomp] leaves it open to gain new insights into vertex-transitive 4-connectivity. This is where we apply Theorem 1: to obtain a characterisation of vertex-transitive 4-connectivity. Recall that a graph $G$ is arc-transitive if for every two ordered pairs $(u_{i},v_{i})$ with $u_{i}v_{i}\in E(G)$ for $i=1,2$ there is an automorphism $\varphi$ of $G$ such that $\varphi(u_{1})=u_{2}$ and $\varphi(v_{1})=v_{2}$. The $K_{4}$-expansion of a 4-regular graph $G$ is obtained from $G$ by expanding every vertex to a $K_{4}$, and adding a perfect matching such that two $K_{4}$’s are joined by a matching-edge if and only if their original vertices in $G$ are joined by a $G$-edge. For other definitions related to 2, see Section 12.
{restatable}
mainresultmaintransitive A 4-connected finite graph $G$ is vertex-transitive if and only if $G$ either is
  1. (i)
    a quasi-5-connected vertex-transitive graph,
  2. (ii)
    the $K_{4}$-expansion of a quasi-5-connected arc-transitive graph,
  3. (iii)
    a cycle of $K_{4}$-bags of length $\geqslant 4$ (Figure 1.7),
  4. (iv)
    a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 4$ (Figure 1.8),
  5. (v)
    a cycle of triangle-bags of length $\geqslant 6$ (Figure 1.9),
or $G$ is one of the following graphs or a $K_{4}$-expansion thereof:
  1. (vi)
    a cycle of $K_{2,2}$-bags of length $\geqslant 3$ (Figure 1.10),
  2. (vii)
    the line-graph of the cube (Figure 1.11).
2 strengthens a classical theorem of Godsil and Royle from geometric group theory in a special case [Godsil, Theorem 3.4.2].
Figure 1.7.
A cycle of $K_{4}$-bags
Figure 1.8.
A cycle alternating between $K_{4}$-bags and $C_{4}$-bags
Figure 1.9.
A cycle of triangle-bags
Figure 1.10. A cycle of $K_{2,2}$-bags and its $K_{4}$-expansion
Figure 1.11. The line-graph of the cube and its $K_{4}$-expansion
Applications that require canonicity and explicitness. There are applications for partial solutions to 1.1, and in particular for Theorem 1 and [Tridecomp], that absolutely require canonicity and explicitness, often both. Examples of these applications include:
  1. $\cdot$
    Computing excluded minors. A major challenge at the interface of computer science and graph-minor theory is to determine the excluded minors for explicit minor-closed graph-classes. Partial solutions to 1.1 for $k\leqslant 3$ can be used to determine the excluded minors for planar graphs and series-parallel graphs. We have conviction that solutions for larger $k$ imply exact structure theorems for other graph-classes, such as the class of apex-graphs, and that these structure theorems can be used to determine the excluded minors of these classes.
  2. $\cdot$
    Connectivity augmentation. Carmesin and Sridharan used the canonicity and explicitness of the tri-separation decomposition [Tridecomp] to construct the first FPT-algorithm for connectivity augmentation from 0 to 4 [conny_aug]. We would be excited to see if Theorem 1 can be used to extend this to 5-connectivity.
  3. $\cdot$
[…]
  4. $\cdot$
    Parallel computing. Total-nestedness makes computing the decomposition in parallel a feasible option, as all partial solutions can be merged without conflict; see for example [Sarah].
  5. $\cdot$
    Cops and robber. In the cops and robber game of Robertson and Seymour, $k$ cops have a winning strategy for catching a robber in a graph iff that graph has a tree-decomposition of width $<k$ [GMcopsRobber]. When the robber is invisible, however, the probability of the cops catching the robber depends on the minimum number of leaf-bags attained by any tree-decomposition of width $<k$. While with other approaches to constructing small-width tree-decompositions, it is not even clear how to verify that they minimize the number of leaf-bags, our theorem gives such tree-decompositions right away in the case of $k$ cops searching $k$-connected graphs for $k\leqslant 4$ [CopsInvisibleRobber].
  6. $\cdot$
    Groups. The characterisation of vertex-transitivity for 4-connectivity (2) crucially relies on canonicity and total-nestedness. In a different direction, local separators have recently been used to prove a low-order Stallings-type theorem for finite nilpotent groups [StallingsNilpotent]. The proof heavily relies on the Tutte-type decomposition for local 2-separators [Local2sep] and exploits the combination of canonicity and total-nestedness. A next step could be using Theorem 1 and 2, or the tri-separation decomposition [Tridecomp].
Comparison to the tri-separation decomposition. Carmesin and Kurkofka introduced the first Tutte-type canonical decomposition for 3-connected graphs [Tridecomp]. They used tri-separations instead of strict tri-separations to obtain their decomposition. The main difference between tri-separations on the one hand, and strict tri-separations and tetra-separations on the other hand, is how their degree-conditions handle neighbours in the separator: tri-separations count them, but strict tri-separations and tetra-separations ignore them. This seemingly subtle difference has a surprisingly big impact: the verbatim extension of the tri-separation approach to 4-connectivity fails 2.
For 3-connected graphs, the strict tri-separation decomposition can be viewed as a refinement of the tri-separation decomposition, but the details are a bit technical; see Appendix B.
1.1. Overview of the proof
To get started, let $G$ be a 4-connected graph. Let $N$ denote its set of totally-nested tetra-separations, and let $\mathcal{T}=(T,\mathcal{V})$ denote the mixed-tree-decomposition of $G$ that is uniquely determined by $N$. Let $\tau$ denote the torso of $\mathcal{T}$ at an arbitrary node $t$ of $T$. It will be straightforward to check that $\tau$ is 4-connected or a $K_{4}$.
In order to show that $\tau$ satisfies one of the outcomes of Theorem 1, we will study how $\tau$ interacts with the tetra-separations of $G$. Indeed, the outcome will be determined by this. Let $(A,B)$ be a tetra-separation of $G$, which may possibly be crossed. We say that $(A,B)$ interlaces $\tau$ if, roughly speaking, both $A\smallsetminus B$ and $B\smallsetminus A$ intersect $\tau$. Here may think of $(A,B)$ as ‘cutting through’ $\tau$. Whether $\tau$ is interlaced or not already yields a first outcome:
(1) If $\tau$ is not interlaced by any tetra-separation of $G$, then $\tau$ is quasi-5-connected.
(See 10.5.) A technical lemma from [Tridecomp] does most of the work of proving 1.
By 1, we may assume from now on that some tetra-separation $(A,B)$ of $G$ interlaces $\tau$. Then $(A,B)$ is not in $N$, so some tetra-separation $(C,D)$ crosses $(A,B)$. The remaining outcomes will be determined by the way in which $(A,B)$ and $(C,D)$ cross. More specifically, the outcome will be determined by the size of the vertex-centre $Z$, which consists of all vertices in the intersection of the separators of $(A,B)$ and $(C,D)$. We will show that
(2) $|Z|\in\{0,2,4\}.$
(See the (Crossing Lemma). (4.1).) So there will be exactly three outcomes. Here we focus only on the cycle-of-graphs outcome, which we get for $|Z|=0$. This outcome is the hardest to prove, and we use all the new methods introduced in this paper in the proof of this outcome. It is also different from all situations encountered in [Tridecomp].
Assume now that $|Z|=0$. As our first step, we will use $(A,B)$ and $(C,D)$ to cut $G$ into four 2-connected pieces, arranged in the shape of a 4-cycle, and apply the Tutte-decomposition to each piece; see Figure 1.12. Since $G$ is 4-connected, the Tutte-decompositions will be path-decompositions. We will then merge the four Tutte-decompositions into a cycle-decomposition $\mathcal{O}=(O,(G_{o})_{o\in V(O)})$. So all adhesion-sets of $\mathcal{O}$ have size two. We then use the 4-connectivity of $G$ to note that the torsos of $\mathcal{O}$ are triangles, 4-cycles, or 3-connected. Moreover, we will find some rules on how these types of torsos may be arranged on $O$. For instance, a torso that is a 4-cycle must always have two 3-connected torsos as neighbours, and if two triangle-torsos are neighbours then their adhesion-set is spanned by an edge.
Figure 1.12. Constructing a cycle-decomposition from four Tutte-decompositions
Every node $o$ of $O$ defines a ($\leqslant 4$)-separation of $G$: its sides are $V(G_{o})$ and $\bigcup_{o^{\prime}\in V(O-o)}V(G_{o^{\prime}})$, so its separator is the union of the two adhesion-sets at $G_{o}$. From this ($\leqslant 4$)-separation, we then obtain a candidate $(U_{o},W_{o})$ for a totally-nested tetra-separation by performing slight alterations to the separator, replacing some vertices with edges in a systematic way. The first key result will be a characterisation of those torsos $H_{o}$ that give rise to a totally-nested tetra-separation $(U_{o},W_{o})$ in terms of an efficiently testable property: being good. For example, only 3-connected torsos can be good, and having $\geqslant 6$ vertices in addition to being 3-connected suffices for being good.
(3) $(U_{o},W_{o})$ is a totally-nested tetra-separation of $G$ if and only if $H_{o}$ is good.
(See 9.41). In order to show the hard backward implication of 3, we will introduce a
(4) characterisation of total-nestedness in terms of connectivity-properties of separators.
(See 6.16.) Examples of connectivity-properties of a tetra-separator $S(A,B)$ that are necessary for $(A,B)$ to be totally-nested include, roughly speaking, that for every bipartition $(\pi_{1},\pi_{2})$ of $S(A,B)$ into two classes $\pi_{i}$ of size 2 there must be
  • $\cdot$
    5 independent $\pi_{1}$$\pi_{2}$ paths in $G$ on the one hand, and
  • $\cdot$
    3 independent paths between the two elements of $\pi_{1}$ in $G-\pi_{2}$ on the other hand.
A major difference to tri-separations is that the connectivity-properties required of the separators are much more complex, as tetra-separations can cross in significantly more complicated ways than tri-separations; see Figure 1.13 for an example.
Figure 1.13. Three tetra-separations cross in a way that tri-separations cannot.
With 4 at hand, the problem of proving the hard backward implication of 3 becomes a problem about finding many paths in many configurations, roughly speaking. To this end, we develop a number of lemmata for finding such paths systematically, and then combine them to prove 3. Finding the paths takes a lot of work, especially when $O$ is short; see Figure 1.14 for example.
Figure 1.14. The bag $G_{o}$ is the top left quarter. Its candidate $(U_{o},W_{o})$ for a totally-nested tetra-separation is indicated in yellow. The sets $\pi_{1}$ and $\pi_{2}$ consist of the blue and the green vertices, respectively. As a small part of proving 3 via 4, we have to find 5 independent $\pi_{1}$$\pi_{2}$ paths (red). We find two such paths in $G_{o}$ using Menger as $G_{o}$ is good. To find the other three, we essentially go around $O$ in both directions starting from $o$, constructing one path in the one direction and two in the other. The example shows that the construction must be highly efficient to avoid collision of the paths when the decomposition-cycle $O$ is short.
By 3, each tetra-separation $(U_{o},W_{o})$ for a good torso $H_{o}$ corresponds to a directed edge $e_{o}$ of $T$. The second key result about $\mathcal{O}$ will be that
(5) The edges $e_{o}$ are precisely the directed edges of the form $(s,t)$ with $st\in E(T)$,
where $t$ is the node of $T$ with $\tau$ as torso. Phrased differently, the tetra-separations $(U_{o},W_{o})$ from good torsos $H_{o}$ are precisely the elements of $N$ that ‘carve out’ $\tau$. (See Lemma 9.49.)
Once we have proved 5, we are almost done. Indeed, it now follows that the torso $\tau$ can be obtained from $G$ by essentially replacing the good torsos of $\mathcal{O}$ with $K_{4}$’s and contracting a few edges. Then it will be fairly immediate to see that $\tau$ is a cycle of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices. (See 9.50.)
Proving 5, however, is a little tricky. We will assume for a contradicting that 5 fails, so some totally-nested tetra-separation $(E,F)$ will ‘cut through’ $\mathcal{O}$. Since $(E,F)$ is nested with both $(A,B)$ and $(C,D)$, we have $E\subseteq A\cap C$, say; see Figure 1.15. We then start adjusting both $(A,B)$ and $(C,D)$: by moving bags of $\mathcal{O}$ from $A\smallsetminus B$ to $B\smallsetminus A$ and from $C\smallsetminus D$ to $D\smallsetminus C$, one bag at a time, while using the total-nestedness of $(E,F)$ to maintain the inclusion $E\subseteq A\cap C$. Eventually, there will be essentially only one bag $G_{o}$ left, and we will see that its torso $H_{o}$ is good. We will then show that $(E,F)=(U_{o},W_{o})$ is the only possibility. Since $(U_{o},W_{o})$ does not ‘cut through’ $\mathcal{O}$, this will yield a contradiction.
Figure 1.15. $(E,F)$ (blue) is nested with $(A,B)$ (yellow) and $(C,D)$ (purple). We use the structure of $\mathcal{O}$ to move $(A,B)$ closer to $(E,F)$ by moving one bag to the other side. This replaces the top half of yellow with red.
Outlook. We would be excited to see an extension of Theorem 1 to higher connectivity. However, the verbatim extension of tetra-separations fails 2 for $k=5$.
Open Problem 1.4.
Extend Theorem 1 to $k$-connectivity for $k\geqslant 5$.
Both the strict tri-separation decomposition and the tri-separation decomposition may have quasi-4-connected torsos. To decompose these torsos further, we can apply the tetra-separation decomposition, but only to those quasi-4-connected torsos that are even 4-connected.
Open Problem 1.5.
Extend Theorem 1 from 4-connectivity to quasi-4-connectivity.
Simply using the tetra-separation decomposition fails: see Figure 1.16 for an illustrative example that essentially generalises 2 to quasi-4-connectivity.
Figure 1.16. A quasi-4-connected version of 2: all tetra-separations are crossed, yet the graph is far from $5$-connected and each $K_{10}$ can be replaced with an arbitrary 5-connected graph
The tri-separation decomposition has been used to prove a characterisation of vertex-transitivity for 3-connectivity [Tridecomp, Corollary 2]. One of the outcomes is that the graph is quasi-4-connected or a triangle-expansion of a quasi-4-connected graph. [^2] The actual outcome is slightly more specific.
Open Problem 1.6.
Extend 2 from 4-connectivity to quasi-4-connectivity.
We believe that a solution to 1.5 can be used to approach 1.6.
Open Problem 1.7.
Extend the FPT-algorithm of [conny_aug] for connectivity augmentation from 0 to 4 to 5-connectivity.
Since the tri-separation decomposition [Tridecomp] has been heavily exploited in [conny_aug], we hope that Theorem 1 can be used to approach 1.7.
[…]
For directed graphs, the case $k=1$ of 1.1 has recently been solved almost completely by Bowler, Gut, Hatzel, Kawarabayashi, Muzi and Reich [Flo].
[…]
Organisation of the paper. Section 2 introduces or recalls all essential terminology. Section 3 provides stronger but more technical versions of Theorem 1 and 2, and proves all while assuming a number of key lemmata that remain to be proved later in the paper. Section 4 provides a characterisation of how tetra-separations cross. Section 5 introduces a systematic way for obtaining tetra-separations from mixed-4-separations. Section 6 gives a characterisation of total-nestedness for tetra-separations in terms of connectivity-properties of their separators. The next sections investigate the different outcomes of Theorem 1: Section 7 deals with the $K_{4,m}$; Section 8 deals with the double-wheel; Section 9 deals with the cycle-of-graphs; and Section 10 deals with quasi-5-connectivity. This will conclude the proof of Theorem 1. Section 11 introduces an ‘angry theorem’ for 4-connectivity: a characterisation of the 4-connected graphs all whose tetra-separations are crossed (hence the name). Section 12 uses Theorem 1 to obtain a characterisation of the vertex-transitive 4-connected finite graphs. Appendix A proves that the circular saws from 1 really exhibit the properties we used. Appendix B compares the strict tri-separation-decomposition (2) with the tri-separation decomposition [Tridecomp].
2. Essential terminology
This section introduces all terminology used by Theorem 1. For graph-theoretic terminology we follow [bibel].
2.1. Tetra-separations
A mixed-separation+ of a graph $G$ is a pair $(A,B)$ such that $A\cup B=V(G)$. The separator $)𝑆𝐴𝐵{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{S(A,B)}}$ of $(A,B)$ is the disjoint union of the vertex set $A\cap B$ and the edge-set $E(A\smallsetminus B,B\smallsetminus A)$. The order of $(A,B)$ is the size of $S(A,B)$. A mixed-separation+ of order $k$ is also called mixed-$k$-separation+. A mixed-separation+ is proper if both $A\smallsetminus B$ and $B\smallsetminus A$ are nonempty. In this case, we refer to $(A,B)$ as a mixed-separation without the plus-sign. A mixed-separation of order $k$ is also called mixed-$k$-separation. A mixed-separation with $k$ vertices in its separator but no edges is called a separation, or $k$-separation.
Definition 2.1.
A tetra-separation of a graph $G$ is a mixed-4-separation $(A,B)$ of $G$ that satisfies the following two conditions:
  1. degree-condition:
    every vertex in $A\cap B$ has $\geqslant 2$ neighbours in both $A\smallsetminus B$ and $B\smallsetminus A$;
  2. matching-condition:
    the edges in $S(A,B)$ form a matching.
Definition 2.2.
A strict tri-separation of a graph $G$ is a mixed-3-separation of $G$ that satisfies both the degree-condition and the matching-condition.
Recall that a graph $G$ is quasi-$k$-connected if it is $(k-1)$-connected and every $(k-1)$-separation $(A,B)$ of $G$ satisfies $|A\smallsetminus B|=1$ or $|B\smallsetminus A|=1$.
Lemma 2.3.
Let $G$ be a 4-connected graph.
  1. (i)
    If $G$ has no tetra-separation, then $G$ is quasi-5-connected.
  2. (ii)
    If $G$ is quasi-5-connected and $|G|\geqslant 8$, then $G$ has no tetra-separation.
We will prove Lemma 2.3 in Section 5.2.
Example 2.4.
The lower-bound $|G|\geqslant 8$ in Lemma 2.3 (ii) is tight: the graph depicted in Figure 2.1 is quasi-5-connected, has $7$ vertices, but has a tetra-separation.
Figure 2.1. A small quasi-5-connected graph with a tetra-separation (red)
The mixed-separations+ of a graph come with the usual partial ordering:
$(A,B)\;{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{\leqslant}}\;(C,D):\Leftrightarrow A\subseteq C\text{ and }B\supseteq D.$
Two mixed-separations+ $(A,B)$ and $(C,D)$ are nested if, after possibly switching the name $A$ with $B$ or the name $C$ with $D$, we have $(A,B)\leqslant(C,D)$. Otherwise, they cross. A set of mixed-separations+ is nested if all its elements are pairwise nested. In the context of $(A,B)$ and $(C,D)$ crossing, the vertex-centre is the vertex-set $A\cap B\cap C\cap D$.
A tetra-separation of a graph $G$ is totally-nested if it is nested with every tetra-separation of $G$.
2.2. Stars, mixed-tree-decompositions and torsos
The following definition is supported by Figure 2.2. Let $G$ be a graph. A pair $\mathcal{T}=(T,(V_{t})_{t\in T})$ of a tree $T$ and a family of vertex sets $V_{t}\subseteq V(G)$ indexed by the nodes $t\in T$ is a mixed-tree-decomposition of $G$ if it satisfies the following two conditions:
  1. (M1)
    $V(G)=\bigcup_{t\in T}V_{t}$;
  2. (M2)
    the subgraph of $T$ induced by $\{\,t\in T:v\in V_{t}\,\}$ is connected for every vertex $v\in G$.
We refer to the vertex sets $V_{t}$ as bags. Mixed-tree-decompositions generalise both tree-decompositions and tree-cut decompositions. Every oriented edge $(t_{1},t_{2})$ of $T$ induces a mixed-separation+ $(A_{1},A_{2})$ where $A_{i}=\bigcup_{t\in V(T_{i})}V_{t}$ for the component $T_{i}$ of $T-t_{1}t_{2}$ with $t_{i}\in T_{i}$. This defines a map ${\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{\alpha_{\mathcal{T}}(t_{1},t_{2})}}:=(A_{1},A_{2})$ with domain the set ${\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{\vec{E}(T)}}$ of oriented edges of $T$. The set of induced mixed-separations+ of $\mathcal{T}$ is nested. It is symmetric: for each element $(A_{1},A_{2})$ it also contains $(A_{2},A_{1})$. We call $\mathcal{T}$ proper if all its induced mixed-separations+ are proper. The adhesion-set of an edge $t_{1}t_{2}$ of $T$ is the separator of $\alpha_{\mathcal{T}}(t_{1},t_{2})$, which is well-defined.
Figure 2.2. A mixed-tree-decomposition. The decomposition-tree is blue. Bags are grey. Edges and vertices in adhesion-sets are red.
A star of mixed-separations+, or star for short, is a set $\sigma=\{\,(A_{i},B_{i}):i\in I\,\}$ of mixed-separations+ $(A_{i},B_{i})$ such that $(A_{i},B_{i})\leqslant(B_{j},A_{j})$ for all distinct $i,j\in I$. For every node $t$ of $T$, the set $\sigma_{t}:=\{\,\alpha_{\mathcal{T}}(s,t):st\in E(T)\,\}$ is a star. Now let $S$ be a set of mixed-separations+ with $\sigma\subseteq S$. We call $\sigma$ a splitting star of $S$ if $\sigma$ is a star and for every $(C,D)\in S$ there is $i\in I$ with either $(C,D)\leqslant(A_{i},B_{i})$ or $(D,C)\leqslant(A_{i},B_{i})$. If $\mathcal{T}$ is proper, then the splitting stars of the image of $\alpha_{\mathcal{T}}$ are precisely the stars $\sigma_{t}$. Conversely, if $S$ is a nested symmetric set of mixed-separations of a finite graph $G$, then there is a unique (up to isomorphisms for decomposition-trees) mixed-tree-decomposition $\mathcal{T}(S)$ of $G$ such that $\alpha_{\mathcal{T}(S)}$ is a bijection $\vec{E}(T)\to S$. The mixed-tree-decomposition $\mathcal{T}(S)=(T,\mathcal{V})$ has an explicit description (see [Tridecomp, §3.1]):
  1. (i)
    the nodes of $T$ are precisely the splitting stars of $S$;
  2. (ii)
    two nodes $\sigma,\sigma^{\prime}$ form an edge iff there is $(A,B)\in\sigma$ with $(B,A)\in\sigma^{\prime}$;
  3. (iii)
    $V_{\sigma}:=\bigcap\,\{\,B:(A,B)\in\sigma\,\}$
where $V_{\emptyset}:=V(G)$ if necessary. Then $S$ can be recovered from $\mathcal{T}(S)$: it equals the image of $\alpha_{\mathcal{T}(S)}$. In this sense, $S$ and $\mathcal{T}(S)$ contain the same information, and we can switch at will between the two perspectives that they offer.
When $G$ is infinite and $S$ is the set of totally-nested tetra-separations of $G$, then $\mathcal{T}(S)$ is a mixed-tree-decomposition of $G$ by standard arguments as in [Tridecomp, Lemma 3.3.10].
A mixed-separation $(C,D)$ interlaces a star $\sigma$ of mixed-separations if for every $(A,B)\in\sigma$ we have either $(A,B)<(C,D)$ or $(A,B)<(D,C)$. The notion of interlacing yields a characterisation of splitting stars: a star $\sigma$ included in a nested set $S$ of mixed-separations of a graph $G$ is a splitting star of $S$ if and only if no element of $S$ interlaces $\sigma$ [Tridecomp, Lemma 2.2.3].
Let $\mathcal{T}=(T,\mathcal{V})$ be a mixed-tree-decomposition of a graph $G$. For a mixed-separation+ $(A,B)$, we let the vertex-set $)𝜆𝐴𝐵{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{\lambda(A,B)}}$ consist of all vertices in $S(A,B)$ and the endvertex in $A\smallsetminus B$ for each edge in $S(A,B)$. The expanded-torso of $t$, or of $V_{t}$, is obtained from the subgraph of $G$ induced by
$V_{t}\cup\bigcup_{st\in E(T)}\lambda(\alpha_{\mathcal{T}}(s,t))$
by making each of the vertex-sets $\lambda(\alpha_{\mathcal{T}}(s,t))$ into a clique. From this expanded-torso we obtain the compressed-torso, or torso for short, by contracting all edges contained in adhesion-sets of the edges $st\in E(T)$; see Figure 2.3.
More generally, let $\sigma=\{\,(A_{i},B_{i}):i\in I\,\}$ be a star of mixed-separations+. The expanded-torso of $\sigma$ is obtained from the subgraph $V_{\sigma}\cup\bigcup_{i\in I}\lambda(A_{i},B_{i})$ by making each of the vertex-sets $\lambda(A_{i},B_{i})$ into a clique. The compressed-torso, or torso, of $\sigma$ is obtained from this expanded torso by contracting all edges contained in any separators $S(A_{i},B_{i})$. The two notions of torsos are compatible: the torso of a node $t$ of $T$ equals the torso of the star $\sigma_{t}$.
Figure 2.3. The torsos of the mixed-tree-decomposition depicted in Figure 2.2.
2.3. Graph-decompositions
Let $G$ and $H$ be graphs. Let $\mathcal{G}=(G_{h}:h\in V(H))$ be a family of subgraphs $G_{h}\subseteq G$ indexed by the nodes of $H$. The pair $(H,\mathcal{G})$ is an $H$-decomposition or graph-decomposition of $G$ if
  1. (H1)
    $G=\bigcup_{h\in V(H)}G_{h}$, and
  2. (H2)
    for every $v\in V(G)$, the subgraph $H_{v}\subseteq H$ induced on $\{h\in V(H):v\in V(G)\}$ is connected.
We refer to the $G_{h}$ as its parts and to $H$ as its decomposition graph. When $H$ is a cycle, we call $(H,\mathcal{G})$ a cycle-decomposition
The adhesion-graphs of $(H,\mathcal{G})$ are the intersections $G_{h}\cap G_{h^{\prime}}$ for the edges $hh^{\prime}\in E(H)$. The adhesion-sets of $(H,\mathcal{G})$ are the vertex-sets of its adhesion-graphs. The torso of a part $G_{h}$ is obtained from $G_{h}$ by making every adhesion-graph $G_{h}\cap G_{h^{\prime}}$ with $hh^{\prime}\in E(G)$ into a complete graph by adding all possible edges. An edge of a torso with both ends contained in the same adhesion-set is a torso-edge, even if it already exists in $G$. The interior ${\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{\mathring{G}_{h}}}$ of $G_{h}$ is the graph $G_{h}-\bigcup_{h^{\prime}\in H-h}G_{h^{\prime}}$. The interior of a torso is defined similarly. The interior vertices of $G_{h}$ are the vertices of the interior of $G_{h}$, and the interior vertices of a torso are defined similarly.
Let $X$ be a graph. We call $X$ a generalised wheel with centre $v$ if $X$ is 3-connected with $v\in V(X)$ such that $X-v$ has a cycle-decomposition with all adhesion-sets of size 1 and into $K_{2}$’s and triangles. We call $X$ a generalised double-wheel with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that $X-u-v$ has a cycle-decomposition with all adhesion-sets of size 1 and into $K_{2}$’s and triangles. We call $X$ a cycle of $X_{1}$-torsos, …, $X_{k}$-torsos where the $X_{i}$ are graphs, if $X$ is 4-connected and has a cycle-decomposition with all adhesion-sets of size two such that every torso is isomorphic to some $X_{i}$. A thickened $K_{4,m}$ is obtained from a $K_{4,m}$ by turning the left side into a $K_{4}$. A sprinkled $K_{4,m}$ is obtained from a $K_{4,m}$ by adding any number of edges to the left side.
3. Proofs of Theorem 1 and 2 and strengthenings thereof
Instead of proving Theorem 1 and 2 directly, we will first prove stronger but slightly more technical versions of them, Theorem 3.1 and Corollary 3.2 below, and then immediately obtain Theorem 1 and 2 from their strengthenings. The proof of Theorem 3.1 uses a number of lemmas that have not yet been introduced. Stating and proving these lemmas will be the main objectives of later sections.
Theorem 3.1 (Technical strengthening of Theorem 1).
Let $G$ be a 4-connected graph. Let $N$ denote the set of all totally-nested tetra-separations of $G$. Let $\sigma$ be a splitting star of $N$ with torso $\tau$.
  1. (i)
    If $\sigma$ is not interlaced by any tetra-separation, then $\tau$ is quasi-5-connected.
Assume now that $\sigma$ is interlaced by a tetra-separation $(A,B)$, which is crossed by another tetra-separation $(C,D)$ with vertex-centre $Z$. Then $|Z|\in\{0,2,4\}$.
  1. (ii)
    If $|Z|=0$, then $\tau$ is a cycle of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices.
  2. (iii)
    If $|Z|=2$, then $\tau$ is a generalised double-wheel with centre $Z$.
  3. (iv)
    If $|Z|=4$, then $\tau$ is a thickened $K_{4,m}$ with $Z$ as the left side and $m\geqslant 0$, or the entire graph $G$ is a sprinkled $K_{4,m}$ with $Z$ as the left side and $m\geqslant 4$.
Proof.
If no tetra-separation of $G$ interlaces $\sigma$, then $\tau$ is quasi-5-connected by 10.5. So we may assume that some tetra-separation $(A,B)$ of $G$ interlaces $\sigma$. Then $(A,B)$ is crossed by some tetra-separation $(C,D)$, with $Z$ denoting the vertex-centre. By the (Crossing Lemma). (4.1), we have $|Z|\in\{0,2,4\}$.
(ii). Assume $|Z|=0$. Then $\tau$ is a cycle of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices by 9.50.
(iii). Assume $|Z|=2$. Then $\tau$ is a generalised double-wheel with centre $Z$ by 8.37.
(iv). Assume $|Z|=4$. If $\sigma$ is non-empty, then $\tau$ is a thickened $K_{4,m}$ with $Z$ as left side and $m\geqslant 0$ by 7.7. Otherwise the entire graph $G$ is a sprinkled $K_{4,m}$ with $Z$ as left side and $m\geqslant 4$, also by 7.7. ∎
Proof of Theorem 1.
Immediate from Theorem 3.1. ∎
Corollary 3.2 (Technical strengthening of 2).
Let $G$ be a 3-connected graph. Let $N$ be the set of all strict tri-separations of $G$ that are totally-nested (with regard to the strict tri-separations of $G$). Let $\sigma$ be a splitting star of $N$ with torso $\tau$.
  1. (i)
    If $\sigma$ is not interlaced by any strict tri-separation, then $\tau$ is quasi-4-connected.
Assume now that $\sigma$ is interlaced by a strict tri-separation $(A,B)$, which is crossed by another strict tri-separation $(C,D)$ with vertex-centre $Z$.
  1. (ii)
    If $Z$ consists of exactly one vertex $v$, then $\tau$ is a generalised wheel with centre $v$.
  2. (iii)
    Otherwise $Z$ has size three. If $\sigma\neq\emptyset$, then $\tau$ is a thickened $K_{3,m}$ with $m\geqslant 0$, and otherwise $G$ is a sprinkled $K_{3,m}$ with $m\geqslant 4$.
Proof.
Let $\hat{G}$ be the graph obtained from $G$ by adding an apex-vertex $\alpha$. Every tetra-separation of $\hat{G}$ contains $\alpha$ in its separator. Hence the map $\varphi$ that sends $(A,B)$ to $(A\cup\{\alpha\},B\cup\{\alpha\})$ is a bijection between the set of strict tri-separations of $G$ and the set of tetra-separations of $\hat{G}$. Two strict tri-separations of $G$ are nested if and only if their images under $\varphi$ are nested. Hence $\varphi$ restricts to a bijection between the set $N$ of totally-nested strict tri-separations of $G$ and the set $\hat{N}$ of totally-nested tetra-separations of $\hat{G}$. Let $\hat{\sigma}:=\varphi(\sigma)$, which is a splitting star of $\hat{N}$. Let $\hat{\tau}$ denote the torso of $\hat{\sigma}$. Since $\alpha$ lies in all separators of tetra-separations of $\hat{G}$, no edge incident to $\alpha$ is contracted in the definition of the torso $\hat{\tau}$. Thus $\tau=\hat{\tau}-\alpha$.
(i). Assume that $\sigma$ is not interlaced by any strict tri-separation. Then $\hat{\sigma}$ is not interlaced by any tetra-separation, so $\hat{\tau}$ is quasi-5-connected by Theorem 3.1 (i). If $\tau$ is not quasi-4-connected, this is witnessed by some separation $(X,Y)$ of $\tau$, and then $(X\cup\{\alpha\},Y\cup\{\alpha\})$ witnesses that $\hat{\tau}$ is not quasi-5-connected. Hence $\tau$ is quasi-4-connected.
So we may assume next that $\sigma$ is interlaced by a strict tri-separation $(A,B)$, which is crossed by another strict tri-separation $(C,D)$ with $Z$ denoting the centre. Then $\varphi(A,B)$ interlaces $\sigma$ and is crossed by $\varphi(C,D)$ with vertex-centre $\hat{Z}:=Z\cup\{\alpha\}$.
(ii). Assume that $Z$ consists of exactly one vertex $v$. Then $\hat{\tau}$ is a generalised double-wheel with centre $\{v,\alpha\}$ by Theorem 3.1 (iii). Hence $\hat{\tau}$ is a generalised wheel with centre $v$.
(iii). The vertex-centre $Z$ cannot have sizes 0 or 2, since then $\hat{Z}$ would have size 1 or 3, which is excluded in Theorem 3.1. Thus $|Z|=3$, so $|\hat{Z}|=4$. Since $\alpha\in\hat{Z}$, both outcomes of Theorem 3.1 (iv) translate to the two outcomes of (iii). ∎
Proof of 2.
Immediate from Corollary 3.2. ∎
Example 3.3.
The outcome in Theorem 3.1 is not in general unique. For example, consider the graph $G$ depicted in Figure 3.1 on the left. Every tetra-separation of $G$ is crossed, hence $\sigma=\emptyset$ is a splitting star of the set $N(G)$ of all totally-nested tetra-separations of $G$. On the left of Figure 3.1, we see two tetra-separations that interlace $\sigma$ and cross with empty vertex-centre. From their perspective, $G$ is a cycle of $K_{4}$-torsos and $4$-wheel torsos, see Theorem 3.1 (ii). On the right of Figure 3.1, we see two tetra-separations that interlace $\sigma$ and cross with a vertex-centre of size 2. From their perspective, $G$ is a generalised wheel, see Theorem 3.1 (iii).
For another example, consider the graph $G$ depicted in Figure 3.2 on the left. Again, $\sigma=\emptyset$ is a splitting star of $N(G)$. On the left of Figure 3.1, we see two tetra-separations that interlace $\sigma$ and cross with empty vertex-centre. From their perspective, $G$ is a cycle of $K_{4}$-torsos, see Theorem 3.1 (ii). On the right of Figure 3.1, we see two tetra-separations that interlace $\sigma$ and cross with a vertex-centre of size 4. From their perspective, $G$ is a $K_{4,4}$, see Theorem 3.1 (iv).
Figure 3.1. The drawings show the same graph, which can be seen with the help of the red cycle and the blue vertices. Two tetra-separations cross with empty vertex-centre (left) and with vertex-centre of size 2 (right).
Figure 3.2. The drawings show the same graph. Two tetra-separations cross with empty vertex-centre (left) and with vertex-centre of size 4 (right).
4. How tetra-separations cross
In this section, we characterise the ways in which tetra-separations can cross. The main result here is the (Crossing Lemma). (4.1).
We require the following terminology for corner diagrams from [Tridecomp, §1.3], which is supported by Figure 4.1. Let $(A,B)$ and $(C,D)$ be mixed-separations+ of a graph $G$. All of the following definitions depend on $(A,B)$ and $(C,D)$ being clear from context. The corner for $\{A,C\}$, or the corner $AC$ for short, is the vertex $(A\smallsetminus B)\cap(C\smallsetminus D)$. The corners for the other three combinations of one side from $(A,B)$ and one side from $(C,D)$ are defined analogously. Two corners are adjacent when their pair share a side, and opposite otherwise. An edge of $G$ is diagonal if its endvertices lie in opposite corners. The centre consists of all vertices in the vertex-centre and all diagonal edges.
The link for $C$, or $C$-link for short, consists of
  • $\cdot$
    all non-diagonal edges $f$ that have an end in the corner for a pair containing $C$ (these being the corners $AC$ or $BC$) and satisfy $f\in S(A,B)$, and
  • $\cdot$
    all vertices in $(C\smallsetminus D)\cap S(A,B)$.
The links for the other three sides $A,B,D$ are defined analogously. Each $L$-link is adjacent to the two corners for the pairs containing $L$. Two links are adjacent if they are adjacent to the same corner, and opposite otherwise. An edge of $G$ is jumping if its endvertices lie in opposite links.
Let $X\in\{A,B\}$ and $Y\in\{C,D\}$. An edge $e$ of $G$ dangles from the $Y$-link through the $X$-link if $e$ has an endvertex in the $Y$-link and $e$ is contained in the $X$-link. We then call $e$ dangling for short.
Figure 4.1. A crossing-diagram. Corners are yellow. Links are blue. The centre is red. The red edge is diagonal. The purple edge is jumping. The left blue edge dangles from the $A$-link through the $C$-link.
The following definition is supported by Figure 4.2. The corner for $AC$ is potter if it is empty and its adjacent links have size two and there exists a path $u_{1}u_{2}u_{3}u_{4}$ in $G$ such that $u_{1}u_{2}$ dangles from the vertex $u_{2}$ in the $C$-link through the $A$-link, and $u_{3}u_{4}$ dangles from the vertex $u_{3}$ in the $A$-link through the $C$-link.
Figure 4.2. The corner $AC$ is potter.
In the remainder of this section, we prove the following:
Key Lemma 4.1 (Crossing Lemma).
Let $(A,B)$ and $(C,D)$ be crossing tetra-separations in a 4-connected graph $G$. Then exactly one of the following assertions holds.
  1. (i)
    All links are empty and the centre has size four.
  2. (ii)
    All links have size one and the centre has size two.
  3. (iii)
    All links have size two and the centre is empty.
Moreover, there are no diagonal edges and no jumping edges. If there is an edge that dangles from one link through another, then we are in case (iii) and the corner in between the two links is potter.
A mixed-separation $(A,B)$ is an edge-cut if $S(A\cap B)\subseteq E(G)$. An edge-cut $(A,B)$ is atomic if $|A|=1$ or $|B|=1$.
Lemma 4.2.
Let $(A,B)$ be a mixed-4-separation of a 4-connected graph $G$ that satisfies the degree-condition. Then the following assertions are equivalent:
  1. (i)
    $(A,B)$ is not a tetra-separation;
  2. (ii)
    $(A,B)$ is an atomic cut in $G$;
  3. (iii)
    $|A\smallsetminus B|=1$ or $|B\smallsetminus A|=1$.
Proof.
(ii) $\Rightarrow$ (i) is trivial.
(i) $\Rightarrow$ (iii). If $(A,B)$ is not a tetra-separation, then there exist two edges in $S(A,B)$ that share an endvertex $u$, say $u\in A\smallsetminus B$. Since $G$ is 4-connected and since $(A,B\cup\{u\})$ is a mixed-separation+ of order at most 3, $(A,B)$ cannot be proper. Hence, $A\smallsetminus(B\cup\{u\})=\emptyset$ and $A\smallsetminus B=\{u\}$.
(iii) $\Rightarrow$ (ii). If say $|A\smallsetminus B|=1$ then $A\cap B=\emptyset$ by the degree-condition. It follows that $(A,B)$ is an atomic cut in $G$. ∎
Let $(A,B)$ and $(C,D)$ be mixed-separations+ of a graph $G$. The corner-separator $L(A,C)$ for the corner $AC$ is the union of the $A$-link, the $C$-link, the vertex-centre, and the diagonal edges that have an end in the corner $AC$. The corner-separators for the other three corners are defined analogously.
Lemma 4.3.
[Tridecomp, Lemma 1.3.4] For two mixed-separations $(A,B)$ and $(C,D)$ of $G$, we have
$|L(A,C)|+|L(B,D)|\leqslant|S(A,B)|+|S(C,D)|\leavevmode\nobreak\ .$
Moreover, if we have equality, there are no jumping edges, and every diagonal edge has its endvertices in the corners for $AC$ and $BD$.
Lemma 4.4.
Let two tetra-separations $(A,B)$ and $(C,D)$ cross in a 4-connected graph $G$. Then there exist two opposite corner-separators of size at least 4.
Proof.
Otherwise, two adjacent corner-separators have size at most 3, say the corner-separator for $AC$ and the corner-separator for $BC$. By 4-connectivity, the corners for $AC$ and $BC$ are empty. Since $C\smallsetminus D\neq\emptyset$, the link $L$ for $C$ is nonempty. The link $L$ only contains vertices since its adjacent corners are empty, so $L=(A\cap B)\smallsetminus D$. If $|L|\geqslant 2$ then the vertices in $L$ have at most one neighbour in $A\smallsetminus B$ (and $B\smallsetminus A$), contradicting the degree-condition for $(A,B)$. Otherwise, if $|L|=1$ then $(C,D)$ is not a tetra-separation by Lemma 4.2, a contradiction. ∎
The following lemma may be regarded as an analogue of [Tridecomp, Lemma 1.3.5].
Lemma 4.5.
Let $(A,B)$ and $(C,D)$ be two mixed-$k$-separations in a $k$-connected graph $G$ that cross so that two opposite corner-separators have size at least $k$. Then either
  1. (i)
    all links have the same size $\ell$ and the centre has size $k-2\ell$ for some $\ell\leqslant\lfloor k/2\rfloor$; or
  2. (ii)
    two adjacent links have size $i$, the other two links have size $j$, and the centre has size $k-i-j$, where $1\leqslant i\leqslant\lfloor(k-1)/2\rfloor$ and $i<j\leqslant k-i$.
Proof.
By Lemma 4.3, two opposite corner-separators have size exactly $k$, say the corner-separators for $AC$ and $BD$. Let $a,b,c,d$ denote the sizes of the links for $A,B,C,D$, respectively. Let $x$ denote the size of the centre. By Lemma 4.3, every diagonal edge has its endvertices in the corners for $AC$ and $BD$; and there are no jumping edges. We get $a+c+x=k$ and $b+d+x=k$ and $a+b+x=k$ and $c+d+x=k$. It follows that $b=c$ and $a=d$. Without loss of generality $a=d\leqslant b=c$.
Since $a+c\leqslant k$ and $a\leqslant c$ we have $a=d\in\{0,1,\ldots,\lfloor k/2\rfloor\}$. If $a=d=0$ then the corner for $AD$ must be nonempty, and hence $x=k$ since $G$ is $k$-connected. Then $b=c=0$ and we get outcome (i).
Otherwise $1\leqslant a=d\leqslant\lfloor k/2\rfloor$. If $a=b=c=d=:\ell$ then we get $x=k-a-c=k-2\ell$ and so get outcome (i). Otherwise $i:=a=d<b=c=:j$. Note that $i\leqslant\lfloor(k-1)/2\rfloor$. With $x=k-a-c=k-i-j$, we get outcome (ii). ∎
Lemma 4.6.
Let $(A,B)$ and $(C,D)$ be two crossing tetra-separations in a 4-connected graph $G$. Then all links have the same size $\ell$ and the centre has size $4-2\ell$, for some $\ell\in\{0,1,2\}$.
Proof.
Suppose that not all links have size $\ell$, for some $\ell\in\{0,1,2\}$. By Lemma 4.4 and Lemma 4.5, we get that two adjacent links have size one, the other two links have size at least two and the centre has size at most one. Without loss of generality, the corner $K$ of $AC$ is the corner whose adjacent links have size one. The corner-separator of $K$ has size at most 3, so $K$ is empty since $G$ is 4-connected. Then one of the links of size one contains a vertex $u$, say $u\in(A\cap B)\smallsetminus D$. But then $u$ has at most one neighbour in $A\smallsetminus B$, contradicting the degree-condition for $(A,B)$. Hence all links have the same size $\ell$ for some $\ell\in\{0,1,2\}$, and then the centre has size $4-2\ell$ by Lemma 4.4 and Lemma 4.5. ∎
Observation 4.7.
Let $(A,B)$ and $(C,D)$ be crossing tetra-separations in a 4-connected graph $G$. If a corner is potter, then its two adjacent corners are nonempty. In particular, no two adjacent corners are potter. ∎
Lemma 4.8.
Let $(A,B)$ and $(C,D)$ be two crossing tetra-separations in a 4-connected graph $G$.
  1. (i)
    There are no jumping edges and no diagonal edges.
  2. (ii)
    Let $X,Y\in\{A,B,C,D\}$. If some edge dangles from the $Y$-link through the $X$-link, then the corner for $XY$ is potter.
Proof.
(i) By Lemma 4.6, all links have the same size $\ell$ and the centre has size $4-2\ell$. By Lemma 4.3, there do not exist jumping edges nor diagonal edges.
(ii) Without loss of generality, $X=A$ and $Y=C$. Let $e=u_{1}u_{2}$ be a dangling edge such that $u_{2}$ lies in the $C$-link and $e$ lies in the $A$-link. We first show that the corner $K$ for $AC$ is empty. Assume for contradiction that $K\neq\emptyset$. By Lemma 4.6 and 4-connectivity, $|L(A,C)|=4$. Then $L(A,C)-e$ is a mixed-3-separator separating the vertices in the corner for $AC$ from $u_{1}$. This contradicts 4-connectivity by Lemma 6.1.
By the degree-condition of $(A,B)$ and since $K$ is empty, the vertex $u_{2}$ in the $C$-link has a neighbour $u_{3}$ in the $A$-link. By the degree-condition of $(C,D)$ and since $K$ is empty, the vertex $u_{3}$ has a neighbour $u_{4}$ in $(C\smallsetminus D)\cap B$. It remains to show that $u_{4}$ is in the corner for $BC$. Otherwise, $u_{4}$ is in the $C$-link. By the degree-condition of $(A,B)$ and since $K$ is empty, the vertex $u_{4}$ has a neighbour in $A\cap D$ besides $u_{3}$, and therefore the $A$-link contains a third vertex or edge besides $e$ and $u_{3}$. This contradicts Lemma 4.6. ∎
Proof of the (Crossing Lemma). (4.1).
We combine Lemma 4.6 and Lemma 4.8. ∎
5. Shifts: a tool for finding tetra-separations
This section provides a ‘shifting’ method for obtaining tetra-separations from genuine 4-separations, and from mixed-4-separations more generally. There are ‘small’ mixed-4-separations that fail to yield tetra-separations via the shifting method. We characterise these in 5.10.
5.1. Left-shifts and right-shifts
Definition 5.1 (Left-shift, right-shift).
Let $(A,B)$ be a mixed-separation+ in a graph. Let $L$ be the set of vertices in $A\cap B$ with at most one neighbour in $A\smallsetminus B$, and let $R$ be the set of vertices in $A\cap B$ with at most one neighbour in $B\smallsetminus A$. The left-shift of $(A,B)$ is $(A\smallsetminus L,B)$. The right-shift of $(A,B)$ is $(A,B\smallsetminus R)$.
In the following, we prove lemmas for the left-shift. All results have analogues for right-shifts by symmetry.
Observation 5.2.
The left-shift $(A^{\prime},B^{\prime})$ of a mixed-separation+ $(A,B)$ satisfies $A^{\prime}\smallsetminus B^{\prime}=A\smallsetminus B$ and $B^{\prime}\smallsetminus A^{\prime}\supseteq B\smallsetminus A$.∎
Lemma 5.3.
Let $(A,B)$ be a mixed-$k$-separation+ of a graph $G$ with left-shift $(A^{\prime},B^{\prime})$.
  1. (i)
    $(A^{\prime},B^{\prime})$ is a $-\ell$-separation+ with $\ell\leqslant k$. We have $\ell=k$ if every vertex in $S(A,B)$ has a neighbour in $A\smallsetminus B$.
  2. (ii)
    Every vertex in $S(A^{\prime},B^{\prime})$ has $\geqslant 2$ neighbours in $A^{\prime}\smallsetminus B^{\prime}$.
Proof.
(i). By definition, the left-shift $(A^{\prime},B^{\prime})$ is a mixed-separation+. The separator of $(A^{\prime},B^{\prime})$ is obtained from the separator of $(A,B)$ by replacing every vertex $v\in S(A,B)$ that has exactly one neighbour $u\in A\smallsetminus B$ with the edge $uv$ and by removing every vertex in $S(A,B)$ that has no neighbour in $A\smallsetminus B$. Hence the order of $(A^{\prime},B^{\prime})$ is at most that of $(A,B)$, with equality if every vertex in $S(A,B)$ has a neighbour in $A\smallsetminus B$.
(ii). Assume for a contradiction that there is a vertex $v\in S(A^{\prime},B^{\prime})$ with at most one neighbour in $A^{\prime}\smallsetminus B^{\prime}$. Since $A^{\prime}\smallsetminus B^{\prime}=A\smallsetminus B$ by 5.2, the vertex $v$ also has at most one neighbour in $A\smallsetminus B$. But then, by the definition of left-shift, we have $v\notin A^{\prime}$, contradicting $v\in S(A^{\prime},B^{\prime})$. ∎
Lemma 5.4.
Let $(A,B)$ be a mixed-4-separation+ in a 4-connected graph $G$. Then the following assertions are equivalent:
  1. (i)
    no two edges in $S(A,B)$ share an endvertex in $A\smallsetminus B$;
  2. (ii)
    $|A\smallsetminus B|\geqslant 2$ or $|S(A,B)\cap E(G)|\leqslant 1$.
Proof.
(i) $\Rightarrow$ (ii) is immediate.
(ii) $\Rightarrow$ (i). Assume that (i) fails, so there are two edges $ab_{1},ab_{2}\in S(A,B)$ with $a\in A\smallsetminus B$ and both $b_{i}\in B\smallsetminus A$. Let $X$ be the vertex-set that consists of $a$ together with all vertices in $S(A,B)$ and all endvertices in $B$ of edges in $S(A,B)$ other than the $ab_{i}$. Since $|A\smallsetminus B|\geqslant 2$ by (ii), there is $\alpha\in A\smallsetminus B$ distinct from $a$. If some $b_{i}$ is not in $X$, we let $\beta:=b_{i}$. Otherwise both $b_{i}$ are in $X$, each has exactly two neighbours in $A\smallsetminus B$, and $S(A,B)$ contains only edges. Since $b_{1}$ has degree four, it has a neighbour $\beta$ in $B\smallsetminus A$ that is distinct from $b_{2}$. In either case, $X$ is a 3-separator that separates $\alpha$ from $\beta$, contradicting that $G$ is 4-connected. ∎
Lemma 5.5.
Let $(A,B)$ be a mixed-4-separation+ in a 4-connected graph $G$ such that every vertex in $S(A,B)$ has $\geqslant 2$ neighbours in $A\smallsetminus B$. Then the following assertions are equivalent:
  1. (i)
    no two edges in $S(A,B)$ share an end in $A\smallsetminus B$;
  2. (ii)
    $(A,B)$ is not an atomic edge-cut with $|A|=1$;
  3. (iii)
    $|A\smallsetminus B|\geqslant 2$.
Proof.
(i) $\Rightarrow$ (ii) is immediate.
(ii) $\Rightarrow$ (iii). Assume $|A\smallsetminus B|\leqslant 1$. Since every vertex in $S(A,B)$ has $\geqslant 2$ neighbours in $A\smallsetminus B$, we get that $S(A,B)$ is an edge-cut. Since $S(A,B)$ is nonempty, $|A|=1$ follows.
(iii) $\Rightarrow$ (i). This follows from Lemma 5.4 (ii) $\Rightarrow$ (i). ∎
Lemma 5.6.
Let $(A,B)$ be a mixed-4-separation+ in a 4-connected graph $G$ with $A\smallsetminus B\neq\emptyset$ and such that no two edges in $S(A,B)$ share an end in $B\smallsetminus A$. Let $(A^{\prime},B^{\prime})$ denote the left-shift of $(A,B)$. Then the following assertions are equivalent:
  1. (i)
    $(A^{\prime},B^{\prime})$ satisfies the matching-condition;
  2. (ii)
    $|A\smallsetminus B|\geqslant 2$.
Proof.
By Lemma 5.3, every vertex in $S(A^{\prime},B^{\prime})$ has $\geqslant 2$ neighbours in $A^{\prime}\smallsetminus B^{\prime}$.
(i) $\Rightarrow$ (ii). Suppose $|A\smallsetminus B|\leqslant 1$. So $|A\smallsetminus B|=1$ since $A\smallsetminus B\neq\emptyset$. Then $|A^{\prime}\smallsetminus B^{\prime}|=1$ since $A\smallsetminus B=A^{\prime}\smallsetminus B^{\prime}$. Applying Lemma 5.5 to $(A^{\prime},B^{\prime})$ yields that $(A^{\prime},B^{\prime})$ is an atomic edge-cut.
(ii) $\Rightarrow$ (i). From $A\smallsetminus B=A^{\prime}\smallsetminus B^{\prime}$ we get $|A^{\prime}\smallsetminus B^{\prime}|\geqslant 2$. So no two edges in $S(A^{\prime},B^{\prime})$ share an end in $A^{\prime}\smallsetminus B^{\prime}$ by Lemma 5.5 applied to $(A^{\prime},B^{\prime})$. So suppose for a contradiction that there are two distinct edges $a_{1}b,a_{2}b$ in $S(A^{\prime},B^{\prime})$ with $a_{i}\in A^{\prime}\smallsetminus B^{\prime}$ and $b\in B^{\prime}\smallsetminus A^{\prime}$. Then $a_{i}\in A^{\prime}\smallsetminus B^{\prime}=A\smallsetminus B$ and $b\in B\cap A$. Hence $b$ has $\geqslant 2$ neighbours in $A\smallsetminus B$, contradicting that $b$ was replaced with edges when taking the left-shift. ∎
Lemma 5.7.
Let $G$ be a 4-connected graph. Let $(A,B)$ be a mixed-4-separation of $G$. Let $(C,D)$ be a 4-separation of $G$ with right-shift $(C^{\prime},D^{\prime})$. Assume $A\subseteq C$. Then $(A,B)\leqslant(C^{\prime},D^{\prime})$.
Proof.
We obtain $A\subseteq C^{\prime}$ from $A\subseteq C=C^{\prime}$. Let $X$ consist of all endvertices in $A$ of edges in $S(A,B)$. Then $N(B\smallsetminus A)=(A\cap B)\cup X$. From $A\subseteq C$ we obtain $D\smallsetminus C\subseteq B\smallsetminus A$. Hence $D\subseteq(B\smallsetminus A)\cup N(B\smallsetminus A)=B\cup X$. Recall that $D^{\prime}\subseteq D$. Let $v\in D^{\prime}$. It suffices to show that $v\notin X$. If $v\in X$, then $v$ has exactly one neighbour in $B\smallsetminus A$. Since $B\smallsetminus A\supseteq D\smallsetminus C$, the definition of right-shift yields $v\notin D^{\prime}$, a contradiction. So $v\notin X$. ∎
5.2. Left-right-shifts and right-left-shifts
Definition 5.8.
Let $(A,B)$ be a mixed-separation+. The left-right-shift of $(A,B)$ is obtained from $(A,B)$ by first taking the left-shift and then the right-shift. The right-left-shift of $(A,B)$ is obtained from $(A,B)$ by first taking the right-shift and then the left-shift.
Observation 5.9.
Let $(A^{\prime},B^{\prime})$ be the left-right-shift of a mixed-4-separation+ $(A,B)$ in a graph $G$. Then $A^{\prime}\smallsetminus B^{\prime}\supseteq A\smallsetminus B$ and $B^{\prime}\smallsetminus A^{\prime}\supseteq B\smallsetminus A$.
Proof.
We perform the two shifts one after the other, applying 5.2 after each shift. ∎
Key Lemma 5.10.
Let $(A,B)$ be a mixed-4-separation+ in a 4-connected graph $G$. The left-right-shift $(A^{\prime\prime},B^{\prime\prime})$ of $(A,B)$ is a tetra-separation if and only if the following three conditions are satisfied:
  1. (i)
    $|A\smallsetminus B|\geqslant 2$;
  2. (ii)
    the number of vertices in $S(A,B)$ with $\leqslant 1$ neighbour in $A\smallsetminus B$ is $\geqslant 2-|B\smallsetminus A|$;
  3. (iii)
    $|B\smallsetminus A|\geqslant 2$ or $|S(A,B)\cap E(G)|\leqslant 1$.
Proof.
Let $(A^{\prime},B^{\prime})$ denote the left-shift of $(A,B)$.
Backward implication. Assume (i)(iii). By (iii) and Lemma 5.4 (applied to $(B,A)$ instead of $(A,B)$), no two edges in $S(A,B)$ share an endvertex in $B\smallsetminus A$. By (i) and Lemma 5.6, $(A^{\prime},B^{\prime})$ satisfies the matching-condition. By Lemma 5.3, every vertex in $S(A^{\prime},B^{\prime})$ has $\geqslant 2$ neighbours in $A^{\prime}\smallsetminus B^{\prime}$. By (i) and since $G$ is 4-connected, every vertex in $S(A,B)$ has a neighbour in $A\smallsetminus B$. Hence $(A^{\prime},B^{\prime})$ has order four by Lemma 5.3. We further have $|A^{\prime}\smallsetminus B^{\prime}|=|A\smallsetminus B|\geqslant 2$ by 5.9 and we have $|B^{\prime}\smallsetminus A^{\prime}|\geqslant 2$ by (ii).
By Lemma 5.6 (applied for the right-shift), $(A^{\prime\prime},B^{\prime\prime})$ satisfies the matching-condition. By Lemma 5.3 (applied for the right-shift), every vertex in $S(A^{\prime\prime},B^{\prime\prime})$ has $\geqslant 2$ neighbours in $B^{\prime\prime}\smallsetminus A^{\prime\prime}$. Every vertex in $S(A^{\prime\prime},B^{\prime\prime})$ has $\geqslant 2$ neighbours in $A^{\prime\prime}\smallsetminus B^{\prime\prime}$ since $A^{\prime\prime}\smallsetminus B^{\prime\prime}=A^{\prime}\smallsetminus B^{\prime}$ by 5.9 (applied for the right-shift). Hence $(A^{\prime\prime},B^{\prime\prime})$ satisfies the degree-condition. Since $B^{\prime}\smallsetminus A^{\prime}$ is nonempty and $G$ is 4-connected, every vertex in $S(A^{\prime},B^{\prime})$ has a neighbour in $B^{\prime}\smallsetminus A^{\prime}$, so $(A^{\prime\prime},B^{\prime\prime})$ has order four by Lemma 5.3 (applied for the right-shift). Finally, $A^{\prime\prime}\smallsetminus B^{\prime\prime}=A^{\prime}\smallsetminus B^{\prime}$ and $B^{\prime\prime}\smallsetminus A^{\prime\prime}\supseteq B^{\prime}\smallsetminus A^{\prime}$ by 5.9, so $(A^{\prime\prime},B^{\prime\prime})$ is proper.
Forward implication. Suppose first that (iii) fails. Then, by Lemma 5.4 for $(B,A)$, there are two edges $ab_{1},ab_{2}$ in $S(A,B)$ with $a\in A\smallsetminus B$ and $b_{i}\in B\smallsetminus A$. Then $a\in A^{\prime\prime}\smallsetminus B^{\prime\prime}$ and $b_{i}\in B^{\prime\prime}\smallsetminus A^{\prime\prime}$ by 5.9. Hence $(A^{\prime\prime},B^{\prime\prime})$ fails the matching-condition.
Suppose next that (iii) holds, but (i) or (ii) fails. By (iii) and Lemma 5.4 for $(B,A)$, no two edges in $S(A^{\prime},B^{\prime})$ share an endvertex in $B^{\prime}\smallsetminus A^{\prime}$. If (i) fails, we have $|A\smallsetminus B|\leqslant 1$, so Lemma 5.6 gives that $(A^{\prime},B^{\prime})$ violates the matching-condition, and hence $(A^{\prime\prime},B^{\prime\prime})$ also violates the matching-condition. Next, assume that (i) holds, but (ii) fails. As $|A\smallsetminus B|\geqslant 2$, Lemma 5.6 gives that $(A^{\prime},B^{\prime})$ satisfies the matching-condition. The failure of (ii) yields $|B^{\prime}\smallsetminus A^{\prime}|\leqslant 1$. Hence applying Lemma 5.6 to $(B^{\prime},A^{\prime})$ (instead of $(A,B)$) yields that $(A^{\prime\prime},B^{\prime\prime})$ violates the matching-condition. ∎
Proof of Lemma 2.3..
(i). If $G$ is not quasi-5-connected, it has a 4-separation $(A,B)$ with $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$. Hence the right-left-shift of $(A,B)$ is a tetra-separation of $G$ by 5.10.
(ii). Let $(A,B)$ be a tetra-separation of $G$. Let $X:=A\smallsetminus B$ and $Y:=B\smallsetminus A$. Let $F$ denote the set of edges in $S(A,B)$. We have
$8\leqslant|G|=|X|+|Y|+4-|F|$
and therefore
$|F|\leqslant(\,|X|-2)+(\,|Y|-2).$
Both $X$ and $Y$ have size $\geqslant 2$ by Lemma 4.2 since $(A,B)$ is a tetra-separation. Hence both terms $|X|-2$ and $|Y|-2$ are $\geqslant 0$. We may therefore bipartition $F$ into classes $F_{X}$ and $F_{Y}$ of $\leqslant|X|-2$ and $\leqslant|Y|-2$ edges, respectively. We obtain $X^{\prime}$ from $X$ by adding for each edge in $F_{X}$ its endvertex in $A$ to $X$, and we obtain $Y^{\prime}$ similarly. Then $(X^{\prime},Y^{\prime})$ is a 4-separation of $G$. We have $|X^{\prime}\smallsetminus Y^{\prime}|\geqslant|X|-|F_{X}|\geqslant 2$, and similarly $|Y^{\prime}\smallsetminus X^{\prime}|\geqslant 2$. Thus $G$ is not quasi-5-connected. ∎
6. Understanding nestedness through connectivity
The purpose of this section is to characterise the property of total-nestedness for tetra-separations $(A,B)$ in terms of connectivity-properties of their separators $S(A,B)$; see 6.16. This translates the problem of proving total-nestedness of $(A,B)$ to a problem about constructing various families of paths that link up the vertices and edges in the separator $S(A,B)$ in different ways.
For an element $x\in V(G)\cup E(G)$ we let $\hat{x}$ denote the singleton $\{x\}$ if $x$ is a vertex, and the set of endvertices of $x$ if $x$ is an edge. For a set $X\subseteq V(G)\cup E(G)$ we write ${\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{\hat{X}}}:=\bigcup_{x\in X}\hat{x}$.
A bipartition of a set is balanced if its two classes have the same size.
Lemma 6.1.
Let $G$ be a $k$-connected graph. Then $G$ has no $-\ell$-separation for any $\ell<k$.
Proof.
Assume for a contradiction that $(A,B)$ is a $-\ell$-separation of $G$ with $\ell<k$. Then $S(A,B)$ contains an edge $ab$ with $a\in A\smallsetminus B$ and $b\in B\smallsetminus A$. We claim that $A\smallsetminus(B\cup\{a\})$ or $B\smallsetminus(A\cup\{b\})$ is non-empty. Indeed, otherwise $V(G)=(A\cap B)\cup\{a,b\}$, so $|G|=\ell+1\leqslant k$, contradicting that $k$-connected graphs such as $G$ have $>k$ vertices by definition. So suppose that there is a vertex $v\in A\smallsetminus(B\cup\{a\})$. Let $X$ be obtained from $S(A,B)$ by keeping all vertices, replacing $ab$ with $a$, and replacing each other edge with its endvertex in $B\smallsetminus A$. Then $X$ is an $\ell$-separator of $G$ that separates $v$ from $b$, contradicting that $G$ is $k$-connected. ∎
Let $X,Y$ be two vertex-sets in a graph $G$. A mixed-separation $(A,B)$ of $G$ strongly separates $X$ and $Y$ if $X\subseteq A\smallsetminus B$ and $Y\subseteq B\smallsetminus A$. Note that if $(A,B)$ strongly separates $X$ and $Y$, then $X$ and $Y$ are disjoint. The following lemma is folklore; we include a proof for convenience.
Lemma 6.2 (Menger for independent paths).
Let $X,Y$ be two disjoint vertex-sets in a graph $G$. Then the following two numbers are equal:
  1. (i)
    the maximum number of independent $X$$Y$ paths in $G$;
  2. (ii)
    the minimum order of a mixed-separation in $G$ that strongly separates $X$ and $Y$.
Moreover, if $G$ is $k$-connected, then (i), (ii)${}\geqslant k$.
Proof.
As usual, (i) $\leqslant$ (ii), so it remains to show (i) $\geqslant$ (ii). Since every $X$$Y$ edge forms a path in every maximum set of independent $X$$Y$ paths in $G$, we may assume without loss of generality that $G$ contains no $X$$Y$ edges. Let $X^{\prime}:=N_{G}(X)$ and $Y^{\prime}:=N_{G}(Y)$. Note that $X^{\prime}\cup Y^{\prime}$ is disjoint from $X\cup Y$. Let $G^{\prime}:=G-(X\cup Y)$.
[…]
The ‘moreover’-part follows from Lemma 6.1. ∎
6.1. Half-connected
A mixed-separation $(A,B)$ of a graph $G$ is half-connected if $G[A\smallsetminus B]$ or $G[B\smallsetminus A]$ is connected [Tridecomp].
Lemma 6.3.
Let $(A,B)$ be a tetra-separation of a 4-connected graph $G$. Then the following assertions are equivalent:
  1. (i)
    $(A,B)$ is half-connected;
  2. (ii)
    no tetra-separation of $G$ crosses $(A,B)$ so that all links are empty.
Proof.
$\neg$ (ii) $\Rightarrow$ $\neg$ (i). If a tetra-separation of $G$ crosses $(A,B)$ so that all links are empty, then all corners are nonempty. Hence, neither $G[A\smallsetminus B]$ nor $G[B\smallsetminus A]$ is connected.
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). Assume that $(A,B)$ is not half-connected; that is, $G[A\smallsetminus B]$ and $G[B\smallsetminus A]$ are not connected. Then $S(A,B)$ consists of vertices only. Let $C_{A}$ be a component of $G[A\smallsetminus B]$ and let $C_{B}$ be a component of $G[B\smallsetminus A]$. Let $C:=V(C_{A})\cup V(C_{B})\cup S(A,B)$ and $D:=V(G-C_{A}-C_{B})$. Then $(C,D)$ is a genuine 4-separation with $S(C,D)=S(A,B)$ that crosses $(A,B)$ with all links empty. To see that $(C,D)$ is a tetra-separation, note first that both $G[C\smallsetminus D]$ and $G[D\smallsetminus C]$ are disconnected. Since $G$ is 4-connected, every vertex in $S(A,B)$ has a neighbour in every component of $G-S(A,B)$. So $(C,D)$ satisfies the degree-condition, and it trivially satisfies the matching-condition. ∎
6.2. 3-linked
Definition 6.4 (3-linked).
Let $X=\{x_{1},x_{2},x_{3},x_{4}\}$ be the separator of a tetra-separation $(A,B)$ of a graph $G$. We say that $\{x_{1},x_{2}\}$ is 3-linked around $X$ if at least one of the following holds:
  1. (6.4.1)
    $x_{3}$ or $x_{4}$ is an edge;
  2. (6.4.2)
    $x_{1}$ or $x_{2}$ is a vertex and there is an $\hat{x}_{1}$$\hat{x}_{2}$ edge;
  3. (6.4.3)
    there are three independent $\hat{x}_{1}$$\hat{x}_{2}$ paths in $G-x_{3}-x_{4}$.
Similarly, we define that $\{x_{i},x_{j}\}$ is 3-linked around $X$ for other choices of distinct indices $i,j\in[4]$. We say that $(A,B)$ is 3-linked if every pair of distinct elements in $S(A,B)$ is 3-linked around $X$.
Lemma 6.5.
Let $X=\{x_{1},x_{2},x_{3},x_{4}\}$ be the separator of a tetra-separation $(A,B)$ in a 4-connected graph $G$. The following assertions are equivalent:
  1. (i)
    $\{x_{1},x_{2}\}$ is 3-linked around $X$;
  2. (ii)
    no tetra-separation of $G$ crosses $(A,B)$ so that $x_{1}$ and $x_{2}$ are in opposite links while the centre has size 2.
Proof.
$\neg$ (ii) $\Rightarrow$ $\neg$ (i). Assume that a tetra-separation $(C,D)$ crosses $(A,B)$ so that $x_{1}$ and $x_{2}$ are in opposite links while the centre has size 2. By the (Crossing Lemma). (4.1), all links have size 1 and there are no jumping edges, no diagonal edges and no dangling edges. So, $x_{3}$ and $x_{4}$ are both vertices, which shows that (6.4.1) does not hold. Since there are no dangling edges and no jumping edges, if there exists an $\hat{x}_{1}$$\hat{x}_{2}$ edge then both $x_{1}$ and $x_{2}$ are edges. This shows that (6.4.2) does not hold. Since there are no diagonal edges, no jumping edges and no dangling edges, every $\hat{x}_{1}$$\hat{x}_{2}$ path that avoids $x_{3}$ and $x_{4}$ has an inner vertex or an edge in the link for $A$ or the link for $B$. Hence, there are at most two independent $\hat{x}_{1}$$\hat{x}_{2}$ paths in $G-x_{3}-x_{4}$. This shows that (6.4.3) does not hold. So, $\{x_{1},x_{2}\}$ is not 3-linked around $X$.
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). Assume that $\{x_{1},x_{2}\}$ is not 3-linked around $X$. Then $x_{3}$ and $x_{4}$ are both vertices and there are at most two independent $\hat{x}_{1}$$\hat{x}_{2}$ paths in $G-x_{3}-x_{4}$. By (Menger for independent paths). (6.2), there is a mixed-2-separation $(C,D)$ of $G-x_{3}-x_{4}$ that strongly separates $\hat{x}_{1}$ from $\hat{x}_{2}$, say with $\hat{x}_{1}\subseteq C\smallsetminus D$ and $\hat{x}_{2}\subseteq D\smallsetminus C$. Let $(C^{\prime},D^{\prime}):=(C\cup\{x_{3},x_{4}\},D\cup\{x_{3},x_{4}\})$, which is a mixed-4-separation. Observe that $(C^{\prime},D^{\prime})$ crosses $(A,B)$ since the $C^{\prime}$-link and the $D^{\prime}$-link are non-empty.
We claim that both $C^{\prime}\smallsetminus D^{\prime}$ and $D^{\prime}\smallsetminus C^{\prime}$ have size $\geqslant 2$. Assume otherwise, say $x_{1}$ is a vertex and it is the only vertex in $C^{\prime}\smallsetminus D^{\prime}$. By the degree-condition for $x_{1}\in S(A,B)$, it has at least four neighbours in $V(G)\smallsetminus(A\cap B)$. This, however, contradicts that $(C,D)$ has order two.
Let $(C^{\prime\prime},D^{\prime\prime})$ be the left-right-shift of $(C^{\prime},D^{\prime})$, which is a tetra-separation by 5.10. First, note that $\hat{x}_{1}\subseteq C^{\prime}\smallsetminus D^{\prime}\subseteq C^{\prime\prime}\smallsetminus D^{\prime\prime}$ and $\hat{x}_{2}\subseteq D^{\prime}\smallsetminus C^{\prime}\subseteq D^{\prime\prime}\smallsetminus C^{\prime\prime}$ by 5.9. So, $x_{1}$ and $x_{2}$ are in the $C^{\prime\prime}$-link and the $D^{\prime\prime}$-link, respectively. In particular, $(C^{\prime\prime},D^{\prime\prime})$ crosses $(A,B)$.
We claim that $x_{3},x_{4}\in S(C^{\prime\prime},D^{\prime\prime})$. Assume not, say $x_{3}\in C^{\prime\prime}\smallsetminus D^{\prime\prime}$. Let $e\in S(C^{\prime\prime},D^{\prime\prime})$ be the edge that $x_{3}$ sends to $D^{\prime\prime}\smallsetminus C^{\prime\prime}$. By the (Crossing Lemma). (4.1), $e$ is not a jumping edge. So there is an $X\in\{A,B\}$ such that $e$ dangles from the $C^{\prime\prime}$-link through the $X$-link, say $X=A$. By the (Crossing Lemma). (4.1), the corner $AC^{\prime\prime}$ is potter. In particular, $x_{1}$ is an edge that dangles from the $A$-link through the $C^{\prime\prime}$-link. Then, however, $\hat{x}_{1}\nsubseteq C^{\prime\prime}\smallsetminus D^{\prime\prime}$, a contradiction. Hence, $(C^{\prime\prime},D^{\prime\prime})$ crosses $(A,B)$ so that $x_{1}$ and $x_{2}$ are in opposite links while the centre has size 2. ∎
6.3. Potter-linked
Let $(A,B)$ be a mixed-separation of a graph $G$. Let $\{x_{1},x_{2}\}\subseteq S(A,B)$ consist of an edge and a vertex, say $x_{1}$ is the edge and $x_{2}$ is the vertex. Write $x_{1}=ab$ with $a\in A\smallsetminus B$ and $b\in B\smallsetminus A$. In the context of $(A,B)$, we say that $\{x_{1},x_{2}\}$ is $A$-weird if $x_{2}$ has exactly one neighbour in $A\smallsetminus\{a\}$. If $(A,B)$ is a tetra-separation, this is equivalent to saying that $x_{2}$ has exactly two neighbours in $A$ and $a$ is one of them.
Example 6.6.
Let $(A,B)$ and $(C,D)$ be crossing tetra-separations in a graph $G$. Assume that the corner for $AC$ is potter, witnessed by an edge $u_{1}u_{2}$ that dangles from the vertex $u_{2}$ in the $C$-link through the $A$-link, and by an edge $u_{3}u_{4}$ that dangles from the vertex $u_{3}$ in the $A$-link through the $C$-link; see Figure 4.2. Then $\{u_{2},u_{3}u_{4}\}$ is $A$-weird and $\{u_{3},u_{1}u_{2}\}$ is $C$-weird.
Let $X,Y\subseteq V(G)\cup E(G)$. An $X$$Y$ edge means an edge $xy$ with $x\in X$ and $y\in Y$.
Definition 6.7 (0-potter-linked).
Let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of the separator of a tetra-separation $(A,B)$ in a graph $G$. We say that $\{\pi_{1},\pi_{2}\}$ is 0-potter-linked with respect to $(A,B)$ if at least one of the following assertions holds:
  1. (6.7.1)
    there is a $\pi_{1}$$\hat{\pi}_{2}$ edge or a $\hat{\pi}_{1}$$\pi_{2}$ edge;
  2. (6.7.2)
    there are five independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G$.
Note that this is symmetric in $\pi_{1}$ and $\pi_{2}$. A tetra-separation $(A,B)$ is 0-potter-linked if $\{\pi_{1},\pi_{2}\}$ is 0-potter-linked for every balanced bipartition $\{\pi_{1},\pi_{2}\}$ of $S(A,B)$.
Lemma 6.8.
Let $(A,B)$ be a tetra-separation of a 4-connected graph $G$. Let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of $S(A,B)$. Then the following assertions are equivalent:
  1. (i)
    $\{\pi_{1},\pi_{2}\}$ is 0-potter-linked;
  2. (ii)
    no tetra-separation of $G$ crosses $(A,B)$ so that $\pi_{1}$ and $\pi_{2}$ form opposite links and no corner is potter.
Proof.
$\neg$ (ii) $\Rightarrow$ $\neg$ (i). Assume that a tetra-separation $(C,D)$ of $G$ crosses $(A,B)$ so that $\pi_{1}$ and $\pi_{2}$ form opposite links and no corner is potter. Hence by the (Crossing Lemma). (4.1), no edge is dangling, jumping, or diagonal. It follows that if there exists a $\hat{\pi}_{1}$$\hat{\pi}_{2}$ edge $uv$, then both $u$ and $v$ are endvertices of edges in $S(A,B)$. This shows that (6.7.1) fails.
It remains to show that there are at most four $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G$. As no edge is dangling, jumping, or diagonal, every $\hat{\pi}_{1}$$\hat{\pi}_{2}$ path in $G$ must have an internal vertex or an edge in the $A$-link or in the $B$-link. The sum of the sizes of the $A$-link and the $B$-link is at most 4 since the size of $S(C,D)$ is at most 4. Hence (6.7.2) fails.
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). Assume that $\{\pi_{1},\pi_{2}\}$ is not 0-potter-linked. Then there are at most four independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G$. By (Menger for independent paths). (6.2), there is a mixed-4-separation $(C,D)$ of $G$ that strongly separates $\hat{\pi}_{1}$ from $\hat{\pi}_{2}$, say with $\hat{\pi}_{1}\subseteq C\smallsetminus D$ and $\hat{\pi}_{2}\subseteq D\smallsetminus C$. In particular, $|C\smallsetminus D|\geqslant 2$ and $|D\smallsetminus C|\geqslant 2$. Observe that $(C,D)$ crosses $(A,B)$ since the $C$-link and the $D$-link are non-empty. Let $(C^{\prime},D^{\prime})$ be the left-right-shift of $(C,D)$, which is a tetra-separation by 5.10.
[…]
The following definition is supported by Figure 6.1.
Definition 6.9 (1-potter-linked).
Let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of the separator of a tetra-separation $(A,B)$ of a graph $G$. Let $(X,Y)=(A,B)$ or $(X,Y)=(B,A)$. We say that $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $X$ with respect to $(A,B)$ if at least one of the following holds:
  1. (6.9.1)
    $\pi_{1}$ is not $X$-weird;
  2. (6.9.2)
    there is a $\pi_{1}$$(\hat{\pi}_{2}\cap Y)$ edge or a $(\hat{\pi}_{1}\cap Y)$$\pi_{2}$ edge;
  3. (6.9.3)
    there are three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[Y]$.
Note that this is not symmetric in $\pi_{1},\pi_{2}$. We say that $\{\pi_{1},\pi_{2}\}$ is 1-potter-linked with respect to $(A,B)$ if for both $X\in\{A,B\}$$\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $X$ and $\pi_{2}$ is 1-potter-linked to $\pi_{1}$ around $X$ with respect to $(A,B)$. A tetra-separation $(A,B)$ is 1-potter-linked if $\{\pi_{1},\pi_{2}\}$ is 1-potter-linked for every balanced bipartition $\{\pi_{1},\pi_{2}\}$ of $S(A,B)$.
Figure 6.1. $\pi_{1}$ is red and $X$-weird. $\pi_{2}$ is blue. Both pictures include how a mixed-separation $(C,D)$ might exploit the weirdness of $\pi_{1}$ to cross $(A,B)$. Left: the yellow edges satisfy (6.9.2). Right: the yellow paths satisfy (6.9.3).
Lemma 6.10.
Let $(A,B)$ be a tetra-separation of a 4-connected graph $G$ and let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of $S(A,B)$. Then the following assertions are equivalent:
  1. (i)
    $\{\pi_{1},\pi_{2}\}$ is 1-potter-linked with respect to $(A,B)$;
  2. (ii)
    no tetra-separation of $G$ crosses $(A,B)$ so that $\pi_{1}$ and $\pi_{2}$ form opposite links and exactly one corner is potter.
Proof.
$\neg$ (ii) $\Rightarrow$ $\neg$ (i). Assume that a tetra-separation $(C,D)$ crosses $(A,B)$ so that $\pi_{1}$ and $\pi_{2}$ form opposite links, say $\pi_{1}$ is the $C$-link and $\pi_{2}$ is the $D$-link, and so that exactly one corner is potter, say the $AC$-corner. We show that $\pi_{1}$ is not 1-potter-linked to $\pi_{2}$ with respect to $(A,B)$. For this, we show that for $(X,Y)=(A,B)$ in Definition 6.9, all of (6.9.1)(6.9.3) fail.
Observe that $\pi_{1}$ is $A$-weird by Example 6.6, so (6.9.1) fails. By assumption, the corner $AC$ is the only corner that is potter. By the (Crossing Lemma). (4.1), every link has size two while the centre is empty. Moreover, there do not exist jumping or diagonal edges, and the $B$-link contains no dangling edge. It follows that there is no $\pi_{1}$$(\hat{\pi}_{2}\cap B)$ edge and no $(\hat{\pi}_{1}\cap B)$$\pi_{2}$ edge, so (6.9.2) fails. Moreover, every $\hat{\pi}_{1}$$\hat{\pi}_{2}$ path in $G[B]$ has an internal vertex or an edge in the $B$-link. Thus, there are at most two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[B]$, so (6.9.3) fails.
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). Assume that $\{\pi_{1},\pi_{2}\}$ is not 1-potter-linked with respect to $(A,B)$. Without loss of generality, $\pi_{1}$ is not 1-potter-linked to $\pi_{2}$ and (6.9.1)(6.9.3) fail for $(X,Y)=(A,B)$. By $\neg$ (6.9.1), $\pi_{1}$ is $A$-weird. Write $\pi_{1}=:\{v_{1},e_{1}\}$ where $v_{1}$ is a vertex and $e_{1}=a_{1}b_{1}$ is an edge with $a_{1}\in A\smallsetminus B$ and $b_{1}\in B\smallsetminus A$. Let $a^{\prime}_{1}$ be the unique neighbour of $v_{1}$ in $A$ besides $a_{1}$. Note that $a^{\prime}_{1}$ lies in $A\smallsetminus B$ by the degree-condition for $v_{1}$.
By $\neg$ (6.9.3), there are at most two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[B]$. By (Menger for independent paths). (6.2), there is a mixed-separation $(C_{1},D_{1})$ of $G[B]$ of order $\leqslant 2$ that strongly separates $\hat{\pi}_{1}$ from $\hat{\pi}_{2}$, say with $\hat{\pi}_{1}\cap B\subseteq C_{1}\smallsetminus D_{1}$ and $\hat{\pi}_{2}\cap B\subseteq D_{1}\smallsetminus C_{1}$. In particular, $|C_{1}\smallsetminus D_{1}|\geqslant 2$ and $|D_{1}\smallsetminus C_{1}|\geqslant 2$. See Figure 6.2 for the next definition. We define $(C_{2},D_{2}):=(C_{1}\cup a_{1},(D_{1}\cup A)\smallsetminus\{v_{1}\})$. We have $S(C_{2},D_{2})=S(C_{1},D_{1})\cup\{a^{\prime}_{1}v_{1},a_{1}\}$ since $v_{1}$ does not have a neighbour in $A$ besides $a_{1}^{\prime}$. By 4-connectivity of $G$ and Lemma 6.1, $(C_{2},D_{2})$ has order exactly 4. Let $(C,D)$ be the left-right-shift of $(C_{2},D_{2})$. Since $C_{1}\smallsetminus D_{1}\subseteq C_{2}\smallsetminus D_{2}$ we have $|C_{2}\smallsetminus D_{2}|\geqslant 2$, and similarly $|D_{2}\smallsetminus C_{2}|\geqslant 2$. Hence $(C,D)$ is a tetra-separation by 5.10.
Consider the crossing-diagram for $(A,B)$ and $(C,D)$. We claim that the $C$-link includes $\pi_{1}$. The $C$-link clearly contains $v_{1}$. The vertex $a_{1}\in S(C_{2},D_{2})$ has two neighbours $v_{1},b_{1}$ in $C_{2}\smallsetminus D_{2}$, so $a_{1}\in S(C,D)$. Hence $e_{1}\notin S(C,D)$. Since the end $b_{1}$ of $e_{1}$ lies in $C_{1}\smallsetminus D_{1}\subseteq C\smallsetminus D$ by 5.9, the edge $e_{1}$ lies in the $C$-link. Next, we claim that the $D$-link includes $\pi_{2}$. Every vertex in $\pi_{2}$ clearly lies in the $D$-link. Now let $f$ be an edge in $\pi_{2}$. The endvertex of $f$ in $B\smallsetminus A$ lies in $\hat{\pi}_{2}\cap B\subseteq D_{1}\smallsetminus C_{1}\subseteq D\smallsetminus C$ by 5.9. The endvertex of $f$ in $A\smallsetminus B$ must be distinct from $a_{1}$ by the matching-condition, and hence is contained in $D_{2}\smallsetminus C_{2}\subseteq D\smallsetminus C$ by 5.9. Therefore, both ends of $f$ lie in $D\smallsetminus C$ while $f\in S(A,B)$, so $f$ lies in the $D$-link.
Since the $C$-link includes $\pi_{1}$ and the $D$-link includes $\pi_{2}$, the two tetra-separations $(A,B)$ and $(C,D)$ cross, and the inclusions are equalities by the (Crossing Lemma). (4.1). It remains to show that the $AC$-corner is the only corner that is potter. The edge $v_{1}a_{1}^{\prime}$ dangles from the $C$-link through the $A$-link, so the corner $AC$ is potter by the (Crossing Lemma). (4.1). By 4.7, the corners for $AD$ and $BC$ are not potter. Assume for contradiction that the $BD$-corner is potter. Then $\pi_{2}$ is $B$-weird and hence contains an edge $f$. But then the endvertex of $f$ in $B\smallsetminus A$ is contained in $D\smallsetminus C$ as shown above. In particular, the corner for $BD$ is not empty, contradicting that $BD$ is potter. ∎
Figure 6.2. The situation in the second half of the proof of Lemma 6.10. $\pi_{1}$ is red and $\pi_{2}$ is blue.
Definition 6.11 (2-potter-linked).
Let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of the separator of a tetra-separation $(A,B)$ of a graph $G$. We say that $\{\pi_{1},\pi_{2}\}$ is 2-potter-linked with respect to $(A,B)$ if for both $(X,Y)=(A,B)$ and $(X,Y)=(B,A)$, at least one of the following holds:
  1. (6.11.1)
    $\pi_{1}$ is not $X$-weird;
  2. (6.11.2)
    $\pi_{2}$ is not $Y$-weird.
A tetra-separation $(A,B)$ is 2-potter-linked if $\{\pi_{1},\pi_{2}\}$ is 2-potter-linked for every balanced bipartition $\{\pi_{1},\pi_{2}\}$ of $S(A,B)$.
Observation 6.12.
Let $(A,B)$ be a tetra-separation of a graph $G$. Let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of $S(A,B)$. Assume that $\pi_{1}$ is $A$-weird. Then there is no $\pi_{1}$$\pi_{2}$ edge.
Proof.
Assume that there is a $\pi_{1}$$\pi_{2}$ edge $uv$ with $u\in\pi_{1}$ and $v\in\pi_{2}$. Let $u^{\prime}$ be the endvertex in $A\smallsetminus B$ of the edge in $\pi_{1}$. Then $v$ is the only neighbour of $u$ in $A\smallsetminus\{u^{\prime}\}$ by definition of $A$-weird. Hence, $u\in S(A,B)$ has at most one neighbour in $A\smallsetminus B$, contradicting the degree-condition. ∎
Lemma 6.13.
Let $(A,B)$ be a tetra-separation of a 4-connected graph $G$ and let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of $S(A,B)$. Then the following assertions are equivalent:
  1. (i)
    $\{\pi_{1},\pi_{2}\}$ is 2-potter-linked with respect to $(A,B)$;
  2. (ii)
    no tetra-separation of $G$ crosses $(A,B)$ so that $\pi_{1}$ and $\pi_{2}$ form opposite links and exactly two corners are potter.
Proof.
$\neg$ (ii) $\Rightarrow$ $\neg$ (i). Assume that a tetra-separation $(C,D)$ crosses $(A,B)$ so that $\pi_{1}$ and $\pi_{2}$ form opposite links, say $\pi_{1}$ is the $C$-link and $\pi_{2}$ is the $D$-link, and so that exactly two corners are potter. By 4.7, no two adjacent corners are potter. So we can assume without loss of generality that the $AC$-corner and the $BD$-corner are potter while the other two corners are not potter. Then $\pi_{1}$ is $A$-weird and $\pi_{2}$ is $B$-weird by Example 6.6. So, (6.11.1) and (6.11.2) fail for $(X,Y)=(A,B)$.
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). Assume that $\{\pi_{1},\pi_{2}\}$ is not 2-potter-linked with respect to $(A,B)$. Say (6.11.1) and (6.11.2) fail for $(X,Y)=(A,B)$; that is to say, $\pi_{1}$ is $A$-weird and $\pi_{2}$ is $B$-weird. For $i\in\{1,2\}$, write $\pi_{i}=:\{v_{i},e_{i}\}$ where $v_{i}$ is a vertex and $e_{i}=a_{i}b_{i}$ is an edge with $a_{i}\in A\smallsetminus B$ and $b_{i}\in B\smallsetminus A$; see Figure 6.3. Let $a^{\prime}_{1}$ be the unique neighbour of $v_{1}$ in $A\smallsetminus\{a_{1}\}$, and let $b_{2}^{\prime}$ be the unique neighbour of $v_{2}$ in $B\smallsetminus\{b_{1}\}$. By the degree-condition, we have $a^{\prime}_{1}\in A\smallsetminus B$ and $b^{\prime}_{1}\in B\smallsetminus A$.
Let $C:=B\smallsetminus\{v_{2}\})\cup\{a_{1}\}$ and $D:=(A\smallsetminus\{v_{1}\})\cup\{b_{2}\}$. Since $\pi_{1}$ is $A$-weird, $\pi_{2}$ is $B$-weird, and $v_{1}v_{2}\notin E(G)$ by 6.12, we have $S(C,D)=\{a^{\prime}_{1}v_{1},a_{1},b_{2},v_{2}b^{\prime}_{2}\}$. Note that $(C,D)$ crosses $(A,B)$ since $v_{1}$ is in the $C$-link and $v_{2}$ is in the $D$-link. The vertex $a_{1}$ has two neighbours in $C\smallsetminus D$, namely $v_{1}$ and $b_{1}$. Since $a_{1}$ has degree at least four and it is not adjacent to $b_{2}$, the only other vertex in $S(C,D)$, it follows that $a_{1}$ also has at least two neighbours in $D\smallsetminus C$. Similarly, $b_{2}$ also has two neighbours in both $C\smallsetminus D$ and $D\smallsetminus C$. Hence $(C,D)$ satisfies the degree-condition. It satisfies the matching-condition since $v_{1}\neq v_{2}$ means that $a^{\prime}_{1}v_{1}$ and $v_{2}b^{\prime}_{2}$ do not share endvertices. Hence, $(C,D)$ is a tetra-separation that crosses $(A,B)$ with $\pi_{1}$ and $\pi_{2}$ as opposite links. The corners $AC$ and $BD$ are potter by construction. By 4.7, no other corner is potter. ∎
Figure 6.3. The situation in the second half of the proof of Lemma 6.13. $\pi_{1}$ is red and $\pi_{2}$ is blue.
Corollary 6.14.
Let $(A,B)$ be a tetra-separation of a 4-connected graph $G$ and let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of $S(A,B)$. Then the following assertions are equivalent:
  1. (i)
    $\{\pi_{1},\pi_{2}\}$ is $h$-potter-linked with respect to $(A,B)$ for every $h\in\{0,1,2\}$;
  2. (ii)
    no tetra-separation of $G$ crosses $(A,B)$ so that $\pi_{1}$ and $\pi_{2}$ form opposite links.
Proof.
By 4.7, every crossing pair of tetra-separations has at most two corners that are potter. So the corollary immediately follows from Lemma 6.8, Lemma 6.10 and Lemma 6.13. ∎
Summary.
Definition 6.15.
A tetra-separation $(A,B)$ is externally 5-connected if it is half-connected, 3-linked and $h$-potter-linked for every $h\in\{0,1,2\}$.
Key Lemma 6.16.
For every tetra-separation $(A,B)$ of a 4-connected graph $G$, the following assertions are equivalent:
  1. (i)
    $(A,B)$ is totally-nested;
  2. (ii)
    $(A,B)$ is externally 5-connected.
Proof.
We combine Lemma 6.3, Lemma 6.5 and Corollary 6.14 with the (Crossing Lemma). (4.1). ∎
7. Thickened $K_{4,m}$’s
This section deals with the thickened $K_{4,m}$ outcome of Theorem 1; see 7.8.
A sprinkled $K_{4,m}$ is a graph obtained from a $K_{4,m}$ with a side $X$ of size four by possibly adding edges between vertices in $X$. We refer to $X$ as a sprinkled side.
Let $Z\subseteq V(G)$ and $K$ a component of $G-Z$. When $Z$ and $K$ are clear from context, we write $(U^{\prime}_{K},W^{\prime}_{K}):=(V(K)\cup Z,V(G-K))$. Let $Z$ be a set of four vertices in a 4-connected graph $G$ such that $G-Z$ has $\geqslant 4$ components. For each component $K$ of $G-Z$, let $(U_{K},W_{K})$ denote the left-right-shift of $(U^{\prime}_{K},W^{\prime}_{K})$.
Lemma 7.1.
Let $Z$ be a set of four vertices in a 4-connected graph $G$, such that $G-Z$ has $\geqslant 4$ components. Let $K$ be a component of $G-Z$. Then $(U_{K},W_{K})$ equals the left-shift of $(U^{\prime}_{K},W^{\prime}_{K})$. In particular, $(U_{K},W_{K})\leqslant(U^{\prime}_{K},W^{\prime}_{K})$ and $W_{K}=W^{\prime}_{K}$.
Proof.
Let $(A,B)$ denote the left-shift of $(U^{\prime}_{K},W^{\prime}_{K})$. Since every vertex in $S(U^{\prime}_{K},W^{\prime}_{K})$ has $\geqslant 2$ neighbours in $W^{\prime}_{K}\smallsetminus U^{\prime}_{K}$, and since $W^{\prime}_{K}\smallsetminus U^{\prime}_{K}\subseteq B\smallsetminus A$ by 5.2, it follows that every vertex in $S(A,B)$ has $\geqslant 2$ neighbours in $B\smallsetminus A$. Hence $(A,B)=(U_{K},W_{K})$ and $W^{\prime}_{K}=B=W_{K}$. ∎
Lemma 7.2.
Let $Z$ be a set of four vertices in a 4-connected graph $G$ such that $G-Z$ has $\geqslant 4$ components. Let $K$ be a component of $G-Z$. Then the following assertions are equivalent:
  1. (i)
    $K$ has $\geqslant 2$ vertices;
  2. (ii)
    the left-right-shift of $(U^{\prime}_{K},W^{\prime}_{K})$ is a totally-nested tetra-separation.
Proof.
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). If $K$ has only one vertex, then $|U_{K}\smallsetminus W_{K}|=1$, so $(U^{\prime}_{K},W^{\prime}_{K})$ fails to be a tetra-separation by 5.10.
(i) $\Rightarrow$ (ii). We abbreviate $(U,W):=(U_{K},W_{K})$ and $(U^{\prime},W^{\prime}):=(U^{\prime}_{K},W^{\prime}_{K})$. Since $K$ has $\geqslant 2$ vertices, $(U,W)$ is a tetra-separation by 5.10. To show that $(U,W)$ is totally-nested, by 6.16 it suffices to show that $(U,W)$ is externally 5-connected.
Half-connected: We have $G[U\smallsetminus W]=K$ by Lemma 7.1.
To check the remaining aspects of external 5-connectivity, for each $z\in Z$ we denote by $z^{*}$ the element of $S(U,W)$ that equal or incident to $z$. By Lemma 7.1, $z^{*}$ is either equal to $z$ or $z^{*}$ is an edge between $K$ and $z$. Denote by $z^{U}$ the vertex in $\hat{z}\cap U$. We can use the three components of $G-Z$ besides $K$ to find paths:
  1. ($\dagger$)
    There are three independent $z_{1}$$z_{2}$ paths in $G$ with all internal vertices outside $K$, for every two distinct vertices $z_{1},z_{2}\in Z$.
3-linked: follows from ($\dagger$)Section 7.
0-potter-linked: Let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of $S(U,W)$. We find three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths through $W\smallsetminus U$ by ($\dagger$)Section 7. We claim that there are two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths through $K$. Let $\{\pi^{\prime}_{1},\pi^{\prime}_{2}\}$ denote the bipartition of $Z$ that corresponds to $\{\pi_{1},\pi_{2}\}$. For both $i$, let $\pi_{i}^{K}$ denote the set of neighbours of $\pi^{\prime}_{i}$ in $K$. Then $|\pi_{i}^{K}|\geqslant 2$, since otherwise $\pi_{i}^{K}\cup\pi^{\prime}_{3-i}$ is a 3-separator of $G$ as $|K|\geqslant 2$, contradicting 4-connectivity. So it suffices to find two disjoint $\pi_{1}^{K}$$\pi_{2}^{K}$ paths in $K$. If these are missing, then by Menger’s theorem $\pi_{1}^{K}$ and $\pi_{2}^{K}$ are separated by a single vertex $v$, and then $\pi^{\prime}_{1}\cup\{v\}$ is a 3-separator of $G$, contradicting 4-connectivity.
1-potter-linked: Let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of $S(U,W)$. Assume that $\pi_{1}$ is weird, say. Then $\pi_{1}$ is $U$-weird since each vertex in $Z$ has $\geqslant 3$ neighbours in $W\smallsetminus U$. By ($\dagger$)Section 7, there exist three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[W]$. Hence $\{\pi_{1},\pi_{2}\}$ is 1-potter-linked with respect to $(U,W)$.
2-potter-linked: Let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of $S(U,W)$. No $\pi_{i}$ is $W$-weird, since each vertex in $Z$ has $\geqslant 3$ neighbours in $W\smallsetminus U$. So, (6.11.1) or (6.11.2) always holds. ∎
We write $)𝜎𝑍{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{\sigma(Z)}}$ for the star $\{\,(U_{K},W_{K}):K$ is a component of $G-Z$ with $|K|\geqslant 2\,\}$.
Corollary 7.3.
Let $Z$ be a set of four vertices in a 4-connected graph $G$ such that $G-Z$ has $\geqslant 4$ components. Then $\sigma(Z)$ is a star of totally-nested tetra-separations.
Proof.
By Lemma 7.2, every element in $\sigma(Z)$ is a totally-nested tetra-separation. Let $K\neq L$ be components of $G-Z$, each of size $\geqslant 2$. By Lemma 7.1, $(U_{K},W_{K})\leqslant(U_{K}^{\prime},W_{K}^{\prime})$ and $(U_{L},W_{L})\leqslant(U_{L}^{\prime},W_{L}^{\prime})$. With $(U_{K}^{\prime},W_{K}^{\prime})\leqslant(W_{L}^{\prime},U_{L}^{\prime})$, it follows that $(U_{K},W_{K})\leqslant(W_{L},U_{L})$. So, $\sigma(Z)$ is a star. ∎
Lemma 7.4.
Let $(A,B)$ be a 4-separation in a 4-connected graph $G$ such that $|A\smallsetminus B|\geqslant 2$ and every vertex in $S(A,B)$ has $\geqslant 2$ neighbours in $B\smallsetminus A$. Let $(A^{\prime\prime},B^{\prime\prime})$ be the left-right-shift of $(A,B)$. Then no tetra-separation $(C,D)$ of $G$ satisfies $(A^{\prime\prime},B^{\prime\prime})<(C,D)\leqslant(A,B)$.
Proof.
Suppose for a contradiction that there is a tetra-separation $(C,D)$ with $(A^{\prime\prime},B^{\prime\prime})<(C,D)\leqslant(A,B)$. By 5.10, $(A^{\prime\prime},B^{\prime\prime})$ is a tetra-separation. By Lemma 7.1, we have $B=B^{\prime\prime}$. So both inclusions $B^{\prime\prime}\supseteq D\supseteq B$ are actually equalities. We therefore have $A^{\prime\prime}\subsetneq C\subseteq A$. Let $v\in C\smallsetminus A^{\prime\prime}$. Then $v$ is a vertex in $S(A,B)$ with at most one neighbour in $A\smallsetminus B$. But since $D=B$ we have $C\smallsetminus D\subseteq A\smallsetminus B$, so $v$ also has at most one neighbour in $C\smallsetminus D$, contradicting that $(C,D)$ is a tetra-separation. ∎
Lemma 7.5.
Let $G$ be a 4-connected graph. Let $(A,B)$ be a mixed-4-separation in $G$ that is not half-connected. Then $(A,B)$ is a tetra-separation.
Proof.
By assumption, both $G[A\smallsetminus B]$ and $G[B\smallsetminus A]$ have two components. In particular, $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$. Then by Lemma 5.4$(A,B)$ satisfies the matching-condition. Let $v$ be a vertex in $S(A,B)$. By 4-connectivity of $G$ and by Lemma 6.1, $v$ has a neighbour in every component of $G-S(A,B)$. In particular, $v$ satisfies the degree-condition. Hence, $(A,B)$ is a tetra-separation. ∎
Lemma 7.6.
Let $Z$ be a set of four vertices in a 4-connected graph $G$ such that $G-Z$ has $\geqslant 4$ components. Let $(E,F)$ be a totally-nested tetra-separation of $G$. Then there exists a component $K$ of $G-Z$ with $|K|\geqslant 2$ such that $(E,F)\leqslant(U^{\prime}_{K},W^{\prime}_{K})$ or $(F,E)\leqslant(U^{\prime}_{K},W^{\prime}_{K})$.
Proof.
We denote by $\mathcal{K}$ the set of components of $G-Z$. First, we show that there is a component $K\in\mathcal{K}$ such that either $E\smallsetminus Z\subseteq V(K)$ or $F\smallsetminus Z\subseteq V(K)$. Assume otherwise. Then we can bipartition $\mathcal{K}$ into two sets $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ such that for both $i\in\{1,2\}$, $|\mathcal{K}_{i}|\geqslant 2$ and $\bigcup_{K\in\mathcal{K}_{i}}K$ intersects both $E$ and $F$. We define $A_{i}:=Z\cup\bigcup_{K\in\mathcal{K}_{i}}V(K)$. Then $(A_{1},A_{2})$ is a tetra-separation by Lemma 7.5. Now $E\nsubseteq A_{i}$ and $F\nsubseteq A_{i}$ for both $i\in\{1,2\}$ implies that $(A_{1},A_{2})$ crosses $(E,F)$, contradicting that $(E,F)$ is totally-nested.
Hence we have $E\subseteq V(K)\cup Z$, say, for a component $K\in\mathcal{K}$. Since every vertex in $Z$ has $\geqslant 2$ neighbours outside $V(K)\cup Z$, it follows that $(U^{\prime}_{K},W^{\prime}_{K})$ equals its right-shift. Hence $(E,F)\leqslant(U^{\prime}_{K},W^{\prime}_{K})$ by Lemma 5.7. Since $(E,F)$ is a tetra-separation, we have $|E\smallsetminus F|\geqslant 2$ by Lemma 5.5. Therefore, $|K|\geqslant 2$. ∎
Key Lemma 7.7.
Let $G$ be a 4-connected graph, and let $N$ denote the set of totally-nested tetra-separations of $G$. Assume that two tetra-separations $(A,B)$ and $(C,D)$ of $G$ cross so that the vertex-centre $Z$ has size four. Let $m$ denote the number of trivial components of $G-Z$. Then:
  1. (i)
    $\sigma(Z)$ is a splitting star of $N$.
  2. (ii)
    $\sigma(Z)$ is interlaced by both $(A,B)$ and $(C,D)$.
  3. (iii)
    If $\sigma(Z)\neq\emptyset$, then $G-Z$ has a component with $\geqslant 2$ vertices, and the torso of $\sigma(Z)$ is a thickened $K_{4,m}$ where $Z$ is the left side.
  4. (iv)
    If $\sigma(Z)=\emptyset$, then every component of $G-Z$ is trivial, and $G=\tau$ is a sprinkled $K_{4,m}$ with $Z$ as left side and $m\geqslant 4$.
Proof.
(i). By Corollary 7.3, $\sigma(Z)$ is a star in $N$. To show that $\sigma(Z)$ is splitting, assume for a contradiction that it is interlaced by a totally-nested tetra-separation $(E,F)$. By Lemma 7.6, there is a component $K$ of $G-Z$ with $|K|\geqslant 2$ and $(E,F)\leqslant(U^{\prime}_{K},W^{\prime}_{K})$, say. Since $K$ has at least two vertices, we have $(U_{K},W_{K})\in\sigma(Z)$. Note that $(U_{K},W_{K})\leqslant(U^{\prime}_{K},W^{\prime}_{K})$ by Lemma 7.1, which implies $(U_{K},W_{K})\not<(F,E)$. By Lemma 7.4, we cannot have $(U_{K},W_{K})<(E,F)$. Since $(U_{K},W_{K})\in N$, we therefore have $(E,F)\leqslant(U_{K},W_{K})$ or $(F,E)\leqslant(U_{K},W_{K})$, contradicting that $(E,F)$ interlaces $\sigma(Z)$.
(ii). We denote by $\sigma^{\prime}(Z)$ the star $\{\,(U^{\prime}_{K},W^{\prime}_{K}):K$ is a component of $G-Z$ with $|K|\geqslant 2\,\}$. Both $(A,B)$ and $(C,D)$ interlace $\sigma^{\prime}(Z)$, hence they also interlace $\sigma(Z)$ by Lemma 7.1.
(iii) and (iv) are straightforward. ∎
A graph $G$ is 4-angry if it is 4-connected and every tetra-separation of $G$ is crossed by a tetra-separation.
Key Lemma 7.8.
Let $G$ be a 4-angry graph. Then the following assertions are equivalent:
  1. (i)
    two tetra-separations of $G$ cross with all links empty;
  2. (ii)
    $G$ is a sprinkled $K_{4,m}$ with $m\geqslant 4$.
Proof.
(ii) $\Rightarrow$ (i). Let $X$ be a sprinkled side of $G$. Let $\{P_{1},\ldots,P_{4}\}$ be a partition of the other side $V(G-X)$. Let $(A,B)$ be the mixed-4-separation $(P_{1}\cup P_{2}\cup X,X\cup P_{3}\cup P_{4})$ and let $(C,D)$ be the mixed-4-separation $(P_{1}\cup P_{3}\cup X,X\cup P_{2}\cup P_{4})$. Then $(A,B)$ and $(C,D)$ are tetra-separations with separators equal to $X$, and they cross.
(i) $\Rightarrow$ (ii) follows from 7.7 (iii) and (iv). ∎
8. Generalised double-wheels
To deal with the generalised double-wheel outcome of Theorem 1, we first introduce a particular cycle-decomposition in Section 8.1 for analysing the wheel-structure obtained from two tetra-separations that cross with vertex-centre of size two. Section 8.2 introduces machinery that exploits the structure of this cycle-decomposition to find independent paths. Section 8.3 employs this machinery to show that all non-trivial bags of the cycle-decomposition give rise to totally-nested tetra-separations (8.27). Section 8.4 shows that these totally-nested tetra-separations form a splitting star as in Theorem 1 and deduces that its torso is a generalised double-wheel (8.37). Finally, Section 8.5 characterises the special case where additionally all tetra-separations of the graph are crossed (8.41).
8.1. Block-bagel
Let $G$ be a 2-connected graph. A Block-bagel of $G$ is a cycle-decomposition $\mathcal{O}=(O,\mathcal{G})$ of $G$ with $\mathcal{G}=(G_{t}:t\in V(O))$ such that
  1. (i)
    every bag $G_{t}$ of $\mathcal{O}$ is either 2-connected or a $K_{2}$, and
  2. (ii)
    all adhesion sets have size one.
The vertex in the adhesion-set of an edge $e$ of $O$ is the adhesion-vertex of $e$. When $\mathcal{O}$ as introduced as $(O,\mathcal{G})$, we tacitly assume that $\mathcal{G}=(G_{t}:t\in V(O))$.
Observation 8.1.
Let $G$ be a 2-connected graph with a block-bagel $\mathcal{O}=(O,\mathcal{G})$. Then every two edges of $O$ have distinct adhesion-vertices.
Proof.
Otherwise, two edges $rs$ and $st$ have the same adhesion-vertex $v$. But then $v$ is a cutvertex of $G$ that separates $V(G_{s}-v)$ from $V(G-G_{s})$. ∎
Assume now that $G$ is a 4-connected graph with two tetra-separations $(A,B)$ and $(C,D)$ that cross with all links of size one. Then the centre $Z^{\prime}$ consists of two vertices by the (Crossing Lemma). (4.1). We will construct a block-bagel of $G-Z^{\prime}$. The $AC$-shrimp is obtained from $G[A\cap C]-Z^{\prime}$ by adding any edges in the $A$-link or the $C$-link (including the endvertices of these edges). Note that the $AC$-shrimp is connected because $G-Z^{\prime}$ is 2-connected. The $AC$-block-decomposition means the block-decomposition of the $AC$-shrimp. Note that the $AC$-block-decomposition is a path-decomposition because $G$ is 4-connected. Its decomposition-path is called the $AC$-block-path. We obtain similarly definitions for the other three corners.
Lemma 8.2.
Let $G$ be a 4-connected graph with two tetra-separations $(A,B)$ and $(C,D)$ that cross with all links of size one. Let $K$ be one of the four corners, and consider an adjacent $L$-link $\Lambda$.
  1. (i)
    If $\Lambda$ consists of a vertex, then this vertex is contained in a unique bag $V_{t}$ of the $K$-block-decomposition, and $G[V_{t}]$ is 2-connected.
  2. (ii)
    If $\Lambda$ consists of an edge $e=uv$, then there is a unique bag $V_{t}$ of the $K$-block-decomposition with $V_{t}=\{u,v\}$ and $G[V_{t}]=K_{2}$. Moreover, $u$ and $v$ lie in the two corners adjacent to the $L$-link.
Proof.
(i). The bag $V_{t}$ is unique since $G$ is 4-connected. It follows from the degree-condition that $G[V_{t}]$ must be 2-connected.
(ii). The first part is immediate. For the ‘moreover’-part we use that $e$ is not dangling by the (Crossing Lemma). (4.1). ∎
Corollary 8.3.
Let $G$ be a 4-connected graph with two tetra-separations $(A,B)$ and $(C,D)$ that cross with all links of size one. Then all four corners are non-empty.∎
The bag $V_{t}$ and node $t$ in Lemma 8.2 are said to represent the $L$-link in the $K$-block-decomposition and in the $K$-block-path.
For each corner $K$ let $\mathcal{T}^{K}=(T^{K},\mathcal{V}^{K})$ denote the $K$-block-decomposition. We obtain $O$ from the disjoint union $\bigsqcup_{K}T^{K}$, where $K$ ranges over the four corners, by doing the following for each link $L$: for both corners $K$ and $K^{\prime}$ adjacent to $L$ we consider the nodes $t$ and $t^{\prime}$ that represent the $L$-link in $T^{K}$ and $T^{K^{\prime}}$, respectively, and
  • $\cdot$
    we join the nodes $t$ and $t^{\prime}$ by an edge $)𝑜𝐿{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{o(L)}}$ in $O$ if their bags $G[V_{t}]$ and $G[V_{t^{\prime}}]$ are 2-connected;
  • $\cdot$
    we otherwise identify the nodes $t$ and $t^{\prime}$ to a single node $)𝑜𝐿{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{o(L)}}$ of $O$.
We say that $o(L)$ represents the $L$-link in $O$. For each node $t$ of $O$ that is not of the form $o(L)$, we define $G_{t}:=G[V^{K}_{t}]$ where $K$ is the unique corner with $t\in T^{K}$. For each node $t$ of $O$ with $t=o(L)$ for some link $L$, we let $G_{t}$ be the $K_{2}$-subgraph of $G$ from Lemma 8.2 (ii) that represents the $L$-link in the $K$-block-decompositions for the two corners $K$ adjacent to $L$. The pair $(O,(G_{t})_{t\in V(O)})$ is the block-bagel induced by $(A,B)$ and $(C,D)$. It really is a block-bagel of $G-Z^{\prime}$.
The following setting is supported by Figure 8.1.
Setting 8.4 (Block-bagel setting).
Let $(A,B)$ and $(C,D)$ be two tetra-separations of a 4-connected graph $G$ that cross with all links of size one. Let $Z^{\prime}$ be the centre, which consists of two vertices $z^{\prime}_{1},z^{\prime}_{2}$. Let $\mathcal{O}=(O,(G_{t})_{t\in V(O)})$ be the Block-bagel induced by $(A,B)$ and $(C,D)$. Write $O=t(0),t(1),\ldots,t(m),t(0)$ where indices are treated as elements of $\mathbb{Z}_{m+1}$. For every node $t\in V(O)$, we let $U_{t}^{\prime}:=V(G_{t})\cup Z^{\prime}$ and $W_{t}^{\prime}:=Z^{\prime}\cup\bigcup_{s\in V(O-t)}V(G_{s})$, and we let $(U_{t},W_{t})$ be the right-left-shift of $(U_{t}^{\prime},W_{t}^{\prime})$.
Note that $(U_{t}^{\prime},W_{t}^{\prime})$ is a mixed-4-separation+.
Figure 8.1. The (Block-bagel setting). (8.4). The vertices $z^{\prime}_{1}$ and $z^{\prime}_{2}$ are in the centre. Left: the arrangement of the bags. Center: The 4-separation+ $(U^{\prime},W^{\prime})$. The vertex $x^{\prime}$ is red, the vertex $y^{\prime}$ is green. Right: The right-left-shift $(U,W)$ of $(U^{\prime},W^{\prime})$. Here $x$ is the red edge and $y$ is the green vertex. One of the $z_{i}$ is a vertex, the other is the blue edge.
8.2. N-Paths lemmata
Definition 8.5 (Helpful).
Let $\mathcal{O}=(O,\mathcal{G})$ be a block-bagel of a graph $G$. An element $x\in V(O)\cup E(O)$ is helpful if
  • $\cdot$
    $x$ is an edge with $K_{2}$-bags at both ends, or
  • $\cdot$
    $x$ is a node with a 2-connected bag.
We define ${\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{h(\mathcal{O})}}$ to be the number of helpful elements of the decomposition-cycle $O$.
For every helpful $x$ there is a unique corner $K$ such that $x$ stems from the $K$-block-path. We then say that $x$ belongs to the corner $K$.
Lemma 8.6 (Helpful Lemma).
Assume the (Block-bagel setting). (8.4). Then for every corner there is a helpful element that belongs to that corner. In particular, $h(\mathcal{O})\geqslant 4$.
Proof.
Let $K$ be an arbitrary corner. If the $K$-block-decomposition has a 2-connected bag, then we are done. Otherwise, the $K$-shrimp is a path. The corner $K$ contains a vertex $v$ by Corollary 8.3. Then $v$ is the adhesion-vertex of a helpful edge that belongs to the corner $K$. ∎
Lemma 8.7.
Assume the  (Block-bagel setting). (8.4). The adhesion-vertex of a helpful edge of $O$ sends edges to both vertices $z^{\prime}_{1},z^{\prime}_{2}$ in the centre.
Proof.
Otherwise the adhesion-vertex has degree $\leqslant 3$, contradicting 4-connectivity. ∎
The following two lemmata are supported by Figure 8.2.
Lemma 8.8.
Assume the  (Block-bagel setting). (8.4). Assume that $t$ is a helpful node of $O$. Then each vertex $z^{\prime}_{i}$ in the centre has a neighbour $w_{i}$ in the interior of the bag $G_{t}$. Moreover, if $|G_{t}|\geqslant 4$, then we can find the $w_{i}$ so that $w_{1}\neq w_{2}$.
Proof.
Since $t$ is helpful, its bag $G_{t}$ is 2-connected. Hence $G_{t}$ has non-empty interior. Each $z^{\prime}_{i}$ in the centre has a neighbour $w_{i}$ in this interior, since otherwise $z^{\prime}_{3-i}$ together with the two adhesion-vertices of $G_{t}$ forms a 3-separator of $G$, contradicting 4-connectivity.
Assume now that $G_{t_{1}}$ contains at least four vertices. If $w_{1}=w_{2}$ is the only possibility, then $w_{1}$ together with the two adhesion-vertices of $G_{t}$ forms a 3-separator, contradicting 4-connectivity. ∎
Figure 8.2. The neighbours $w_{i}$ found by Lemma 8.8 are blue. The paths found by Lemma 8.9 are red.
Lemma 8.9 (Bag-Path Lemma).
Assume the  (Block-bagel setting). (8.4). Let $t$ be a helpful node of $O$. Then there is a $z^{\prime}_{1}$$z^{\prime}_{2}$ path in $G-z^{\prime}_{1}z^{\prime}_{2}$ with interior in $\mathring{G}_{t}$.
Proof.
Otherwise the two adhesion-vertices of $G_{t}$ together with one of the $z^{\prime}_{i}$ forms a 3-separator of $G$, contradicting 4-connectivity. ∎
Lemma 8.10 (One-Path Lemma).
Assume the  (Block-bagel setting). (8.4). Let $P\subseteq O$ be a path of length $\geqslant 2$ with end-nodes $r$ and $s$. Let $u\in V(G_{r})$ and $v\in V(G_{s})$. Then there is a $u$$v$ path in $\bigcup_{t\in V(P)}G_{t}$ that internally avoids the adhesion-vertices of the edges not in $P$.
Proof.
Let $u_{i}$ denote the adhesion-vertex of $t(i)\,t(i+1)$. We assume without loss of generality that $P=t(1)\,t(2)\ldots t(k)$. For each $i\in\{1,2,\ldots,k\}$, we construct a path $P_{i}$ in $G_{t(i)}$, as follows. For $i=1$ we claim that $P_{i}$ can be chosen to be a $u$$u_{1}$ path with interior in $G_{t(1)}-u_{0}$. This is clearly possible if the bag $G_{t(1)}$ is a $K_{2}$. Otherwise $G_{t(1)}$ is 2-connected. Then we find two independent $u$$u_{1}$ paths in $G_{t(1)}$, and at least one of them internally avoids $u_{0}$. Similarly, if $i=k$ then we let $P_{i}$ be a $u_{k-1}$$v$ path with interior in $G_{t(k)}-u_{k}$. Otherwise, if $2\leqslant i\leqslant k$, then we let $P_{i}$ be a $u_{i-1}$$u_{i}$ path in $G_{t(i)}$. Then the concatenation $P:=P_{1}P_{2}\ldots P_{k}$ is a desired $u$$v$ path. ∎
Lemma 8.11 (Two-Fan Lemma).
Assume the  (Block-bagel setting). (8.4). Let $u$ and $v$ be the two adhesion-vertices in a bag $G_{t}$. Assume that either $t$ is helpful or $u$ is the adhesion-vertex of a helpful edge. Then there exist two independent paths $P_{1},P_{2}$ with all internal vertices in $\mathring{G}_{t}$ such that $P_{1}$ is a $u$$z_{1}^{\prime}$ path and $P_{2}$ is a $u$$v$ path.
Proof.
If $u$ is the adhesion-vertex of a helpful edge, then $uz_{1}^{\prime}\in E(G)$ by Lemma 8.7, and $uv\in E(G)$ since $G_{t}$ is a $K_{2}$. Then we let $P_{1}:=uz_{1}^{\prime}$ and $P_{2}:=uv$. Otherwise, $t$ is helpful. By Lemma 8.8, $z^{\prime}_{1}$ has a neighbour $w$ in $\mathring{G}_{t}$. Since $G_{t}$ is 2-connected, it contains two independent paths $Q_{1},Q_{2}$ such that $Q_{1}$ is a $u$$w$ path and $Q_{2}$ is a $u$$v$ path. Then we obtain $P_{1}$ from $Q_{1}$ by appending $z^{\prime}_{1}$, and we let $P_{2}:=Q_{2}$. ∎
The following lemma is supported by Figure 8.3.
Lemma 8.12 (Two-Paths Lemma).
Assume the  (Block-bagel setting). (8.4). Let $P$ be an $r$$s$ path in $O$. Assume that either the first edge $e$ or the second node of $P$ is helpful. Assume that $P$ contains at least two helpful elements distinct from $r$. Let $u$ be the adhesion-vertex of the edge $e$. Then there exist two independent $u$$z_{1}^{\prime}$ paths in $G$ with all interior vertices in $\bigcup_{t\in V(\mathring{P})}G_{t}\cup\mathring{G}_{s}$.
Proof.
Without loss of generality, $P=t(0)\,t(1)\ldots t(k)$ with $r=t(0)$ and $s=t(k)$. By assumption, either the edge $e=t(0)\,t(1)$ or the second node $t(1)$ is helpful. Let $u_{i}$ denote the adhesion-vertex of $t(i)\,t(i+1)$. In particular, $u=u_{0}$. By the (Two-Fan Lemma). (8.11), there are two independent paths $P_{1},Q_{1}$ with interior in $\mathring{G}_{t(1)}$ such that $P_{1}$ is a $u_{0}$$z_{1}^{\prime}$ path and $Q_{1}$ is a $u_{0}$$u_{1}$ path.
If $t(1)t(2)$ or $t(2)$ is helpful, then $k=2$. In this case, we find a $u_{1}$$z_{1}^{\prime}$ path $Q_{2}$ with interior in $\mathring{G}_{t(2)}$ by the (Two-Fan Lemma). (8.11). Then the two paths $P_{1}$ and $u_{0}Q_{1}u_{1}Q_{2}z_{1}^{\prime}$ are as desired. Otherwise, $t(1)$ is helpful, $t(2)$ has a $K_{2}$-bag and at least one of $t(2)t(3)$ or $t(3)$ is helpful. So $k=3$ in this case. Then $u_{1}u_{2}\in E(G)$ since $t(2)$ has a $K_{2}$-bag. We find a $u_{2}$$z_{1}^{\prime}$ path $Q_{2}$ with interior in $\mathring{G}_{t(3)}$ by the (Two-Fan Lemma). (8.11). Then the paths $P_{1}$ and $u_{0}Q_{1}u_{1}u_{2}Q_{2}z_{1}^{\prime}$ are as desired. ∎
Figure 8.3. The paths constructed in the (Two-Paths Lemma). (8.12). Left: the first edge on $P$ is helpful. Right: the second node on $P$ is helpful.
8.3. Totally-nested tetra-separations from good bags
Definition 8.13 (Good bag, bad bag).
Assume the (Block-bagel setting). (8.4). A bag $G_{t}$ of $\mathcal{O}$ is good if $G_{t}$ is 2-connected and either
  • $\cdot$
    $|G_{t}|\geqslant 4$, or
  • $\cdot$
    $G_{t}$ is a triangle and a neighbouring bag is a $K_{2}$.
Otherwise, $G_{t}$ is bad. We also call $t$ good or bad, respectively.
Note that $G_{t}$ is bad if and only if either $G_{t}$ is a $K_{2}$ or $G_{t}$ is a triangle and both neighbouring bags are 2-connected.
The following setting is supported by Figure 8.4.
Setting 8.14 (Block-bagel bag setting).
Assume the (Block-bagel setting). (8.4). Let $U^{\prime}:=U_{t(0)}^{\prime}$ and $W^{\prime}:=W_{t(0)}^{\prime}$. Let $x^{\prime}$ be the adhesion-vertex of $t(-1)\,t(0)$ and let $y^{\prime}$ be the adhesion-vertex of $t(0)\,t(1)$. Observe that $S(U^{\prime},W^{\prime})=\{x^{\prime},y^{\prime},z_{1}^{\prime},z_{2}^{\prime}\}$. Let $(U,W)$ be the right-left-shift of $(U^{\prime},W^{\prime})$. Let $x,y,z_{1},z_{2}\in S(U,W)$ correspond to $x^{\prime},y^{\prime},z^{\prime}_{1},z^{\prime}_{2}$, respectively. [^3] So $x$ is incident or equal to $x^{\prime}$, and so on. Let $x^{W}$ denote the unique vertex in $\hat{x}\cap W$, and define $y^{W}$ and $z_{i}^{W}$ similarly.
Figure 8.4. The (Block-bagel setting). (8.4). $S(U^{\prime},W^{\prime})$ is blue, $S(U,W)$ is red. We have $y^{W}=y^{\prime}$ and $z^{W}_{i}=z^{\prime}_{i}$, so these three vertices are blue. The vertex $x^{W}$ is green.
Observation 8.15.
Assume the  (Block-bagel bag setting). (8.14). If $G_{t(1)}$ is a $K_{2}$, then $y$ is the edge in $G_{t(1)}$. Symmetrically, if $G_{t(-1)}$ is a $K_{2}$, then $x$ is the edge in $G_{t(-1)}$.∎
Lemma 8.16.
Assume the  (Block-bagel bag setting). (8.14). Then each vertex $z^{\prime}_{i}$ in the centre has $\geqslant 3$ neighbours outside $Z^{\prime}$ contained in $W^{\prime}\smallsetminus U^{\prime}$ , and in particular in $W\smallsetminus U$. In particular, $z_{i}^{W}=z_{i}^{\prime}$.
Proof.
We combine the (Helpful Lemma). (8.6) with Lemma 8.7 and Lemma 8.8, and we use 5.9 for $W^{\prime}\smallsetminus U^{\prime}\subseteq W\smallsetminus U$. ∎
Lemma 8.17.
Assume the  (Block-bagel bag setting). (8.14). Then the following assertions are equivalent:
  1. (i)
    $G_{t(0)}$ is good, or all three bags $G_{t(-1)},G_{t(0)},G_{t(1)}$ are $K_{2}$’s;
  2. (ii)
    $(U,W)$ is a tetra-separation.
Proof.
(i) $\Rightarrow$ (ii). Suppose first that $G_{t(0)}$ is good. By Corollary 8.3, each of the four corners contains a vertex. Since $G_{t(0)}$ is 2-connected, it is included in $A\cap C$, say, so every corner except $AC$ has its vertex contained in $W^{\prime}\smallsetminus U^{\prime}$. Hence $|W^{\prime}\smallsetminus U^{\prime}|\geqslant 3$. The 4-separation $(U^{\prime},W^{\prime})$ trivially satisfies the matching-condition. If the bag $G_{t(0)}$ is not a triangle, then it has $\geqslant 4$ vertices, so $|U^{\prime}\smallsetminus W^{\prime}|\geqslant 2$. In this case, $(U,W)$ is a tetra-separation by 5.10. Otherwise, $G_{t(0)}$ is a triangle. Then some neighbouring bag, say $G_{t(1)}$, is a $K_{2}$. Hence $|U^{\prime}\smallsetminus W^{\prime}|=1$ and the vertex $y^{\prime}$ has exactly one neighbour in $W^{\prime}\smallsetminus U^{\prime}$. Again by 5.10, $(U,W)$ is a tetra-separation.
Suppose now that all three bags $G_{t(-1)},G_{t(0)},G_{t(1)}$ are $K_{2}$’s. Then $G_{t(0)}$ meets at most two corners, so the vertices in the other two corners lie in $W^{\prime}\smallsetminus U^{\prime}$, giving $|U^{\prime}\smallsetminus W^{\prime}|\geqslant 2$. Both $x^{\prime}$ and $y^{\prime}$ have exactly one neighbour in $W^{\prime}\smallsetminus U^{\prime}$ since both $G_{t(-1)}$ and $G_{t(1)}$ are $K_{2}$’s. Hence $(U,W)$ is a tetra-separation by 5.10.
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). As (i) fails, $G_{t(0)}$ either is a $K_{2}$ neighboured by at least one 2-connected bag, or $G_{t(0)}$ is a triangle neighboured by 2-connected bags. Suppose first that $G_{t(0)}$ is a $K_{2}$ neighboured by a 2-connected bag, say $G_{t(1)}$. Then $y^{\prime}$ has two neighbours in $G_{t(1)}$, which are also contained in $W^{\prime}\smallsetminus U^{\prime}$. The two vertices $z^{\prime}_{i}$ in the centre have $\geqslant 3$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$ by Lemma 8.16. So only $x^{\prime}$ might have $\leqslant 1$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$. But $|U^{\prime}\smallsetminus W^{\prime}|=0$, so $(U,W)$ is no tetra-separation by 5.10.
Suppose now that $G_{t(0)}$ is a triangle with both neighbouring bags 2-connected. Then, as above, all vertices in $S(U^{\prime},W^{\prime})$ have $\geqslant 2$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$. But $|U^{\prime}\smallsetminus W^{\prime}|=1$, so $(U,W)$ is no tetra-separation by 5.10. ∎
Observation 8.18.
Assume the  (Block-bagel bag setting). (8.14). Let $t(i)\,t(i+1)$ be the edge of $O$ that has $y^{W}$ as adhesion-vertex. Then either the edge $t(i)\,t(i+1)$ or the node $t(i+1)$ is helpful.∎
Lemma 8.19.
Assume the  (Block-bagel bag setting). (8.14). Then there exist three independent $z_{1}^{\prime}$$z_{2}^{\prime}$ paths in $G[W]-\{x^{\prime},y^{\prime}\}-z^{\prime}_{1}z^{\prime}_{2}$.
Proof.
For each helpful $h\in V(O)\cup E(O)$ we construct a $z^{\prime}_{1}$$z^{\prime}_{2}$ path $P_{h}$, as follows. If $h$ is an edge, then its adhesion-vertex send edges to both $z^{\prime}_{i}$ by Lemma 8.7, and we let $P_{h}$ be the path defined by these two edges; see Figure 8.5. If $h$ is a helpful node, then we use the (Bag-Path Lemma). (8.9) to let $P_{h}$ by a $z^{\prime}_{1}$$z^{\prime}_{2}$ path avoiding the edge $z^{\prime}_{1}z^{\prime}_{2}$ (if it exists) and with interior in $\mathring{G}_{h}$.
All paths $P_{h}$ are independent. By the (Helpful Lemma). (8.6), there are at least four of them. At most one of them meets the bag $G_{t(0)}$. By Lemma 8.16, both vertices $z^{\prime}_{i}$ in the centre are contained in $W$. Hence at least three of the paths $P_{h}$ are independent $z^{\prime}_{1}$$z^{\prime}_{2}$ paths in $G[W]-\{x^{\prime},y^{\prime}\}-z^{\prime}_{1}z^{\prime}_{2}$. ∎
Figure 8.5. The red paths are constructed in Lemma 8.19.
The following lemma is supported by Figure 8.6.
Lemma 8.20.
Assume the  (Block-bagel bag setting). (8.14) with $G_{t(0)}$ 2-connected. Then there are three independent $\{x^{W},y^{W}\}$$z_{1}^{\prime}$ paths $P_{1},P_{2},P_{3}$ in $G[W]$, such that:
  1. (i)
    both $P_{1}$ and $P_{2}$ are $y^{W}$$z_{1}^{\prime}$ paths, and
  2. (ii)
    $P_{3}$ is an $x^{W}$$z_{1}^{\prime}$ path.
Proof.
Since $t(0)$ is helpful, (Helpful Lemma). (8.6) implies that
(6) $\geqslant 3$ elements of $O-t(0)$ are helpful.
Let $f=t(j)\,t(j+1)$ be the edge on $O$ that has $y^{W}$ as adhesion-vertex. Then $f$ or $t(j+1)$ is helpful by 8.18. If $t(j)$ is helpful, then $j=0$, and $j=1$ otherwise. Let $P$ be the path in $O$ that starts with $t(j),t(j+1)$ and has two helpful elements distinct from $t(j)$ and is minimal with these properties. The path $P$ is well-defined since $O-t(0)$ has $\geqslant 3$ helpful elements by 6. Moreover, the end $s$ of $P$ is distinct from $t(0)$. By the (Two-Paths Lemma). (8.12), we find two independent $y^{W}$$z_{1}^{\prime}$ paths $Q_{1},Q_{2}$ with all interior vertices in $\bigcup_{t\in V(\mathring{P})}G_{t}\cup\mathring{G}_{s}$. By construction, the paths $Q_{1}$ and $Q_{2}$ are contained in $G[W]$.
Now let $e=t(i)\,t(i-1)$ be the edge on $O$ that has $x^{W}$ as adhesion-vertex. Then $e$ or $t(i-1)$ is helpful by 8.18. And if $t(i)$ is helpful, then $i=0$. By the (One-Path Lemma). (8.10), we find an $x^{W}$$z_{1}^{\prime}$ path $Q_{3}$ with interior in $\mathring{G}_{t(i-1)}$. Once more, $Q_{3}\subseteq G[W]$ by construction.
It remains to show that the three paths $Q_{i}$ are independent. We already know that $Q_{1}$ and $Q_{2}$ are independent, so it remains to show that $Q_{3}$ is independent from $Q_{1}$ and $Q_{2}$. If the edge $e$ is helpful, then it is the first helpful element on $O$ following $t(0)$ in the direction of $e$. In particular, $e$ does not lie on $P$ by 6. Since the bag $G_{t(i-1)}$ is a $K_{2}$, it has empty interior. So $Q_{3}=x^{W}z^{\prime}_{1}$ avoids $\mathring{G}_{s}$, even if $s=t(i-1)$, so $Q_{3}$ is independent from $Q_{1}$ and $Q_{2}$. Otherwise $t(i-1)$ is helpful. Then $t(i-1)$ is the first helpful element on $O$ following $t(0)$ in the direction of $e$. In particular, both edges on $O$ incident to $t(i-1)$ are unhelpful. So $t(i-1)$ does not lie on $P$ by 6. Hence $Q_{3}$ is independent from $Q_{1}$ and $Q_{2}$. ∎
Figure 8.6. The paths constructed in Lemma 8.20. $P_{1}$ is yellow, $P_{2}$ is blue, $P_{3}$ is green.
Lemma 8.21.
Let $G$ be a graph, let $u\in U\subseteq V(G)$ and $v\in V\subseteq V(G)$. Every $u$$v$ path in $G$ contains a $U$$V$ path.∎
In the following, we characterise when $(U,W)$ is a totally-nested tetra-separation.
Half-connected
Lemma 8.22.
Assume the  (Block-bagel bag setting). (8.14). Then $(U,W)$ is half-connected.
Proof.
We show that $G[W\smallsetminus U]$ is connected. Without loss of generality, assume that the interior of $G_{t(0)}$ is contained in the corner $AC$. Let $v$ be a vertex in the corner $BD$, which exists by Corollary 8.3. Let $t(i)$ be a node of $O$ with $v\in G_{t(i)}$.
We claim that every vertex $w$ in $W$ is connected to $v$ in $G[W\smallsetminus U]$. Since $G_{t(i)}\subseteq G[W\smallsetminus U]$, every vertex in $G_{t(i)}$ is connected to $v$ in $G[W\smallsetminus U]$. If $w$ does not lie in the centre $Z^{\prime}$, then we are done by the (One-Path Lemma). (8.10). Otherwise $w=z^{\prime}_{1}$, say. Then $w$ has a neighbour in $W\smallsetminus Z^{\prime}$ by Lemma 8.19, which we already know is linked to $v$ in $G[W\smallsetminus U]$. Hence $(U,W)$ is half-connected. ∎
3-linked
Lemma 8.23.
Assume the  (Block-bagel bag setting). (8.14) with $G_{t(0)}$ 2-connected. Then $(U,W)$ is 3-linked.
Proof.
Let $\pi$ consist of two elements of $S(U,W)$. We show that $\pi$ is 3-linked around $S(U,W)$.
Case $\pi=\{x,y\}$. We may assume that both $z_{1},z_{2}$ are vertices, by (6.4.1). Since the bag $G_{t(0)}$ is 2-connected, it contains two independent $x^{\prime}$$y^{\prime}$ paths. By Lemma 8.21, these paths contain two independent $\hat{x}$$\hat{y}$ paths which are also included in the bag $G_{t(0)}$. We find a third $\hat{x}$$\hat{y}$ path $G$ that avoids both the bag $G_{t(0)}$ and the centre, by the (One-Path Lemma). (8.10). So we are done by (6.4.3).
Case $\pi=\{z_{1},z_{2}\}$. We find three independent $z^{\prime}_{1}$$z^{\prime}_{2}$ paths in $G-x-y$ by Lemma 8.19. These contain three independent $\hat{z}_{1}$$\hat{z}_{2}$ paths by Lemma 8.21. So we are done by (6.4.3).
Case $\pi=\{x,z_{1}\}$. We find an $x^{\prime}$$z^{\prime}_{1}$ path with interior in $\mathring{G}_{t(0)}$, by the (Two-Fan Lemma). (8.11). This path contains a $\hat{x}$$\hat{z}_{1}$ path $P_{1}$, by Lemma 8.21. Next, we find two independent $x^{W}$$z^{\prime}_{1}$ paths $P_{2}$ and $P_{3}$ in $G$ that internally avoid $G_{t(0)}$, by Lemma 8.20. Both paths $P_{2}$ and $P_{3}$ are $\hat{x}$$\hat{z}_{1}$ paths by Lemma 8.16. Hence the $P_{i}$ witness (6.4.3).
All remaining cases are symmetric to the case $\pi=\{x,z_{1}\}$. ∎
0-potter-linked
Lemma 8.24.
Assume the  (Block-bagel bag setting). (8.14) with $G_{t(0)}$ good. Then $(U,W)$ is 0-potter-linked.
Proof.
Let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of $S(U,W)$.
Case $\pi_{1}=\{x,y\}$ and $\pi_{2}=\{z_{1},z_{2}\}$. Suppose first that $G_{t(0)}$ is a triangle with a neighbouring bag that is a $K_{2}$; say $G_{t(1)}$ is a $K_{2}$. Let $v$ denote the vertex of the triangle $G_{t(0)}$ besides $x^{\prime}$ and $y^{\prime}$. Since $G$ is 4-connected, $v$ sends edges to both vertices $z^{\prime}_{i}$ in the centre. Similarly, the adhesion-vertex $y^{\prime}$ in between the triangle-bag $G_{t(0)}$ and the $K_{2}$-bag $G_{t(1)}$ sends at least one edge to some vertex in the centre, say to $z^{\prime}_{1}$.
We claim that $z^{\prime}_{1}=z_{1}$. On the one hand, $z^{\prime}_{1}$ has $\geqslant 3$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$ by Lemma 8.16, so $z^{\prime}_{1}$ remains in the separator when taking the right-shift $(U^{\prime\prime},W^{\prime\prime})$ of $(U^{\prime},W^{\prime})$. However, $y^{\prime}$ becomes the edge $y$ of the $K_{2}$-bag $G_{t(1)}$ in the separator $S(U^{\prime\prime},W^{\prime\prime})$ by 8.15. Hence $z^{\prime}_{1}$ has two neighbours $y^{\prime}$ and $v$ in $U^{\prime\prime}\smallsetminus W^{\prime\prime}$. Thus $z^{\prime}_{1}$ also remains in the separator when taking the left-shift $(U,W)$ of $(U^{\prime\prime},W^{\prime\prime})$. In particular, $z^{\prime}_{1}=z_{1}$.
So $z_{1}$ is a vertex. Recall that $z_{1}$ sends an edge to $y^{\prime}\in\hat{\pi}_{1}$. Hence we have (6.7.1).
Suppose otherwise that $G_{t(0)}$ is 2-connected and has at least four vertices. We will find five independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G$ for (6.7.2). There exist two disjoint $\{z^{\prime}_{1},z^{\prime}_{2}\}$$\mathring{G}_{t(0)}$ paths $P_{1},P_{2}$ by Lemma 8.8. Since $G_{t(0)}$ is 2-connected, we can extend the paths $P_{1},P_{2}$ to two disjoint $\{z^{\prime}_{1},z^{\prime}_{2}\}$$\{x^{\prime},y^{\prime}\}$ paths with interior in $\mathring{G}_{t(0)}$. We find three independent $\{x_{1}^{W},x_{2}^{W}\}$$\{z_{1}^{\prime},z_{2}^{\prime}\}$ paths in $G[W]$ by Lemma 8.20. The five paths are independent, and they contain five $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths by Lemma 8.21.
Case $\pi_{1}=\{x,z_{1}\}$ and $\pi_{2}=\{y,z_{2}\}$. Here it suffices to find five independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G$ for (6.7.2). There are two independent $x^{\prime}$$y^{\prime}$ paths in $G_{t(0)}$ by 2-connectivity. We find three independent $z^{\prime}_{1}$$z^{\prime}_{2}$ paths in $G[W]-\{x^{\prime},y^{\prime}\}$ by Lemma 8.19. The latter three paths are internally disjoint from the former two. By Lemma 8.21, the five paths contain five independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths.
All remaining cases are symmetric to one of the two cases above. ∎
1-potter-linked and 2-potter-linked
Lemma 8.25.
Assume the  (Block-bagel bag setting). (8.14). Then no pair of two elements of $S(U,W)$ is $W$-weird.
Proof.
Let $\pi$ consist of two elements of $S(U,W)$. Assume for a contradiction that $\pi$ is $W$-weird.
Assume $\pi=\{x,z_{1}\}$. The vertex $z^{\prime}_{1}$ in the centre has $\geqslant 3$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$ by Lemma 8.16. Hence $z^{W}_{1}=z^{\prime}_{1}$. Since the $\geqslant 3$ neighbours also lie in $W\smallsetminus U$ by 5.9, we further get that $z_{1}$ must be the edge in $\pi$. Since $z_{1}$ is an edge, $x$ is a vertex which has exactly one neighbour in $G[W]-z_{1}^{W}$. Hence $G_{t(-1)}$ is a $K_{2}$. But then $x$ is the edge of $G_{t(-1)}$ by 8.15, a contradiction.
Assume $\pi=\{x,y\}$. Say $x$ is a vertex and $y$ is an edge. Then $x=x^{\prime}$, so $G_{t(-1)}$ is 2-connected by 8.15. Hence $x$ has two neighbours $u,v$ in $G_{t(-1)}$. By the (Helpful Lemma). (8.6), neither $u$ nor $v$ is an endvertex of the edge $y$. Hence $\{x,y\}$ is not $W$-weird, a contradiction.
All other cases are covered by symmetry. ∎
Lemma 8.26.
Assume the  (Block-bagel bag setting). (8.14) with $G_{t(0)}$ 2-connected. Then $(U,W)$ is 1-potter-linked and 2-potter-linked.
Proof.
Let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of $S(U,W)$. Neither $\pi_{i}$ is $W$-weird by Lemma 8.25. In particular, $\{\pi_{1},\pi_{2}\}$ is 2-potter-linked.
To show that $\{\pi_{1},\pi_{2}\}$ is 1-potter-linked, it suffices to find three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[W]$. If one of the $\pi_{i}$ is of the form $\{x,y\}$, then we find three such paths by Lemma 8.20 and Lemma 8.21. Otherwise $z_{1}\in\pi_{1}$ and $z_{2}\in\pi_{2}$, say. Then we find three desired paths by Lemma 8.19 and Lemma 8.21. ∎
Summary
Key Lemma 8.27.
Assume the  (Block-bagel setting). (8.4). Let $t$ be a node of $O$. If the bag $G_{t}$ is good, then $(U_{t},W_{t})$ is a totally-nested tetra-separation.
Proof.
We may assume the (Block-bagel bag setting). (8.14) with $t=t(0)$. Then $(U_{t},W_{t})$ is a tetra-separation by Lemma 8.17. It suffices to show that $(U_{t},W_{t})$ is externally 5-connected, by 6.16. This follows from Lemma 8.22 (half-connected), Lemma 8.23 (3-linked), Lemma 8.24 (0-potter-linked) and Lemma 8.26 (1-potter-linked and 2-potter-linked). ∎
8.4. Splitting stars and torsos
Definition 8.28 (Tetra-star $\sigma(\mathcal{O})$).
Assume the (Block-bagel setting). (8.4). The tetra-star of $\mathcal{O}$ is the set of all $(U_{t},W_{t})$ for good bags $G_{t}$ of $\mathcal{O}$. We denote the tetra-star of $\mathcal{O}$ by ${\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{\sigma(\mathcal{O})}}$.
Lemma 8.29.
Assume the (Block-bagel setting). (8.4). Then $\sigma(\mathcal{O})$ is a star of totally-nested tetra-separations.
Proof.
[…]
The following two lemmata are supported by Figure 8.7.
Lemma 8.30 (Link-Moving Lemma for $K_{2}$-bags).
Assume the (Block-bagel setting). (8.4). Let $P$ and $Q$ be the two paths that are the components of $O-o(C)-o(D)$. Assume $o(A)\in P$ and $o(B)\in Q$. Let $c$ denote the end of $P$ that is adjacent or incident to $o(C)$ in $O$. Define $d$ similarly. Let $s$ be a node in the same component of $P-o(A)$ as $c$ such that $G_{s}=K_{2}$. Let $P^{\prime}:=dP\mathring{s}$ and $Q^{\prime}:=\mathring{s}Pc\cup Q$.
  1. (i)
    $A^{\prime}:=\bigcup_{t\in V(P^{\prime})}V(G_{t})\cup Z^{\prime}$ and $B^{\prime}:=\bigcup_{t\in V(Q^{\prime})}V(G_{t})\cup Z^{\prime}$ are the two sides of a tetra-separation of $G$.
  2. (ii)
    $S(A^{\prime},B^{\prime})=Z^{\prime}\cup(D\text{-link})\cup\{e\}$, where $e$ is the edge in $G_{s}$.
  3. (iii)
    $(A^{\prime},B^{\prime})$ separates some two vertices in $B\cup D$.
  4. (iv)
    $(A^{\prime},B^{\prime})$ and $(C,D)$ cross with all links of size one, and the block-bagel induced by $(A^{\prime},B^{\prime})$ and $(C,D)$ is equal to $\mathcal{O}$.
Proof.
The pair $(A^{\prime},B^{\prime})$ is a mixed-separation by construction. Moreover, (ii) also holds by construction. We show (i) next. Let $d$ be the element in the $D$-link.
Matching-condition: We have to show that, if $d$ in an edge, then it does not share an endvertex with $e$. The endvertices of $d$ lie in the corners for $AD$ and $BD$ since there do not exist dangling edges in the crossing-diagram of $(A,B)$ and $(C,D)$ by the (Crossing Lemma). (4.1). Both endvertices of $e$ lie in the corner for $AC$ by Lemma 8.2 (i). Hence $d$ and $e$ share no ends.
Degree-condition: All vertices in $S(A^{\prime},B^{\prime})$ are also contained in $S(A,B)$. Since $B\smallsetminus A\subseteq B^{\prime}\smallsetminus A^{\prime}$, every vertex in $S(A^{\prime},B^{\prime})$ has at least two neighbours in $B^{\prime}\smallsetminus A^{\prime}$.
If $d$ is a vertex, then we claim that every neighbour $u$ of $d$ in $A\smallsetminus B$ also also lies in $A^{\prime}\smallsetminus B^{\prime}$. Indeed, since $ud$ is not a dangling edge, $u$ is contained in the corner $AD$ or the $A$-link with respect to $(A,B),(C,D)$. By Lemma 8.2 (i), no endvertex of $e$ lies in the $A$-link with respect to $(A,B),(C,D)$. So, we have $u\in A^{\prime}\smallsetminus B^{\prime}$.
It remains to show that both vertices $z_{1}^{\prime}$ and $z_{2}^{\prime}$ in the centre have at least two neighbours in $A^{\prime}\smallsetminus B^{\prime}$. Let $s^{\prime}$ be the neighbour of $s$ on $P^{\prime}$. Observe that either $s^{\prime}$ or $ss^{\prime}$ is helpful and belongs to the corner $AC$. Moreover, by the (Helpful Lemma). (8.6) there is a helpful element that belongs to the $AD$-corner. Hence, there are at least two helpful elements that belong to $(ss^{\prime})\cup T_{AD}$, where $T_{AD}$ is the $AD$-block-path. By Lemmas 8.7 and 8.8, we are done.
We also have (iii) since there is a vertex in $A\cap(D\smallsetminus C)\subseteq A^{\prime}\smallsetminus B^{\prime}$ that is separated from every vertex in $B\smallsetminus A\subseteq B^{\prime}\smallsetminus A^{\prime}$ by $(A^{\prime},B^{\prime})$. Next, we show (iv). Clearly, $Z^{\prime}$ is the centre of the crossing-diagram of $(A^{\prime},B^{\prime})$ with $(C,D)$. So, $(A^{\prime},B^{\prime})$ and $(C,D)$ cross with all links of size one. It is straightforward to check that the block-bagel induced by $(A^{\prime},B^{\prime})$ and $(C,D)$ is equal to $\mathcal{O}$. ∎
Let $P$ be a path. Let $u$ be a vertex on $P$, and let $e$ be an edge on $P$. Then $uP\mathring{e}$ denotes the subpath of $P$ that starts in $u$ and ends in the first end of $e$ it encounters. In particular, $e$ is not an edge of $uP\mathring{e}$.
Lemma 8.31 (Link-Moving Lemma for adhesion-vertices).
Assume the (Block-bagel setting). (8.4). Let $P$ and $Q$ be the two paths that are the components of $O-o(C)-o(D)$. Assume $o(A)\in P$ and $o(B)\in Q$. Let $c$ denote the end of $P$ that is adjacent or incident to $o(C)$ in $O$. Define $d$ similarly. Let $e$ be an edge of $O$ with 2-connected bags at both ends. Assume that $e$ stems from the $AC$-block-path. Let $P^{\prime}:=dP\mathring{e}$ and $Q^{\prime}:=\mathring{e}Pc\cup Q$. Then:
  1. (i)
    $A^{\prime}:=\bigcup_{t\in V(P^{\prime})}V(G_{t})\cup Z^{\prime}$ and $B^{\prime}:=\bigcup_{t\in V(Q^{\prime})}V(G_{t})\cup Z^{\prime}$ are the two sides of a tetra-separation of $G$.
  2. (ii)
    $S(A^{\prime},B^{\prime})=Z^{\prime}\cup(D\text{-link})\cup\{v\}$, where $v$ is the adhesion-vertex of the edge $e$.
  3. (iii)
    $(A^{\prime},B^{\prime})$ separates some two vertices in $B\cup D$.
  4. (iv)
    $(A^{\prime},B^{\prime})$ and $(C,D)$ cross with all links of size one, and the block-bagel induced by $(A^{\prime},B^{\prime})$ and $(C,D)$ is equal to $\mathcal{O}$.
Proof.
The proof is analogous to the proof of the (Link-Moving Lemma for $K_{2}$-bags). (8.30), except that the matching-condition is trivially fulfilled, and that the vertex $v$ satisfies the degree condition since it is contained in two 2-connected bags $G_{s},G_{s^{\prime}}$ with $V(G_{s}-v)\subseteq A^{\prime}\smallsetminus B^{\prime}$ and $V(G_{s^{\prime}}-v)\subseteq B^{\prime}\smallsetminus A^{\prime}$. ∎
Figure 8.7. The tetra-separation $(A^{\prime},B^{\prime})$ constructed in the (Link-Moving Lemma for $K_{2}$-bags). (8.30) (left) and in the (Link-Moving Lemma for adhesion-vertices). (8.31) (right).
Lemma 8.32.
Assume the (Block-bagel setting). (8.4). Let $T$ denote the $AC$-block-path. Let $T^{\prime}:=T-o(A)-o(C)$. Then either $T^{\prime}$ is non-empty or $|A\cap C|\leqslant 3$.
Proof.
If the $A$-link or the $C$-link contains a vertex, then this vertex is contained in a 2-connected bag $G_{t}$ with $t\in V(T^{\prime})$, by Lemma 8.2. Otherwise both the $A$-link and the $C$-link consist of edges. Then $G_{o(A)}$ and $G_{o(C)}$ are $K_{2}$-bags. Each of the two bags $G_{o(A)},G_{o(C)}$ has exactly one endvertex $v_{A},v_{C}$ in the corner for $AC$ by Lemma 8.2, respectively. If $T^{\prime}$ is empty, then $v_{A}=v_{C}$, and $A\cap C=\{v_{A}\}\cup Z^{\prime}$ has size three. ∎
Lemma 8.33.
Assume the (Block-bagel setting). (8.4). Let $(E,F)$ be a tetra-separation of $G$ that interlaces the tetra-star $\sigma(\mathcal{O})$. Then $(E,F)$ is crossed by another tetra-separation of $G$, or there is a 2-connected bag $G_{t}$ of $\mathcal{O}$ such that $E\subseteq U^{\prime}_{t}$ or $F\subseteq U^{\prime}_{t}$.
Proof.
Assume that $(E,F)$ is totally-nested. Then $(E,F)\leqslant(A,B)$ and $(E,F)\leqslant(C,D)$, say. We choose two tetra-separations $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$ with $|A^{\prime}\cap C^{\prime}|$ minimum such that
  1. (i)
    $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$ cross with all links of size one; and
  2. (ii)
    the block-bagel induced by $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$ is equal to $\mathcal{O}$; and
  3. (iii)
    $(E,F)\leqslant(A^{\prime},B^{\prime})$ and $(E,F)\leqslant(C^{\prime},D^{\prime})$.
Note that the existence of $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$ follows from $(A,B)$ and $(C,D)$ exhibiting properties (i)(iii). Let $T$ denote the $A^{\prime}C^{\prime}$-block-path, and let $T^{\prime}:=T-o(A^{\prime})-o(C^{\prime})$. By Lemma 8.32, $T^{\prime}$ is non-empty. Let $\mathcal{O}^{\prime}=\mathcal{O}$ be the block-bagel induced by $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$.
We claim that every bag $G_{t}$ with $t\in T^{\prime}$ is 2-connected. Assume otherwise that $G_{t}=K_{2}$ for some $t\in T^{\prime}$, and let $e$ denote the edge of $G_{t}$. In the context of $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$, we apply the (Link-Moving Lemma for $K_{2}$-bags). (8.30) to $t$ and obtain a tetra-separation $(A^{\prime\prime},B^{\prime\prime})$ such that
  • $\cdot$
[…]
  • $\cdot$
    $(A^{\prime\prime},B^{\prime\prime})$ separates two vertices of $B^{\prime}\cup D^{\prime}$.
Since $B^{\prime}\cup D^{\prime}\subseteq F$, the tetra-separation $(A^{\prime\prime},B^{\prime\prime})$ separates two vertices in $F$. If $(A^{\prime\prime},B^{\prime\prime})$ also separates two vertices in $E$, then it crosses $(E,F)$, a contradiction. Hence we have either $(E,F)\leqslant(A^{\prime\prime},B^{\prime\prime})$ or $(E,F)\leqslant(B^{\prime\prime},A^{\prime\prime})$. In the former case, we have that $(A^{\prime\prime},B^{\prime\prime})$ and $(C^{\prime},D^{\prime})$ satisfy all of (i)-(iii), but have $|A^{\prime\prime}\cap C^{\prime}|<|A^{\prime}\cap C^{\prime}|$, contradicting the minimality of $|A^{\prime}\cap C^{\prime}|$. In the latter case, we apply the (Link-Moving Lemma for $K_{2}$-bags). (8.30) to $t$ and obtain a tetra-separation $(C^{\prime\prime},D^{\prime\prime})$ such that
  • $\cdot$
[…]
  • $\cdot$
    $(C^{\prime\prime},D^{\prime\prime})$ separates two vertices of $B^{\prime}\cup D^{\prime}$.
Moreover, we have
$E\subseteq B^{\prime\prime}\cap A^{\prime}\cap C^{\prime}\subseteq C^{\prime\prime}\quad\text{and}\quad F\supseteq A^{\prime\prime}\cup D^{\prime}\cup B^{\prime}\supseteq D^{\prime\prime}.$
Hence $(E,F)\leqslant(C^{\prime\prime},D^{\prime\prime})$. Therefore, we have that $(A^{\prime},B^{\prime})$ and $(C^{\prime\prime},D^{\prime\prime})$ satisfy all of (i)-(iii), but have $|A^{\prime}\cap C^{\prime\prime}|<|A^{\prime}\cap C^{\prime}|$, contradicting the minimality of $|A^{\prime}\cap C^{\prime}|$.
We claim that $T^{\prime}$ has no edge. Assume otherwise that $tt^{\prime}$ is an edge of $T^{\prime}$. Then both bags $G_{t},G_{t^{\prime}}$ at its ends are 2-connected by the above claim. In the context of $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$, we apply the (Link-Moving Lemma for adhesion-vertices). (8.31) to the edge $tt^{\prime}$ and obtain a tetra-separation $(A^{\prime\prime},B^{\prime\prime})$ with the properties stated in the lemma. Then we derive a contradiction in the same way as in the previous claim.
By the above, every bag $G_{t}$ with $t\in T^{\prime}$ is 2-connected, and $T^{\prime}$ has no edge. Since $T^{\prime}$ is non-empty, it has a unique node $t$, and $G_{t}$ is 2-connected. Then $A^{\prime}\cap C^{\prime}=V(G_{t})\cup Z^{\prime}$. By (iii), $E\subseteq U^{\prime}_{t}$. ∎
Lemma 8.34.
Assume the (Block-bagel setting). (8.4). Every tetra-separation of $G$ that interlaces the tetra-star $\sigma(\mathcal{O})$ is crossed by another tetra-separation of $G$.
Proof.
Assume for contradiction that $\sigma(\mathcal{O})$ is interlaced by a totally-nested tetra-separation $(E,F)$ of $G$. By Lemma 8.33, there is a 2-connected bag $G_{t}$ of $\mathcal{O}$ with $E\subseteq U_{t}^{\prime}$, say. By Lemma 5.7, we have $(E,F)\leqslant(U_{t}^{\prime\prime},W_{t}^{\prime\prime})$ where $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$ is the right-shift of $(U^{\prime}_{t},W^{\prime}_{t})$.
Assume first that $G_{t}$ is good. Then $(U_{t},W_{t})\in\sigma(\mathcal{O})$ by definition of $\sigma(\mathcal{O})$. Since $(E,F)$ interlaces $\sigma(\mathcal{O})$, we have $(U_{t},W_{t})<(E,F)$ or $(U_{t},W_{t})<(F,E)$. We cannot have $(U_{t},W_{t})<(F,E)$ since otherwise $\emptyset\neq E\smallsetminus F\subseteq(W_{t}\smallsetminus U_{t})\cap(U_{t}^{\prime\prime}\smallsetminus W_{t}^{\prime\prime})$ contradicts that $U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t}\subseteq U_{t}\smallsetminus W_{t}$ by 5.2. So $(U_{t},W_{t})<(E,F)\leqslant(U_{t}^{\prime\prime},W_{t}^{\prime\prime})$. Since $(U_{t},W_{t})$ is the left-shift of $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, some vertex in $S(E,F)$ must have $\leqslant 1$ neighbour in $E\smallsetminus F$, contradicting the degree-condition.
Otherwise, $G_{t}$ is bad. Since $G_{t}$ is 2-connected, it is a triangle, and since $G_{t}$ is bad, both neighbouring bags in $\mathcal{O}$ are 2-connected. Hence both adhesion-vertices of $G_{t}$ are contained in the separator of $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$. Therefore, the unique interior vertex of the triangle-bag $G_{t}$ is the only vertex of $U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t}$. Since $(E,F)\leqslant(U_{t}^{\prime\prime},W_{t}^{\prime\prime})$, we have $E\smallsetminus F\subseteq U_{t}^{\prime\prime}\smallsetminus W_{t}^{\prime\prime}$. Hence $|E\smallsetminus F|\leqslant 1$, which by Lemma 4.2 contradicts that $(E,F)$ is a tetra-separation. ∎
Corollary 8.35.
Assume the (Block-bagel setting). (8.4). Let $t$ be a node of $O$. Then the following assertions are equivalent:
  1. (i)
    the bag $G_{t}$ is good;
  2. (ii)
    $(U_{t},W_{t})$ is a totally-nested tetra-separation of $G$.
Proof.
(i) $\Rightarrow$ (ii): this is 8.27.
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). Assume that $G_{t}$ is bad. Then $(U_{t},W_{t})$ fails to be a tetra-separation, unless $G_{t}$ and both neighbouring bags are $K_{2}$’s, by 8.27. In this case, both vertices of $G_{t}$ lie in $U_{t}\smallsetminus W_{t}$, so $(U_{t},W_{t})$ is distinct from all tetra-separations in the tetra-star $\sigma(\mathcal{O})$. Therefore, $(U_{t},W_{t})$ interlaces $\sigma(\mathcal{O})$, and so $(U_{t},W_{t})$ is crossed by another tetra-separation, by Lemma 8.34. ∎
Lemma 8.36.
Let $G$ be a 4-connected graph. Let $\sigma$ be a star of mixed-4-separations of $G$. Then the torso of $\sigma$ is 4-connected or a $K_{4}$.
Proof.
We combine Lemmas 2.6.3 and 2.6.4 from [Tridecomp]. ∎
Key Lemma 8.37.
Let $G$ be a 4-connected graph, and let $N$ denote the set of totally-nested tetra-separations of $G$. Let $\sigma$ be a splitting star of $N$ that is interlaced by two tetra-separations $(A,B)$ and $(C,D)$ of $G$ that cross with vertex-centre $Z$ of size two. Then the torso $\tau$ of $\sigma$ is a generalised double-wheel with centre $Z$.
Proof.
By the (Crossing Lemma). (4.1), the centre $Z$ equals the vertex-centre and all links have size one. Let $\mathcal{O}$ be the Block-bagel induced by $(A,B)$ and $(C,D)$, which has centre $Z$. By Lemma 8.29, the tetra-star $\sigma(\mathcal{O})$ is a star of totally-nested tetra-separations of $G$. By Lemma 8.34, the tetra-star $\sigma(\mathcal{O})$ is a splitting star of $N$. Since both $(A,B)$ and $(C,D)$ interlace both $\sigma$ and $\sigma(\mathcal{O})$, we obtain $\sigma=\sigma(\mathcal{O})$. Hence it remains to compute the torso $\tau$ of $\sigma(\mathcal{O})$. The torso $\tau$ is 4-connected by Lemma 8.36. We can obtain $\tau$ from $\mathcal{O}$ by replacing each good bag $G_{t}$ by a $K_{4}$ on the vertex set $Z\cup\{u,v\}$ where $u$ and $v$ are the adhesion-vertices of $G_{t}$, leaving only $K_{2}$-bags and triangle-bags, and contracting some edges in $K_{2}$-bags. Hence $\tau$ is a generalised wheel with centre $Z$. ∎
8.5. Angry block-bagels
Let $X$ be a graph. We call $X$ a double-wheel with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that $X-u-v$ is a cycle. So $u$ and $v$ send edges to all vertices of $X-u-v$, and the edge $uv$ may or may not exist. The rim-length of $X$ then means the length of the cycle $X-u-v$. We call $X$ a double-wheel of triangles with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that $X-u-v$ has a cycle-decomposition $(O,\mathcal{G})$ with all adhesion-sets of size 1 and into triangles. The rim-length of $X$ means the length of the cycle $O$.
Lemma 8.38.
Let $G$ be a 4-connected graph with a block-bagel $\mathcal{O}=(O,\mathcal{G})$ with centre $Z$ such that all bags of $\mathcal{O}$ are bad. Then either $G$ is a double-wheel with centre $Z$ or $G$ is a double-wheel of triangles with centre $Z$. In either case, the rim-length equals $h(\mathcal{O})$.
Proof.
Since all bags of $\mathcal{O}$ are bad, all bags are either $K_{2}$’s or triangles. Moreover, if there is a bag that is a triangle, then its neighbouring bags are triangles, too. It follows that if some bag is a triangle, then all bags are triangles. In this case, $G$ is a double-wheel of triangles with centre $Z$. Then $h(\mathcal{O})$ equals the number of triangle-bags of $\mathcal{O}$, and hence equals the rim-length. Otherwise, every bag is a $K_{2}$. In this case, $G$ is a double-wheel with centre $Z$. Then $h(\mathcal{O})$ equals the number $K_{2}$-bags of $\mathcal{O}$, and hence equals the rim-length. ∎
Lemma 8.39.
Let $G$ be a double-wheel with rim-length $\geqslant 4$ and centre $Z$. Then $G$ has two tetra-separations that cross with all links of size one and centre equal to $Z$.
Proof.
Let $O$ denote the cycle $G-Z$. Let $e_{1},f_{1},e_{2},f_{2}$ be the first four edges on $O$. Let $(A_{i},B_{i})$ be the tetra-separation of $G$ with separator $\{e_{i},f_{i}\}\cup Z$. Then $(A_{1},B_{1})$ crosses $(A_{2},B_{2})$ with the four edges as links and with centre equal to $Z$. ∎
Lemma 8.40.
Let $G$ be a double-wheel of triangles with rim-length $\geqslant 4$ and centre $Z$. Then $G$ has two tetra-separations that cross with all links of size one and centre equal to $Z$.
Proof.
Since $G$ is a double-wheel of triangles with rim-length $\geqslant 4$, the graph $G-Z$ has a cycle-decomposition $\mathcal{O}=(O,(G_{t})_{t\in O})$ with $|O|\geqslant 4$ such that all bags are triangles. Let $(A,B)$ and $(C,D)$ be crossing 2-separations of $O$. Let $\hat{A}:=Z\cup\bigcup_{t\in A}V(G_{t})$ and define $\hat{B},\hat{C},\hat{D}$ similarly. Then $(\hat{A},\hat{B})$ and $(\hat{C},\hat{D})$ are tetra-separations of $G$ that cross as desired. ∎
Key Lemma 8.41.
Let $G$ be a 4-angry graph. Then the following assertions are equivalent:
  1. (i)
    $G$ has two tetra-separations that cross with all links of size one;
  2. (ii)
    $G$ has a block-bagel $\mathcal{O}$ with $h(\mathcal{O})\geqslant 4$ all whose bags are bad;
  3. (iii)
    $G$ is a double-wheel with rim-length $\geqslant 4$ or a double-wheel of triangles with rim-length $\geqslant 4$.
Proof.
(i) $\Rightarrow$ (ii). Assume that $G$ has two tetra-separations $(A,B)$ and $(C,D)$ that cross with all links of size one. Let $Z$ be the centre of $(A,B)$ and $(C,D)$, which is equal to the vertex-centre and has size two, by the (Crossing Lemma). (4.1). Let $\mathcal{O}$ be the block-bagel induced by $(A,B)$ and $(C,D)$. Then we have the (Block-bagel setting). (8.4). By the (Helpful Lemma). (8.6), we have $h(\mathcal{O})\geqslant 4$. If $\mathcal{O}$ has a good bag, then $G$ has a totally-nested tetra-separation by 8.27, contradicting that $G$ is 4-angry. Hence, every bag of $\mathcal{O}$ is bad.
(ii) $\Rightarrow$ (iii). Follows from Lemma 8.38.
(iii) $\Rightarrow$ (i). Follows from Lemma 8.39 and Lemma 8.40. ∎
9. Cycles of 2-connected torsos
One of the outcomes of Theorem 1 is a cycle of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices. To investigate how this outcome arises, we assume that we are given two tetra-separations $(A,B)$ and $(C,D)$ of a 4-connected graph $G$ that cross with empty vertex-centre. In order to construct a cycle-decomposition for analysing the cyclic structure that is determined by $(A,B)$ and $(C,D)$, we first use the corner-diagram to divide $G$ into four smaller pieces (Section 9.1). We then apply the Tutte-decomposition to all four pieces and merge them to obtain a cycle-decomposition of $G$ (Section 9.2). In Section 9.3, we summarise assumptions shared by a fair share of the lemmas to come. Section 9.4 provides machinery that exploits the structure of the cycle-decomposition to find independent paths. We use this machinery in Section 9.5 to characterise which bags of the cycle-decomposition give rise to totally-nested tetra-separations (9.41). Section 9.6 shows that these totally-nested tetra-separations form a splitting star as in Theorem 1, and determines the structure of its torso (9.50). Finally, Section 9.7 characterises the special case where additionally all tetra-separations of the graph are crossed.
9.1. Bananas
Let $G$ be a 4-connected graph. Assume that two tetra-separations $(A,B)$ and $(C,D)$ of $G$ cross with all links of size two. Recall that the centre is empty by the (Crossing Lemma). (4.1). We obtain the graph ${\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{G^{*}}}$ from $G$ by adding edges, as follows. For each $L$-link $\{x,y\}$ and each adjacent corner $K=LS$ (where $L,S$ are sides), we join the unique vertices in $\hat{x}\cap S$ and $\hat{y}\cap S$ by an edge, which we call the $K$-edge for the $L$-link; see Figure 9.1. We will only use $G^{*}$ when $(A,B)$ and $(C,D)$ are clear from context. The expanded corner for $AC$ is obtained from the corner $(A\smallsetminus B)\cap(C\smallsetminus D)$ for $AC$ by adding the endvertices of all edges contained in the two adjacent links. The banana for $AC$, or $AC$-banana, is the subgraph of $G^{*}$ induced by the expanded corner for $AC$. The expanded corners and bananas are defined similarly for the other three corners.
Figure 9.1. The edges added to $G$ to obtain $G^{*}$ are coloured red and green. The red edge is the $AC$-edge for the $A$-link.
Lemma 9.1.
Let $G$ be a 4-connected graph. Let $(A,B)$ and $(C,D)$ be tetra-separations of $G$ that cross with all links of size two. Let $K$ be a corner. Then the expanded corner for $K$ has size $\geqslant 4$.
Proof.
Assume $K=AC$ without loss of generality. Let $L$ be the union of the $A$-link and the $C$-link, and note that $|L|=4$. For each element $x$ in $L$, we define $v_{x}:=x$ if $x$ is a vertex, and $v_{x}:=u$ if $x$ is an edge and $u$ is the endvertex of $x$ in $(B\smallsetminus A)\cup(D\smallsetminus C)$. By the matching-condition for $(A,B)$ and $(C,D)$, $x\mapsto v_{x}$ is an injection that maps elements in $L$ to vertices of the expanded corner for $AC$. So, the expanded corner for $AC$ has size $\geqslant 4$. ∎
Lemma 9.2.
Let $G$ be a 4-connected graph. Let $(A,B)$ and $(C,D)$ be tetra-separations of $G$ that cross with all links of size two. Let $K$ be a corner. Then the following assertions are equivalent:
  1. (i)
    the $K$-edges for the links adjacent to $K$ are distinct;
  2. (ii)
    the corner $K$ is not potter.
Proof.
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). Say $K$ is the corner for $AC$. Let $\beta$ denote the $AC$-banana. Let $e(A)$ denote the $AC$-edge for the $A$-link, and let $e(C)$ denote the $AC$-edge for the $C$-link. Assume that $e(A)=e(C)=:e$. We will show that $AC$ is potter. If both endvertices of $e$ lie in the corner for $AC$, then each link adjacent to $AC$ consists of a matching of size two with the endvertices of $e$ as one side of the matching; see Figure 9.2. As $G$ is 4-connected, the corner for $AC$ cannot contain any other vertices besides the endvertices of $e$. But then both endvertices of $e$ have degree $\leqslant 3$ in $G$, contradicting that $G$ is 4-connected. So we may assume that $e$ has an endvertex $v$ that lies in the $A$-link, say. Then the $C$-link contains an edge $f$ that ends in $v$. Hence $f$ dangles from the $A$-link through the $C$-link, which implies that the corner for $AC$ is potter by the (Crossing Lemma). (4.1).
$\neg$ (ii) $\Rightarrow$ $\neg$ (i) holds by definition. ∎
Figure 9.2. The situation in the proof of Lemma 9.2
Lemma 9.3.
Let $G$ be a 4-connected graph. Let $(A,B)$ and $(C,D)$ be tetra-separations of $G$ that cross with all links of size two. The bananas of the four corners are 2-connected.
Proof.
[…]
Figure 9.3. The construction of $(\hat{X},\hat{Y})$ in the proof of Lemma 9.3. The extensions of $X$ to $\hat{X}$ and of $Y$ to $\hat{Y}$ are indicated by the dashed lines.
9.2. Tutte-bananas and Tutte-bagels
A Tutte-bagel of $G$ is a cycle-decomposition $(O,\mathcal{G})$ of $G$ with $\mathcal{G}=(G_{t}:t\in V(O))$ that satisfies all of the following conditions:
  1. (i)
    every torso is 3-connected, a 4-cycle or a triangle;
  2. (ii)
    all adhesion sets have size two;
  3. (iii)
    whenever two triangle-torsos are neighbours, the adhesion-graph in between them is a $K_{2}$.
Lemma 9.4.
Let $G$ be a 4-connected graph, and let $(O,\mathcal{G})$ be a Tutte-bagel of $G$.
  1. (i)
    If a torso is 3-connected or a 4-cycle, then its adhesion sets are disjoint.
  2. (ii)
    If a torso is a triangle, then its adhesion sets are distinct but share exactly one vertex.
  3. (iii)
    If a torso is a 4-cycle, then the two neighbouring torsos must be 3-connected.∎
Let $G$ be a graph. A 2-separation $(A,B)$ of $G$ is totally-nested if it is nested with every 2-separation of $G$. This is equivalent to saying that $(A,B)$ is half-connected and at least one of $G[A]$ and $G[B]$ is 2-connected [Tridecomp, Corollary 3.3.3]. Assume now that $G$ is 2-connected and $N$ denotes its set of totally-nested 2-separations. Then the tree-decomposition $\mathcal{T}=\mathcal{T}(N)$ defined by $N$ is the Tutte-decomposition of $G$. Every torso of $\mathcal{T}$ is 3-connected, a cycle or a $K_{2}$. Tutte introduced his decomposition in [TutteCon]. A modern proof is included in the appendix of [Tridecomp]. The Tutte-decomposition $\mathcal{T}=(T,\mathcal{V})$ has the following two properties:
  1. (i)
    Every adhesion-set has size two.
  2. (ii)
    Every adhesion-set between two neighbouring cycle-torsos is spanned by an edge in $G$.
  3. (iii)
    Every $K_{2}$-torso neighbours at least three torsos, none of which are $K_{2}$’s.
These properties give an alternative characterisation of the Tutte-decomposition.
Definition 9.5 (Tutte-bananas).
Let $G$ be a 4-connected graph. Let $(A,B)$ and $(C,D)$ be tetra-separations of $G$ that cross with all links of size two. Let $K$ be one of the four corners. Recall that the $K$-banana is 2-connected by Lemma 9.3. The $K$-Tutte-banana is the Tutte-decomposition of the $K$-banana. The $K$-Tutte-path is the decomposition-tree of the $K$-Tutte-banana.
Lemma 9.6.
Tutte-bananas are path-decompositions all whose torsos are either 3-connected graphs, 4-cycles or triangles.
Proof.
This follows from $G$ being 4-connected. ∎
Lemma 9.7.
Let $G$ be a 4-connected graph. Let $(A,B)$ and $(C,D)$ be tetra-separations of $G$ that cross with all links of size two. Let $K$ be one of the four corners, and consider an adjacent $L$-link $\Lambda$. Let $(T,\mathcal{V})$ be the $K$-Tutte-banana.
  1. (i)
    If $\Lambda$ consists of two vertices, then there is a unique node $t$ of $T$ with $\Lambda\subseteq V_{t}$. Moreover, the torso of $V_{t}$ is 3-connected, and $t$ is an end-node of the path $T$.
  2. (ii)
    If $\Lambda$ contains an edge, then there is unique node $t$ of $T$ with $\hat{\Lambda}=V_{t}$. Moreover, the torso of $V_{t}$ is a 4-cycle or a triangle, and $t$ is an end-node of the path $T$.
Proof.
Say $K$ is the $AC$-corner and $\Lambda$ is the $A$-link.
(i). Let $u,v$ be the two vertices in $\Lambda$. Recall that $uv$ is an edge in $G^{*}$. So there is at least one bag $V_{t}\in\mathcal{V}$ that contains both $u$ and $v$. Let $V_{s}\in\mathcal{V}$ contain the two ends of the $AC$-edge for the $C$-link, and choose $V_{s}$ to minimise the distance from $s$ to $t$ in $T$. Recall that $T$ is a path by Lemma 9.6. If $t$ is not an end of $T$, there is $V_{r}\in\mathcal{V}$ with $t\in\mathring{r}Ts$ and $V_{r}\smallsetminus V_{t}$ non-empty, so every vertex in $V_{r}\smallsetminus V_{t}$ is separated from a vertex outside the $AC$-banana by the adhesion-set of $rt$ which has size two, contradicting 4-connectivity. Hence $t$ is an end of $T$. Similarly, either $u$ or $v$ avoids the adhesion-set of $V_{t}$, say $u$. If the torso of $V_{t}$ is a cycle, then $u\in S(C,D)$ violates the degree-condition. Thus the torso is 3-connected.
(ii). Since $\Lambda$ contains an edge $e$, the subgraph $H$ of $G^{*}$ induced by $\hat{\Lambda}$ is either a triangle or a 4-cycle, depending on whether the other element of $\Lambda$ besides $e$ is a vertex $v$ or an edge $f$, respectively. If $H$ is a triangle, then $v\in D\smallsetminus C$ has degree two in the $AC$-banana. Hence $H$ is a leaf-bag of $(T,\mathcal{V})$, and no other bag of $(T,\mathcal{V})$ includes $\Lambda$. ∎
The node $t$ of $T$ as in Lemma 9.7 is said to represent the $L$-link in $T$, and in the $K$-Tutte-banana.
Lemma 9.8.
Let $G$ be a 4-connected graph. Let $(A,B)$ and $(C,D)$ be tetra-separations of $G$ that cross with all links of size two. Let $K$ be one of the four corners, and let $(T,\mathcal{V})$ be its Tutte-banana. The two links adjacent to $K$ are represented by distinct end-nodes of $T$, unless $T$ has only one node. In the latter case, both links consist of vertices only, and the unique torso of $(T,\mathcal{V})$ is 3-connected.
Proof.
Without loss of generality, $K$ is the corner $AC$. Both the $A$-link and the $C$-link are represented by end-nodes $a$ and $c$ of $T$, respectively, by Lemma 9.7. Assume now that $a=c$. Then $T$ has no edge, since otherwise the adhesion-set of this edge would form a 2-separator of $G$, contradicting 4-connectivity. Hence $|T|=1$, and in particular the corner $AC$ is not potter.
We claim that both the $A$-link and the $C$-link consist of vertices only. Indeed, if the $A$-link contains an edge, then the $AC$-edge $e$ for the $A$-link is distinct from the $AC$-edge for the $C$-link by Lemma 9.2, so the ends of $e$ together form a 2-separator of the $AC$-banana. This 2-separator gives rise to a 2-separation of the $AC$-banana that cuts off the triangle-torso or 4-cycle-torso at $a$ by Lemma 9.7, and this 2-separation is totally-nested as its separator is spanned by the edge $e$. Hence the 2-separation corresponds to an edge of $T$, a contradiction. ∎
Definition 9.9 (Tutte-bagel induced by crossing tetra-separations).
Let $G$ be a 4-connected graph. Let $(A,B)$ and $(C,D)$ be tetra-separations of $G$ that cross with all links of size two. For each corner $K$ let $\mathcal{T}^{K}=(T^{K},\mathcal{V}^{K})$ denote the $K$-Tutte-banana. Recall that $T^{K}$ is a path by Lemma 9.6.
We obtain $O$ from the disjoint union $\bigsqcup_{K}T^{K}$, where $K$ ranges over the four corners, by doing the following for each link $L$: for both corners $K$ and $K^{\prime}$ adjacent to $L$ we consider the nodes $t$ and $t^{\prime}$ that represent the $L$-link in $T^{K}$ and $T^{K^{\prime}}$, respectively, and
  • $\cdot$
    we join the nodes $t$ and $t^{\prime}$ by an edge $)𝑜𝐿{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{o(L)}}$ in $O$ if the torsos associated with these nodes in their respective Tutte-bananas are 3-connected;
  • $\cdot$
    we otherwise identify the nodes $t$ and $t^{\prime}$ to a single node $)𝑜𝐿{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{o(L)}}$ of $O$.
We say that $o(L)$ represents the $L$-link in $O$. For each node $t$ of $O$ that is not of the form $o(L)$, we define $G_{t}:=G[V^{K}_{t}]$ where $K$ is the unique corner with $t\in T^{K}$. For each node of $O$ of the form $o(L)$ for some link $L$, we consider the two nodes $t\in T^{K}$ and $t^{\prime}\in T^{K^{\prime}}$ that were identified to obtain $o(L)$, and we let $G_{o(L)}:=G[V^{K}_{t}]$ which is independent of $K$ by Lemma 9.7. We refer to the pair $(O,(G_{t}:t\in V(O))$ as the Tutte-bagel induced by $(A,B)$ and $(C,D)$. It remains to show that this really is a Tutte-bagel.
Lemma 9.10.
Let $G$ be a 4-connected graph. Let $(A,B)$ and $(C,D)$ be tetra-separations of $G$ that cross with all links of size two. The Tutte-bagel induced by $(A,B)$ and $(C,D)$ really is a Tutte-bagel.
Proof.
Let $\mathcal{O}=(O,(G_{t}:t\in V(O))$ denote the Tutte-bagel induced by $(A,B)$ and $(C,D)$. First, we show that $\mathcal{O}$ is a cycle-decomposition. That $O$ is a cycle follows from Lemma 9.8.
(H1). Recall that there are no jumping edges and no diagonal edges by the (Crossing Lemma). (4.1). Every vertex and edge of $G$ is included in some expanded corner, and hence is included in the bag of some Tutte-banana. Hence every vertex and edge of $G$ appears in some $G_{t}$.
(H2). Let $v\in G$ be a vertex and consider $O_{v}:=O[\,\{t\in V(O):v\in G_{t}\}\,]$.
If $v$ avoids $\hat{\Lambda}$ for every link $\Lambda$, then $v$ lies in a unique banana, and so $O_{v}$ is a subpath of the decomposition-path of the corresponding Tutte-banana. If $v$ lies in $\hat{\Lambda}$ for precisely one $L$-link $\Lambda$, then $v$ lies in the bananas for the two corners adjacent to $L$, and so $O_{v}$ is a path. Otherwise $v$ lies in $\hat{\Lambda}_{1}\cap\hat{\Lambda}_{2}$ for two adjacent $L_{i}$-links $\Lambda_{i}$. Then $v$ lies in the corner $K$ in between the two links. Then $v$ lies in all bags of the $K$-Tutte-banana, and in the unique bags at the nodes representing the $L_{i}$-links in the Tutte-bananas of the corners adjacent to $K$.
Let $H_{t}$ be a torso of $\mathcal{O}$. If $t$ represents some link, then $H_{t}$ is a 4-cycle or a triangle. Otherwise $H_{t}$ stems from a Tutte-banana. Then $H_{t}$ is 3-connected, a 4-cycle or a triangle, by Lemma 9.6.
Since every adhesion-set between two neighbouring triangle-torsos of a Tutte-banana is spanned by an edge in $G$, the same is true for $\mathcal{O}$. ∎
Lemma 9.11.
Let $G$ be a 4-connected graph. Let $(A,B)$ and $(C,D)$ be tetra-separations of $G$ that cross with all links of size two. Let $\mathcal{O}$ be the Tutte-bagel induced by $(A,B)$ and $(C,D)$. Every $K$-edge for an $L$-link spans a unique adhesion-set of $\mathcal{O}$.∎
9.3. Settings
The following settings are supported by Figure 9.4.
Setting 9.12 (Tutte-bagel setting).
Let $G$ be a 4-connected graph. Let $(A,B)$ and $(C,D)$ be tetra-separations of $G$ that cross with all links of size two. Let $\mathcal{O}=(O,(G_{t}:t\in V(O)))$ be the Tutte-bagel induced by $(A,B)$ and $(C,D)$, with torsos being denoted by $H_{t}$. Write $O=:t(0),t(1),\ldots,t(m),t(0)$ with indices in $\mathbb{Z}_{m+1}$. Assume that the representatives of the links appear in this enumeration in the order $o(C),o(B),o(D),o(A)$. For each node $t$ of $O$ let $U^{\prime}_{t}:=V(G_{t})$ and $W^{\prime}_{t}:=\bigcup\,\{\,V(G_{s}):s\in V(O-t)\,\}$. Let $(U_{t},W_{t})$ denote the right-left-shift of $(U^{\prime}_{t},W^{\prime}_{t})$.
Setting 9.13 (Tutte-bagel torso setting).
Assume the (Tutte-bagel setting). (9.12). Assume that the torso $H_{t(0)}$ of $\mathcal{O}$ is 3-connected; so $H_{t(0)}$ stems from the $AC$-banana. Let $\{x^{\prime}_{1},x^{\prime}_{2}\}$ be the adhesion set of $t(-1)t(0)$ and let $\{y^{\prime}_{1},y^{\prime}_{2}\}$ be the adhesion set of $t(0)t(1)$. Let $(U^{\prime},W^{\prime}):=(U^{\prime}_{t(0)},W^{\prime}_{t(0)})$ and $(U,W):=(U_{t(0)},W_{t(0)})$. Note that $S(U^{\prime},W^{\prime})=\{x^{\prime}_{1},x^{\prime}_{2},y^{\prime}_{1},y^{\prime}_{2}\}$. Let $x_{i}\in S(U,W)$ be equal or incident to $x^{\prime}_{i}$, and let $x^{W}_{i}$ be the unique vertex in $\hat{x}_{i}\cap W$. Define $y_{i}$ and $y^{W}_{i}$ similarly.
Figure 9.4. Left: The (Tutte-bagel setting). (9.12). Right: the (Tutte-bagel torso setting). (9.13). The separator $S(U^{\prime},W^{\prime})$ is blue, while $S(U,W)$ is red. We have $x^{W}_{i}=x^{\prime}_{i}$ for both $i=1,2$ and $y^{W}_{1}=y^{\prime}_{1}$ (say), so these three vertices are red. The vertex $y^{W}_{2}$ is green.
Let $\mathcal{O}$ be a Tutte-bagel of a 4-connected graph $G$. Let $H_{t}$ be a torso of $\mathcal{O}$. Assume first that $H_{t}$ is a 4-cycle. Then $H_{t}$ has exactly two edges whose endvertices lie in distinct adhesion-sets of $H_{t}$. We call these edges the free edges of $H_{t}$. Assume now that $H_{t}$ is a triangle. Then the adhesion-sets of $H_{t}$ share exactly one vertex $v$. We call $v$ the tip of the triangle $H_{t}$. There is a unique edge of $H_{t}$ whose endvertices lie in distinct adhedion-sets of $H_{t}$. We call this edge the free edge of the triangle $H_{t}$.
Lemma 9.14.
Assume the (Tutte-bagel torso setting). (9.13). The vertex-set $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at $t(0)\,t(1)$ if the torso $H_{t(1)}$ is 3-connected, and it is the adhesion-set at $t(1)\,t(2)$ otherwise. In the latter case, the torso at $t(2)$ is not a 4-cycle.
Proof.
The former case is straightforward, so it remains to consider the case where $H_{t(1)}$ is not 3-connected. If the torso $H_{t(1)}$ is a 4-cycle, then the $y_{i}$ are the free edges of $H_{t(1)}$, and so $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at $t(1)t(2)$. Hence we may assume that the torso $H_{t(1)}$ is a triangle. The next torso $H_{t(2)}$ cannot be 4-cycle by Lemma 9.4, so it is 3-connected or a triangle.
Assume first that $H_{t(2)}$ is 3-connected. Since $H_{t(2)}$ shares a vertex with $H_{t(1)}$, say $y^{\prime}_{1}$, the vertex-set of $H_{t(2)}$ must be included in $C$. Hence the $\geqslant 2$ neighbours of $y^{\prime}_{1}$ in the bag $G_{t(2)}$ lie in $W^{\prime}\smallsetminus U^{\prime}$. Thus $y^{W}_{1}=y^{\prime}_{1}$. Since $H_{t(1)}$ is a triangle, $y_{2}$ is its free edge, so $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at $t(1)\,t(2)$.
Otherwise it is a triangle, so the adhesion-graph at $t(1)\,t(2)$ is a $K_{2}$-subgraph of $G$. Then the $y^{\prime}_{i}$ incident to the free edge of the triangle $H_{t(1)}$ corresponds to an edge $y_{i}$. The $y^{\prime}_{j}$ that is the tip of $H_{t(1)}$ has $y^{W}_{i}$ as a neighbour and is incident to the free edge of $H_{t(2)}$, so $y^{\prime}_{j}=y_{j}$. Hence $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at $t(1)\,t(2)$. ∎
Recall that $O_{v}=O[\,\{t\in V(O):v\in G_{t}\}\,]$ for a vertex $v\in V(G)$ in the context of a Tutte-bagel $\mathcal{O}$ of $G$.
Lemma 9.15.
Let $G$ be a 4-connected graph, and let $\mathcal{O}=(O,\mathcal{G})$ be a Tutte-bagel of $G$. Then both of the following assertions hold:
  1. (i)
    Every two triangle-torsos of $\mathcal{O}$ have distinct tips.
  2. (ii)
    Let $P\subseteq O_{v}$ be a path with ends $r\neq t$ such that $H_{r},H_{t}$ are triangles. Then $P=rt$ or $P=rst$ where $H_{s}$ is a triangle.
Proof.
(i). Suppose for a contradiction that two triangle-torsos of $\mathcal{O}$ have the same tip. Say the torsos are $H_{r}$ and $H_{t}$, and $v$ is their shared tip. Let $e$ be the free edge of $H_{r}$, and let $e^{\prime}$ be the free edge of $H_{t}$. Then $\{v,e,e^{\prime}\}$ is a mixed-3-separator, contradicting 4-connectivity by Lemma 6.1.
(ii). First, assume that there is a node $p$ on $P$ other than $r,t$ with $v\in G_{p}$ such that the torso $H_{p}$ is a 4-cycle or 3-connected. That the adhesion-sets at $G_{p}$ are disjoint then contradicts that $v$ lies in all three bags $G_{r},G_{p},G_{t}$. Otherwise, all torsos of all nodes on $P$ are triangles. Let $e$ be the edge of $P$ incident with $r$, and let $e^{\prime}$ be the edge of $P$ incident with $t$. By (i), if $|P|\geqslant 4$ then the adhesion sets of $e$ and $e^{\prime}$ are disjoint, which contradicts that $v$ lies in all bags of nodes in $P$ by $P\subseteq O_{v}$. Otherwise $|P|\leqslant 3$ and we get the statement. ∎
9.4. N-Paths lemmata
Lemma 9.16.
Let $G$ be a 4-connected graph with a Tutte-bagel $\mathcal{O}=(O,\mathcal{G})$ and decomposition-cycle $O=t(0),\ldots t(m),t(0)$. Let $0\leqslant i<j\leqslant m$. Let $\pi_{\ell}$ denote the adhesion-set of $t(\ell)t(\ell+1)$ for all relevant indices $\ell$. There exist two disjoint $\pi_{i}$$\pi_{j}$ paths in $G_{t(i+1)}\cup\cdots\cup G_{t(j)}$.
Proof.
For each torso $H_{t(\ell)}$ we use that it is 2-connected to find two disjoint paths $P_{\ell},Q_{\ell}$ between its two adhesion-sets. These paths use no torso-edges, so they are subgraphs of their bags and of $G$. Hence combining the paths in the bags $G_{t(i+1)},\ldots,G_{t(j)}$ yields a pair of desired $\pi_{i}$$\pi_{j}$ paths. ∎
Lemma 9.17.
Let $\mathcal{O}$ be a Tutte-bagel of a 4-connected graph $G$. Let $H_{t}$ be a 3-connected torso of $\mathcal{O}$ with adhesion sets $\pi$ and $\pi^{\prime}$. The two vertices in $\pi$ are linked by a path in $G_{t}-\pi^{\prime}$, unless $H_{t}=K_{4}$.
Proof.
Assume that the two vertices $x_{1},x_{2}$ in $\pi$ lie in distinct components $C_{i}$ of $G_{t}-\pi^{\prime}$. Then $V(C_{i})=\{x_{i}\}$ and $G_{t}-\pi^{\prime}=C_{1}\cup C_{2}$, as $G$ is $4$-connected. Hence $G_{t}$ has only 4 vertices. Since $H_{t}$ is 3-connected, it is a $K_{4}$. ∎
Lemma 9.18.
Assume the (Tutte-bagel setting). (9.12). Let $K$ be a corner. If the $K$-Tutte-banana has no 3-connected torsos, then all its torsos are triangles, all its adhesion-graphs are $K_{2}$-subgraphs of $G$, and the two links adjacent to $K$ are represented by triangle-torsos in the Tutte-banana.
Proof.
Since the Tutte-banana $\mathcal{T}$ for $K$ has no 3-connected torsos, it also has no 4-cycles as torsos by Lemma 9.4. Hence all torsos of $\mathcal{T}$ are triangles; in particular, both links adjacent to $K$ are represented by triangle-torsos of $\mathcal{T}$. Again by Lemma 9.4, it follows that all adhesion-graphs of $\mathcal{T}$ are $K_{2}$-subgraphs of $G$. ∎
Lemma 9.19 (One-Path Lemma).
Assume the (Tutte-bagel setting). (9.12). Let $a_{1}a_{2}$ be the $AD$-edge for the $A$-link. Then there is an $a_{1}$$a_{2}$ path $P$ in $G[A\cap D]$ that uses at most one vertex in the $D$-link. Moreover, if $AD$ is not potter, then for each vertex $d$ in the $D$-link there is such a path $P$ that avoids $d$. And if both $a_{1},a_{2}$ lie in the $A$-link, then $P$ does not use the edge $a_{1}a_{2}$ (if it exists).
Proof.
If the corner $AD$ is potter, then $a_{1}a_{2}$ is an edge in $G$, and the path $a_{1}a_{2}$ is as desired. So we may assume that the corner $AD$ is not potter. Let $\mathcal{T}$ denote the $AD$-Tutte-banana. We distinguish two cases.
[…]
Suppose second that $\mathcal{T}$ has no 3-connected torsos. Then all torsos of $\mathcal{T}$ are triangles, both the $A$-link and the $D$-link are represented by triangle-torsos, and $a_{1}a_{2}\in E(G)$ by Lemma 9.18. In particular, exactly one of $a_{1},a_{2}$ lies in the $A$-link, say it is $a_{1}$. Now $a_{2}$ cannot lie in the $D$-link, since otherwise some edge dangles from the $D$-link through the $A$-link which by the (Crossing Lemma). (4.1) implies that the corner $AD$ is potter, a contradiction. Hence the path $a_{1}a_{2}$ is as desired. ∎
Lemma 9.20 (Two-Paths Lemma).
Assume the (Tutte-bagel torso setting). (9.13). Let $\ell\geqslant 1$ be maximum such that $t(\ell)$ stems from the BD-Tutte-banana. There exist two independent $y^{W}_{1}$$y^{W}_{2}$ paths $P_{1},P_{2}$ in $G_{t(1)}\cup\cdots\cup G_{t(\ell)}$ not using the edge $y^{\prime}_{1}y^{\prime}_{2}$ (if it exists) such that:
  1. (i)
    $P_{1}\cup P_{2}$ uses at most one vertex of the $D$-link.
  2. (ii)
    If the corner $BD$ is not potter, then for every vertex $d$ in the $D$-link we can choose the two paths so that $d\notin P_{1}\cup P_{2}$.
Proof.
By Lemma 9.14, there is $i\in\{1,2\}$ such that $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at $t(i-1)t(i)$. Moreover, the torso $H_{t(i)}$ is not a 4-cycle. We consider two cases, supported by Figure 9.5.
Suppose first that the torso at $t(i)$ is 3-connected. Then we find two $y^{W}_{1}$$y^{W}_{2}$ paths $Q_{1},Q_{2}$ in the torso at $t(i)$ that avoid the torso-edge $y^{W}_{1}y^{W}_{2}$. Note that $G_{t(i)}\subseteq G[C]$. If neither $Q_{j}$ uses the other torso-edge $z_{1}z_{2}$, then we are done. So we may assume that $Q_{1}$, say, uses $z_{1}z_{2}$. Let $e=u_{1}u_{2}$ be the $BD$-edge for the $B$-link. Let $k\geqslant i$ be maximum such that $t(k)$ stems from the $BC$-Tutte-banana. By Lemma 9.16, we find two disjoint $\{z_{1},z_{2}\}$$\{u_{1},u_{2}\}$ paths $R_{1},R_{2}$ in $G_{t(i+1)}\cup\cdots\cup G_{t(k)}$, taking the trivial paths if $k=i$. We apply the (One-Path Lemma). (9.19) (with $B$ in place of $A$) to the $BD$-edge $u_{1}u_{2}$ for the $B$-link to find a $u_{1}$$u_{2}$ path $R_{3}$ in $G[B\cap D]$ that uses at most one vertex in the $D$-link. Moreover, if $BD$ is not potter, then we find $R_{3}$ so that it avoids any prescribed vertex $d$ in the $D$-link. Then $P_{1}:=(Q_{1}-z_{1}z_{2})\cup R_{1}\cup R_{2}\cup R_{3}$ and $P_{2}:=Q_{2}$ are as desired.
Suppose otherwise that the torso at $t(i)$ is a triangle. Then $i=2$ and the torso at $t(1)$ is not 3-connected, by Lemma 9.14. Since the torso at $t(2)$ is a triangle, the torso at $t(1)$ cannot be a 4-cycle by Lemma 9.4, so it must be a triangle as well. Then the adhesion-graph of $t(1)t(2)$ is a $K_{2}$-subgraph of $G$ with vertices $y^{W}_{1},y^{W}_{2}$. Then we let $P_{2}:=Q_{2}:=y^{W}_{1}y^{W}_{2}$ and construct $P_{1}:=R_{1}\cup R_{2}\cup R_{3}$ as in the first case with the (One-Path Lemma). (9.19). ∎
Figure 9.5. Two paths found by Lemma 9.20. The vertices $y^{W}_{i}$ are blue. The path $P_{2}=Q_{2}$ is red. The path $P_{1}$ is green. The green vertices in the $B$-link are $u_{1},u_{2}$.
Lemma 9.21 (Three-Paths Lemma).
Assume the (Tutte-bagel torso setting). (9.13). There exist three independent $\{x^{W}_{1},y^{W}_{1}\}$$\{x^{W}_{2},y^{W}_{2}\}$ paths $P_{1},P_{2}$ and $P_{3}$ in $G[W]-x^{\prime}_{1}x^{\prime}_{2}-y^{\prime}_{1}y^{\prime}_{2}$ such that two $P_{i}$ link $y^{W}_{1}$ to $y^{W}_{2}$ and one $P_{i}$ links $x^{W}_{1}$ to $x^{W}_{2}$.
Proof.
Recall that not both corners $AD,BD$ are potter by 4.7.
Suppose first that the corner $AD$ is not potter. We apply the (Two-Paths Lemma). (9.20) to find two independent $y^{W}_{1}$$y^{W}_{2}$ paths $P_{1},P_{2}$ in $G[W]-y^{\prime}_{1}y^{\prime}_{2}$ such that $P_{1}\cup P_{2}$ uses at most one vertex in the $D$-link. Then we apply the (One-Path Lemma). (9.19) to find an $x^{W}_{1}$$x^{W}_{2}$ path $P_{3}\subseteq G[W]-x^{\prime}_{1}x^{\prime}_{2}$ that avoids $P_{1}\cup P_{2}$.
Suppose otherwise that the corner $BD$ is not potter. We proceed as in the first case, except that we find $P_{3}$ first and then find $P_{1},P_{2}$ afterwards. ∎
9.5. Totally-nested tetra-separations from good bags
Definition 9.22 (Good bag, bad bag).
Let $\mathcal{O}$ be a Tutte-bagel of a 4-connected graph $G$. A torso $H_{t}$ of $\mathcal{O}$ is good if it is 3-connected and it satisfies at least one of the following conditions:
  1. (i)
    $|H_{t}|\geqslant 6$;
  2. (ii)
    some neighbouring torso is a 4-cycle;
  3. (iii)
    both neighbouring torsos are triangles; or
  4. (iv)
    $|H_{t}|=5$ and some neighbouring torso is a triangle.
If $H_{t}$ fails to be good, it is bad.
Lemma 9.23.
Assume the (Tutte-bagel torso setting). (9.13). The following assertions are equivalent:
  1. (i)
    $(U,W)$ is a tetra-separation;
  2. (ii)
    $H_{t(0)}$ is good.
Proof.
Note that $|W^{\prime}\smallsetminus U^{\prime}|\geqslant 2$.
(ii) $\Rightarrow$ (i). Assume that $H_{t(0)}$ is good. If $|H_{t(0)}|\geqslant 6$, then $U^{\prime}\smallsetminus W^{\prime}$ has size $\geqslant 2$, so $(U,W)$ is a tetra-separation by 5.10. If $|H_{t(0)}|=5$ or $|H_{t(0)}|=4$, then $H_{t(0)}$ being good ensures via the neighbouring torsos that at least one or at least two vertices in $S(U^{\prime},W^{\prime})$ have exactly one neighbour in $W^{\prime}\smallsetminus U^{\prime}$, respectively, so $(U,W)$ is a tetra-separation by 5.10.
(i) $\Rightarrow$ (ii). Assume that $(U,W)$ fails to be a tetra-separation. Then by 5.10, the number of vertices in $S(U^{\prime},W^{\prime})$ with $\leqslant 1$ neighbour in $W^{\prime}\smallsetminus U^{\prime}$ fails to be $\geqslant 2-|U^{\prime}\smallsetminus W^{\prime}|$. Hence all sufficient conditions for $H_{t(0)}$ being good fail. ∎
Half-connected
Lemma 9.24.
Assume the (Tutte-bagel torso setting). (9.13) with $(U,W)$ a tetra-separation. Then $(U,W)$ is half-connected.
Proof.
If $S(U,W)$ contains an edge, then $G-S(U,W)$ has exactly two components by 4-connectivity, so $(U,W)$ is half-connected. Otherwise $(U,W)=(U^{\prime},W^{\prime})$. We will show that $G[W\smallsetminus U]$ is connected. Let $\pi$ denote the adhesion-set of $\mathcal{O}$ that is spanned by the $BD$-edge for the $D$-link by Lemma 9.11.
We claim that every vertex $v\in W\smallsetminus U$ not in $\pi$ sends a fan of two paths to $\pi$ in $G[W\smallsetminus U]$. Indeed, $\pi$ together with $\{x^{\prime}_{1},x^{\prime}_{2}\}$, say, separates $v$ from $G_{t(0)}$. Since $G$ is 4-connected, there is a fan of four paths from $v$ to $\{x^{\prime}_{1},x^{\prime}_{2}\}\cup\pi$ by Lemma 6.2. The two paths ending in $\pi$ are as desired.
So it remains to show that there is at least one vertex in $W\smallsetminus U$ besides the two vertices in $\pi$. The vertex $x_{1}$ has at least two neighbours in $W\smallsetminus U$, at least one of which must be in $A\smallsetminus B$. The neighbour in $A\smallsetminus B$ avoids $\pi\subseteq B$. ∎
3-linked
Lemma 9.25.
Let $\mathcal{O}$ be a Tutte-bagel of a 4-connected graph $G$. Let $H_{t}$ be a 3-connected torso of $\mathcal{O}$ with adhesion sets $\{a_{1},a_{2}\}$ and $\{b_{1},b_{2}\}$. Then there is an $a_{1}$$b_{1}$ path in $G_{t}-\{a_{2},b_{2}\}$.
Proof.
Assume for a contradiction that the two vertices $a_{1},b_{1}$ lie in distinct components $C,D$ of $G_{t}-\{a_{2},b_{2}\}$. Then $V(C)=\{a_{1}\}$ and $V(D)=\{b_{1}\}$, as $G$ is $4$-connected. Hence $G_{t}$ has only 4 vertices. Since $H_{t}$ is 3-connected, it is a $K_{4}$. But then $a_{1}b_{1}$ is an edge in $G_{t}$, a contradiction. ∎
Lemma 9.26.
Assume the (Tutte-bagel torso setting). (9.13). Let $P^{\prime}$ be a path in $G_{t(0)}$ that joins two vertices $v^{\prime}_{1},v^{\prime}_{2}\in S(U^{\prime},W^{\prime})$ and avoids the other two vertices in $S(U^{\prime},W^{\prime})$. Let $v_{i}\in S(U,W)$ be equal or incident to $v^{\prime}_{i}$. Then $P^{\prime}$ contains a $\hat{v}_{1}$$\hat{v}_{2}$ path $P$ that avoids the two vertices or edges in $S(U,W)$ besides $v_{1}$ and $v_{2}$.
Proof.
If $v_{i}$ is an edge, then all neighbours of $v^{\prime}_{i}$ in $G_{t(0)}$ lie in $S(U^{\prime},W^{\prime})$ except the one that is joined to $v^{\prime}_{i}$ via the edge $v_{i}$, so $P^{\prime}$ uses $v_{i}$. Hence $P^{\prime}$ contains the desired subpath. ∎
Lemma 9.27.
Assume the (Tutte-bagel torso setting). (9.13) with $(U,W)$ a tetra-separation. Let $1\leqslant k\leqslant m$ be maximum with $y^{W}_{1}\in G_{t(k)}$. Assume that $y_{2}$ is a vertex. Then exactly one of the following two assertions holds:
  1. (i)
    The torso $H_{t(k)}$ is 3-connnected, there are $\leqslant 2$ torsos in between $H_{t(0)}$ and $H_{t(k)}$, and they are triangles.
  2. (ii)
    $H_{t(k)}$ is a triangle, the two edges in $H_{t(k)}$ incident to $y^{W}_{1}$ exists in $G$ and $y^{W}_{2}\notin H_{t(k)}$. Moreover, there are exactly two triangle-torsos between $H_{t(0)}$ and $H_{t(k)}$.
Proof.
If $H_{t(1)}$ is 3-connected, we have (i). Otherwise $H_{t(1)}$ is a 4-cycle or a triangle. Since $y_{2}$ is a vertex by assumption, $H_{t(1)}$ is a triangle and $y_{2}=y^{W}_{2}$ is its tip, by Lemma 9.14. Hence $y^{W}_{1}$ is incident to the free edge of $H_{t(1)}$. If $H_{t(2)}$ is 3-connected, then we have (i). Otherwise $H_{t(2)}$ is a triangle, by Lemma 9.4. By Lemma 9.15, $y^{W}_{1}$ is the tip of $H_{t(2)}$. Hence $y^{W}_{2}$ is not contained in the adhesion-set between $H_{t(2)}$ and $H_{t(3)}$, and in particular $y^{W}_{2}$ is not contained in $H_{t(3)}$. If $H_{t(3)}$ is 3-connected, we have (i). Otherwise $H_{t(3)}$ also is a triangle. By Lemma 9.15, $y^{W}_{1}$ is incident to the free edge of $H_{t(3)}$. By Lemma 9.4, the adhesion-graph between the triangle-torsos $H_{t(2)}$ and $H_{t(3)}$ is a $K_{2}$-subgraph of $G$. Hence the two edges of $H_{t(3)}$ incident to $y^{W}_{1}$ exist in $G$. Thus we have (ii). ∎
Lemma 9.28.
Let $\mathcal{O}$ be a Tutte-bagel of a 4-connected graph $G$. Let $H_{t}$ be a 3-connected torso of $\mathcal{O}$ with adhesion-sets $\{a_{1},a_{2}\}$ and $\{b_{1},b_{2}\}$. Then either
  1. (i)
    there exist two independent $a_{1}$$b_{1}$ paths in the bag $G_{t}$, or
  2. (ii)
    the torso $H_{t}$ is a 4-wheel with rim $a_{1}a_{2}b_{1}b_{2}$ and both edges $a_{1}a_{2},b_{1}b_{2}$ are missing from $G$.
Proof.
Assume that (i) fails. Then there exists a cutvertex $v$ of $G_{t}$ that separates $a_{1}$ from $b_{1}$. Since $H_{t}$ is 3-connected and $G_{t}$ misses at most two of the torso-edges, $v$ is distinct from $a_{2}$ and $b_{2}$. Adding the edges $a_{1}a_{2}$ and $b_{1}b_{2}$ makes $v$ into the separator of a mixed-3-separation $(X,Y)$ of $H_{t}$ with $a_{1}\in X\smallsetminus Y$ and $b_{1}\in Y\smallsetminus X$. Since $S(X,Y)$ contains edges, $H_{t}-S(X,Y)$ has exactly two components. These components include the vertex-sets $\{a_{1},v,b_{2}\}$ and $\{a_{2},v,b_{1}\}$, respectively. Since $G$ is 4-connected, these inclusions are equalities, and $v$ has degree four in $H_{t}$. As $H_{t}$ is 3-connected, $H_{t}$ is a 4-wheel with rim $a_{1}a_{2}b_{1}b_{2}$. Since (i) fails by assumption, both torso-edges $a_{1}a_{2}$ and $b_{1}b_{2}$ must be missing from $G$. ∎
Lemma 9.29.
Assume the (Tutte-bagel torso setting). (9.13) with $(U,W)$ a tetra-separation. Assume that both $x_{2}$ and $y_{2}$ are vertices. There are two independent $x^{W}_{1}$$y^{W}_{1}$ paths in $G[W]-x_{2}-y_{2}$.
Proof.
Let $1\leqslant i\leqslant m$ be maximum with $y^{W}_{1}\in G_{t(i)}$. Let $1\leqslant j\leqslant m$ be minimum with $x^{W}_{1}\in G_{t(j)}$. Let $\pi_{\ell}$ denote the adhesion-set at $t(\ell)t(\ell+1)$ for all $\ell$. We consider two cases.
Case $i<j$. This case is supported by Figure 9.6. If $H_{t(i)}$ is 3-connected, we let $F_{y}$ by a $y^{W}_{1}$$\pi_{i}$ fan of two paths in $G_{t(i)}$ avoiding the vertex in $\pi_{i-1}-y^{W}_{1}$. Otherwise, by Lemma 9.27, $H_{t(i)}$ is a triangle avoiding $y^{W}_{2}$ such that both edges incident to $y^{W}_{1}$ in $H_{t(i)}$ exist in $G$, and we let $F_{y}$ denote the $y^{W}_{1}$$\pi_{i}$ fan of two paths defined by these two edges. In both cases, $F_{y}$ avoids $y_{2}$. Similarly, we obtain an $x^{W}_{1}$$\pi_{j-1}$ fan $F_{x}\subseteq G_{t(j)}$ avoiding $x_{2}$. By Lemma 9.16, we find two disjoint $\pi_{i}$$\pi_{j-1}$ paths in $G_{t(i+1)}\cup\cdots\cup G_{t(j-1)}$. Call these paths $P$ and $Q$. Then $F_{x}\cup P\cup Q\cup F_{y}$ is a union of two independent $x^{W}_{1}$$y^{W}_{1}$ paths in $G[W]-x_{2}-y_{2}$.
Figure 9.6. Case $i<j$ in the proof of Lemma 9.29. Here, $H_{t(i)}$ is 3-connected and $H_{t(j)}$ is a triangle. The unique vertex in $\pi_{i-1}-y^{W}_{1}$ is black.
Case $i\geqslant j$. This case is supported by Figure 9.7. We claim that the corner $BD$ is not potter, and assume for a contradiction that it is. Let $\pi$ denote the unique adhesion-set of the $BD$-Tutte-banana. Since $BD$ is potter, neither corner $BC$ nor $AD$ is potter by 4.7. Hence if $y_{1}$ is an edge, it neither dangles from the $C$-link through the $B$-link nor from the $B$-link through the $C$-link, by the (Crossing Lemma). (4.1). So $y^{W}_{1}\in C\smallsetminus D$, and hence $y^{W}_{1}\notin\pi$. Similarly, $x^{W}_{1}\notin\pi$. But this contradicts $i\geqslant j$, so the corner $BD$ is not potter.
We claim that $y^{W}_{1}$ lies in the $B$-link and $x^{W}_{1}$ lies in the $D$-link. Assume for a contradiction that $y^{W}_{1}$ avoids the $B$-link, say. Then $y^{W}_{1}\in C\smallsetminus D$. Let $\pi$ denote the adhesion-set defined by the $BD$-edge for the $B$-link. Since $BD$ is not potter, no edge dangles from the $D$-link through the $B$-link. Hence $\pi\subseteq(B\smallsetminus A)\cap D$. Now $y^{W}_{1}\notin\pi$ since $y^{W}_{1}\notin D$, and $x^{W}_{1}\notin\pi$ since $x^{W}_{1}\notin B\smallsetminus A$, contradicting $i\geqslant j$. So $y^{W}_{1}$ lies in the $B$-link and $x^{W}_{1}$ lies in the $D$-link.
By (Crossing Lemma). (4.1), the corners $BC$ and $AD$ are potter. Let $T$ denote the $BD$-Tutte-path. Then $o(B)$ and $o(D)$ are the ends of $T$. Note that $T$ has $\geqslant 3$ nodes.
Assume first that $o(B)$ or $o(D)$ has a neighbour in $T$ with a triangle-torso in $\mathcal{O}$. Say this is a neighbour $t$ of $o(B)$. Since $y^{W}_{1}$ is the tip of the triangle $H_{o(B)}$, the vertex $y^{W}_{1}$ is an end of the free edge $f$ of $H_{t}$ by Lemma 9.15. Let $u$ denote the other end of $f$. Let $v$ denote the tip of $H_{t}$. Note that $\{u,v\}$ is an adhesion-set, so $x^{W}_{1}\in\{u,v\}$. Hence $x^{W}_{1}$ lies in the two triangle-torsos from the potter corner $AD$ and in the triangle-torso $H_{t}$. Thus $o(D)$ and $t$ are neighbours in $T$ by Lemma 9.15. In particular, $T$ has exactly three nodes. Since $v$ lies in the two triangle-torsos $H_{t}$ and $H_{o(B)}$ while $x^{W}_{1}$ lies in the two triangle-torsos for the potter corner $AD$, we cannot have $v=x^{W}_{1}$ by Lemma 9.15. Hence $x^{W}_{1}=u$. By Lemma 9.4, the two adhesion-sets of $H_{t}$ are spanned by edges in $G$. Hence $x^{W}_{1}y^{W}_{1}$ and $x^{W}_{1}vy^{W}_{1}$ are desired paths.
Figure 9.7. Case $i\geqslant j$ in the proof of Lemma 9.29. Left: $H_{t}$ is a triangle.
Right: $H_{t}$ is 3-connected.
Assume otherwise that neither $o(B)$ nor $o(D)$ has a neighbour in $T$ with a triangle-torso. Then the torsos neighbouring $o(B)$ and $o(D)$ via $T$ are 3-connected. Since 3-connected torsos have disjoint adhesion-sets and $i\geqslant j$, we have $|T|=3$. Let $t$ denote the internal node of $T$. Let $\{x^{W}_{1},u\}$ denote the adhesion-set of the edge $o(D)t$, and let $\{y^{W}_{1},v\}$ denote the adhesion-set of the edge $o(B)t$. We find two independent $y^{W}_{1}$$x^{W}_{1}$ paths in $G_{t}$, and are done, unless $H_{t}$ is a 4-wheel with rim $x^{W}_{1}uy^{W}_{1}v$ and $x^{W}_{1}u,y^{W}_{1}v\notin E(G)$, by Lemma 9.28. But in the latter case, $u$ and $v$ have degree three in $G$, contradicting 4-connectivity. ∎
Lemma 9.30.
Assume the (Tutte-bagel torso setting). (9.13) with $(U,W)$ a tetra-separation. Then $\{x_{i},y_{j}\}$ is 3-linked around $S(U,W)$ for all $i,j\in[2]$.
Proof.
Say $i=j=1$. If $x_{2}$ or $y_{2}$ is an edge, then we are done by (6.4.1), so assume that both are vertices. By Lemma 9.25, there is an $x^{\prime}_{1}$$y^{\prime}_{1}$ path $P^{\prime}$ in $G_{t(0)}-\{x^{\prime}_{2},y^{\prime}_{2}\}$. Then $P^{\prime}$ contains a $\hat{x}_{1}$$\hat{y}_{1}$ path $P$ that avoids the vertices $x_{2}$ and $y_{2}$, by Lemma 9.26. We find two independent $x^{W}_{1}$$y^{W}_{1}$ paths in $G[W]-x_{2}-y_{2}$ using Lemma 9.29. Hence $\{x_{1},y_{1}\}$ is 3-linked around $S(U,W)$ by (6.4.3). ∎
Lemma 9.31.
Assume the (Tutte-bagel torso setting). (9.13) with $(U,W)$ a tetra-separation. Then $\{x_{1},x_{2}\}$ is 3-linked around $S(U,W)$. Similarly, $\{y_{1},y_{2}\}$ is 3-linked around $S(U,W)$.
Proof.
If $y_{1}$ or $y_{2}$ is an edge, we are done, so assume that both are vertices. There exist two independent $x^{W}_{1}$$x^{W}_{2}$ paths in $G[W]-\{y_{1},y_{2}\}$ by the (Three-Paths Lemma). (9.21). So it remains to find an $x^{\prime}_{1}$$x^{\prime}_{2}$ path in $G^{\prime}:=G_{t_{0}}-\{y^{\prime}_{1},y^{\prime}_{2}\}$ (as any such path contains a suitable $\hat{x}_{1}$$\hat{x}_{2}$ path by Lemma 9.26). Assume for a contradiction that such a path is missing. Then $H_{t(0)}$ is a $K_{4}$ by Lemma 9.17. Since both $y_{1}$ and $y_{2}$ are vertices, the torso $H_{t(1)}$ is 3-connected by Lemma 9.14. Since $H_{t(0)}$ is good, it follows that $H_{t(-1)}$ is a 4-cycle. So $x_{1}$ and $x_{2}$ are the two free edges on this 4-cycle. By 4-connectivity of $G$, both $x^{\prime}_{1}$ and $x^{\prime}_{2}$ have degree at least four. Since $H_{t(-1)}$ is a 4-cycle and $H_{t(0)}$ is a $K_{4}$, it follows that the edge $x^{\prime}_{1}x^{\prime}_{2}$ must exist in $G$, contradicting that there is no $x^{\prime}_{1}$$x^{\prime}_{2}$ path in $G^{\prime}:=G_{t_{0}}-\{y^{\prime}_{1},y^{\prime}_{2}\}$. ∎
Potter-linked
Lemma 9.32.
Assume the (Tutte-bagel torso setting). (9.13) with $(U,W)$ a tetra-separation. Then $(U,W)$ is 0-potter-linked.
Proof.
Let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of $S(U,W)$. It suffices to find five independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G$.
Suppose first that $\pi_{1}=\{x_{1},x_{2}\}$ and $\pi_{2}=\{y_{1},y_{2}\}$. We find two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths $P_{1},P_{2}$ in $G[W]$ using Lemmas 9.16 and 9.14. The torso $H_{t(0)}$ is 3-connected by assumption, and hence it contains three independent $\{x^{\prime}_{1},x^{\prime}_{2}\}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ paths by (Menger for independent paths). (6.2). None of these paths use torso-edges, so all three paths are subgraphs of the bag $G_{t(0)}$. The three $\{x^{\prime}_{1},x^{\prime}_{2}\}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ paths can be shortened to three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths $Q_{1},Q_{2},Q_{3}$ by Lemma 9.26. Hence $\pi_{1}$ is 0-potter-linked to $\pi_{2}$.
Suppose otherwise that $\pi_{1}=\{x_{1},y_{1}\}$ and $\pi_{2}=\{x_{2},y_{2}\}$, say. We find three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths $P_{1},P_{2},P_{3}$ in $G[W]-x^{\prime}_{1}x^{\prime}_{2}-y^{\prime}_{1}y^{\prime}_{2}$ using the (Three-Paths Lemma). (9.21). By Lemma 9.28, we find two independent $\{x^{\prime}_{1},y^{\prime}_{1}\}$$\{x^{\prime}_{2},y^{\prime}_{2}\}$ paths in the bag $G_{t(0)}$, which shorten to two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths $P_{4},P_{5}$ in $G_{t(0)}$ by Lemma 9.26, unless $H_{t(0)}$ is a 4-wheel and both torso-edges $x^{\prime}_{1}x^{\prime}_{2},y^{\prime}_{1}y^{\prime}_{2}$ are missing in $G$. But the latter is impossible: since $H_{t(0)}$ is good, some neighbouring torso would be a triangle or a 4-cycle, but then a vertex in the adhesion-set would have degree three in $G$, contradicting 4-connectivity. ∎
Assume the (Tutte-bagel torso setting). (9.13). We call $x_{i}$ an inward edge if $x_{i}$ is an edge and $x^{W}_{i}=x^{\prime}_{i}$; that is, if $x_{i}$ is an edge with both endvertices in $G_{t(0)}$. We call $x_{i}$ an outward edge if $x_{i}$ is an edge but not an inward edge; that is, if the endvertex $x^{W}_{i}$ of $x_{i}$ lies outside of $G_{t(0)}$.
Lemma 9.33.
Assume the (Tutte-bagel torso setting). (9.13) with $(U,W)$ a tetra-separation. Assume that $x_{1}$ is an inward edge. Then all of the following hold:
  1. (i)
    Some $y^{\prime}_{i}$ is a neighbour of $x^{\prime}_{1}$ and $y^{W}_{i}=y^{\prime}_{i}$.
  2. (ii)
    $\{x_{1},x_{2}\}$ is not $W$-weird.
Proof.
(i). Since $H_{t(0)}$ is 3-connected, the vertex $x^{\prime}_{1}$ has $\geqslant 3$ neighbours in $H_{t(0)}$. Let $u$ be a neighbour of $x^{\prime}_{1}$ in $H_{t(0)}$ distinct from $x^{\prime}_{2}$ and the end of the edge $x_{1}$ besides $x^{\prime}_{1}$. All interior vertices of $H_{t(0)}$ are contained in $U\smallsetminus W$, so $u\in\{y^{\prime}_{1},y^{\prime}_{2}\}$. Say $u=y^{\prime}_{1}$. Assume for a contradiction that $y^{W}_{1}\notin H_{t(0)}$. Then $x^{\prime}_{1}y^{\prime}_{1}$ is a $(W\smallsetminus U)$$(U\smallsetminus W)$ edge, so $(U,W)$ violates the matching-condition, a contradiction.
(ii). Assume for a contradiction that $\{x_{1},x_{2}\}$ is $W$-weird. Then $x_{2}=x^{\prime}_{2}$ since $x_{1}$ is an edge by assumption. Moreover, $x^{\prime}_{2}$ has exactly one neighbour outside $H_{t(0)}$ by weirdness. But then $x_{2}$ should be an edge in $S(U,W)$ by the definition of right-left-shift of $(U^{\prime},W^{\prime})$, contradicting that $x_{2}$ is a vertex. ∎
Lemma 9.34.
Assume the (Tutte-bagel torso setting). (9.13) with $(U,W)$ a tetra-separation. Let $\pi_{1}:=\{x_{1},x_{2}\}$ and $\pi_{2}:=\{y_{1},y_{2}\}$. Assume that $x_{1}$ is an outward edge. Assume that $\pi_{1}$ is $U$-weird. Then $x_{2}=x^{\prime}_{2}$ and there is an $x_{2}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ edge $x_{2}y_{i}^{\prime}$ in $G$ with $y_{i}^{\prime}=y_{i}^{W}$.
Proof.
Since $\pi_{1}$ is $U$-weird and $x_{1}$ is an outward edge, we get that $x_{2}=x^{\prime}_{2}$, that $x_{2}x^{\prime}_{1}$ is an edge, and that $x_{2}$ has at most one neighbour in $U\smallsetminus\{x^{\prime}_{1}\}$. However, $H_{t(0)}$ is 3-connected, so $x_{2}$ has some two neighbours $u,v$ in $H_{t(0)}$ besides $x^{\prime}_{1}$. Hence one of these neighbours, say $v$, is not in $U$. Since all interior vertices of $H_{t(0)}$ lie in $U\smallsetminus W$, the vertex $v$ must be one of the $y^{\prime}_{i}$, say $v=y^{\prime}_{1}$. Since $v\notin U$, we get that $y_{1}$ is an inward edge. In particular, $y^{W}_{1}=y^{\prime}_{1}$. ∎
Lemma 9.35.
Assume the (Tutte-bagel torso setting). (9.13) with $(U,W)$ a tetra-separation. Let $\pi_{1}:=\{x_{1},x_{2}\}$ and $\pi_{2}:=\{y_{1},y_{2}\}$. Assume that $\pi_{1}$ is $W$-weird and $x_{1}$ is an outward edge. Then no $y_{i}$ is an inward edge.
Proof.
Assume otherwise, say $y_{1}$ is an inward edge. Then by Lemma 9.33 we have $y^{W}_{1}=y^{\prime}_{1}$ and some $x^{\prime}_{j}$ is a neighbour of $y^{W}_{1}$. Then $j\neq 1$ since otherwise $x^{\prime}_{1}y^{W}_{1}$ is a $(U\smallsetminus W)$$(W\smallsetminus U)$ edge, contradicting the matching-condition. So $j=2$. Then $y_{1}^{\prime}=y_{1}^{W}$ is the only neighbour of $x_{2}$ in $W\smallsetminus\{x_{1}^{W}\}$. It follows that $x_{1}^{W}$ is the only neighbour of $x_{2}^{\prime}$ in $W^{\prime}\smallsetminus U^{\prime}$. This contradicts that $x_{2}$ is a vertex by definition of right-left-shift. ∎
Lemma 9.36.
Assume the (Tutte-bagel torso setting). (9.13) with $(U,W)$ a tetra-separation. Let $\pi_{1}:=\{x_{1},x_{2}\}$ and $\pi_{2}:=\{y_{1},y_{2}\}$. Then $\{\pi_{1},\pi_{2}\}$ is 1-potter-linked.
Proof.
By symmetry, it suffices to show that $\pi_{1}$ is 1-potter-linked to $\pi_{2}$. If $\pi_{1}$ is not weird for any side of $(U,W)$, then $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ by (6.9.1). So we may assume that $\pi_{1}$ is weird for some side; in particular, $x_{1}$ is an edge and $x_{2}=x^{\prime}_{2}$ is a vertex, say.
Case: $x_{1}$ is an inward edge. By Lemma 9.33, there is a neighbour $y^{\prime}_{i}$ of $x^{\prime}_{1}$ with $y^{\prime}_{i}=y^{W}_{i}$, and $\pi_{1}$ is not $W$-weird. So $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $W$ by (6.9.1). Since we assumed earlier that $\pi_{1}$ is weird for some side of $(U,W)$, but showed that $\pi_{1}$ is not $W$-weird, it follows that $\pi_{1}$ is $U$-weird. By Lemma 9.16, there exist two disjoint $\{x^{W}_{1},x^{W}_{2}\}$$\{y^{W}_{1},y^{W}_{2}\}$ paths in $\bigcup_{t\in O-t(0)}G_{t}$. Since $x_{1}$ is an inward edge and $y^{\prime}_{i}=y^{W}_{i}$, the edge $x^{\prime}_{1}y^{\prime}_{i}$ can also be written as $x^{W}_{1}y^{W}_{i}$. This edge together with the two paths forms a set of three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[W]$, so $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $U$ by (6.9.3).
Case: $x_{1}$ is an outward edge. If $\pi_{1}$ is $U$-weird, then we find an $x_{2}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ edge $e=x_{2}y_{i}^{\prime}$ with $y_{i}^{\prime}=y_{i}^{W}$ by Lemma 9.34. Since $x_{2}\in\pi_{1}$ and $y_{i}^{\prime}\in\hat{\pi}_{2}\cap W$, the edge $e$ is a $\pi_{1}$$(\hat{\pi}_{2}\cap W)$ edge, which witnesses that $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $U$ by (6.9.2). So it remains to show that $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $W$. For this, we assume that $\pi_{1}$ is $W$-weird by (6.9.1). By Lemma 9.35, no $y_{i}$ is an inward edge. Since $H_{t(0)}$ is 3-connected, we find three independent $\{x^{\prime}_{1},x^{\prime}_{2}\}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ paths in $H_{t(0)}$ by (Menger for independent paths). (6.2). All three paths are subgraphs of $G[U]$ since no $y_{i}$ is an inward edge. Moreover, they are three $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[U]$. This means $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $W$ by (6.9.3). ∎
Lemma 9.37.
Assume the (Tutte-bagel torso setting). (9.13) with $(U,W)$ a tetra-separation. Let $\pi_{1}:=\{x_{1},x_{2}\}$ and $\pi_{2}:=\{y_{1},y_{2}\}$. Then $\{\pi_{1},\pi_{2}\}$ is 2-potter-linked.
Proof.
By symmetry and by 6.12, it suffices to show the following: if $\pi_{1}$ is $U$-weird and $\pi_{2}$ is $W$-weird, then there is a $\pi_{1}$$\pi_{2}$ edge in $G$. So we assume that $\pi_{1}$ is $U$-weird and $\pi_{2}$ is $W$-weird. Then no $y_{i}$ is an inward edge by Lemma 9.33. Without loss of generality, $y_{1}$ is an outward edge. Then by Lemma 9.35, no $x_{i}$ is an inward edge. So we can assume that $x_{1}$ is an outward edge. Then by Lemma 9.34, we find an $x_{2}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ edge $x_{2}y_{i}^{\prime}$ with $y_{i}^{\prime}=y_{i}^{W}$. Note that $y_{i}$ is a vertex, so $y_{i}^{\prime}=y_{i}\in\pi_{2}$. Then $x_{2}y_{i}^{\prime}$ is the desired edge. ∎
Lemma 9.38.
Assume the (Tutte-bagel torso setting). (9.13) with $(U,W)$ a tetra-separation. If $x_{1}$ is an edge, then $x^{W}_{1}\in A$.
Proof.
Suppose for a contradiction that $x^{W}_{1}\in B\smallsetminus A$. By the (Crossing Lemma). (4.1), there are no jumping edges and no diagonal edges. Then $x_{1}$ dangles from the $A$-link through the $D$-link. Hence the corner $AD$ is potter by the (Crossing Lemma). (4.1). But then $x^{\prime}_{1}$ has a neighbour in $D$-link, so $x^{\prime}_{1}$ has $\geqslant 2$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$, contradicting that $x_{1}$ is an edge rather than the vertex $x^{\prime}_{1}$. ∎
Lemma 9.39.
Assume the (Tutte-bagel torso setting). (9.13) with $(U,W)$ a tetra-separation. Then $\{x_{1},y_{1}\}$ is not $W$-weird.
Proof.
Assume for a contradiction that $\{x_{1},y_{1}\}$ is $W$-weird. Then $x_{1}$ is an edge and $y_{1}=y^{\prime}_{1}$ is a vertex, say.
Assume first that $x_{1}$ is an inward edge. Then $y^{\prime}_{1}$ has $\leqslant 1$ neighbour in $W\smallsetminus\{x^{\prime}_{1}\}\supseteq V(G-G_{t(0)})$. But then, by the definition of $(U,W)$ via right-left-shift, $y_{1}$ should be an edge rather than the vertex $y^{\prime}_{1}$, a contradiction.
Assume otherwise that $x_{1}$ is an outward edge. By Lemma 9.38, $x^{W}_{1}\in A$. However, since $y^{\prime}_{1}$ is contained in $A\cap C$ and the Tutte-bagel induces a path-decomposition of $G[A]$, the vertex $y^{\prime}_{1}=y_{1}$ does not neighbour $x^{W}_{1}$, contradicting that $\{x_{1},y_{1}\}$ is $W$-weird. ∎
Lemma 9.40.
Assume the (Tutte-bagel torso setting). (9.13) with $(U,W)$ a tetra-separation. Let $\pi_{1}:=\{x_{1},y_{1}\}$ and $\pi_{2}:=\{x_{2},y_{2}\}$. Then $\{\pi_{1},\pi_{2}\}$ is both 1-potter-linked and 2-potter-linked.
Proof.
By Lemma 9.39, neither $\pi_{1}$ nor $\pi_{2}$ is $W$-weird. Hence $\{\pi_{1},\pi_{2}\}$ is 2-potter-linked. It remains to show that $\pi_{1}$ is 1-potter-linked to $\pi_{2}$, say. Since $\pi_{1}$ is not $W$-weird, it remains to consider the case where $\pi_{1}$ is $U$-weird. We find three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[W]$ with the (Three-Paths Lemma). (9.21). Hence $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ via (6.9.3). ∎
Key Lemma 9.41.
Assume the  (Tutte-bagel setting). (9.12). Then the following two assertions are equivalent for every node $t$ of $O$:
  1. (i)
    $(U_{t},W_{t})$ is a totally-nested tetra-separation;
  2. (ii)
    the torso of $\mathcal{O}$ at $t$ is good.
Proof.
If $H_{t}$ is bad, then $(U_{t},W_{t})$ is not even a tetra-separation by Lemma 9.23, let alone totally-nested. So we may assume now that $H_{t}$ is good. In particular, we may assume the (Tutte-bagel torso setting). (9.13) with $t=t(0)$.
By Lemma 9.23, $(U,W)$ is a tetra-separation. It suffices to show that $(U,W)$ is externally 5-connected by 6.16. By Lemma 9.24, $(U,W)$ is half-connected. To see that $(U,W)$ is 3-linked, we combine Lemma 9.30 with Lemma 9.31. By Lemma 9.32, $(U,W)$ is 0-potter-linked. We get that $(U,W)$ is 1-potter-linked and 2-potter-linked from Lemmas 9.36, 9.40 and 9.37. ∎
9.6. Splitting stars and torsos
Lemma 9.42.
Assume the (Tutte-bagel setting). (9.12). Let $H_{s}$ be a torso of $\mathcal{O}$. Assume that $s$ stems from the $AC$-Tutte-path $T$. Assume that the $A$-link is represented by a triangle-torso, and that $s$ is neither equal nor adjacent to the node $o(A)$ in $O$. Then $G_{s}\subseteq G[C\smallsetminus D]$.
Proof.
Let $r$ be the unique neighbour of $o(A)$ in $T$. If $H_{r}$ is 3-connected or a 4-cycle, then its adhesion sets are disjoint, and we get $G_{s}\subseteq G[C\smallsetminus D]$. Otherwise $H_{r}$ is a triangle. The tips of $H_{o(A)}$ and $H_{r}$ are the two vertices in the adhesion-set of the edge $o(A)r$. It follows that the other adhesion-set of $H_{r}$ is included in $C\smallsetminus D$, so $G_{s}\subseteq G[C\smallsetminus D]$. ∎
Lemma 9.43.
Assume the (Tutte-bagel setting). (9.12). Let $H_{s}$ be a triangle-torso with tip $u$. Assume that $s$ stems from the $AC$-Tutte-path. Assume that $s\neq o(A)$, and that if the $A$-link is represented by a triangle-torso then $s$ is not adjacent to $o(A)$. Then $u$ has no neighbour in the $D$-link.
Proof.
Assume not for a contradiction, and let $v$ be a neighbour of $u$ in the $D$-link. Recall that $u\notin D$ by Lemma 9.42. So the edge $uv$ dangles from the $D$-link through the $A$-link. Hence the corner $AD$ is potter by the (Crossing Lemma). (4.1), and the triangle-torso $H_{o(A)}$ has $uv$ as free edge. By assumption, $o(A)$ and $s$ are not adjacent in $O$. By Lemma 9.15, there is precisely one torso $H_{r}$ in between the two triangle-torsos $H_{o(A)}$ and $H_{s}$, which also contains $u$. But $H_{r}$ has $u$ as tip since $uv$ is free in $H_{o(A)}$, which by Lemma 9.15 contradicts that $H_{s}$ also has $u$ as tip. ∎
Figure 9.8. The situation in Lemma 9.44.
Let $P$ be a path, and let $u,v$ be vertices on $P$. Recall that $uP\mathring{v}$ denotes the subpath of $P$ that starts at $u$ and ends just before the vertex $v$. The statement of the following lemma is supported by Figure 9.8.
Lemma 9.44 (Link-Moving Lemma for vertices).
Assume the (Tutte-bagel setting). (9.12). Let $P$ and $Q$ be the two paths that are the two components of $O-o(C)-o(D)$. Assume that $o(A)\in P$ and $o(B)\in Q$. Let $c$ denote the end of $P$ that is adjacent or incident to $o(C)$ in $O$. Define $d$ similarly. Let $s$ be a node in the same component of $P-o(A)$ as $c$ such that $H_{s}$ is a 4-cycle or a triangle. Assume that if the $A$-link is represented by a triangle-torso, then $s$ is not adjacent to $o(A)$ in $P$. Let $P^{\prime}:=dP\mathring{s}$ and $Q^{\prime}:=\mathring{s}Pc\cup Q$. Then:
  1. (i)
    $A^{\prime}:=\bigcup_{t\in V(P^{\prime})}V(G_{t})$ and $B^{\prime}:=\bigcup_{t\in V(Q^{\prime})}V(G_{t})$ are the two sides of a tetra-separation of $G$.
  2. (ii)
    $S(A^{\prime},B^{\prime})=S\cup(D$-link)))), where $S$ consists of the two free edges of $H_{s}$ if $H_{s}$ is a 4-cycle and of the free edge and the tip of $H_{s}$ if $H_{s}$ is a triangle.
  3. (iii)
    $(A^{\prime},B^{\prime})$ separates some two vertices in $B\cup D$.
Proof.
(ii) holds by construction, and gives that the mixed-separation $(A^{\prime},B^{\prime})$ has order four. We show (i) next.
Matching-condition. Let $ab$ be an edge in the $D$-link with $a\in A\smallsetminus B$ and $b\in B\smallsetminus A$. Let $f$ be an edge in $S$. Assume for a contradiction that $ab$ and $f$ share an endvertex, which can only be $a$. Since $a\in(A\smallsetminus B)\cap D$ and $V(G_{s})\subseteq A\cap C$, we find that $a$ is in the $A$-link. So $ab$ dangles from the $A$-link through the $D$-link. Then the corner $AD$ is potter by the (Crossing Lemma). (4.1). Thus the $A$-link is represented by a triangle-torso, so $a\notin D$ by Lemma 9.42, a contradiction.
Degree-condition. It is straightforward to check that every vertex in $S(A^{\prime},B^{\prime})$ has $\geqslant 2$ neighbours in $B^{\prime}\smallsetminus A^{\prime}$. Now let $v\in S(A^{\prime},B^{\prime})$ be a vertex; we claim that it has $\geqslant 2$ neighbours in $A^{\prime}\smallsetminus B^{\prime}$. If $v$ is in $S$, then $H_{s}$ is a triangle with tip $v$. Let $H_{r}$ denote the torso neighbouring $H_{s}$ in the direction of $o(A)$ on the path $P$. If $H_{r}$ is 3-connected, then $v$ has two neighbours in $G_{r}$, both of which lie in $A^{\prime}\smallsetminus B^{\prime}$ since $G_{s}\subseteq G[A\cap C]$. Otherwise $H_{r}$ is a triangle, and the adhesion-graph in between $H_{r}$ and $H_{s}$ is a $K_{2}$-subgraph of $G$ by Lemma 9.4. Hence $v$ has two neighbours in $G_{r}$, both of which lie in $A^{\prime}\smallsetminus B^{\prime}$ since $r\neq o(A)$ by assumption. So we may assume next that $v$ lies in the $D$-link. Every neighbour of $v$ in $A\smallsetminus B$ also lies in $A^{\prime}\smallsetminus B^{\prime}$, except possibly when a neighbour of $v$ lies in $S$ or an edge incident to $v$ lies in $S$. The former exception is impossible by Lemma 9.43, while the latter is excluded by Lemma 9.42. This conclude the proof of (i).
(iii). We have to find two vertices in $B\cup D$ that are contained in $A^{\prime}\smallsetminus B^{\prime}$ and in $B^{\prime}\smallsetminus A^{\prime}$, respectively. Every vertex in $B\smallsetminus A$ lies in the intersection of $B\cup D$ with $B^{\prime}\smallsetminus A^{\prime}$, and $B\smallsetminus A$ is non-empty. Any vertex in $D\smallsetminus B$ lies in $A^{\prime}\smallsetminus B^{\prime}$ by Lemma 9.42, so it remains to show that $D\smallsetminus B$ is non-empty. If $D\smallsetminus B$ is empty, then the $D$-link consists of two edges. But both these edges have ends in the corner for $AD$ by the (Crossing Lemma). (4.1), a contradiction. ∎
Lemma 9.45 (Link-Moving Lemma for edges).
Assume the (Tutte-bagel setting). (9.12). Let $P$ and $Q$ be the two paths that are the two components of $O-o(C)-o(D)$. Assume that $o(A)\in P$ and $o(B)\in Q$. Let $c$ denote the end of $P$ that is adjacent or incident to $o(C)$ in $O$. Define $d$ similarly. Let $e$ be an edge of $O$ such that the torsos at the ends of $e$ are 3-connected. Assume that $e$ stems from the $AC$-Tutte-path. Let $P^{\prime}:=dP\mathring{e}$ and $Q^{\prime}:=\mathring{e}Pc\cup Q$. Then:
  1. (i)
    $A^{\prime}:=\bigcup_{t\in V(P^{\prime})}V(G_{t})$ and $B^{\prime}:=\bigcup_{t\in V(Q^{\prime})}V(G_{t})$ are the two sides of a tetra-separation of $G$.
  2. (ii)
    $S(A^{\prime},B^{\prime})=\pi\cup(D$-link)))), where $\pi$ is the adhesion-set of $ac$.
  3. (iii)
    $(A^{\prime},B^{\prime})$ separates some two vertices in $B\cup D$.
Proof.
The proof is analogous to the proof of Lemma 9.44, except that everything is straightforward because the two 3-connected torsos at the ends of $e$ both have disjoint adhesion-sets and have all their vertices contained in $A\cap C$. ∎
Lemma 9.46.
Assume the (Tutte-bagel setting). (9.12). Let $i:=1$ if $H_{t(1)}$ is 3-connected, and $i:=2$ otherwise. Similarly, let $j:=-1$ if $H_{t(-1)}$ is 3-connected, and $j:=-2$ otherwise. The right-shift of $(U^{\prime}_{t},W^{\prime}_{t})$ has sides
$V(G_{t(0)})\quad\text{and}\quad\bigcup\,\{\,V(G_{t(\ell)}):i\leqslant\ell\leqslant j\text{ in the cyclic ordering of }\mathbb{Z}_{m+1}\,\}.$
Proof.
This is just a different perspective on Lemma 9.14. ∎
Definition 9.47 (Tetra-star $\sigma(\mathcal{O})$).
Assume the (Tutte-bagel setting). (9.12). The tetra-star of $\mathcal{O}$ is the set of all $(U_{t},W_{t})$ for good torsos $H_{t}$ of $\mathcal{O}$. We denote the tetra-star of $\mathcal{O}$ by ${\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{\sigma(\mathcal{O})}}$.
Lemma 9.48.
Assume the (Tutte-bagel setting). (9.12). Let $(E,F)$ be a totally-nested tetra-separation of $G$ that interlaces $\sigma$. Then there is a 3-connected torso $H_{t}$ of $\mathcal{O}$ such that $(E,F)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$ or $(F,E)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, where $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$ denotes the right-shift of $(U^{\prime}_{t},W^{\prime}_{t})$.
Proof.
Since $(E,F)$ is totally-nested, we have $(E,F)\leqslant(A,B)$ and $(E,F)\leqslant(C,D)$, say. In particular, $E\subseteq A\cap C$. Let $\beta$ denote the $AC$-banana, and let $(T,\mathcal{V})$ denote the $AC$-Tutte-banana.
An element $x\in V(T)\cup E(T)$ is admissible if
  • $\cdot$
    $x$ is an edge with 3-connected torsos at its end-nodes; or
  • $\cdot$
    $x$ is an interior node of $T$ with a triangle or 4-cycle as torso.
A mixed-1-separation $(M,N)$ of $T$ is admissible if the vertex or edge in $S(M,N)$ is admissable. The $\beta$-lift of an admissible $(M,N)$ is
  • $\cdot$
    $(\bigcup_{t\in M}V_{t},\bigcup_{t\in N}V_{t})$ if $S(M,N)$ consists of an edge, and
  • $\cdot$
    $(\bigcup_{t\in M\smallsetminus N}V_{t},\bigcup_{t\in N\smallsetminus M}V_{t})$ if $S(M,N)$ consists of a vertex.
Note that the $\beta$-lift is a mixed-2-separation of the banana $\beta$. A $G$-lift of an admissible $(M,N)$ is a tetra-separation of $G$ that induces the $\beta$-lift of $(M,N)$ on $\beta$ and separates some two vertices in $F$.
We claim that:
(7) Every admissible $(M,N)$ has a $G$-lift.
Given an admissible $(M,N)$, let $x$ denote the element of its separator. If $x$ is an edge, then the (Link-Moving Lemma for edges). (9.45) produces a $G$-lift. So we may assume that $x$ is an interior node of $T$. Let $T^{\prime}$ be obtained from $T$ by removing any endvertex with a triangle or 4-cycle torso that represents a link. Any vertex removed in this way is not admissible, and neither is its incident edge in $T$. Hence $x\in T^{\prime}$.
Let us show that $x$ is not the only vertex of $T^{\prime}$. Otherwise the torso $H_{x}$ is a triangle or 4-cycle. By Lemma 9.18, $H_{x}$ is a triangle and both the $A$-link and the $C$-link are represented by triangle-torsos. In particular, $A\cap C=3$. But $E$, as a side of a tetra-separation, has size $\geqslant 4$, and $E$ is included in $A\cap C$, a contradiction. So $x$ is not the only vertex of $T^{\prime}$.
Let $y$ be an endvertex of $T^{\prime}$ that is distinct from $x$. By symmetry, we may assume that $y$ is adjacent or incident to $o(A)$ when viewed as a node of $O$. Then the (Link-Moving Lemma for vertices). (9.44) produces a $G$-lift. This concludes the proof that 7 holds.
For every admissible $(M,N)$ we use 7 to choose a $G$-lift $(\bar{M},\bar{N})$ so that for $(N,M)$ we choose $(\bar{N},\bar{M})$. We claim that exactly one of $E\subseteq\bar{M}$ and $E\subseteq\bar{N}$ holds, for all $(M,N)$. Indeed, otherwise $(\bar{M},\bar{N})$ separates $E$, and since it also separates $F$ by definition, $(\bar{M},\bar{N})$ crosses $(E,F)$, a contradiction. We let
$\omega$ :=${}:=\{\,(M,N):(M,N)\text{ is admissable and }\bar{N}\supseteq E\,\}$
$\Omega$ :=${}:=\bigcap\,\{\,N\smallsetminus M:(M,N)\in\omega\,\}\cap V(T^{\prime}).$
We claim that $\Omega\neq\emptyset$. Assume for a contradiction that $\Omega$ is empty.
Suppose first that $\omega$ is empty. Then $V(T^{\prime})$ is empty. Hence $T$ is a $K_{2}$ both whose endvertices have triangle or 4-cycle torsos that represent links. By Lemma 9.4, both torsos are triangles. Hence the corner $AC$ is potter, so $A\cap C=2$, contradicting that $E\subseteq A\cap C$ where $E$ has size $\geqslant 4$.
Suppose otherwise that $\omega$ is non-empty. So $\omega$ has one or two maximal elements. Assume first that $\omega$ has a unique maximal element $(M,N)$. Since $\Omega$ is empty but $N\smallsetminus M$ is not, $N\smallsetminus M$ must consist of an end of $T$ that misses in $T^{\prime}$. Hence $N\smallsetminus M=\{o(C)\}$, say, and the unique element $x$ of $S(M,N)$ is not an edge. So $x$ is a neighbour of $o(C)$, and both $x$ and $o(C)$ have a triangle or 4-cycle as torso. Hence $\bar{N}\cap A$ equals the adhesion-set of the edge $xo(C)$, contradicting that $E$ has size $\geqslant 4$ and $E\subseteq\bar{N}\cap A$.
Assume otherwise that $\omega$ has two maximal elements $(M_{1},N_{1})$ and $(M_{2},N_{2})$. Let $x_{i}$ denote the unique element of $S(M_{i},N_{i})$ for both $i$. We claim that both $x_{i}$ are vertices. Otherwise $x_{1}$ is an edge, say. Then both endvertices of $x_{1}$ have 3-connected torsos. Hence $x_{2}$ cannot be an endvertex of $x_{1}$. Therefore, an endvertex of $x_{1}$ is in $\Omega$, a contradiction. So both $x_{i}$ are vertices. Since $\Omega$ is empty, the vertices $x_{i}$ are neighbours. Hence $E$ is included in the adhesion-set of $x_{1}x_{2}$, contradicting that $E$ has size $\geqslant 4$. Thus $\Omega\neq\emptyset$.
Every node in $\Omega$ has a 3-connected torso in $\mathcal{O}$. Since $\Omega$ spans a subpath of $T$ that contains no admissible edges, we have $|\Omega|=1$. Hence $\Omega$ consists of exactly one node $t^{*}$, and its torso $H_{t^{*}}$ is 3-connected.
To show that $(E,F)\leqslant$ the right-shift of $(U^{\prime}_{t},W^{\prime}_{t})$, by Lemma 9.46 and Lemma 5.7 it remains to show $E\subseteq V(G_{t^{*}})$. Assume for a contradiction that there is a vertex $v\in E\smallsetminus V(G_{t^{*}})$. Recall that $E\subseteq A\cap C$, so in particular $E\subseteq V(\beta)$. Hence there is a node $s$ of $T$ with $v\in V_{s}$, and $s\neq t^{*}$ by assumption. Let $r$ be the penultimate node on the path $sTt^{*}$.
Assume first that the torso $H_{r}$ is 3-connected. Then the edge $rt^{*}$ is admissible. Let $(M,N)\in\omega$ with $S(M,N)=\{rt^{*}\}$. Note that $r\in M\smallsetminus N$ and $t^{*}\in N\smallsetminus M$ since $\{t^{*}\}=\Omega$. Let $(\hat{M},\hat{N})$ be the $\beta$-lift of $(M,N)$. Observe that $v\in\hat{M}\smallsetminus\hat{N}$. Therefore, $v\in\bar{M}\smallsetminus\bar{N}$, contradicting that $E\subseteq\bar{N}$.
Assume otherwise that the torso $H_{r}$ is either a triangle or a 4-cycle. Then $r$ is admissible or $r$ is one of $o(A)$ and $o(C)$. Assume first that $r$ is admissible. Let $(M,N)\in\omega$ with $S(M,N)=\{r\}$. Note that $r,s\in M$ and $t^{*}\in N\smallsetminus M$. Let $(\hat{M},\hat{N})$ be the $\beta$-lift of $(M,N)$. Then $v\in\hat{M}\smallsetminus\hat{N}$, so $v\in\bar{M}\smallsetminus\bar{N}$, contradicting that $E\subseteq\bar{N}$. Assume otherwise that $r=o(A)$, say. Then $H_{r}$ is a triangle or 4-cycle representing the $A$-link. Since $v$ does not lie in $H_{t^{*}}$, we in particular have that $v$ does not lie in the adhesion-set of the edge $rt^{*}$. But then $v$ lies in $D\smallsetminus C$, contradicting that $E\subseteq C$. ∎
Lemma 9.49.
Assume the (Tutte-bagel setting). (9.12). Every tetra-separation of $G$ that interlaces the tetra-star $\sigma$ of $\mathcal{O}$ is crossed.
Proof.
Assume for a contradiction that $\sigma$ is interlaced by a totally-nested tetra-separation $(E,F)$ of $G$. By Lemma 9.48, there is a 3-connected torso $H_{t}$ of $\mathcal{O}$ such that $(E,F)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, say, where $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$ denotes the right-shift of $(U^{\prime}_{t},W^{\prime}_{t})$.
Assume first that $H_{t}$ is good. Then $(U_{t},W_{t})$ is a totally-nested tetra-separation by 9.41. In particular, $(U_{t},W_{t})\in\sigma$. Since $(E,F)$ interlaces $\sigma$, we have one of $(U_{t},W_{t})<(E,F)$ and $(U_{t},W_{t})<(F,E)$. We cannot have $(U_{t},W_{t})<(F,E)$, since otherwise $\emptyset\neq E\smallsetminus F\subseteq(W_{t}\smallsetminus U_{t})\cap(U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t})$ contradicts the fact that $U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t}\subseteq U_{t}\smallsetminus W_{t}$ by 5.2. Hence $(U_{t},W_{t})<(E,F)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$. Since $(U_{t},W_{t})$ is the left-shift of $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, some vertex in $S(E,F)$ must have $\leqslant 1$ neighbour in $E\smallsetminus F$, contradicting the degree-condition.
Assume otherwise that $G_{t}$ is bad. Hence $G_{t}$ has four or five vertices. If $H_{t}$ has four vertices, then all but at most one of the four vertices in $S(U^{\prime}_{t},W^{\prime}_{t})$ has $\geqslant 2$ neighbours in $W^{\prime}_{t}\smallsetminus U^{\prime}_{t}$. If $H_{t}$ has five vertices, then all four vertices in $S(U^{\prime}_{t},W^{\prime}_{t})$ have $\geqslant 2$ neighbours in $W^{\prime}_{t}\smallsetminus U^{\prime}_{t}$. In either case, we have $|U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t}|\leqslant 1$. Since $(E,F)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, we have $E\smallsetminus F\subseteq U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t}$. Therefore, $|E\smallsetminus F|\leqslant 1$, which by Lemma 5.5 contradicts that $(E,F)$ is a tetra-separation. ∎
Key Lemma 9.50.
Let $G$ be a 4-connected graph, and let $N$ denote the set of totally-nested tetra-separations of $G$. Let $\sigma$ be a splitting star of $N$ that is interlaced by two tetra-separations $(A,B)$ and $(C,D)$ of $G$ that cross with empty vertex-centre. Then the torso $\tau$ of $\sigma$ is a cycle of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices.
Proof.
By the (Crossing Lemma). (4.1), all links have size two. By Lemma 9.10, $(A,B)$ and $(C,D)$ induce a Tutte-bagel $\mathcal{O}$. By 9.41, the tetra-star $\sigma(\mathcal{O})$ is a star of totally-nested tetra-separations of $G$. By Lemma 9.49, the tetra-star $\sigma(\mathcal{O})$ is a splitting star of $N$. Since both $(A,B)$ and $(C,D)$ interlace both $\sigma$ and $\sigma(\mathcal{O})$, we obtain $\sigma=\sigma(\mathcal{O})$. Hence it remains to compute the torso $\tau$ of $\sigma(\mathcal{O})$. The torso $\tau$ is 4-connected by Lemma 8.36. We can obtain $\tau$ from $\mathcal{O}$ by replacing each good torso with a $K_{4}$, leaving only 3-connected torsos with five or four vertices, and contracting some free edges in torsos that are 4-cycles or triangles. Each torso that is a 4-cycle neighbours two 3-connected torsos by Lemma 9.4, and the 4-cycle-torso witnesses that both 3-connected torsos are good. Hence the free edges of the 4-cycle-torsos are contained in the separators of the elements of $\sigma(\mathcal{O})$ that stem from the good neighbouring torsos, and so they are contracted. Hence $\tau$ is a cycle of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices. ∎
9.7. Angry Tutte-bagels
Let $G$ be a 4-connected graph with a Tutte-bagel $\mathcal{O}=(O,\mathcal{G})$. A triangle-strip of $\mathcal{O}$ is a maximal connected subgraph $\Delta$ of $O$ such that every node in $\Delta$ has a triangle-torso with regard to $\mathcal{O}$. Phrased differently, a triangle-strip is a component of
[…]
Triangle-strips are paths unless all torsos of $O$ are triangles. An element of $V(O)\cup E(O)$ is furious if either
  • $\cdot$
    it is an edge with 3-connected torsos at both ends, or
  • $\cdot$
    it is a vertex with a 4-cycle as torso.
We define
${\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{\alpha(\mathcal{O})}}:=\#(\text{furious edges and vertices of }\mathcal{O})+\sum_{\Delta}\left\lceil\tfrac{2}{3}|\Delta|\right\rceil$
where the sum ranges over all triangle-strips $\Delta$ of $\mathcal{O}$.
Lemma 9.51.
Assume the (Tutte-bagel setting). (9.12). Then $\alpha(\mathcal{O})\geqslant 4$.
Proof.
The four links injectively correspond to furious edges or vertices or nodes in triangle-strips. Hence it suffices to show that there is no path $rst$ in $O$ with all three nodes $r,s,t$ having triangle-torsos representing links. Assume otherwise. Then two adjacent corners are potter, contradicting 4.7. ∎
Lemma 9.52.
Let $G\neq K_{5}$ be a 4-connected graph with a Tutte-bagel $\mathcal{O}=(O,\mathcal{G})$ such that all torsos of $\mathcal{O}$ are bad and $\alpha(\mathcal{O})\geqslant 4$. Then $G$ has two tetra-separations that cross with all links of size two.
Proof.
We claim that we can select four elements $x(i)$ of $V(O)\cup E(O)$, where $i\in\mathbb{Z}_{4}$, such that
  • $\cdot$
    every selected element is furious or belongs to some triangle-strip;
  • $\cdot$
    no three consecutive nodes are chosen from any triangle-strip.
If all torsos of $\mathcal{O}$ are triangles, then we use $|O|\geqslant 6$, which we have since $G\neq K_{5}$ and $\alpha(\mathcal{O})\geqslant 4$. Otherwise, we use $\alpha(\mathcal{O})\geqslant 4$ and that the triangle-strips of $\mathcal{O}$ are induced paths in $O$.
We may assume that $O$ traverses the $x(i)$ in the cyclic ordering of $\mathbb{Z}_{4}$. Let $P_{A},P_{B}$ denote the components of $O-x(0)-x(2)$, let $A:=\bigcup_{t\in V(P_{A})}V(G_{t})$ and define $B$ similarly. We use $x(1),x(3)$ to define $P_{C},P_{D}$ and $C,D$ analogously.
We define $S(i)$ for $i\in\mathbb{Z}_{4}$ to consist of
  • $\cdot$
    the two vertices in the adhesion-set of $x(i)$ if $x(i)$ is a furious edge;
  • $\cdot$
    the two free edges of the 4-cycle-torso at $x(i)$ if $x(i)$ is a furious vertex;
  • $\cdot$
    the tip and the free edge of the triangle-torso at $x(i)$ if $x(i)$ stems from a triangle-strip.
We claim that $(A,B)$ and $(C,D)$ are tetra-separations with separators $S(A,B)=S(0)\cup S(2)$ and $S(C,D)=S(1)\cup S(3)$. By symmetry, it suffices to show that
  1. (i)
    every vertex in $S(0)$ has $\geqslant 2$ neighbours in $A\smallsetminus B$, and
  2. (ii)
    every edge in $S(0)$ has an end in $A\smallsetminus B$ avoiding $\widehat{S(2)}$.
For this, let $t(1),t(2),t(3)$ denote the first three vertices on $P_{A}$ that come just after $x(0)$ on $O$.
Case: $x(0)$ is a furious vertex. So $H_{x(0)}$ is a 4-cycle and $S(0)$ consists of two edges, and we only have to show (ii). The neighbouring torso $H_{t(1)}$ must be 3-connected by Lemma 9.4, so the adhesion-set of $x(0)t(1)$ is included in $A\smallsetminus B$ and avoids $\widehat{S(2)}$. Hence (ii) holds.
Case: $x(0)$ is a furious edge. Then $S(0)$ is the adhesion-set of the edge $x(0)$, and we only have to show (i). Note that $t(1)$ is an end of $x(0)$, so the torso $H_{t(1)}$ is 3-connected. Let $U:=V(H_{t(1)}-S(0))$. Every vertex in $S(0)$ has two neighbours in $G$ that are contained in $U$. To ensure that these neighbours lie in $A\smallsetminus B$, it suffices to show that the adhesion-set $\pi$ of $t(1)t(2)$ is included in $A\smallsetminus B$. Since $t(1)t(2)$ is the first candidate for $x(1)$, with $x(2)$ coming later, $\pi$ can only meet $B$ if $H_{t(2)}$ is a triangle. But then $t(2)$ is the first candidate for $x(1)$, so $x(2)$ comes after $t(2)$. If $H_{t(3)}$ is not a triangle, we are done, and otherwise the tip of the triangle $H_{t(3)}$ avoids $\pi$ by Lemma 9.15. Hence $\pi\subseteq A\smallsetminus B$ as desired.
Case: $x(0)$ is in a triangle-strip. Assume first that $H_{t(1)}$ is 3-connected. Since $H_{t(1)}$ is bad and neighboured by the triangle-torso $H_{x(0)}$, its other neighbouring torso $H_{t(2)}$ must be 3-connected. Hence the edge $t(1)t(2)$ is the first candidate for $x(1)$, and so the edge $t(2)t(3)$ is the first candidate for $x(2)$. Therefore, the entire vertex-set of $H_{t(1)}$ is included in $A\smallsetminus B$, which ensures (i) and (ii).
Assume otherwise that $H_{t(1)}$ is a triangle. Let $\pi$ denote the adhesion-set of the edge $t(1)t(2)$. Note that $\pi$ contains an end of the free edge in $S(0)$ and both vertices in $\pi$ are neighbours of the tip of the triangle $H_{x(0)}$. Hence to show both (i) and (ii), it suffices to show that $\pi$ is included in $A\smallsetminus B$. The node $t(1)$ is the first candidate for $x(1)$, so $t(2)$ is the first candidate for $x(2)$. By Lemma 9.4, $H_{t(2)}$ is either a triangle or 3-connected. If $H_{t(2)}$ is 3-connected, then $\pi$ is included in $A\smallsetminus B$. Otherwise $H_{t(2)}$ is a triangle. Then $H_{x(0)}$, $H_{t(1)}$ and $H_{t(2)}$ are three consecutive triangles. The way we have selected the $x(i)$ then ensures that $x(2)$ comes after $t(2)$. Again, $H_{t(3)}$ is either a triangle or 3-connected. If $H_{t(3)}$ is 3-connected, then $\pi$ is included in $A\smallsetminus B$. Otherwise, if $H_{t(3)}$ is a triangle, then its tip is distinct from the tip of $H_{t(2)}$ by Lemma 9.15, and $t(3)$ is the first candidate for $x(2)$. So $\pi$ is included in $A\smallsetminus B$.
This completes the proof of (i) and (ii). Hence $(A,B)$ and $(C,D)$ are tetra-separations with $S(A,B)=S(0)\cup S(2)$ and $S(C,D)=S(1)\cup S(3)$. The tetra-separations $(A,B)$ and $(C,D)$ cross with all links of size two as the links are precisely the four non-empty sets $S(i)$. ∎
Key Lemma 9.53.
Let $G$ be a 4-angry graph. Then the following two assertions are equivalent:
  1. (i)
    $G$ has two tetra-separations that cross with all links of size two;
  2. (ii)
    $G\neq K_{5}$ and $G$ has a Tutte-bagel $\mathcal{O}$ with $\alpha(\mathcal{O})\geqslant 4$ all whose torsos are bad.
Proof.
(i) $\Rightarrow$ (ii). Clearly, $G\neq K_{5}$ since $K_{5}$ does not have a tetra-separation. By Lemma 9.10, $(A,B)$ and $(C,D)$ induce a Tutte-bagel $\mathcal{O}$. By 9.41, the tetra-star $\sigma(\mathcal{O})$ is a star of totally-nested tetra-separations of $G$. Since $G$ has no totally-nested tetra-separations by assumption, all torsos of $\mathcal{O}$ are bad. We get $\alpha(\mathcal{O})\geqslant 4$ from Lemma 9.51.
(ii) $\Rightarrow$ (i) is Lemma 9.52. ∎
10. Quasi-5-connected torsos
This section deals with the quasi-5-connected outcome of Theorem 1; see 10.5.
A mixed-separation+ $(A,B)$ of a graph $G$ almost interlaces a star $\sigma$ of mixed-separations+ of $G$ if $(C,D)\leqslant(A,B)$ or $(C,D)\leqslant(B,A)$ for all $(C,D)\in\sigma$. The difference between ‘interlaces’ and ‘almost interlaces’ is ‘$<$’ versus $\leqslant$’ in the definitions.
Lemma 10.1.
(Analogue of [Tridecomp, Lemma 2.6.7]) Let $G$ be a 4-connected graph. Let $\sigma$ be a star of mixed-4-separations of $G$ satisfying the matching-condition. Let $(A,B)$ be a separation of the torso $\tau$ of $\sigma$. Then there exists a mixed-separation $(\hat{A},\hat{B})$ of $G$ such that:
  1. (i)
    the order of $(\hat{A},\hat{B})$ is at most the order of $(A,B)$;
  2. (ii)
    $(\hat{A},\hat{B})$ almost interlaces $\sigma$;
  3. (iii)
    $|\hat{A}\smallsetminus\hat{B}|\geqslant|A\smallsetminus B|$ and $|\hat{B}\smallsetminus\hat{A}|\geqslant|B\smallsetminus A|$; and
  4. (iv)
    for every $(U,W)\in\sigma$ we have $|W\cap(\hat{A}\smallsetminus\hat{B})|\geqslant|A\smallsetminus B|$ and $|W\cap(\hat{B}\smallsetminus\hat{A})|\geqslant|B\smallsetminus A|$.
Proof.
The proof is the same as the proof of [Tridecomp, Lemma 2.6.7], with one exception: we replace Lemma 1.2.3 with the matching-condition. ∎
Lemma 10.2.
Let $(A,B)$ be a tetra-separation in a 4-connected graph $G$. Let $(C,D)$ be a mixed-4-separation+ in $G$ with $(A,B)\leqslant(C,D)$. Let $(C^{\prime},D^{\prime})$ denote either the left-right-shift or the right-left-shift of $(C,D)$. Then $(A,B)\leqslant(C^{\prime},D^{\prime})$.
Proof.
We have $B\supseteq D\supseteq D^{\prime}$. Since $A\subseteq C$, it remains to show that every vertex in $S(C,D)\cap A$ also lies in $S(C^{\prime},D^{\prime})$. Every vertex $v$ in $S(C,D)\cap A$ also lies in $S(A,B)$. Hence every such $v$ has $\geqslant 2$ neighbours in $A\smallsetminus B\subseteq C\smallsetminus D$. Thus $v\in S(C^{\prime},D^{\prime})$. ∎
Lemma 10.3.
Let $G$ be a 4-connected graph. Let $(U,W)$ be a tetra-separation of $G$, and let $(A,B)$ be a mixed-4-separation of $G$ such that $(U,W)\leqslant(A,B)$. Suppose that $|W\cap(A\smallsetminus B)|\geqslant 1$. Let $(A^{\prime},B^{\prime})$ be either the left-right-sift or the right-left-shift of $(A,B)$. Then $(U,W)<(A^{\prime},B^{\prime})$.
Proof.
We have $(U,W)\leqslant(A^{\prime},B^{\prime})$ by Lemma 10.2. Recall that $A\smallsetminus B\subseteq A^{\prime}\smallsetminus B^{\prime}$ by 5.9. Hence $W\cap(A^{\prime}\smallsetminus B^{\prime})$ is nonempty, so in particular the inclusion $W\supseteq B^{\prime}$ is proper. ∎
Lemma 10.4.
Let $G$ be a 4-connected graph. Let $\sigma$ be a star of tetra-separations of $G$. Suppose that the torso $\tau$ of $\sigma$ has a 4-separation $(A,B)$ such that both sides have size $\geqslant 6$. Then $\sigma$ is interlaced by a tetra-separation of $G$.
Proof.
By Lemma 10.1, there is a mixed-$k$-separation+ $(\hat{A},\hat{B})$ of $G$ such that
  • $\cdot$
    $k\leqslant 4$ by (i);
  • $\cdot$
    $(\hat{A},\hat{B})$ almost interlaces $\sigma$ by (ii);
  • $\cdot$
    $|\hat{A}\smallsetminus\hat{B}|\geqslant 2$ and $|\hat{B}\smallsetminus\hat{A}|\geqslant 2$ by (iii); and
  • $\cdot$
    for every $(U,W)\in\sigma$ we have $|W\cap(\hat{A}\smallsetminus\hat{B})|\geqslant 2$ and $|W\cap(\hat{B}\smallsetminus\hat{A})|\geqslant 2$ by (iv).
Since $G$ is 4-connected, $(\hat{A},\hat{B})$ is a mixed-4-separation by Lemma 6.1. Let $(\bar{A},\bar{B})$ denote the left-right-shift of $(\hat{A},\hat{B})$. Then $(\bar{A},\bar{B})$ is a tetra-separation by 5.10. By Lemma 10.3, $(\bar{A},\bar{B})$ interlaces $\sigma$. ∎
Key Lemma 10.5.
Let $G$ be a 4-connected graph, and let $N$ denote the set of totally-nested tetra-separations of $G$. Let $\sigma$ be a splitting star of $N$ that is not interlaced by any tetra-separation of $G$. Then the torso $\tau$ of $\sigma$ is quasi-5-connected.
Proof.
By Lemma 8.36, $\tau$ is 4-connected or a $K_{4}$, and we are done in the latter case. By the contrapositive of Lemma 10.4, every 4-separation of $\tau$ has a side of size five. ∎
11. An Angry Theorem for 4-connectivity
A key step in proving Tutte’s decomposition theorem is to characterise the graphs that are 2-angry: those that are 2-connected and all whose 2-separators are crossed – hence the name. Tutte proved the first angry theorem: the 2-angry graphs are precisely the 3-connected graphs and the cycles [TutteCon]. For a modern proof, see [Tridecomp, Theorem 3.3.4]. An angry theorem for 3-connectivity has been proved using tri-separations [Tridecomp, Theorem 1.1.5]. However, like here, the angry theorem was not used in the proof of the decomposition theorem. Still, angry theorems have recently found an application in group theory: the local version of the 2-angry theorem [Local2sep] was used as a key ingredient in [StallingsNilpotent], and versions for higher connectivity are expected to be needed to continue in the direction of [StallingsNilpotent]. Here we prove an angry theorem for 4-connectivity. A graph $G$ is 4-angry if it is 4-connected and every tetra-separation of $G$ is crossed by a tetra-separation.
Theorem 11.1 (Angry Theorem for 4-connectivity).
A 4-connected graph $G$ with $|G|\geqslant 8$ is 4-angry if and only if it is one of the following:
  1. (i)
    quasi-5-connected;
  2. (ii)
    a cycle $\mathcal{O}$ of triangles, 4-cycles, 3-connected graphs on $\leqslant 5$ vertices, with only bad torsos and $\alpha(\mathcal{O})\geqslant 4$;
  3. (iii)
    a double-wheel with rim of length $\geqslant 4$ or a double-wheel of triangles with rim of length $\geqslant 4$;
  4. (iv)
    a sprinkled $K_{4,m}$ with $m\geqslant 4$.
Proof.
Forward implication. Assume that $G$ is 4-angry. If $G$ has no tetra-separation, then (i) by Lemma 2.3. So $G$ has a tetra-separation $(A,B)$, which is crossed by a tetra-separation $(C,D)$ as $G$ is 4-angry. By the (Crossing Lemma). (4.1), all links have the same size $\ell$, and $\ell\in\{0,1,2\}$.
  • $\cdot$
    If $\ell=0$, then (iv) by 7.8.
  • $\cdot$
    If $\ell=1$, then (iii) by 8.41.
  • $\cdot$
    If $\ell=2$, then (ii) by 9.53 and since $G\neq K_{5}$ by $|G|\geqslant 8$.
Backward implication. We use one of Lemma 2.3, 7.8, 8.41 and 9.53. Here Lemma 2.3 needs $|G|\geqslant 8$. ∎
12. Vertex-transitive graphs
[…]
Recall that a graph $G$ is arc-transitive if for every two ordered pairs $(u_{1},v_{1}),(u_{2},v_{2})$ of vertices forming edges $u_{1}v_{1},u_{2}v_{2}\in E(G)$, there is an automorphism $\varphi$ of $G$ such that $(\varphi(u_{1}),\varphi(v_{1}))=(u_{2},v_{2})$.
Theorem 12.1.
[Godsil, Theorem 3.4.2] Every finite vertex-transitive $d$-regular graph is $k$-connected for $k:=\lceil\frac{2}{3}(d+1)\rceil$.
Corollary 12.2.
Every finite $d$-regular graph with $d\geqslant 4$ is 4-connected.∎
A graph $G$ is essentially 5-connected if $G$ is 4-connected and every tetra-separation $(A_{1},A_{2})$ of $G$ has only edges in its separator and at least one side $A_{i}$ with $G[A_{i}]=K_{4}$.
Definition 12.3.
Let $G$ be a graph with a cycle-decomposition $\mathcal{O}=(O,(G_{t})_{t\in V(O)})$. Let $\ell$ denote the length of the cycle $O$. We say that $G$ is a
  • $\cdot$
    cycle of $K_{4}$-bags of length $\ell$ if all bags of $\mathcal{O}$ are $K_{4}$’s and all adhesion-sets of $\mathcal{O}$ are disjoint (Figure 1.7);
  • $\cdot$
    cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\ell$ if the bags along $O$ alternate between being $K_{4}$’s and being 4-cycles, and all adhesion-sets of $\mathcal{O}$ are disjoint (Figure 1.8);
  • $\cdot$
    cycle of triangle-bags of length $\ell$ if all bags are triangles, and every two non-consecutive edges on $O$ have disjoint adhesion-sets (Figure 1.9);
  • $\cdot$
    cycle of $K_{2,2}$-bags of length $\ell$ if all bags of $\mathcal{O}$ are $K_{2,2}$’s, all torsos of $\mathcal{O}$ are $K_{4}$’s and all adhesion-sets of $\mathcal{O}$ are disjoint (Figure 1.10).
Observation 12.4.
  • $\cdot$
    Every cycle of $K_{4}$-bags of length $\ell$ is isomorphic to $K_{2}\boxtimes C_{\ell}$.
  • $\cdot$
    Every cycle of $K_{4}$-bags of length 3 is isomorphic to $K_{6}$.
  • $\cdot$
    Every cycle of $K_{2,2}$-bags of length 3 is isomorphic to the octahedron.
  • $\cdot$
    Every cycle of $K_{2,2}$-bags of length 4 is isomorphic to $K_{4,4}$.
  • $\cdot$
    Every cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\ell\in\{4,6\}$ is essentially 5-connected.
  • $\cdot$
    Every cycle of triangle-bags of length $\ell$ is isomorphic to $C_{\ell}^{2}$ (the graph obtained from the $\ell$-cycle $C_{\ell}$ by joining every vertex $v$ to all vertices of distance at most 2 from $v$ on $C_{\ell}$).
  • $\cdot$
    Every cycle of triangle-bags of length 5 is isomorphic to $K_{5}$. ∎
Lemma 12.5.
Let $G$ be a vertex-transitive 4-connected finite graph, and let $(A_{1},A_{2})$ be a totally-nested tetra-separation in $G$. Then $S(A_{1},A_{2})$ consists of four edges, and $(A_{1},A_{2})$ has a side $A_{i}$ with $G[A_{i}]=K_{4}$.
Proof.
Let $M$ be the union of the $\operatorname{Aut}(G)$-orbits of $(A_{1},A_{2})$ and $(A_{2},A_{1})$. Let $(C,D)$ be $\leqslant$-minimal in $M$; that is, $(C^{\prime},D^{\prime})\not<(C,D)$ for every $(C^{\prime},D^{\prime})\in M$.
We claim that $S(C,D)$ contains no vertex. Assume otherwise that $u\in S(C,D)$ is a vertex. Then let $v$ be an arbitrary vertex in $C\smallsetminus D$ and let $\varphi\in\operatorname{Aut}(G)$ such that $\varphi(u)=v$. Then $(\varphi(C),\varphi(D))<(C,D)$ or $(\varphi(D),\varphi(C))<(C,D)$, contradicting the $\leqslant$-minimality of $(C,D)$. Hence $S(C,D)$ consists of four edges.
We claim that every vertex in $C$ is incident to an edge in $S(C,D)$. Assume for a contradiction that $v\in C$ is not. Let $u\in C\smallsetminus D$ be an endvertex of one of the four edges in $S(C,D)$ and let $\varphi\in\operatorname{Aut}(G)$ such that $\varphi(u)=v$. Then $(\varphi(C),\varphi(D))<(C,D)$ or $(\varphi(D),\varphi(C))<(C,D)$, contradicting the $\leqslant$-minimality of $(C,D)$.
Finally, since $G$ is 4-connected, every vertex has degree $\geqslant 4$, so $G[C]$ is a $K_{4}$. The tetra-separation $(A_{1},A_{2})$ shares the above properties of $(C,D)$ since $(C,D)\in M$. ∎
Corollary 12.6.
Let $G$ be a vertex-transitive 4-connected finite graph such that every tetra-separation is totally-nested. Then $G$ is essentially 5-connected.∎
Lemma 12.7.
Let $G$ be a vertex-transitive 4-connected finite graph that has two tetra-separations $(A,B),(C,D)$ that cross with all links empty. Then $G=K_{4,4}$, which is a cycle of $K_{2,2}$-bags of length 4.
Proof.
Let $Z$ be the centre of the crossing-diagram, which consists of four vertices by the (Crossing Lemma). (4.1). Note that $G-Z$ has at least four components.
[…]
By 4-connectivity of $G$, each of the $\geqslant 4$ vertices of $G-Z$ has neighbourhood equal to $Z$. So, $G$ is 4-regular by vertex-transitivity. In particular, $G-Z$ has exactly four vertices, and $G[Z]$ is independent. So, $G=K_{4,4}$. ∎
Lemma 12.8.
Let $G$ be a vertex-transitive 4-connected finite graph that has two tetra-separations $(A,B),(C,D)$ that cross with all links of size one. Let $\mathcal{O}$ be the block-bagel induced by $(A,B)$ and $(C,D)$. Then every bag of $\mathcal{O}$ is bad.
Proof.
Assume for a contradiction that $\mathcal{O}=:(O,\mathcal{G})$ has a good bag $G_{t}$. Then we have the (Block-bagel bag setting). (8.14) with $t=t(0)$, which we assume in the following. By 8.27, $(U,W)$ is a totally-nested tetra-separation of $G$. By Lemma 12.5, the separator $S(U,W)$ consists of four edges, and either $G[U]=K_{4}$ or $G[W]=K_{4}$. It follows with vertex-transitivity that $G$ is 4-regular and that the neighbourhood around each vertex induces $K_{1}\sqcup K_{3}$ in $G$.
By the (Helpful Lemma). (8.6), $O$ has at least four helpful edges or vertices. By 4-regularity, there are exactly four helpful edges or nodes of $O$. For each helpful edge, we colour its adhesion-vertex red. For each helpful node $t$, there exists a unique interior vertex of $G_{t}$ that is a neighbour of $z^{\prime}_{1}$ (by 4-regularity), and we colour this vertex red. Then exactly four vertices have been coloured red. By the above, the red vertices induced a $K_{1}\sqcup K_{3}$ in $G-Z^{\prime}$, where $Z^{\prime}$ is the centre of the crossing-diagram. None of the red vertices in the $K_{3}$ can be an interior vertex of a bag, so all three are adhesion-vertices. But then $O$ would have to be a triangle, contradicting that $O$ has four helpful edges or nodes. ∎
Lemma 12.9.
Let $G$ be a vertex-transitive 4-connected finite graph that has two tetra-separations $(A,B),(C,D)$ that cross with all links of size one. Let $\mathcal{O}$ be the Block-bagel induced by $(A,B)$ and $(C,D)$. Then every bag of $\mathcal{O}$ is a $K_{2}$.
Proof.
Write $\mathcal{O}=:(O,(G_{t})_{t\in V(O)})$, and let $u,v$ be the two vertices in the centre. By Lemma 12.8, every bag of $\mathcal{O}$ is bad. In particular, every bag of $\mathcal{O}$ is a $K_{2}$ or a triangle. Assume for a contradiction that $\mathcal{O}$ has a bag that is a triangle. Since every bag of $\mathcal{O}$ is bad, it follows that all bags of $\mathcal{O}$ are triangles.
Let $G_{t}$ be an arbitrary bag of $\mathcal{O}$, and let $x$ be the vertex in $G_{t}$ that is not an adhesion-vertex. By 4-connectivity of $G$, we have $xu,xv\in E(G)$ and $x$ has degree 4. Hence $G$ is 4-regular, and since $G_{t}$ was arbitrary, we get $|O|\leqslant 4$. As every node of $O$ is helpful, we get $|O|\geqslant 4$ from the (Helpful Lemma). (8.6), so $O$ has length 4. Hence each of $u,v$ has four neighbours, one in the interior of each bag, so by 4-regularity the edge $uv$ cannot exist in $G$. Therefore, $u$ is not contained in a triangle. But all bags are triangles, contradicting vertex-transitivity. ∎
Lemma 12.10.
Let $G$ be a vertex-transitive 4-connected finite graph that has two tetra-separations $(A,B),(C,D)$ that cross with all links of size one. Then $G$ is an octahedron, which is a cycle of $K_{2,2}$-bags of length 3.
Proof.
By Lemma 12.9, every bag of the block-bagel $\mathcal{O}=(O,\mathcal{G})$ induced by $(A,B)$ and $(C,D)$ is a $K_{2}$. Hence every edge of $O$ is helpful, so $|O|\geqslant 4$ by the (Helpful Lemma). (8.6). Since each vertex of $G$ contained in $\bigcup_{t\in V(O)}G_{t}$ has degree $\leqslant 4$ in $G$, the graph $G$ is 4-regular. Let $u,v$ be the vertices in the centre. By Lemma 8.7 it follows that $|O|=4$ and $uv\notin E(G)$. Then $G$ is an octahedron. ∎
Lemma 12.11.
Let $G$ be a vertex-transitive 4-connected finite graph that has two tetra-separations $(A,B),(C,D)$ that cross with all links of size two. Let $\mathcal{O}$ be the Tutte-bagel induced by $(A,B)$ and $(C,D)$. Let $H_{t}$ be a good torso of $\mathcal{O}$. Then the bag $G_{t}$ is a $K_{4}$ and its two neighbouring torsos are 4-cycles.
Proof.
We may assume the (Tutte-bagel torso setting). (9.13) with $t=t(0)$. By 9.41$(U,W)$ is a totally-nested tetra-separation. By Lemma 12.5, $S(U,W)$ consists of four edges, and we have either $G[U]=K_{4}$ or $G[W]=K_{4}$.
We claim that $G[U]=K_{4}$. Assume otherwise, so $G[W]=K_{4}$. Then $G[W]=G_{s}$ for some bag $G_{s}\neq G_{t}$. Hence $O$ has length $\leqslant 4$, two 3-connected bags (namely $G_{s},G_{t}$), and all other bags have triangles or 4-cycles as torsos by Lemma 9.14. But then $\alpha(\mathcal{O})<4$ contradicts Lemma 9.51.
Hence $G[U]=K_{4}$. Next, we claim that $S(U,W)$ contains no inward edge, and assume for a contradiction that $x_{1}\in S(U,W)$ is an inward edge with endvertex $u$ in the interior of $G_{t}$. By Lemma 9.33, some $y^{\prime}_{i}$, say $y^{\prime}_{1}$, is a neighbour of $x^{\prime}_{1}$ and $y^{W}_{1}=y^{\prime}_{1}$. In particular, $y_{1}$ cannot be an outward edge. Since $S(U,W)$ consists of four edges and $y_{1}$ is one of them, $y_{1}$ must be an inward edge. Let $u$ be the endvertex of $y_{1}$ in the interior of $G_{t}$. Then $u\neq v$ by the matching-condition for $(U,W)$. Let $w$ be a neighbour of $x^{\prime}_{1}$ in $G$ but outside of $G_{t}$. Then the vertex-set $\{u,v,w\}$ is independent in $G$ and consists entirely of neighbours of $x^{\prime}_{1}$. Since $G[U]=K_{4}$ and $S(U,W)$ consists of four edges, $G$ is 4-regular and every vertex lies in a $K_{4}$ by vertex-transitivity. But $\{u,v,w\}$ witnesses that $x^{\prime}_{1}$ does not lie in a $K_{4}$, a contradiction. Therefore, $S(U,W)$ contains no inward edge.
So all edges in $S(U,W)$ are outward edges. Hence $G_{t}=K_{4}$ and both neighbouring torsos are 4-cycles. ∎
Lemma 12.12.
Let $G$ be a vertex-transitive 4-connected finite graph that has two tetra-separations $(A,B),(C,D)$ that cross with all links of size two. Let $\mathcal{O}$ be the Tutte-bagel induced by $(A,B)$ and $(C,D)$. If some torso of $\mathcal{O}$ is good, then $G$ is a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 8$.
Proof.
Write $\mathcal{O}=:(O,(G_{t})_{t\in V(O)})$. Write $O=t(0)\,t(1)\ldots t(m)\,t(0)$ with indices in $\mathbb{Z}_{m+1}$. By Lemma 12.11, if a torso such as $H_{t(0)}$ is good, then its bag is a $K_{4}$ and the two neighbouring torsos $H_{t(-1)}$ and $H_{t(1)}$ are 4-cycles. Then, by Lemma 9.4, the torsos $H_{t(-2)}$ and $H_{t(2)}$ are 3-connected. Hence, both of these torsos are good as they neighbour torsos that are 4-cycles. By induction, we get that $O$ has even length, that bags $G_{t(i)}$ are $K_{4}$’s for even $i$, and that torsos $H_{t(i)}$ are 4-cycles for odd $i$. By Lemma 9.51, there are $\geqslant 4$ bags that are $K_{4}$’s. It follows that $G$ is a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 8$. ∎
Lemma 12.13.
Let $\mathcal{O}$ be a Tutte-bagel of a 4-connected graph. Let $G_{t}$ be a bag of $\mathcal{O}$ on five vertices. Let $\pi,\pi^{\prime}$ be the two adhesion-sets at $G_{t}$, and let $v$ be the interior vertex of $G_{t}$. Then the bag $G_{t}$ contains a full matching between $\pi$ and $\pi^{\prime}$, and $v$ sends edges to all four vertices in $\pi\cup\pi^{\prime}$.
Proof.
Since $G$ is 4-connected, $v$ sends edges to all four vertices in $\pi\cup\pi^{\prime}$. Since the torso $H_{t}$ is 3-connected with $|H_{t}|=5$, it is either a $4$-wheel, a $K_{5}^{-}$ or a $K_{5}$. In each case, $H_{t}$ contains a full matching between $\pi$ and $\pi^{\prime}$, and both edges in the matching also exist in the bag $G_{t}$. ∎
Lemma 12.14.
Let $G$ be a vertex-transitive 4-connected finite graph that has two tetra-separations $(A,B),(C,D)$ that cross with all links of size two. Let $\mathcal{O}$ be the Tutte-bagel induced by $(A,B)$ and $(C,D)$. Assume that $\mathcal{O}$ has a torso on five vertices. Then $G$ is the line graph of the cube.
Proof.
Write $\mathcal{O}=:(O,\mathcal{G})$. By Lemma 12.12, all bags of $\mathcal{O}$ are bad. Let $H_{t}$ be a torso of $\mathcal{O}$ on five vertices. Since the unique interior vertex $x$ of $H_{t}$ has degree 4, the graph $G$ is 4-regular. Since $H_{t}$ is bad, it is neighboured by a 3-connected torso $H_{s}$. Since $H_{s}$ is bad as well, it has $\leqslant 5$ vertices. That is, the torso of $H_{s}$ either is a $K_{4}$ or has exactly five vertices.
We claim that $H_{s}$ has exactly five vertices. Otherwise $H_{s}=K_{4}$. Let $\{u,v\}$ denote the adhesion-set in between the torsos $H_{t}$ and $H_{s}$. Then $uv\notin E(G)$ by 4-regularity of $G$. Let $\{u^{\prime},v^{\prime}\}$ denote the other adhesion-set at $H_{s}$. Since $v$ is contained in at least two triangles in $G$, we must have $u^{\prime}v^{\prime}\in E(G)$. The vertex $v$ is contained in exactly two triangles: one from $H_{s}$ and the other is $vu^{\prime}v^{\prime}$. The vertex $v^{\prime}$, however, is contained in at least two triangles (namely $vu^{\prime}v^{\prime}$ and $uu^{\prime}v^{\prime}$), two of which intersect in the edge $u^{\prime}v^{\prime}$. This contradicts the vertex-transitivity of $G$. Hence $H_{s}$ has exactly five vertices.
It follows by induction that all torsos of $\mathcal{O}$ have exactly five vertices. By Lemma 12.13, each bag contains a 2-matching between its two adhesion-sets, and the interior vertex of the bag sends edges to all four other vertices in the bag. By 4-regularity and since every bag looks like this, these are all the edges that there are in the bags.
Since each bag has a unique interior vertex, the set $X$ of all these vertices has size $|O|$. Since all torsos are 3-connected, we have $|O|\geqslant 4$ by Lemma 9.51. If $|O|\geqslant 5$, then every vertex in $X$ is contained in exactly two induced 4-cycles in $G$, while every vertex in an adhesion set is contained in exactly one induced 4-cycle in $G$, contradicting vertex-transitivity. Therefore, we have $|O|=4$. Consider the graph $H:=G-X$, which is either an 8-cycle or a disjoint union of two 4-cycles. If $H$ is an 8-cycle, then every vertex in $X$ is contained in two induced 4-cycles of $G$ while every vertex in $V(G-X)$ is only contained in one induced 4-cycle of $G$, contradicting vertex-transitivity. Therefore $H$ is a disjoint union of two 4-cycles, and it follows that $G$ is the line graph of the cube. ∎
Lemma 12.15.
Let $G$ be a vertex-transitive 4-connected finite graph that has two tetra-separations $(A,B),(C,D)$ that cross with all links of size two. Let $\mathcal{O}$ be the Tutte-bagel induced by $(A,B)$ and $(C,D)$. Assume that every bag of $\mathcal{O}$ has $\leqslant 4$ vertices. Then $G$ is one of the following graphs:
  • $\cdot$
    a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 8$;
  • $\cdot$
    a cycle of triangle-bags of length $\geqslant 6$;
  • $\cdot$
    a cycle of $K_{2,2}$-bags of length $\geqslant 4$;
  • $\cdot$
    a cycle of $K_{4}$-bags of length $\geqslant 4$.
Proof.
If $\mathcal{O}$ has a good torso, then $G$ is a cycle alternating between $K_{4}$-bags and $C_{4}$-bags by Lemma 12.12, and it is of length $\geqslant 8$ by Lemma 9.51. Hence, we may assume that all torsos of $\mathcal{O}$ are bad.
We claim that $\mathcal{O}$ has no 4-cycle as torso. Indeed, if $\mathcal{O}$ has a 4-cycle as a torso, then the neighbouring torsos are 3-connected by Lemma 9.4, and they are good, a contradiction. Thus,
(8) all torsos are either triangles or $K_{4}$’s.
We claim that if some torso of $\mathcal{O}$ is a triangle, then also its neighbouring torsos are triangles. For this, let $H_{t}$ and $H_{s}$ be neighbouring torsos with $H_{t}$ a triangle. By 8, the torso $H_{s}$ is either a triangle or a $K_{4}$. Assume for a contradiction that $H_{s}=K_{4}$. Let $\{u,v\}$ be the adhesion-set in between $H_{t}$ and $H_{s}$, so that $v$ is the tip of the triangle-torso $H_{t}$, say. Since every vertex has degree $\geqslant 4$, the edge $uv$ must exist in $G$, and $G$ must be 4-regular. Let $w$ be the vertex of the triangle $H_{t}$ besides $u$ and $v$. Let $G_{r}$ be the bag neighbouring $G_{t}$ besides $G_{s}$. Then $v$ has exactly one neighbour in the bag $G_{r}$. By 8, the torso $H_{r}$ can only be a triangle or a $K_{4}$. As $v$ has exactly one neighbour in the bag $G_{r}$, the torso $H_{r}$ cannot be a $K_{4}$, so it must be a triangle. But then the edge $vw$ exists by Lemma 9.4, so $v$ has two neighbours in $G_{r}$, giving it degree five in $G$, contradicting 4-regularity. Therefore, every triangle-torso of $\mathcal{O}$ is neighboured by two triangle-torsos.
So if $\mathcal{O}$ has at least one triangle-torsos, then all torsos are triangles, and all adhesion-sets are spanned by edges by Lemma 9.4. Hence $G$ is a cycle of triangle-bags, and it has length $\geqslant 6$ by Lemma 9.51 and since $G\neq K_{5}$.
So we may assume that $\mathcal{O}$ has no triangle-torsos. Then by 8, all torsos of $\mathcal{O}$ are $K_{4}$’s. Then either all adhesion-sets are spanned by edges in $G$, in which case $G$ is a cycle of $K_{4}$-bags; or none of them are, in which case $G$ is a cycle of $K_{2,2}$-bags. In both cases, the length is $\geqslant 4$ by Lemma 9.51. ∎
Lemma 12.16.
Let $G$ be a vertex-transitive 4-connected finite graph that has two tetra-separations $(A,B),(C,D)$ that cross with all links of size two. Then $G$ is one of the following:
  • $\cdot$
    a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 8$;
  • $\cdot$
    a cycle of triangle-bags of length $\geqslant 6$;
  • $\cdot$
    a cycle of $K_{2,2}$-bags of length $\geqslant 4$;
  • $\cdot$
    a cycle of $K_{4}$-bags of length $\geqslant 4$;
  • $\cdot$
    line graph of the cube.
Proof.
We combine Lemmas 12.14 and 12.15. ∎
Corollary 12.17.
Let $G$ be a vertex-transitive 4-connected finite graph that has two tetra-separations $(A,B),(C,D)$ that cross. Then $G$ is one of the following:
  • $\cdot$
    a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 8$;
  • $\cdot$
    a cycle of triangle-bags of length $\geqslant 6$;
  • $\cdot$
    a cycle of $K_{2,2}$-bags of length $\geqslant 3$;
  • $\cdot$
    a cycle of $K_{4}$-bags of length $\geqslant 4$;
  • $\cdot$
    line graph of the cube.
Proof.
By the (Crossing Lemma). (4.1), all links have the same size 00, $1$ or $2$. If all links have size 00, then $G$ is a cycle of $K_{2,2}$-bags of length $4$ by Lemma 12.7. If all links have size $1$, then $G$ is a cycle of $K_{2,2}$-bags of length $3$ by Lemma 12.10. Otherwise all links have size $2$, and we apply Lemma 12.16. ∎
Observation 12.18.
Let $G^{\prime}$ be the expansion of a graph $G$. Let $v$ be a vertex of degree $d$ in $G$, and let $C_{v}$ be the clique of $G^{\prime}$ that corresponds to $v$. Then every vertex in $C_{v}$ has degree $d$ in $G^{\prime}$.∎
Observation 12.19.
Let $G$ be a graph of minimum degree at least $2$, and let $G^{\prime}$ be the expansion of $G$. Then there is a unique partition $\mathcal{P}$ of $V(G^{\prime})$ into (the vertex-sets of) $|\mathcal{P}|=|V(G)|$ cliques. Moreover, the contraction-minor $G^{\prime}/\mathcal{P}$ is isomorphic to $G$.∎
Observation 12.20.
The expansion of a graph $G$ is isomorphic to the graph $H$ with
$V(H)$ =${}=\{(u,v),(v,u):uv\in E(G)\}\text{ and}$
$E(H)$ =${}=\Big{\{}\{(u,v),(v,u)\}:uv\in E(G)\Big{\}}\,\cup\,\Big{\{}\{(u,v_{1}),(u,v_{2})\}:uv_{1},uv_{2}\in E(G)\Big{\}}.$
Lemma 12.21.
Let $G$ be an $r$-regular graph with $r\geqslant 3$. Then the following assertions are equivalent:
  1. (i)
    the $K_{r}$-expansion of $G$ is vertex-transitive;
  2. (ii)
    $G$ is arc-transitive.
Proof.
[…]
(i) $\Rightarrow$ (ii). Let $(u_{1},v_{1}),(u_{2},v_{2})$ be ordered pairs of vertices in $G$ with $u_{i}v_{i}\in E(G)$. We will find an automorphism $\varphi$ of $G$ such that $\varphi(u_{1})=u_{2}$ and $\varphi(v_{1})=v_{2}$. For this, we abbreviate $x_{i}:=(u_{i},v_{i})$. Let $P_{i}\in\mathcal{P}$ be the unique partition-class of $\mathcal{P}$ that contains $x_{i}$. Since $G^{\prime}$ is vertex-transitive by assumption, there is an automorphism $\varphi^{\prime}$ of $G^{\prime}$ that maps $x_{1}$ to $x_{2}$. Note that $\varphi$ then maps $P_{1}$ to $P_{2}$. We define $\varphi:=\gamma\circ\varphi^{\prime}\circ\gamma^{-1}$ by abuse of notation, using that $\operatorname{Aut}(G^{\prime})$ acts on $\mathcal{P}$. Clearly, $\varphi$ is a bijection $V(G)\to V(G)$. Moreover, we have
[…]
$\Leftrightarrow$ there is a $\gamma^{-1}(x)$$\gamma^{-1}(y)$ edge in $G^{\prime}$
$\Leftrightarrow$ there is a $\varphi^{\prime}(\gamma^{-1}(x))$$\varphi^{\prime}(\gamma^{-1}(y))$ edge in $G^{\prime}$
$\Leftrightarrow$ $\varphi(x)\varphi(y)\in E(G)$.
So, $\varphi$ is an automorphism of $G$. Moreover, we have
$\varphi(u_{1})=\gamma(\varphi^{\prime}(\gamma^{-1}(u_{1})))=\gamma(\varphi^{\prime}(P_{1}))=\gamma(P_{2})=u_{2}.$
Similarly, $\varphi(v_{1})=v_{2}$.
(ii) $\Rightarrow$ (i). We assume without loss of generality that $G^{\prime}$ takes the form described in 12.20. Let $x_{1},x_{2}\in V(G^{\prime})$ be two vertices. We will find an automorphism $\varphi^{\prime}$ of $G^{\prime}$ that maps $x_{1}$ to $x_{2}$. We write $x_{i}=:(u_{i},v_{i})$ using 12.20, so $u_{i}v_{i}$ is an edge of $G$. Since $G$ is arc-transitive, there is an automorphism $\varphi$ of $G$ such that $\varphi(u_{1})=u_{2}$ and $\varphi(v_{1})=v_{2}$. Now $\varphi$ defines an automorphism $\varphi^{\prime}$ of $G^{\prime}$ that sends every vertex $x=(u,v)$ to the vertex $\varphi^{\prime}(x):=(\varphi(u),\varphi(v))$. The automorphism $\varphi^{\prime}$ satisfies
$\varphi^{\prime}(x_{1})=(\varphi(u_{1}),\varphi(v_{1}))=(u_{2},v_{2})=x_{2}.\qed$
Corollary 12.22.
Let $G$ be an $r$-regular graph with $r\geqslant 3$. Let $(A,B)$ be a mixed-$r$-separation of $G$ such that $S(A,B)$ consists only of edges and $S(A,B)$ is a matching and $G[A]=K_{r}$. Then $G$ is not arc-transitive. Moreover, the $K_{r}$-expansion of $G$ is not vertex-transitive.
Proof.
By Lemma 12.21, it suffices to show that $G$ is not arc-transitive. Indeed, if $u$ is a vertex in $A$, then $G[N(u)]=K_{1}\sqcup K_{r-1}$ with $r-1\geqslant 2$. Let $v_{1}$ be the isolated vertex in $G[N(u)]$ and let $v_{2}\neq v_{1}$ be another vertex in $N(u)$. Then there is no automorphism of $G$ that maps $(u,v_{1})$ to $(u,v_{2})$. ∎
Corollary 12.23.
Let $r\geqslant 3$. The $K_{r}$-expansion of an $r$-regular graph $G$ is not arc-transitive. In particular, the $K_{r}$-expansion of the $K_{r}$-expansion of an $r$-regular graph $G$ is not vertex-transitive.
Proof.
Follows from Corollary 12.22. ∎
Corollary 12.24.
A cycle of triangle-bags of length $\geqslant 6$ is not arc-transitive. In particular, the $K_{4}$-expansion of a cycle of triangle-bags of length $\geqslant 6$ is not vertex-transitive.
Proof.
Let $G$ be a cycle of triangle-bags of length $\geqslant 6$, and let $\mathcal{O}$ be the corresponding cycle-decomposition. Note that every free edge of $\mathcal{O}$ is contained in exactly one triangle of $G$, while every other edge of $G$ is contained in exactly two triangles. So, $G$ is not edge-transitive. In particular, it is not arc-transitive. By Lemma 12.21, the $K_{4}$-expansion of a cycle of triangle-bags of length $\geqslant 6$ is not vertex-transitive. ∎
Corollary 12.25.
Let $G$ be a finite graph that is arc-transitive, 4-connected, 4-regular and has two tetra-separations that cross. Then $G$ is the line graph of the cube or a cycle of $K_{2,2}$-bags of length $\geqslant 3$.
Proof.
This follows from Corollary 12.17 together with the following observations. First, a cycle of $K_{4}$-bags of length $\geqslant 4$ is not 4-regular. Second, a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 8$ is not arc-transitive by Corollary 12.22. Finally, a cycle of triangle-bags of length $\geqslant 6$ is not arc-transitive by Corollary 12.24. ∎
Lemma 12.26.
Let $G$ be a vertex-transitive essentially 5-connected graph. Then either $G$ is quasi-5-connected or there is a partition $\mathcal{P}$ of $V(G)$ into vertex-sets of $K_{4}$’s such that the edges in $G$ with ends in different classes of $\mathcal{P}$ form a perfect matching in $G$.
Proof.
If $G$ does not have a tetra-separation, then $G$ is quasi-5-connected by Lemma 2.3 (i). Hence, we may assume that $G$ has a tetra-separation $(A,B)$. Since $G$ is essentially 5-connected, the separator of $(A,B)$ consists of four edges and there is a side, say $A$, that induces a $K_{4}$ in $G$. Let $u\in A$ and observe that $G[N(u)]=K_{1}\sqcup K_{3}$; in particular, $u$ is contained in exactly one $K_{4}$. By vertex-transitivity, so is every vertex of $G$. It follows that $G$ admits a partition of its vertex-set into the vertex-sets of $K_{4}$’s. Moreover, $G$ is 4-regular since $u$ has degree 4 in $G$. Then it follows that the edges between distinct partition-classes of $\mathcal{P}$ form a perfect matching in $G$. ∎
Lemma 12.27.
Let $G$ be a vertex-transitive essentially 5-connected graph. Then $G$ is one of the following:
  • $\cdot$
    quasi-5-connected;
  • $\cdot$
    a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\ell\in\{4,6\}$;
  • $\cdot$
    the $K_{4}$-expansion of an arc-transitive 4-connected 4-regular graph.
Proof.
Assume that $G$ is not quasi-5-connected. Then by Lemma 12.26, there is a partition $\mathcal{P}$ of $V(G)$ into the vertex-sets of $K_{4}$’s such that edges between distinct partition-classes of $\mathcal{P}$ form a perfect matching in $G$. We distinguish the following cases.
[…]
Corollary 12.28.
Let $G$ be an arc-transitive essentially 5-connected graph. Then $G$ is quasi-5-connected.
Proof.
Follows from Lemma 12.27 together with the following observations. First, a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\ell\in\{4,6\}$ is not arc-transitive by Corollary 12.22. Second, the $K_{4}$-expansion of a graph is not arc-transitive by Corollary 12.23. ∎
Proof of 2.
Let $G$ be a 4-connected finite graph.
[…]
($\Leftarrow$). Clearly, if $G$ is one of the graphs in (i), (iii), (iv), (v), then $G$ is vertex-transitive. If (ii) $G$ is the $K_{4}$-expansion of a quasi-5-connected arc-transitive graph, then $G$ is vertex-transitive by Lemma 12.21. Both (vi) a cycle of $K_{2,2}$-bags and (vii) the line graph of the cube are arc-transitive. If $G$ is a $K_{4}$-expansion of (vi) or (vii), then $G$ is vertex-transitive by Lemma 12.21. ∎
Acknowledgement.
We thank Joseph Devine for exciting discussions. We thank Agelos Georgakopoulos for telling us about Theorem 12.1 of Godsil and Royle.
Appendix A Circular saws
Here we verify that the circular saws from the introduction have the desired properties.
Lemma A.1.
Let $G$ be a $k$-regular circular saw on $\mathbb{Z}_{n}\times\mathbb{Z}_{2}$ with $n\gg k$. Let $aa^{\prime}$ and $bb^{\prime}$ be two edges of $G$ such that $\{a,a^{\prime}\}$ and $\{b,b^{\prime}\}$ are disjoint and there is no $\{a,a^{\prime}\}$$\{b,b^{\prime}\}$ edge. Let $X$ denote the neighbourhood of $\{a,a^{\prime}\}$, and let $Y$ denote the neighbourhood of $\{b,b^{\prime}\}$. There are $2(k-1)$ disjoint $X$$Y$ paths in $G$.
Proof.
For each vertex $x\in X$ we construct two $x$$Y$ paths $P^{+}_{x},P^{-}_{x}$ inductively, as follows. See Figure A.1 for an example.
In the first step of the construction of $P^{+}_{x}$, we let $P^{+}_{x}$ start and end with $x$. Assume now that $P^{+}_{x}$ ends with the vertex $(\ell,i)$. If $(\ell,i)\in Y$, then we terminate the construction. Otherwise $(\ell,i)\notin Y$. If $i=0$ we append the vertex $(\ell+k-1,1)$, and if $i=1$ we add the vertex $(\ell,0)$. This completes the construction of $P^{+}_{x}$.
In the first step of the construction of $P^{-}_{x}$, we let $P^{-}_{x}$ start and end with $x$. Assume now that $P^{-}_{x}$ ends with the vertex $(\ell,i)$. If $(\ell,i)\in Y$, then we terminate the construction. Otherwise $(\ell,i)\notin Y$. If $i=0$ we append the vertex $(\ell,1)$, and if $i=1$ we add the vertex $(\ell-k+1,0)$. This completes the construction of $P^{-}_{x}$.
Exactly one of the paths $P^{+}_{x},P^{-}_{x}$ has an internal vertex in $X$. Let $P_{x}$ denote the other path. Then $(P_{x}:x\in X)$ is a family of disjoint $X$$Y$ paths. There are $|X|=2(k-1)$ of these paths. ∎
Figure A.1. Paths constructed in the proof of Lemma A.1. The edges $aa^{\prime}$ and $bb^{\prime}$ are coloured blue, the paths are coloured red.
Lemma A.2.
Let $G$ be a $k$-regular circular saw on $\mathbb{Z}_{n}\times\mathbb{Z}_{2}$ with $n\gg k\geqslant 3$. Then $G$ is $k$-connected and the neighbourhoods of its vertices are precisely its $k$-separators.
Proof.
We show first that $G$ is $k$-connected, and assume for a contradiction that some two vertices $a,b$ of $G$ are separated by a set $S$ of $k-1$ vertices. Let $a^{\prime}$ be a neighbour of $a$ outside $S$, and let $b^{\prime}$ be a neighbour of $b$ outside $S$. Note that $12(k-1)\geqslant k+$ 1 since $k\geqslant 3$. Hence by Lemma A.1, there are $k+1$ disjoint $N(\{a,a^{\prime}\})$$N(\{b,b^{\prime}\})$ paths in $G$, contradicting that $|S|\leqslant k-1$. Hence $G$ is $k$-connected.
Next, we assume for a contradiction that $G$ has a $k$-separation $(A,B)$ such that $S(A,B)$ fails to be the neighbourhood of some vertex of $G$. Let $a\in A\smallsetminus B$ and $b\in B\smallsetminus A$. Then there are vertices $a^{\prime}\in N(a)\smallsetminus S(A,B)$ and $b^{\prime}\in N(b)\smallsetminus S(A,B)$. Now Lemma A.1 yields $k+1$ disjoint $N(\{a,a^{\prime}\})$$N(\{b,b^{\prime}\})$ paths in $G$, contradicting that $(A,B)$ has order $k$. ∎
Appendix B Comparison of strict tri-separations and tri-separations
The strict tri-separation-decomposition (2) is different from the tri-separation decomposition [Tridecomp, Theorem 1] as can be seen in Figures B.1 and B.2. However, the strict tri-separation-decomposition can be viewed as a refinement of the tri-separation decomposition via Lemma B.1.
Figure B.1. The totally-nested strict tri-separations
Figure B.2. The totally-nested non-trivial tri-separations
A tri-separation $(A,B)$ is negligible if $|A\smallsetminus B|\leqslant 1$ or $|B\smallsetminus A|\leqslant 1$.
Lemma B.1.
Let $G$ be a 3-connected graph. Let $(A,B)$ be a totally-nested non-trivial tri-separation of $G$. Then the left-right-shift $(A^{\prime},B^{\prime})$ of $(A,B)$ is a totally-nested strict tri-separation, or $(A,B)$ is negligible.
Proof.
Assume that $(A,B)$ is not negligible. Then $(A^{\prime},B^{\prime})$ is a strict tri-separation by 5.10 (which generalises to strict tri-separations via an apex-argument as in the proof of Corollary 3.2). Assume for a contradiction that there is a strict tri-separation $(C,D)$ of $G$ that crosses $(A^{\prime},B^{\prime})$ with $C$-link $\gamma^{\prime}$ and $D$-link $\delta^{\prime}$. We will show that $(C,D)$ crosses $(A,B)$ with non-empty $C$-link $\gamma$ and non-empty $D$-link $\delta$. For this, we consider the crossing-diagram of $(C,D)$ and $(A^{\prime},B^{\prime})$. By [Tridecomp, Lemma 1.3.10], there are two cases.
Case: the centre consists of one vertex and all links have size one. Then $\gamma^{\prime}$ and $\delta^{\prime}$ are non-empty. If $\gamma^{\prime}$ consists of a vertex $v$, then $\gamma$ also consists of $v$. Otherwise $\gamma^{\prime}$ consists of an edge $e$. The edge $e$ cannot be dangling, since then its endvertex in $S(C,D)$ would violate the degree-condition. Hence the endvertices of $e$ are contained in the corners $A^{\prime}C$ and $B^{\prime}C$, respectively. So $e$ or an endvertex of $e$ is contained in $\gamma$. This concludes the proof that $\gamma$ is non-empty, and a similar argumentation shows that $\delta$ is non-empty. Hence $(C,D)$ and $(A,B)$ cross with non-empty opposite links $\gamma$ and $\delta$.
Case: the centre consists of three vertices and all links are empty. Then $S(A^{\prime},B^{\prime})$ consists of three vertices, so $(A^{\prime},B^{\prime})=(A,B)$. But all corners are non-empty, so $(A^{\prime},B^{\prime})$ is not half-connected. However, $(A,B)$ is half-connected because it is totally-nested, by [Tridecomp, Lemma 2.3.4], contradicting that $(A^{\prime},B^{\prime})=(A,B)$. ∎
[…]

### v2

@@
-A Tutte-type canonical decomposition of 4-connected graphs
+A Tutte-type canonical decomposition of 3- and 4-connected graphs
 Abstract.
@@
 As a corollary, we obtain a new Tutte-type canonical decomposition of 3-connected graphs into parts that are either quasi-4-connected, generalised wheels or thickened $K_{3,m}$’s.
-This decomposition is similar yet different from the tri-separation decomposition [Tridecomp].
-As an application of the decomposition for 4-connectivity, we obtain a new theorem characterising all 4-connected vertex-transitive finite graphs as quasi-5-connected, the $K_{4}$-expansion of a quasi-5-connected graph, or on a short explicit list of graphs.
+This decomposition is similar yet different from the tri-separation decomposition [CK23].
+As an application of the decomposition for 4-connectivity, we obtain a new theorem characterising all quasi-4-connected vertex-transitive finite graphs as essentially quasi-5-connected or on a short explicit list of graphs.
 Key words and phrases:
@@
 05C40, 05C75, 05C83, 05E18
-[Uncaptioned image] Funded by DFG, project number 546892829.
+[Uncaptioned image] funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 566118291; 546892829
+funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 546892829
 1. Introduction
 Problem 1.1.
-Is there for every $k\in\mathbb{N}$ an explicit canonical way to decompose every $k$-connected graph along $k$-separators into smaller pieces that are $(k+1)$-connected or ‘basic’?
+Is there for every $k\in\mathbb{N}$ an explicit canonical way to decompose every $k$-connected graph along kk-‘separators’ into smaller pieces that are $(k+1)$-connected or ‘basic’?
 […]
-1.1 has a trivial solution for $k=0$, and for $k=1$ a positive answer is provided by the block-cutvertex-decomposition.
-Already for $k=2$, the solution is a classical theorem of Tutte [TutteCon] from 1966.
-It was not known how to extend Tutte’s approach to $k=3$ until recently, when Carmesin and Kurkofka discovered a new notion of 3-separators which allowed them to find a Tutte-type canonical decomposition of 3-connected graphs [Tridecomp].
+˜1.1 has a trivial solution for $k=0$, and for $k=1$ a positive answer is provided by the block-cutvertex-decomposition.
+Already for $k=2$, the solution is a classical theorem of Tutte [T66] from 1966.
+A non-canonical solution for $k=3$ was found by Grohe [G16] in 2016.
+Still, it was not known how to extend Tutte’s approach to $k=3$ until recently, when Carmesin and Kurkofka discovered a new notion of 3-separators which allowed them to find a Tutte-type canonical decomposition of 3-connected graphs [CK23].
 However, examples show that the obvious extension of their solution to $k=4$ fails.
@@
 In this paper, we introduce a new type of 4-separator and, as our main result, use it to prove a Tutte-type canonical decomposition theorem for 4-connectivity.
-As a corollary, we obtain a new Tutte-type canonical decomposition for 3-connectivity that provides an alternative perspective to [Tridecomp].
-We also obtain a new theorem characterising vertex-transitive 4-connected finite graphs.
+As a corollary, we obtain a new Tutte-type canonical decomposition result for 3-connectivity that provides an alternative perspective to [CK23].
+We also obtain a new theorem that characterises vertex-transitive quasi-4-connected finite graphs.
 Background.
-To solve 1.1 for $k=0$ and $k=1$, one simply cuts along all $k$-separators of the graph.
+To solve ˜1.1 for $k=0$ and $k=1$, one simply cuts along all $k$-separators of the graph.
 But already for $k=2$, this approach fails.
@@
 The resulting pieces of Tutte’s decomposition are 3-connected, cycles or $K_{2}$’s.
-All three decompositions for $k\leqslant 3$ can be described in terms of tree-decompositions.
+All decompositions for $k\leqslant 2$ can be described in terms of tree-decompositions.
 For instance, Tutte’s selection of 2-separators gives a recipe for a tree-decomposition with all adhesion-sets of size $2$ and all whose torsos [^1] A torso of a tree-decomposition is obtained from a bag by making each of its adhesion-sets into a clique.
@@
 For example, if $G$ is $k$-connected and the neighbourhoods of its vertices are precisely the $k$-separators of $G$, then the $k$-separators around neighbouring vertices cross, but $G$ fails to be $(k+1)$-connected.
-Such $G$ exist for all $k\geqslant 3$: take a circular saw as in Figure 1.1, for instance.
-Figure 1.1.
-A circular saw has vertex-set $\mathbb{Z}_{n}\times\mathbb{Z}_{2}$ with $n\gg k$ and every vertex $(v,0)$ sends edges to the $k$ vertices $(v,1),\ldots,(v+k-1,1)$.
-It is $k$-connected, and when $k\geqslant 3$ its neighbourhoods of vertices are precisely its $k$-separators (Lemma A.2).
+Such $G$ exist for all $k\geqslant 3$: take a circular saw as in Figure˜1.1, for instance.
+We believe that such graphs should not have to be decomposed any further at level $k$.
+Figure $1.$ A circular saw has vertex-set $\mathbb{Z}_{n}\times\mathbb{Z}_{2}$ with $n\gg k$ and every vertex $(v,0)$ sends edges to the $k$ vertices $(v,1),\ldots,(v+k-1,1)$.
+It is $k$-connected, and when $k\geqslant 3$ its neighbourhoods of vertices are precisely its $k$-separators (Lemma˜14.2).
 Challenge 2.
-We consider the $k$-connected graph $G$ depicted in Figure 1.2 if $k$ is odd and in Figure 1.3 if $k$ is even.
+We consider the $k$-connected graph $G$ depicted in Figure˜1.2 if $k$ is odd and in Figure˜1.3 if $k$ is even.
 Every $k$-separator of $G$ is formed from a red set, a non-adjacent blue set, and if $k$ is even also the apex-vertex in the centre.
@@
 Figure 1.3. $k$ is even. Red: $(\tfrac{k}{2}-1)$-cliques. Blue: $\tfrac{k}{2}$-cliques. Grey indicates all possible edges. The central vertex is an apex.
-Decomposing 4-connected and 3-connected graphs. We found a way to overcome the two challenges for $k=4$ and $k=3$ with a twofold approach, which we first explain for $k=4$.
-Ad 1.
-We relax the notion of $5$-connected to quasi-5-connected, an idea that we learned about from Grohe’s work [grohe2016quasi].
+Decomposing 3- and 4-connected graphs. We found a way to overcome the two challenges for $k=3$ and $k=4$ with a twofold approach, which we first explain for $k=4$.
+Ad ˜1.
+We relax the notion of $5$-connected to quasi-5-connected, an idea that we learned about from Grohe’s work [G16].
 A graph $G$ is quasi-5-connected if it is 4-connected and every 4-separation $(A,B)$ of $G$ satisfies $|A\smallsetminus B|=1$ or $|B\smallsetminus A|=1$.
-Ad 2.
+We will not decompose quasi-5-connected graphs any further.
+Ad ˜2.
 We introduce the new notion of a tetra-separation, which we use instead of 4-separators.
 There are two key differences.
-On the one hand, tetra-separations may use edges in addition to vertices to cut up the graph, an idea that we learned about from the work of Carmesin and Kurkofka [Tridecomp] (we will discuss the relation later).
+On the one hand, tetra-separations may use edges in addition to vertices to cut up the graph, an idea that we learned about from the work of Carmesin and Kurkofka [CK23] (we will discuss the relation later).
 On the other hand, every vertex used by a tetra-separation to cut up the graph into two sides must have at least two neighbours in each side, not counting neighbours that are also used by the tetra-separation.
 This degree-condition is new.
-Recall that a mixed-separation of a graph $G$ is a pair $(A,B)$ with $A\cup B=V(G)$ and $A\smallsetminus B\neq\emptyset\neq B\smallsetminus A$.
+A mixed-separation of a graph $G$ is a pair $(A,B)$ with $A\cup B=V(G)$ and $A\smallsetminus B\neq\emptyset\neq B\smallsetminus A$.
 The separator of $(A,B)$ is the set of all vertices in $A\cap B$ and all edges in $G$ with one end in $A\smallsetminus B$ and the other in $B\smallsetminus A$.
+In a 4-connected graph, every mixed-separation has a separator of size $\geqslant 4$ (Lemma˜6.1).
 A tetra-separation of $G$ is a mixed-separation $(A,B)$ of $G$ whose separator has size four and which satisfies the following two conditions:
-  1. degree-condition:
-    every vertex in the separator has $\geqslant 2$ neighbours in both $A\smallsetminus B$ and $B\smallsetminus A$;
-  2. matching-condition:
-    the edges in the separator form a matching.
+degree-condition  : every vertex in the separator has $\geqslant 2$ neighbours in both $A\smallsetminus B$ and $B\smallsetminus A$;
+matching-condition  : the edges in the separator form a matching.
 Two tetra-separations $(A,B)$ and $(C,D)$ are nested if, after possibly exchanging the name $A$ with $B$ or the name $C$ with $D$, we have $A\subseteq C$ and $B\supseteq D$.
 A tetra-separation of $G$ is totally-nested if it is nested with every tetra-separation of $G$.
-Figure 1.4 shows the totally-nested tetra-separations of an example-graph similar to the one shown in Figure 1.3.
+Figure˜1.4 shows the totally-nested tetra-separations of an example-graph similar to the one shown in Figure˜1.3.
 Figure 1.4. The totally-nested tetra-separations of a graph. Their separators are coloured.
-While the decompositions for $k$-connectivity for $k\leqslant 2$ such as the Tutte-decomposition can be described as tree-decompositions, the decomposition of 4-connected graphs via tetra-separations requires a notion that sits in between the notions of tree-decomposition (which allows only vertices in adhesion-sets) and Wollan’s notion of tree-cut decomposition (which allows only edges in adhesion-sets) [WollanExcludedImmersion].
-To this end, the notion of mixed-tree-decomposition has been introduced which allows both vertices and edges in adhesion-sets [Tridecomp].
+While the decompositions for $k$-connectivity for $k\leqslant 2$ such as the Tutte-decomposition can be described as tree-decompositions, the decomposition of 4-connected graphs via tetra-separations requires a notion that sits in between the notions of tree-decomposition (which allows only vertices in adhesion-sets) and Wollan’s notion of tree-cut decomposition (which allows only edges in adhesion-sets) [W15].
+To this end, the notion of mixed-tree-decomposition has been introduced which allows both vertices and edges in adhesion-sets [CK23].
 The set of totally-nested tetra-separations of a 4-connected graph $G$ canonically corresponds to a unique mixed-tree-decomposition of $G$.
-See Section 2.2 for details.
+See Section˜2.2 for details.
 In order to state our decomposition-result for 4-connected graphs, we need the following definitions to describe the ‘basic’ parts of the decomposition – those that are not quasi-5-connected.
-Recall that graph-decompositions generalise tree-decompositions in that the decomposition-tree may take the form of an arbitrary graph, such as a cycle, in which case we speak of a cycle-decomposition [canonicalGraphDec].
+Graph-decompositions generalise tree-decompositions in that the decomposition-tree may take the form of an arbitrary graph, such as a cycle, in which case we speak of a cycle-decomposition [DJKK22].
 Let $X$ be a graph.
-We call $X$ a generalised double-wheel with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that $X-u-v$ has a cycle-decomposition of adhesion 1 into $K_{2}$’s and triangles (Figure 1.5).
-We call $X$ a cycle of $X_{1}$-torsos, …, $X_{k}$-torsos where the $X_{i}$ are graphs, if $X$ is 4-connected and has a cycle-decomposition with all adhesion-sets of size 2 such that every torso is isomorphic to some $X_{i}$ (Figure 1.6).
+We call $X$ a generalised double-wheel with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that XuvX-u-v has a cycle-decomposition of adhesion 1 into $K_{2}$’s and triangles (Figure˜1.5).
+We call $X$ a cycle of $X_{1}$-torsos, …, $X_{k}$-torsos, $Y_{1}$-bags, …, $Y_{\ell}$-bags where the $X_{i}$ and $Y_{j}$ are graphs, if $X$ is 4-connected and has a cycle-decomposition with all adhesion-sets of size 2 such that for every bag either its torso is isomorphic to some $X_{i}$ or the bag itself is isomorphic to some $Y_{j}$ (Figure˜1.6).
 Recall that the only 3-connected graphs on $\leqslant 5$ vertices are $K_{4}$, the $4$-wheel, $K_{5}$ minus an edge, and $K_{5}$.
@@
 Figure 1.5. A generalised double-wheel
-Figure 1.6. A cycle of triangle-torsos, $K_{4}$-torsos and $K_{5}$-torsos
+Figure 1.6. A cycle of triangle-bags, $K_{4}$-bags and $K_{5}$-bags
 Theorem 1.
@@
 Remark 1.2.
-Both $N(G)$ and $\mathcal{T}(G)$ in Theorem 1 are canonical: if $\varphi\colon G\to G^{\prime}$ is a graph-isomorphism, then $N(\varphi(G))=\varphi(N(G))$ and $\mathcal{T}(\varphi(G))=\varphi(\mathcal{T}(G))$.
+Both $N(G)$ and $\mathcal{T}(G)$ in Theorem˜1 are canonical: if $\varphi\colon G\to G^{\prime}$ is a graph-isomorphism, then $N(\varphi(G))=\varphi(N(G))$ and $\mathcal{T}(\varphi(G))=\varphi(\mathcal{T}(G))$, and the automorphism group of every $G$ induces a group-action on $N(G)$ and on $\mathcal{T}(G)$.
 Remark 1.3.
-Theorem 1 includes infinite graphs $G$ with no additional restrictions whatsoever. More generally, all graphs in this paper are allowed to be infinite, unless stated otherwise.
-Theorem 1 implies a new Tutte-type canonical decomposition for 3-connectivity.
+Theorem˜1 includes infinite graphs $G$ with no additional restrictions whatsoever. More generally, all graphs in this paper are allowed to be infinite, unless stated otherwise.
+Theorem˜1 implies a new Tutte-type canonical decomposition for 3-connectivity.
 A strict tri-separation is defined like a tetra-separation, except that the separator is required to have size three instead of four.
 A strict tri-separation of a graph $G$ is totally-nested if it is nested with every strict tri-separation of $G$.
-A generalised wheel with centre $v$ is a 3-connected graph $X$ with a vertex $v$ such that $X-v$ has a cycle-decomposition with all adhesion-sets of size 1 and into $K_{2}$’s and triangles.
+A generalised wheel with centre $v$ is a 3-connected graph $X$ with a vertex $v$ such that XvX-v has a cycle-decomposition with all adhesion-sets of size 1 and into $K_{2}$’s and triangles.
 Corollary 2.
@@
     $\tau$ is a thickened $K_{3,m}$ with $m\geqslant 0$, or the entire graph $G$ is a sprinkled $K_{3,m}$ with $m\geqslant 4$.
-Both Remarks 1.2 and 1.3 also apply to 2.
+Both Remarks˜1.2 and 1.3 also apply to ˜2.
+Comparison to the tri-separation decomposition.
+Carmesin and Kurkofka introduced the first Tutte-type canonical decomposition for 3-connected graphs [CK23].
+They used tri-separations instead of strict tri-separations to obtain their decomposition.
+The main difference between tri-separations on the one hand, and strict tri-separations and tetra-separations on the other hand, is how their degree-conditions handle neighbours in the separator: tri-separations count them, but strict tri-separations and tetra-separations ignore them.
+This seemingly subtle difference has a surprisingly big impact: the verbatim extension of the tri-separation approach to 4-connectivity fails ˜2.
+For 3-connected graphs, the strict tri-separation decomposition can be viewed as a refinement of the tri-separation decomposition, but the details are a bit technical; see Section˜15.
+Algorithmic aspects. It is straightforward to come up with a polynomial-time algorithm that produces the decompositions in Theorem˜1 and ˜2 for all 4- and 3-connected graphs.
+Korhonen [K24] recently showed that there is an algorithm that, given an $n$-vertex $m$-edge graph $G$ and an integer $k\geqslant 1$, returns a $k$-lean tree-decomposition of $G$ in time $k^{\mathcal{O}(k^{2})}n+\mathcal{O}(m)$.
+We believe that every 5-lean tree-decomposition of every 4-connected graph $G$ can be made into a tetra-separation decomposition of $G$ as in Theorem˜1 in linear time:
+Conjecture 1.4.
+There is an algorithm that, given a 4-connected $n$-vertex $m$-edge graph $G$, returns the tetra-separation decomposition of $G$ from Theorem˜1 in linear time $\mathcal{O}(n+m)$.
+A positive answer would also provide a linear-time algorithm for the decomposition of 3-connected graphs in ˜2 by following the short proof of ˜2.
+Total-nestedness makes parallel computing a feasible option for computing the tetra-separation decomposition even more efficiently.
+Indeed, all the partial solutions found in parallel, which come in the form of sets of totally-nested tetra-separations, can always be combined without conflict.
+Combining the decompositions for different $k$. Every finite graph $G$ can be decomposed canonically by combining the partial solutions to ˜1.1 inductively, as follows.
+  1. 1)
+    First, we trivially decompose $G$ into its connected components.
+  2. 2)
+    Each connected component we further decompose with the block-cutvertex decomposition into bags that are 2-connected or $K_{2}$’s.
+  3. 3)
+    Each 2-connected bag we further decompose with the Tutte-decomposition into torsos that are 3-connected, cycles or $K_{2}$’s.
+  4. 4)
+    Each 3-connected torso we further decompose with ˜2 into torsos that are quasi-4-connected, generalised wheels or thickened $K_{3,m}$’s. Alternatively, we use [CK23].
+[…]
+Figure 1.7. A quasi-4-connected version of ˜2: all tetra-separations are crossed, yet the graph is far from 55-connected and each $K_{10}$ can be replaced with an arbitrary 5-connected graph
+Let $H$ be an arbitrary graph.
+The following definition of $H^{\Delta}$ is supported by Figure˜1.8.
+Let $U$ be the set of all vertices in $H$ of degree three.
+We obtain $H^{s}$ from $H$ by subdividing every edge that joins two vertices in $U$.
+Then we obtain $H^{s\Delta}$ from $H^{s}$ by making each neighbourhood $N_{H^{s}}(u)$ around a vertex $u\in U$ into a triangle $\Delta(u)$.
+Finally, we let $H^{\Delta}:=H^{s\Delta}-U$.
+If $H$ is quasi-4-connected and has >6>6 vertices, then $H^{\Delta}$ is 4-connected by Lemma˜12.2.
+The operation $H\mapsto H^{\Delta}$ is canonical: every isomorphism $H_{1}\to H_{2}$ induces an isomorphism $H^{\Delta}_{1}\to H^{\Delta}_{2}$.
+Figure 1.8.
+From $H$ (left) over $H^{s}$ and $H^{s\Delta}$ to $H^{\Delta}$ (right).
+Central graphs: vertices in $U$ are green, neighbours of $U$ are circled red, subdivision vertices are fully red, and the triangles $\Delta(u)$ are red.
+[…]
+  1. 5)
+    Each quasi-4-connected torso $\tau$ with >6>6 vertices we make into a 4-connected graph $\tau^{\Delta}$.
+    (Quasi-4-connected graphs on $\leqslant 6$ vertices are treated as ‘basic’ outcome.)
+  2. 6)
+    Finally, we decompose each 4-connected $\tau^{\Delta}$ with Theorem˜1.
+The operation $\tau\mapsto\tau^{\Delta}$ harmonises well with the other decomposition steps. Indeed, we have followed all steps to prove Figure˜1.8 below.
 Vertex-transitive graphs.
-It is well known that all vertex-transitive finite connected graphs are either 3-connected, cycles, or $K_{2}$’s [infiniteSPQR].
-This fact has then been refined in [Tridecomp] by analysing the 3-connected outcome further; see [Tridecomp, Corollary 2].
-However, [Tridecomp] leaves it open to gain new insights into vertex-transitive 4-connectivity.
-This is where we apply Theorem 1: to obtain a characterisation of vertex-transitive 4-connectivity.
 Recall that a graph $G$ is arc-transitive if for every two ordered pairs $(u_{i},v_{i})$ with $u_{i}v_{i}\in E(G)$ for $i=1,2$ there is an automorphism $\varphi$ of $G$ such that $\varphi(u_{1})=u_{2}$ and $\varphi(v_{1})=v_{2}$.
-The $K_{4}$-expansion of a 4-regular graph $G$ is obtained from $G$ by expanding every vertex to a $K_{4}$, and adding a perfect matching such that two $K_{4}$’s are joined by a matching-edge if and only if their original vertices in $G$ are joined by a $G$-edge.
-For other definitions related to 2, see Section 12.
+Let $H$ be an $r$-vertex graph and let $G$ be an $r$-regular graph.
+An $H$-expansion of $G$ is a graph obtained from $G$ by replacing the vertices $v$ of $G$ with pairwise disjoint copies $H_{v}$ of $H$ and by replacing every edge $e=uv$ of $G$ with an $H_{u}$$H_{v}$ edge $\hat{e}$ so that $\hat{e}_{1}$ and $\hat{e}_{2}$ share no ends for every two $e_{1}\neq e_{2}\in E(G)$.
+Thus, the edges $\hat{e}$ form a perfect matching of the $H$-expansion.
+It was mostly222[CK23, Corollary 2] is slightly weaker than the result we mention here.
+The slight strengthening happens in Proposition 13.51 in this paper.
+shown in [CK23] that a finite connected graph $G$ is vertex-transitive if and only if $G$ is either
+  • $\cdot$
+    quasi-4-connected and vertex-transitive,
+  • $\cdot$
+    the $K_{3}$-expansion of a quasi-4-connected 4-regular arc-transitive graph,
+  • $\cdot$
+    a cycle, $K_{2}$ or $K_{1}$.
+[…]
+A graph $G$ is 2-quasi-5-connected if $G$ is quasi-4-connected and every 4-separation $(A,B)$ of $G$ satisfies $|A\smallsetminus B|\leqslant 2$ or $|B\smallsetminus A|\leqslant 2$.
+For other definitions related to Figure˜1.8, see Section˜13.
 {restatable}
-mainresultmaintransitive A 4-connected finite graph $G$ is vertex-transitive if and only if $G$ either is
+mainresultmaintransitive A quasi-4-connected finite graph $G$ is vertex-transitive if and only if $G$ either is
   1. (i)
-    a quasi-5-connected vertex-transitive graph,
+    a $(\geqslant 4)$-regular quasi-5-connected vertex-transitive graph,
   2. (ii)
-    the $K_{4}$-expansion of a quasi-5-connected arc-transitive graph,
+    a 3-regular 2-quasi-5-connected vertex-transitive graph,
   3. (iii)
-    a cycle of $K_{4}$-bags of length $\geqslant 4$ (Figure 1.7),
+    a $K_{4}$-expansion of a quasi-5-connected 4-regular arc-transitive graph, or a vertex-transitive $C_{4}$-expansion of a quasi-5-connected 4-regular arc-transitive graph,
   4. (iv)
-    a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 4$ (Figure 1.8),
+    a cycle of $K_{4}$-bags of length $\geqslant 4$ (Figure˜1.9),
   5. (v)
-    a cycle of triangle-bags of length $\geqslant 6$ (Figure 1.9),
-or $G$ is one of the following graphs or a $K_{4}$-expansion thereof:
+    a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 4$ (Figure˜1.10),
   6. (vi)
-    a cycle of $K_{2,2}$-bags of length $\geqslant 3$ (Figure 1.10),
+    a cycle of triangle-bags of length $\geqslant 6$ (Figure˜1.12),
   7. (vii)
-    the line-graph of the cube (Figure 1.11).
-2 strengthens a classical theorem of Godsil and Royle from geometric group theory in a special case [Godsil, Theorem 3.4.2].
+    a cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length $\geqslant 6$ (Figure˜1.11),
+  8. (viii)
+    a cycle of $C_{4}$-bags of length $\geqslant 4$ (Figure˜1.13) or $K_{3}\square K_{2}$,
+  9. (ix)
+    a cycle of $K_{2,2}$-bags of length $\geqslant 3$, the $K_{4}$-expansion thereof, or a vertex-transitive $C_{4}$-expansion thereof (Figure˜1.14), or
+  10. (x)
+    the line-graph of the cube, its $K_{4}$-expansion, or one of the three $C_{4}$-expansions of it depicted in Figure˜1.15.
 Figure 1.9.
@@
 Figure 1.11.
-A cycle of triangle-bags
-Figure 1.10. A cycle of $K_{2,2}$-bags and its $K_{4}$-expansion
-Figure 1.11. The line-graph of the cube and its $K_{4}$-expansion
-Applications that require canonicity and explicitness.
-There are applications for partial solutions to 1.1, and in particular for Theorem 1 and [Tridecomp], that absolutely require canonicity and explicitness, often both.
+A cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos
+Figure 1.12.
+Cycles of triangle-bags
+Figure 1.13.
+Cycles of $C_{4}$-bags
+Figure 1.14. A cycle of $K_{2,2}$-bags, its $K_{4}$-expansion, and two vertex-transitive $C_{4}$-expansions
+Figure 1.15. The line-graph of the cube, its $K_{4}$-expansion, and its vertex-transitive $C_{4}$-expansions
+When canonicity and an explicit description matter.
+There are applications for partial solutions to ˜1.1, and in particular for Theorem˜1 and [CK23], that require canonicity and explicitness, often both.
 Examples of these applications include:
   1. $\cdot$
-    Computing excluded minors.
-A major challenge at the interface of computer science and graph-minor theory is to determine the excluded minors for explicit minor-closed graph-classes.
-Partial solutions to 1.1 for $k\leqslant 3$ can be used to determine the excluded minors for planar graphs and series-parallel graphs.
-We have conviction that solutions for larger $k$ imply exact structure theorems for other graph-classes, such as the class of apex-graphs, and that these structure theorems can be used to determine the excluded minors of these classes.
+    Connectivity augmentation.
+Carmesin and Sridharan used the canonicity and explicitness of the tri-separation decomposition [CK23] to construct the first FPT-algorithm for connectivity augmentation from 0 to 4 [CS25\$ˆ+\$].
+That is, they showed that there is an algorithm with runtime $C(\ell)\cdot n^{\mathcal{O}(1)}$ and the following specifications:
+    Input  : a graph $G$ on $n$ vertices, a number $\ell\in\mathbb{N}$ and a set $F\subseteq[V(G)]^{2}\smallsetminus E(G)$ of non-edges.
+Output  : a set $X\subseteq F$ of $\leqslant\ell$ edges such that $G+X$ is 4-connected, if such an $X$ exists, and ‘no’ otherwise.
+We strongly believe that Theorem˜1 makes it possible to extend their algorithm to 5-connectivity.
   2. $\cdot$
-    Connectivity augmentation.
-Carmesin and Sridharan used the canonicity and explicitness of the tri-separation decomposition [Tridecomp] to construct the first FPT-algorithm for connectivity augmentation from 0 to 4 [conny_aug].
-We would be excited to see if Theorem 1 can be used to extend this to 5-connectivity.
+    Finding excluded minors.
+A challenge at the interface of computer science and graph-minor theory is to determine the excluded minors for particular minor-closed graph-classes.
+Indeed, while the graph-minor theorem implies the existence of a cubic-time membership test for each minor-closed graph-class, explicit knowledge of the excluded minors is required to obtain an explicit description of that test [RS95].
+Partial solutions to ˜1.1 for $k\leqslant 3$ can be used to determine the excluded minors for planar graphs and series-parallel graphs.
+We have conviction that solutions for larger $k$ imply exact structure theorems for other graph-classes, such as the class of apex-graphs, and that these structure theorems can be used to determine the excluded minors for these classes and thereby provide explicit cubic-time membership tests.
+  3. $\cdot$
+    Graph isomorphism.
+Canonical tree-decompositions offer angles of attack on finding efficient algorithms for the graph isomorphism problem on particular classes of graphs.
+Recently, an FPT-algorithm was found [LPPS22] that solves the graph-isomorphism problem for $H$-minor-free graphs $G,G^{\prime}$ in time $f(H)\cdot n^{\mathcal{O}(1)}$.
+In that paper, the authors wish for a canonical tree-decomposition into bags that are highly cohesive or of bounded size [LPPS22, 1.3], and they mention long cycles as an illustrative example of an obstruction to the existence of such decompositions.
+Our results show that such canonical tree-decompositions exist for low-order connectivity if one trades ‘bounded size’ for ‘basic’ and an explicit description thereof.
+Hence we wonder if our results allow for even more efficient algorithms that solve the graph isomorphism problem for, say, graphs without quasi-5-connected minors.
+  4. $\cdot$
+    Cops and robber.
+In the cops and robber game of Robertson and Seymour, $k$ cops have a winning strategy for catching a robber in a graph iff that graph has a tree-decomposition of width <k<k [ST93].
+Consider now a randomised variation of this game where the robber is invisible, the cops can visit vertices at most once, and both players use mixed strategies.
+[^3] A mixed strategy is a random distribution over the set of all pure (i.e.
+deterministic) strategies for a player.
+Now the maximum winning probability for $k$ cops depends on the minimum number of leaf-bags attained by any tree-decomposition of width <k<k.
+Our results give such tree-decompositions for $k$-connected graphs when $k\leqslant 4$ [P25\$ˆ+\$].
   5. $\cdot$
@@
   6. $\cdot$
-    Parallel computing. Total-nestedness makes computing the decomposition in parallel a feasible option, as all partial solutions can be merged without conflict; see for example [Sarah].
-  5. $\cdot$
-    Cops and robber.
-In the cops and robber game of Robertson and Seymour, $k$ cops have a winning strategy for catching a robber in a graph iff that graph has a tree-decomposition of width $<k$ [GMcopsRobber].
-When the robber is invisible, however, the probability of the cops catching the robber depends on the minimum number of leaf-bags attained by any tree-decomposition of width $<k$.
-While with other approaches to constructing small-width tree-decompositions, it is not even clear how to verify that they minimize the number of leaf-bags, our theorem gives such tree-decompositions right away in the case of $k$ cops searching $k$-connected graphs for $k\leqslant 4$ [CopsInvisibleRobber].
-  6. $\cdot$
-    Groups.
-The characterisation of vertex-transitivity for 4-connectivity (2) crucially relies on canonicity and total-nestedness.
-In a different direction, local separators have recently been used to prove a low-order Stallings-type theorem for finite nilpotent groups [StallingsNilpotent].
-The proof heavily relies on the Tutte-type decomposition for local 2-separators [Local2sep] and exploits the combination of canonicity and total-nestedness.
-A next step could be using Theorem 1 and 2, or the tri-separation decomposition [Tridecomp].
-Comparison to the tri-separation decomposition.
-Carmesin and Kurkofka introduced the first Tutte-type canonical decomposition for 3-connected graphs [Tridecomp].
-They used tri-separations instead of strict tri-separations to obtain their decomposition.
-The main difference between tri-separations on the one hand, and strict tri-separations and tetra-separations on the other hand, is how their degree-conditions handle neighbours in the separator: tri-separations count them, but strict tri-separations and tetra-separations ignore them.
-This seemingly subtle difference has a surprisingly big impact: the verbatim extension of the tri-separation approach to 4-connectivity fails 2.
-For 3-connected graphs, the strict tri-separation decomposition can be viewed as a refinement of the tri-separation decomposition, but the details are a bit technical; see Appendix B.
+    Finite groups.
+The proof of Figure˜1.8 crucially relies on canonicity and total-nestedness.
+In a different direction, Stallings’ theorem applies tree-decomposition methods to Cayley graphs to detect product structure in groups.
+However, Stallings’ approach only works for infinite groups.
+Local separators have recently been used to prove a low-order Stallings-type theorem for finite nilpotent groups [CKKT24].
+The proof heavily relies on the Tutte-type decomposition for local 2-separators [C22] and exploits the combination of canonicity and total-nestedness.
+The next step towards fully generalising Stallings theorem to finite groups will be to use versions for local separators of Theorem˜1 and of either ˜2 or [CK23].
 1.1. Overview of the proof
@@
 Let $N$ denote its set of totally-nested tetra-separations, and let $\mathcal{T}=(T,\mathcal{V})$ denote the mixed-tree-decomposition of $G$ that is uniquely determined by $N$.
-Let $\tau$ denote the torso of $\mathcal{T}$ at an arbitrary node $t$ of $T$.
+Let $\tau$ denote the torso of $\mathcal{T}$ at an arbitrary node $t$ of $T$ (see Section˜2.2 for the definition of torso in the context of mixed-tree-decompositions).
 It will be straightforward to check that $\tau$ is 4-connected or a $K_{4}$.
-In order to show that $\tau$ satisfies one of the outcomes of Theorem 1, we will study how $\tau$ interacts with the tetra-separations of $G$.
+In order to show that $\tau$ satisfies one of the outcomes of Theorem˜1, we will study how $\tau$ interacts with the tetra-separations of $G$.
 Indeed, the outcome will be determined by this.
@@
 (1) If $\tau$ is not interlaced by any tetra-separation of $G$, then $\tau$ is quasi-5-connected.
-(See 10.5.) A technical lemma from [Tridecomp] does most of the work of proving 1.
-By 1, we may assume from now on that some tetra-separation $(A,B)$ of $G$ interlaces $\tau$.
+(See ˜10.5.) A technical lemma from [CK23] does most of the work of proving ˜1.
+By ˜1, we may assume from now on that some tetra-separation $(A,B)$ of $G$ interlaces $\tau$.
 Then $(A,B)$ is not in $N$, so some tetra-separation $(C,D)$ crosses $(A,B)$.
@@
 (2) $|Z|\in\{0,2,4\}.$
-(See the (Crossing Lemma).
+(See (Crossing Lemma).
 (4.1).) So there will be exactly three outcomes.
@@
 This outcome is the hardest to prove, and we use all the new methods introduced in this paper in the proof of this outcome.
-It is also different from all situations encountered in [Tridecomp].
+It is also different from all situations encountered in [CK23].
 Assume now that $|Z|=0$.
-As our first step, we will use $(A,B)$ and $(C,D)$ to cut $G$ into four 2-connected pieces, arranged in the shape of a 4-cycle, and apply the Tutte-decomposition to each piece; see Figure 1.12.
+As our first step, we will use $(A,B)$ and $(C,D)$ to cut $G$ into four 2-connected pieces, arranged in the shape of a 4-cycle, and apply the Tutte-decomposition to each piece; see Figure˜1.16.
 Since $G$ is 4-connected, the Tutte-decompositions will be path-decompositions.
@@
 (3) $(U_{o},W_{o})$ is a totally-nested tetra-separation of $G$ if and only if $H_{o}$ is good.
-(See 9.41). In order to show the hard backward implication of 3, we will introduce a
+(See ˜9.41). In order to show the hard backward implication of ˜3, we will introduce a
 (4) characterisation of total-nestedness in terms of connectivity-properties of separators.
-(See 6.16.) Examples of connectivity-properties of a tetra-separator $S(A,B)$ that are necessary for $(A,B)$ to be totally-nested include, roughly speaking, that for every bipartition $(\pi_{1},\pi_{2})$ of $S(A,B)$ into two classes $\pi_{i}$ of size 2 there must be
+(See ˜6.16.) Examples of connectivity-properties of a tetra-separator $S(A,B)$ that are necessary for $(A,B)$ to be totally-nested include, roughly speaking, that for every bipartition $(\pi_{1},\pi_{2})$ of $S(A,B)$ into two classes $\pi_{i}$ of size 2 there must be
   • $\cdot$
@@
     3 independent paths between the two elements of $\pi_{1}$ in $G-\pi_{2}$ on the other hand.
-A major difference to tri-separations is that the connectivity-properties required of the separators are much more complex, as tetra-separations can cross in significantly more complicated ways than tri-separations; see Figure 1.13 for an example.
+A major difference to tri-separations is that the connectivity-properties required of the separators are much more complex, as tetra-separations can cross in significantly more complicated ways than tri-separations; see Figure˜1.17 for an example.
 Figure 1.17. Three tetra-separations cross in a way that tri-separations cannot.
-With 4 at hand, the problem of proving the hard backward implication of 3 becomes a problem about finding many paths in many configurations, roughly speaking.
-To this end, we develop a number of lemmata for finding such paths systematically, and then combine them to prove 3.
-Finding the paths takes a lot of work, especially when $O$ is short; see Figure 1.14 for example.
+With ˜4 at hand, the problem of proving the hard backward implication of ˜3 becomes a problem about finding many paths in many configurations, roughly speaking.
+To this end, we develop a number of lemmata for finding such paths systematically, and then combine them to prove ˜3.
+Finding the paths takes a lot of work, especially when $O$ is short; see Figure˜1.18 for example.
 Figure 1.18.
@@
 The sets $\pi_{1}$ and $\pi_{2}$ consist of the blue and the green vertices, respectively.
-As a small part of proving 3 via 4, we have to find 5 independent $\pi_{1}$$\pi_{2}$ paths (red).
+As a small part of proving ˜3 via ˜4, we have to find 5 independent $\pi_{1}$$\pi_{2}$ paths (red).
 We find two such paths in $G_{o}$ using Menger as $G_{o}$ is good.
@@
 The example shows that the construction must be highly efficient to avoid collision of the paths when the decomposition-cycle $O$ is short.
-By 3, each tetra-separation $(U_{o},W_{o})$ for a good torso $H_{o}$ corresponds to a directed edge $e_{o}$ of $T$. The second key result about $\mathcal{O}$ will be that
+By ˜3, each tetra-separation $(U_{o},W_{o})$ for a good torso $H_{o}$ corresponds to a directed edge $e_{o}$ of $T$. The second key result about $\mathcal{O}$ will be that
 (5) The edges $e_{o}$ are precisely the directed edges of the form $(s,t)$ with $st\in E(T)$,
@@
 Phrased differently, the tetra-separations $(U_{o},W_{o})$ from good torsos $H_{o}$ are precisely the elements of $N$ that ‘carve out’ $\tau$.
-(See Lemma 9.49.)
-Once we have proved 5, we are almost done.
+(See Lemma˜9.49.)
+Once we have proved ˜5, we are almost done.
 Indeed, it now follows that the torso $\tau$ can be obtained from $G$ by essentially replacing the good torsos of $\mathcal{O}$ with $K_{4}$’s and contracting a few edges.
 Then it will be fairly immediate to see that $\tau$ is a cycle of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices.
-(See 9.50.)
-Proving 5, however, is a little tricky.
-We will assume for a contradicting that 5 fails, so some totally-nested tetra-separation $(E,F)$ will ‘cut through’ $\mathcal{O}$.
-Since $(E,F)$ is nested with both $(A,B)$ and $(C,D)$, we have $E\subseteq A\cap C$, say; see Figure 1.15.
+(See ˜9.50.)
+Proving ˜5, however, is a little tricky.
+We will assume for a contradicting that ˜5 fails, so some totally-nested tetra-separation $(E,F)$ will ‘cut through’ $\mathcal{O}$.
+Since $(E,F)$ is nested with both $(A,B)$ and $(C,D)$, we have $E\subseteq A\cap C$, say; see Figure˜1.19.
 We then start adjusting both $(A,B)$ and $(C,D)$: by moving bags of $\mathcal{O}$ from $A\smallsetminus B$ to $B\smallsetminus A$ and from $C\smallsetminus D$ to $D\smallsetminus C$, one bag at a time, while using the total-nestedness of $(E,F)$ to maintain the inclusion $E\subseteq A\cap C$.
@@
 This replaces the top half of yellow with red.
-Outlook. We would be excited to see an extension of Theorem 1 to higher connectivity. However, the verbatim extension of tetra-separations fails 2 for $k=5$.
-Open Problem 1.4.
-Extend Theorem 1 to $k$-connectivity for $k\geqslant 5$.
-Both the strict tri-separation decomposition and the tri-separation decomposition may have quasi-4-connected torsos.
-To decompose these torsos further, we can apply the tetra-separation decomposition, but only to those quasi-4-connected torsos that are even 4-connected.
-Open Problem 1.5.
-Extend Theorem 1 from 4-connectivity to quasi-4-connectivity.
-Simply using the tetra-separation decomposition fails: see Figure 1.16 for an illustrative example that essentially generalises 2 to quasi-4-connectivity.
-Figure 1.16. A quasi-4-connected version of 2: all tetra-separations are crossed, yet the graph is far from $5$-connected and each $K_{10}$ can be replaced with an arbitrary 5-connected graph
-The tri-separation decomposition has been used to prove a characterisation of vertex-transitivity for 3-connectivity [Tridecomp, Corollary 2].
-One of the outcomes is that the graph is quasi-4-connected or a triangle-expansion of a quasi-4-connected graph.
-[^2] The actual outcome is slightly more specific.
-Open Problem 1.6.
-Extend 2 from 4-connectivity to quasi-4-connectivity.
-We believe that a solution to 1.5 can be used to approach 1.6.
-Open Problem 1.7.
-Extend the FPT-algorithm of [conny_aug] for connectivity augmentation from 0 to 4 to 5-connectivity.
-Since the tri-separation decomposition [Tridecomp] has been heavily exploited in [conny_aug], we hope that Theorem 1 can be used to approach 1.7.
 […]
 […]
-For directed graphs, the case $k=1$ of 1.1 has recently been solved almost completely by Bowler, Gut, Hatzel, Kawarabayashi, Muzi and Reich [Flo].
+For directed graphs, the case $k=1$ of ˜1.1 has recently been solved almost completely by Bowler, Gut, Hatzel, Kawarabayashi, Muzi and Reich [BGHKMR23].
 […]
-Organisation of the paper.
-Section 2 introduces or recalls all essential terminology.
-Section 3 provides stronger but more technical versions of Theorem 1 and 2, and proves all while assuming a number of key lemmata that remain to be proved later in the paper.
-Section 4 provides a characterisation of how tetra-separations cross.
-Section 5 introduces a systematic way for obtaining tetra-separations from mixed-4-separations.
-Section 6 gives a characterisation of total-nestedness for tetra-separations in terms of connectivity-properties of their separators.
-The next sections investigate the different outcomes of Theorem 1: Section 7 deals with the $K_{4,m}$; Section 8 deals with the double-wheel; Section 9 deals with the cycle-of-graphs; and Section 10 deals with quasi-5-connectivity.
-This will conclude the proof of Theorem 1.
-Section 11 introduces an ‘angry theorem’ for 4-connectivity: a characterisation of the 4-connected graphs all whose tetra-separations are crossed (hence the name).
-Section 12 uses Theorem 1 to obtain a characterisation of the vertex-transitive 4-connected finite graphs.
-Appendix A proves that the circular saws from 1 really exhibit the properties we used.
-Appendix B compares the strict tri-separation-decomposition (2) with the tri-separation decomposition [Tridecomp].
+[…]
 2. Essential terminology
-This section introduces all terminology used by Theorem 1. For graph-theoretic terminology we follow [bibel].
+This section introduces all terminology used by Theorem˜1. For graph-theoretic terminology we follow [D24a].
 2.1. Tetra-separations
-A mixed-separation+ of a graph $G$ is a pair $(A,B)$ such that $A\cup B=V(G)$.
-The separator $)𝑆𝐴𝐵{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{S(A,B)}}$ of $(A,B)$ is the disjoint union of the vertex set $A\cap B$ and the edge-set $E(A\smallsetminus B,B\smallsetminus A)$.
-The order of $(A,B)$ is the size of $S(A,B)$.
-A mixed-separation+ of order $k$ is also called mixed-$k$-separation+.
-A mixed-separation+ is proper if both $A\smallsetminus B$ and $B\smallsetminus A$ are nonempty.
-In this case, we refer to $(A,B)$ as a mixed-separation without the plus-sign.
-A mixed-separation of order $k$ is also called mixed-$k$-separation.
-A mixed-separation with $k$ vertices in its separator but no edges is called a separation, or $k$-separation.
+[…]
 Definition 2.1.
 A tetra-separation of a graph $G$ is a mixed-4-separation $(A,B)$ of $G$ that satisfies the following two conditions:
-  1. degree-condition:
-    every vertex in $A\cap B$ has $\geqslant 2$ neighbours in both $A\smallsetminus B$ and $B\smallsetminus A$;
-  2. matching-condition:
-    the edges in $S(A,B)$ form a matching.
+degree-condition: every vertex in $A\cap B$ has $\geqslant 2$ neighbours in both $A\smallsetminus B$ and $B\smallsetminus A$;
+matching-condition: the edges in $S(A,B)$ form a matching.
 Definition 2.2.
 A strict tri-separation of a graph $G$ is a mixed-3-separation of $G$ that satisfies both the degree-condition and the matching-condition.
-Recall that a graph $G$ is quasi-$k$-connected if it is $(k-1)$-connected and every $(k-1)$-separation $(A,B)$ of $G$ satisfies $|A\smallsetminus B|=1$ or $|B\smallsetminus A|=1$.
+Recall that a graph $G$ is quasi-$k$-connected if it is (k1)(k-1)-connected and every (k1)(k-1)-separation $(A,B)$ of $G$ satisfies $|A\smallsetminus B|=1$ or $|B\smallsetminus A|=1$.
 Lemma 2.3.
@@
     If $G$ is quasi-5-connected and $|G|\geqslant 8$, then $G$ has no tetra-separation.
-We will prove Lemma 2.3 in Section 5.2.
+We will prove Lemma˜2.3 in Section˜5.2.
 Example 2.4.
-The lower-bound $|G|\geqslant 8$ in Lemma 2.3 (ii) is tight: the graph depicted in Figure 2.1 is quasi-5-connected, has $7$ vertices, but has a tetra-separation.
+The lower-bound $|G|\geqslant 8$ in Lemma˜2.3 (ii) is tight: the graph depicted in Figure˜2.1 is quasi-5-connected, has $7$ vertices, but has a tetra-separation.
 Figure 2.1. A small quasi-5-connected graph with a tetra-separation (red)
@@
 2.2. Stars, mixed-tree-decompositions and torsos
-The following definition is supported by Figure 2.2.
+The following definition is supported by Figure˜2.2.
 Let $G$ be a graph.
@@
 The adhesion-set of an edge $t_{1}t_{2}$ of $T$ is the separator of $\alpha_{\mathcal{T}}(t_{1},t_{2})$, which is well-defined.
-Figure 2.2. A mixed-tree-decomposition. The decomposition-tree is blue. Bags are grey. Edges and vertices in adhesion-sets are red.
+Figure $2.$ A mixed-tree-decomposition. The decomposition-tree is blue. Bags are grey. Edges and vertices in adhesion-sets are red.
 A star of mixed-separations+, or star for short, is a set $\sigma=\{\,(A_{i},B_{i}):i\in I\,\}$ of mixed-separations+ $(A_{i},B_{i})$ such that $(A_{i},B_{i})\leqslant(B_{j},A_{j})$ for all distinct $i,j\in I$.
@@
 Conversely, if $S$ is a nested symmetric set of mixed-separations of a finite graph $G$, then there is a unique (up to isomorphisms for decomposition-trees) mixed-tree-decomposition $\mathcal{T}(S)$ of $G$ such that $\alpha_{\mathcal{T}(S)}$ is a bijection $\vec{E}(T)\to S$.
-The mixed-tree-decomposition $\mathcal{T}(S)=(T,\mathcal{V})$ has an explicit description (see [Tridecomp, §3.1]):
+The mixed-tree-decomposition $\mathcal{T}(S)=(T,\mathcal{V})$ has an explicit description (see [CK23, §3.1]):
   1. (i)
@@
 In this sense, $S$ and $\mathcal{T}(S)$ contain the same information, and we can switch at will between the two perspectives that they offer.
-When $G$ is infinite and $S$ is the set of totally-nested tetra-separations of $G$, then $\mathcal{T}(S)$ is a mixed-tree-decomposition of $G$ by standard arguments as in [Tridecomp, Lemma 3.3.10].
+When $G$ is infinite and $S$ is the set of totally-nested tetra-separations of $G$, then $\mathcal{T}(S)$ is a mixed-tree-decomposition of $G$ by standard arguments as in [CK23, Lemma 3.3.10].
 A mixed-separation $(C,D)$ interlaces a star $\sigma$ of mixed-separations if for every $(A,B)\in\sigma$ we have either $(A,B)<(C,D)$ or $(A,B)<(D,C)$.
-The notion of interlacing yields a characterisation of splitting stars: a star $\sigma$ included in a nested set $S$ of mixed-separations of a graph $G$ is a splitting star of $S$ if and only if no element of $S$ interlaces $\sigma$ [Tridecomp, Lemma 2.2.3].
-Let $\mathcal{T}=(T,\mathcal{V})$ be a mixed-tree-decomposition of a graph $G$.
-For a mixed-separation+ $(A,B)$, we let the vertex-set $)𝜆𝐴𝐵{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{\lambda(A,B)}}$ consist of all vertices in $S(A,B)$ and the endvertex in $A\smallsetminus B$ for each edge in $S(A,B)$.
-The expanded-torso of $t$, or of $V_{t}$, is obtained from the subgraph of $G$ induced by
+The notion of interlacing yields a characterisation of splitting stars: a star $\sigma$ included in a nested set $S$ of mixed-separations of a graph $G$ is a splitting star of $S$ if and only if no element of $S$ interlaces $\sigma$ [CK23, Lemma 2.2.3].
+[…]
 $V_{t}\cup\bigcup_{st\in E(T)}\lambda(\alpha_{\mathcal{T}}(s,t))$
 by making each of the vertex-sets $\lambda(\alpha_{\mathcal{T}}(s,t))$ into a clique.
-From this expanded-torso we obtain the compressed-torso, or torso for short, by contracting all edges contained in adhesion-sets of the edges $st\in E(T)$; see Figure 2.3.
+From this expanded-torso we obtain the compressed-torso, or torso for short, by contracting all edges contained in adhesion-sets of the edges $st\in E(T)$; see Figure˜2.3.
 More generally, let $\sigma=\{\,(A_{i},B_{i}):i\in I\,\}$ be a star of mixed-separations+.
@@
 The two notions of torsos are compatible: the torso of a node $t$ of $T$ equals the torso of the star $\sigma_{t}$.
-Figure 2.3. The torsos of the mixed-tree-decomposition depicted in Figure 2.2.
+Figure 2.3. The torsos of the mixed-tree-decomposition depicted in Figure˜$2.$
 2.3. Graph-decompositions
@@
 Let $X$ be a graph.
-We call $X$ a generalised wheel with centre $v$ if $X$ is 3-connected with $v\in V(X)$ such that $X-v$ has a cycle-decomposition with all adhesion-sets of size 1 and into $K_{2}$’s and triangles.
-We call $X$ a generalised double-wheel with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that $X-u-v$ has a cycle-decomposition with all adhesion-sets of size 1 and into $K_{2}$’s and triangles.
+We call $X$ a generalised wheel with centre $v$ if $X$ is 3-connected with $v\in V(X)$ such that XvX-v has a cycle-decomposition with all adhesion-sets of size 1 and into $K_{2}$’s and triangles.
+We call $X$ a generalised double-wheel with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that XuvX-u-v has a cycle-decomposition with all adhesion-sets of size 1 and into $K_{2}$’s and triangles.
 We call $X$ a cycle of $X_{1}$-torsos, …, $X_{k}$-torsos where the $X_{i}$ are graphs, if $X$ is 4-connected and has a cycle-decomposition with all adhesion-sets of size two such that every torso is isomorphic to some $X_{i}$.
@@
 A sprinkled $K_{4,m}$ is obtained from a $K_{4,m}$ by adding any number of edges to the left side.
-3. Proofs of Theorem 1 and 2 and strengthenings thereof
-Instead of proving Theorem 1 and 2 directly, we will first prove stronger but slightly more technical versions of them, Theorem 3.1 and Corollary 3.2 below, and then immediately obtain Theorem 1 and 2 from their strengthenings.
-The proof of Theorem 3.1 uses a number of lemmas that have not yet been introduced.
+3. Proofs of Theorem˜1 and ˜2 and strengthenings thereof
+Instead of proving Theorem˜1 and ˜2 directly, we will first prove stronger but slightly more technical versions of them, Theorem˜3.1 and Corollary˜3.2 below, and then immediately obtain Theorem˜1 and ˜2 from their strengthenings.
+The proof of Theorem˜3.1 uses a number of lemmas that have not yet been introduced.
 Stating and proving these lemmas will be the main objectives of later sections.
-Theorem 3.1 (Technical strengthening of Theorem 1).
+Theorem 3.1 (Technical strengthening of Theorem˜1).
 Let $G$ be a 4-connected graph. Let $N$ denote the set of all totally-nested tetra-separations of $G$. Let $\sigma$ be a splitting star of $N$ with torso $\tau$.
@@
 Proof.
-If no tetra-separation of $G$ interlaces $\sigma$, then $\tau$ is quasi-5-connected by 10.5.
+If no tetra-separation of $G$ interlaces $\sigma$, then $\tau$ is quasi-5-connected by ˜10.5.
 So we may assume that some tetra-separation $(A,B)$ of $G$ interlaces $\sigma$.
@@
 (4.1), we have $|Z|\in\{0,2,4\}$.
-(ii). Assume $|Z|=0$. Then $\tau$ is a cycle of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices by 9.50.
-(iii). Assume $|Z|=2$. Then $\tau$ is a generalised double-wheel with centre $Z$ by 8.37.
-(iv).
+˜(ii). Assume $|Z|=0$. Then $\tau$ is a cycle of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices by ˜9.50.
+˜(iii). Assume $|Z|=2$. Then $\tau$ is a generalised double-wheel with centre $Z$ by ˜8.37.
+˜(iv).
 Assume $|Z|=4$.
-If $\sigma$ is non-empty, then $\tau$ is a thickened $K_{4,m}$ with $Z$ as left side and $m\geqslant 0$ by 7.7.
-Otherwise the entire graph $G$ is a sprinkled $K_{4,m}$ with $Z$ as left side and $m\geqslant 4$, also by 7.7.
-∎
-Proof of Theorem 1.
-Immediate from Theorem 3.1. ∎
-Corollary 3.2 (Technical strengthening of 2).
+If $\sigma$ is non-empty, then $\tau$ is a thickened $K_{4,m}$ with $Z$ as left side and $m\geqslant 0$ by ˜$7.$ Otherwise the entire graph $G$ is a sprinkled $K_{4,m}$ with $Z$ as left side and $m\geqslant 4$, also by ˜$7.$ ∎
+Proof of Theorem˜1.
+Immediate from Theorem˜3.1. ∎
+Corollary 3.2 (Technical strengthening of ˜2).
 Let $G$ be a 3-connected graph.
@@
 Thus $\tau=\hat{\tau}-\alpha$.
-(i).
+˜(i).
 Assume that $\sigma$ is not interlaced by any strict tri-separation.
-Then $\hat{\sigma}$ is not interlaced by any tetra-separation, so $\hat{\tau}$ is quasi-5-connected by Theorem 3.1 (i).
+Then $\hat{\sigma}$ is not interlaced by any tetra-separation, so $\hat{\tau}$ is quasi-5-connected by Theorem˜3.1 ˜(i).
 If $\tau$ is not quasi-4-connected, this is witnessed by some separation $(X,Y)$ of $\tau$, and then $(X\cup\{\alpha\},Y\cup\{\alpha\})$ witnesses that $\hat{\tau}$ is not quasi-5-connected.
@@
 Then $\varphi(A,B)$ interlaces $\sigma$ and is crossed by $\varphi(C,D)$ with vertex-centre $\hat{Z}:=Z\cup\{\alpha\}$.
-(ii).
+˜(ii).
 Assume that $Z$ consists of exactly one vertex $v$.
-Then $\hat{\tau}$ is a generalised double-wheel with centre $\{v,\alpha\}$ by Theorem 3.1 (iii).
+Then $\hat{\tau}$ is a generalised double-wheel with centre $\{v,\alpha\}$ by Theorem˜3.1 ˜(iii).
 Hence $\hat{\tau}$ is a generalised wheel with centre $v$.
-(iii).
-The vertex-centre $Z$ cannot have sizes 0 or 2, since then $\hat{Z}$ would have size 1 or 3, which is excluded in Theorem 3.1.
+˜(iii).
+The vertex-centre $Z$ cannot have sizes 0 or 2, since then $\hat{Z}$ would have size 1 or 3, which is excluded in Theorem˜3.1.
 Thus $|Z|=3$, so $|\hat{Z}|=4$.
-Since $\alpha\in\hat{Z}$, both outcomes of Theorem 3.1 (iv) translate to the two outcomes of (iii).
+Since $\alpha\in\hat{Z}$, both outcomes of Theorem˜3.1 ˜(iv) translate to the two outcomes of ˜(iii).
 ∎
-Proof of 2.
-Immediate from Corollary 3.2. ∎
+Proof of ˜2.
+Immediate from Corollary˜3.2. ∎
 Example 3.3.
-The outcome in Theorem 3.1 is not in general unique.
-For example, consider the graph $G$ depicted in Figure 3.1 on the left.
+The outcome in Theorem˜3.1 is not in general unique.
+For example, consider the graph $G$ depicted in Figure˜3.1 on the left.
 Every tetra-separation of $G$ is crossed, hence $\sigma=\emptyset$ is a splitting star of the set $N(G)$ of all totally-nested tetra-separations of $G$.
-On the left of Figure 3.1, we see two tetra-separations that interlace $\sigma$ and cross with empty vertex-centre.
-From their perspective, $G$ is a cycle of $K_{4}$-torsos and $4$-wheel torsos, see Theorem 3.1 (ii).
-On the right of Figure 3.1, we see two tetra-separations that interlace $\sigma$ and cross with a vertex-centre of size 2.
-From their perspective, $G$ is a generalised wheel, see Theorem 3.1 (iii).
-For another example, consider the graph $G$ depicted in Figure 3.2 on the left.
+On the left of Figure˜3.1, we see two tetra-separations that interlace $\sigma$ and cross with empty vertex-centre.
+From their perspective, $G$ is a cycle of $K_{4}$-torsos and $4$-wheel torsos, see Theorem˜3.1 ˜(ii).
+On the right of Figure˜3.1, we see two tetra-separations that interlace $\sigma$ and cross with a vertex-centre of size 2.
+From their perspective, $G$ is a generalised wheel, see Theorem˜3.1 ˜(iii).
+For another example, consider the graph $G$ depicted in Figure˜3.2 on the left.
 Again, $\sigma=\emptyset$ is a splitting star of $N(G)$.
-On the left of Figure 3.1, we see two tetra-separations that interlace $\sigma$ and cross with empty vertex-centre.
-From their perspective, $G$ is a cycle of $K_{4}$-torsos, see Theorem 3.1 (ii).
-On the right of Figure 3.1, we see two tetra-separations that interlace $\sigma$ and cross with a vertex-centre of size 4.
-From their perspective, $G$ is a $K_{4,4}$, see Theorem 3.1 (iv).
+On the left of Figure˜3.1, we see two tetra-separations that interlace $\sigma$ and cross with empty vertex-centre.
+From their perspective, $G$ is a cycle of $K_{4}$-torsos, see Theorem˜3.1 ˜(ii).
+On the right of Figure˜3.1, we see two tetra-separations that interlace $\sigma$ and cross with a vertex-centre of size 4.
+From their perspective, $G$ is a $K_{4,4}$, see Theorem˜3.1 ˜(iv).
 Figure 3.1.
@@
 In this section, we characterise the ways in which tetra-separations can cross. The main result here is the (Crossing Lemma). (4.1).
-We require the following terminology for corner diagrams from [Tridecomp, §1.3], which is supported by Figure 4.1.
+We require the following terminology for corner diagrams from [CK23, §1.3], which is supported by Figure˜4.1.
 Let $(A,B)$ and $(C,D)$ be mixed-separations+ of a graph $G$.
@@
 The left blue edge dangles from the $A$-link through the $C$-link.
-The following definition is supported by Figure 4.2.
+The following definition is supported by Figure˜4.2.
 The corner for $AC$ is potter if it is empty and its adjacent links have size two and there exists a path $u_{1}u_{2}u_{3}u_{4}$ in $G$ such that $u_{1}u_{2}$ dangles from the vertex $u_{2}$ in the $C$-link through the $A$-link, and $u_{3}u_{4}$ dangles from the vertex $u_{3}$ in the $A$-link through the $C$-link.
@@
 Lemma 4.3.
-[Tridecomp, Lemma 1.3.4] For two mixed-separations $(A,B)$ and $(C,D)$ of $G$, we have
+[CK23, Lemma 1.3.4] For two mixed-separations $(A,B)$ and $(C,D)$ of $G$, we have
 $|L(A,C)|+|L(B,D)|\leqslant|S(A,B)|+|S(C,D)|\leavevmode\nobreak\ .$
@@
 If $|L|\geqslant 2$ then the vertices in $L$ have at most one neighbour in $A\smallsetminus B$ (and $B\smallsetminus A$), contradicting the degree-condition for $(A,B)$.
-Otherwise, if $|L|=1$ then $(C,D)$ is not a tetra-separation by Lemma 4.2, a contradiction.
+Otherwise, if $|L|=1$ then $(C,D)$ is not a tetra-separation by Lemma˜4.2, a contradiction.
 ∎
-The following lemma may be regarded as an analogue of [Tridecomp, Lemma 1.3.5].
+The following lemma may be regarded as an analogue of [CK23, Lemma 1.3.5].
 Lemma 4.5.
@@
   2. (ii)
-    two adjacent links have size $i$, the other two links have size $j$, and the centre has size $k-i-j$, where $1\leqslant i\leqslant\lfloor(k-1)/2\rfloor$ and $i<j\leqslant k-i$.
+    two adjacent links have size $i$, the other two links have size $j$, and the centre has size kijk-i-j, where $1\leqslant i\leqslant\lfloor(k-1)/2\rfloor$ and $i<j\leqslant k-i$.
 Proof.
-By Lemma 4.3, two opposite corner-separators have size exactly $k$, say the corner-separators for $AC$ and $BD$.
+By Lemma˜4.3, two opposite corner-separators have size exactly $k$, say the corner-separators for $AC$ and $BD$.
 Let $a,b,c,d$ denote the sizes of the links for $A,B,C,D$, respectively.
 Let $x$ denote the size of the centre.
-By Lemma 4.3, every diagonal edge has its endvertices in the corners for $AC$ and $BD$; and there are no jumping edges.
+By Lemma˜4.3, every diagonal edge has its endvertices in the corners for $AC$ and $BD$; and there are no jumping edges.
 We get $a+c+x=k$ and $b+d+x=k$ and $a+b+x=k$ and $c+d+x=k$.
@@
 Note that $i\leqslant\lfloor(k-1)/2\rfloor$.
-With $x=k-a-c=k-i-j$, we get outcome (ii).
+With x=kac=kijx=k-a-c=k-i-j, we get outcome (ii).
 ∎
@@
 Suppose that not all links have size $\ell$, for some $\ell\in\{0,1,2\}$.
-By Lemma 4.4 and Lemma 4.5, we get that two adjacent links have size one, the other two links have size at least two and the centre has size at most one.
+By Lemma˜4.4 and Lemma˜4.5, we get that two adjacent links have size one, the other two links have size at least two and the centre has size at most one.
 Without loss of generality, the corner $K$ of $AC$ is the corner whose adjacent links have size one.
@@
 But then $u$ has at most one neighbour in $A\smallsetminus B$, contradicting the degree-condition for $(A,B)$.
-Hence all links have the same size $\ell$ for some $\ell\in\{0,1,2\}$, and then the centre has size $4-2\ell$ by Lemma 4.4 and Lemma 4.5.
+Hence all links have the same size $\ell$ for some $\ell\in\{0,1,2\}$, and then the centre has size $4-2\ell$ by Lemma˜4.4 and Lemma˜4.5.
 ∎
@@
 Proof.
-(i) By Lemma 4.6, all links have the same size $\ell$ and the centre has size $4-2\ell$. By Lemma 4.3, there do not exist jumping edges nor diagonal edges.
+(i) By Lemma˜4.6, all links have the same size $\ell$ and the centre has size $4-2\ell$. By Lemma˜4.3, there do not exist jumping edges nor diagonal edges.
 (ii) Without loss of generality, $X=A$ and $Y=C$.
@@
 Assume for contradiction that $K\neq\emptyset$.
-By Lemma 4.6 and 4-connectivity, $|L(A,C)|=4$.
-Then $L(A,C)-e$ is a mixed-3-separator separating the vertices in the corner for $AC$ from $u_{1}$.
-This contradicts 4-connectivity by Lemma 6.1.
+By Lemma˜4.6 and 4-connectivity, $|L(A,C)|=4$.
+Then L(A,C)eL(A,C)-e is a mixed-3-separator separating the vertices in the corner for $AC$ from $u_{1}$.
+This contradicts 4-connectivity by Lemma˜6.1.
 By the degree-condition of $(A,B)$ and since $K$ is empty, the vertex $u_{2}$ in the $C$-link has a neighbour $u_{3}$ in the $A$-link.
@@
 By the degree-condition of $(A,B)$ and since $K$ is empty, the vertex $u_{4}$ has a neighbour in $A\cap D$ besides $u_{3}$, and therefore the $A$-link contains a third vertex or edge besides $e$ and $u_{3}$.
-This contradicts Lemma 4.6.
+This contradicts Lemma˜4.6.
 ∎
 Proof of the (Crossing Lemma). (4.1).
-We combine Lemma 4.6 and Lemma 4.8. ∎
-5. Shifts: a tool for finding tetra-separations
-This section provides a ‘shifting’ method for obtaining tetra-separations from genuine 4-separations, and from mixed-4-separations more generally.
-There are ‘small’ mixed-4-separations that fail to yield tetra-separations via the shifting method.
-We characterise these in 5.10.
-5.1. Left-shifts and right-shifts
-Definition 5.1 (Left-shift, right-shift).
+We combine Lemma˜4.6 and Lemma˜4.8. ∎
+5. Reductions: a tool for finding tetra-separations
+This section provides a ‘reduction’ method for obtaining tetra-separations from genuine 4-separations, and from mixed-4-separations more generally.
+There are ‘small’ mixed-4-separations that fail to yield tetra-separations via the reduction method.
+We characterise these in ˜5.10.
+5.1. Left-reductions and right-reductions
+Definition 5.1 (Left-reduction, right-reduction).
 Let $(A,B)$ be a mixed-separation+ in a graph.
 Let $L$ be the set of vertices in $A\cap B$ with at most one neighbour in $A\smallsetminus B$, and let $R$ be the set of vertices in $A\cap B$ with at most one neighbour in $B\smallsetminus A$.
-The left-shift of $(A,B)$ is $(A\smallsetminus L,B)$.
-The right-shift of $(A,B)$ is $(A,B\smallsetminus R)$.
-In the following, we prove lemmas for the left-shift. All results have analogues for right-shifts by symmetry.
+The left-reduction of $(A,B)$ is $(A\smallsetminus L,B)$.
+The right-reduction of $(A,B)$ is $(A,B\smallsetminus R)$.
+In the following, we prove lemmas for the left-reduction. All results have analogues for right-reductions by symmetry.
 Observation 5.2.
-The left-shift $(A^{\prime},B^{\prime})$ of a mixed-separation+ $(A,B)$ satisfies $A^{\prime}\smallsetminus B^{\prime}=A\smallsetminus B$ and $B^{\prime}\smallsetminus A^{\prime}\supseteq B\smallsetminus A$.∎
+The left-reduction $(A^{\prime},B^{\prime})$ of a mixed-separation+ $(A,B)$ satisfies $A^{\prime}\smallsetminus B^{\prime}=A\smallsetminus B$ and $B^{\prime}\smallsetminus A^{\prime}\supseteq B\smallsetminus A$.∎
 Lemma 5.3.
-Let $(A,B)$ be a mixed-$k$-separation+ of a graph $G$ with left-shift $(A^{\prime},B^{\prime})$.
+Let $(A,B)$ be a mixed-$k$-separation+ of a graph $G$ with left-reduction $(A^{\prime},B^{\prime})$.
   1. (i)
-    $(A^{\prime},B^{\prime})$ is a $-\ell$-separation+ with $\ell\leqslant k$. We have $\ell=k$ if every vertex in $S(A,B)$ has a neighbour in $A\smallsetminus B$.
+    $(A^{\prime},B^{\prime})$ is a mixed-$\ell$-separation+ with $\ell\leqslant k$. We have $\ell=k$ if every vertex in $S(A,B)$ has a neighbour in $A\smallsetminus B$.
   2. (ii)
@@
 Proof.
-(i).
-By definition, the left-shift $(A^{\prime},B^{\prime})$ is a mixed-separation+.
+˜(i).
+By definition, the left-reduction $(A^{\prime},B^{\prime})$ is a mixed-separation+.
 The separator of $(A^{\prime},B^{\prime})$ is obtained from the separator of $(A,B)$ by replacing every vertex $v\in S(A,B)$ that has exactly one neighbour $u\in A\smallsetminus B$ with the edge $uv$ and by removing every vertex in $S(A,B)$ that has no neighbour in $A\smallsetminus B$.
 Hence the order of $(A^{\prime},B^{\prime})$ is at most that of $(A,B)$, with equality if every vertex in $S(A,B)$ has a neighbour in $A\smallsetminus B$.
-(ii).
+˜(ii).
 Assume for a contradiction that there is a vertex $v\in S(A^{\prime},B^{\prime})$ with at most one neighbour in $A^{\prime}\smallsetminus B^{\prime}$.
-Since $A^{\prime}\smallsetminus B^{\prime}=A\smallsetminus B$ by 5.2, the vertex $v$ also has at most one neighbour in $A\smallsetminus B$.
-But then, by the definition of left-shift, we have $v\notin A^{\prime}$, contradicting $v\in S(A^{\prime},B^{\prime})$.
+Since $A^{\prime}\smallsetminus B^{\prime}=A\smallsetminus B$ by ˜5.2, the vertex $v$ also has at most one neighbour in $A\smallsetminus B$.
+But then, by the definition of left-reduction, we have $v\notin A^{\prime}$, contradicting $v\in S(A^{\prime},B^{\prime})$.
 ∎
@@
 Proof.
-(i) $\Rightarrow$ (ii) is immediate.
-(ii) $\Rightarrow$ (i).
-Assume that (i) fails, so there are two edges $ab_{1},ab_{2}\in S(A,B)$ with $a\in A\smallsetminus B$ and both $b_{i}\in B\smallsetminus A$.
+˜(i) $\Rightarrow$ ˜(ii) is immediate.
+˜(ii) $\Rightarrow$ ˜(i).
+Assume that ˜(i) fails, so there are two edges $ab_{1},ab_{2}\in S(A,B)$ with $a\in A\smallsetminus B$ and both $b_{i}\in B\smallsetminus A$.
 Let $X$ be the vertex-set that consists of $a$ together with all vertices in $S(A,B)$ and all endvertices in $B$ of edges in $S(A,B)$ other than the $ab_{i}$.
-Since $|A\smallsetminus B|\geqslant 2$ by (ii), there is $\alpha\in A\smallsetminus B$ distinct from $a$.
+Since $|A\smallsetminus B|\geqslant 2$ by ˜(ii), there is $\alpha\in A\smallsetminus B$ distinct from $a$.
 If some $b_{i}$ is not in $X$, we let $\beta:=b_{i}$.
@@
 Proof.
-(i) $\Rightarrow$ (ii) is immediate.
-(ii) $\Rightarrow$ (iii).
+˜(i) $\Rightarrow$ ˜(ii) is immediate.
+˜(ii) $\Rightarrow$ ˜(iii).
 Assume $|A\smallsetminus B|\leqslant 1$.
@@
 Since $S(A,B)$ is nonempty, $|A|=1$ follows.
-(iii) $\Rightarrow$ (i). This follows from Lemma 5.4 (ii) $\Rightarrow$ (i). ∎
+˜(iii) $\Rightarrow$ ˜(i). This follows from Lemma˜5.4 ˜(ii) $\Rightarrow$ ˜(i). ∎
 Lemma 5.6.
 Let $(A,B)$ be a mixed-4-separation+ in a 4-connected graph $G$ with $A\smallsetminus B\neq\emptyset$ and such that no two edges in $S(A,B)$ share an end in $B\smallsetminus A$.
-Let $(A^{\prime},B^{\prime})$ denote the left-shift of $(A,B)$.
+Let $(A^{\prime},B^{\prime})$ denote the left-reduction of $(A,B)$.
 Then the following assertions are equivalent:
@@
 Proof.
-By Lemma 5.3, every vertex in $S(A^{\prime},B^{\prime})$ has $\geqslant 2$ neighbours in $A^{\prime}\smallsetminus B^{\prime}$.
-(i) $\Rightarrow$ (ii).
+By Lemma˜5.3, every vertex in $S(A^{\prime},B^{\prime})$ has $\geqslant 2$ neighbours in $A^{\prime}\smallsetminus B^{\prime}$.
+˜(i) $\Rightarrow$ ˜(ii).
 Suppose $|A\smallsetminus B|\leqslant 1$.
@@
 Then $|A^{\prime}\smallsetminus B^{\prime}|=1$ since $A\smallsetminus B=A^{\prime}\smallsetminus B^{\prime}$.
-Applying Lemma 5.5 to $(A^{\prime},B^{\prime})$ yields that $(A^{\prime},B^{\prime})$ is an atomic edge-cut.
-(ii) $\Rightarrow$ (i).
+Applying Lemma˜5.5 to $(A^{\prime},B^{\prime})$ yields that $(A^{\prime},B^{\prime})$ is an atomic edge-cut.
+˜(ii) $\Rightarrow$ ˜(i).
 From $A\smallsetminus B=A^{\prime}\smallsetminus B^{\prime}$ we get $|A^{\prime}\smallsetminus B^{\prime}|\geqslant 2$.
-So no two edges in $S(A^{\prime},B^{\prime})$ share an end in $A^{\prime}\smallsetminus B^{\prime}$ by Lemma 5.5 applied to $(A^{\prime},B^{\prime})$.
+So no two edges in $S(A^{\prime},B^{\prime})$ share an end in $A^{\prime}\smallsetminus B^{\prime}$ by Lemma˜5.5 applied to $(A^{\prime},B^{\prime})$.
 So suppose for a contradiction that there are two distinct edges $a_{1}b,a_{2}b$ in $S(A^{\prime},B^{\prime})$ with $a_{i}\in A^{\prime}\smallsetminus B^{\prime}$ and $b\in B^{\prime}\smallsetminus A^{\prime}$.
 Then $a_{i}\in A^{\prime}\smallsetminus B^{\prime}=A\smallsetminus B$ and $b\in B\cap A$.
-Hence $b$ has $\geqslant 2$ neighbours in $A\smallsetminus B$, contradicting that $b$ was replaced with edges when taking the left-shift.
+Hence $b$ has $\geqslant 2$ neighbours in $A\smallsetminus B$, contradicting that $b$ was replaced with edges when taking the left-reduction.
 ∎
@@
 Let $(A,B)$ be a mixed-4-separation of $G$.
-Let $(C,D)$ be a 4-separation of $G$ with right-shift $(C^{\prime},D^{\prime})$.
+Let $(C,D)$ be a 4-separation of $G$ with right-reduction $(C^{\prime},D^{\prime})$.
 Assume $A\subseteq C$.
@@
 If $v\in X$, then $v$ has exactly one neighbour in $B\smallsetminus A$.
-Since $B\smallsetminus A\supseteq D\smallsetminus C$, the definition of right-shift yields $v\notin D^{\prime}$, a contradiction.
+Since $B\smallsetminus A\supseteq D\smallsetminus C$, the definition of right-reduction yields $v\notin D^{\prime}$, a contradiction.
 So $v\notin X$.
 ∎
-5.2. Left-right-shifts and right-left-shifts
+5.2. Left-right-reductions and right-left-reductions
 Definition 5.8.
 Let $(A,B)$ be a mixed-separation+.
-The left-right-shift of $(A,B)$ is obtained from $(A,B)$ by first taking the left-shift and then the right-shift.
-The right-left-shift of $(A,B)$ is obtained from $(A,B)$ by first taking the right-shift and then the left-shift.
+The left-right-reduction of $(A,B)$ is obtained from $(A,B)$ by first taking the left-reduction and then the right-reduction.
+The right-left-reduction of $(A,B)$ is obtained from $(A,B)$ by first taking the right-reduction and then the left-reduction.
 Observation 5.9.
-Let $(A^{\prime},B^{\prime})$ be the left-right-shift of a mixed-4-separation+ $(A,B)$ in a graph $G$.
+Let $(A^{\prime},B^{\prime})$ be the left-right-reduction of a mixed-4-separation+ $(A,B)$ in a graph $G$.
 Then $A^{\prime}\smallsetminus B^{\prime}\supseteq A\smallsetminus B$ and $B^{\prime}\smallsetminus A^{\prime}\supseteq B\smallsetminus A$.
 Proof.
-We perform the two shifts one after the other, applying 5.2 after each shift. ∎
+We perform the two reductions one after the other, applying ˜5.2 after each reduction. ∎
 Key Lemma 5.10.
 Let $(A,B)$ be a mixed-4-separation+ in a 4-connected graph $G$.
-The left-right-shift $(A^{\prime\prime},B^{\prime\prime})$ of $(A,B)$ is a tetra-separation if and only if the following three conditions are satisfied:
+The left-right-reduction $(A^{\prime\prime},B^{\prime\prime})$ of $(A,B)$ is a tetra-separation if and only if the following three conditions are satisfied:
   1. (i)
@@
 Proof.
-Let $(A^{\prime},B^{\prime})$ denote the left-shift of $(A,B)$.
+Let $(A^{\prime},B^{\prime})$ denote the left-reduction of $(A,B)$.
 Backward implication.
-Assume (i)(iii).
-By (iii) and Lemma 5.4 (applied to $(B,A)$ instead of $(A,B)$), no two edges in $S(A,B)$ share an endvertex in $B\smallsetminus A$.
-By (i) and Lemma 5.6, $(A^{\prime},B^{\prime})$ satisfies the matching-condition.
-By Lemma 5.3, every vertex in $S(A^{\prime},B^{\prime})$ has $\geqslant 2$ neighbours in $A^{\prime}\smallsetminus B^{\prime}$.
-By (i) and since $G$ is 4-connected, every vertex in $S(A,B)$ has a neighbour in $A\smallsetminus B$.
-Hence $(A^{\prime},B^{\prime})$ has order four by Lemma 5.3.
-We further have $|A^{\prime}\smallsetminus B^{\prime}|=|A\smallsetminus B|\geqslant 2$ by 5.9 and we have $|B^{\prime}\smallsetminus A^{\prime}|\geqslant 2$ by (ii).
-By Lemma 5.6 (applied for the right-shift), $(A^{\prime\prime},B^{\prime\prime})$ satisfies the matching-condition.
-By Lemma 5.3 (applied for the right-shift), every vertex in $S(A^{\prime\prime},B^{\prime\prime})$ has $\geqslant 2$ neighbours in $B^{\prime\prime}\smallsetminus A^{\prime\prime}$.
-Every vertex in $S(A^{\prime\prime},B^{\prime\prime})$ has $\geqslant 2$ neighbours in $A^{\prime\prime}\smallsetminus B^{\prime\prime}$ since $A^{\prime\prime}\smallsetminus B^{\prime\prime}=A^{\prime}\smallsetminus B^{\prime}$ by 5.9 (applied for the right-shift).
+Assume ˜(i)˜(iii).
+By ˜(iii) and Lemma˜5.4 (applied to $(B,A)$ instead of $(A,B)$), no two edges in $S(A,B)$ share an endvertex in $B\smallsetminus A$.
+By ˜(i) and Lemma˜5.6, $(A^{\prime},B^{\prime})$ satisfies the matching-condition.
+By Lemma˜5.3, every vertex in $S(A^{\prime},B^{\prime})$ has $\geqslant 2$ neighbours in $A^{\prime}\smallsetminus B^{\prime}$.
+By ˜(i) and since $G$ is 4-connected, every vertex in $S(A,B)$ has a neighbour in $A\smallsetminus B$.
+Hence $(A^{\prime},B^{\prime})$ has order four by Lemma˜5.3.
+We further have $|A^{\prime}\smallsetminus B^{\prime}|=|A\smallsetminus B|\geqslant 2$ by ˜5.9 and we have $|B^{\prime}\smallsetminus A^{\prime}|\geqslant 2$ by ˜(ii).
+By Lemma˜5.6 (applied for the right-reduction), $(A^{\prime\prime},B^{\prime\prime})$ satisfies the matching-condition.
+By Lemma˜5.3 (applied for the right-reduction), every vertex in $S(A^{\prime\prime},B^{\prime\prime})$ has $\geqslant 2$ neighbours in $B^{\prime\prime}\smallsetminus A^{\prime\prime}$.
+Every vertex in $S(A^{\prime\prime},B^{\prime\prime})$ has $\geqslant 2$ neighbours in $A^{\prime\prime}\smallsetminus B^{\prime\prime}$ since $A^{\prime\prime}\smallsetminus B^{\prime\prime}=A^{\prime}\smallsetminus B^{\prime}$ by ˜5.9 (applied for the right-reduction).
 Hence $(A^{\prime\prime},B^{\prime\prime})$ satisfies the degree-condition.
-Since $B^{\prime}\smallsetminus A^{\prime}$ is nonempty and $G$ is 4-connected, every vertex in $S(A^{\prime},B^{\prime})$ has a neighbour in $B^{\prime}\smallsetminus A^{\prime}$, so $(A^{\prime\prime},B^{\prime\prime})$ has order four by Lemma 5.3 (applied for the right-shift).
-Finally, $A^{\prime\prime}\smallsetminus B^{\prime\prime}=A^{\prime}\smallsetminus B^{\prime}$ and $B^{\prime\prime}\smallsetminus A^{\prime\prime}\supseteq B^{\prime}\smallsetminus A^{\prime}$ by 5.9, so $(A^{\prime\prime},B^{\prime\prime})$ is proper.
+Since $B^{\prime}\smallsetminus A^{\prime}$ is nonempty and $G$ is 4-connected, every vertex in $S(A^{\prime},B^{\prime})$ has a neighbour in $B^{\prime}\smallsetminus A^{\prime}$, so $(A^{\prime\prime},B^{\prime\prime})$ has order four by Lemma˜5.3 (applied for the right-reduction).
+Finally, $A^{\prime\prime}\smallsetminus B^{\prime\prime}=A^{\prime}\smallsetminus B^{\prime}$ and $B^{\prime\prime}\smallsetminus A^{\prime\prime}\supseteq B^{\prime}\smallsetminus A^{\prime}$ by ˜5.9, so $(A^{\prime\prime},B^{\prime\prime})$ is proper.
 Forward implication.
-Suppose first that (iii) fails.
-Then, by Lemma 5.4 for $(B,A)$, there are two edges $ab_{1},ab_{2}$ in $S(A,B)$ with $a\in A\smallsetminus B$ and $b_{i}\in B\smallsetminus A$.
-Then $a\in A^{\prime\prime}\smallsetminus B^{\prime\prime}$ and $b_{i}\in B^{\prime\prime}\smallsetminus A^{\prime\prime}$ by 5.9.
+Suppose first that ˜(iii) fails.
+Then, by Lemma˜5.4 for $(B,A)$, there are two edges $ab_{1},ab_{2}$ in $S(A,B)$ with $a\in A\smallsetminus B$ and $b_{i}\in B\smallsetminus A$.
+Then $a\in A^{\prime\prime}\smallsetminus B^{\prime\prime}$ and $b_{i}\in B^{\prime\prime}\smallsetminus A^{\prime\prime}$ by ˜5.9.
 Hence $(A^{\prime\prime},B^{\prime\prime})$ fails the matching-condition.
-Suppose next that (iii) holds, but (i) or (ii) fails.
-By (iii) and Lemma 5.4 for $(B,A)$, no two edges in $S(A^{\prime},B^{\prime})$ share an endvertex in $B^{\prime}\smallsetminus A^{\prime}$.
-If (i) fails, we have $|A\smallsetminus B|\leqslant 1$, so Lemma 5.6 gives that $(A^{\prime},B^{\prime})$ violates the matching-condition, and hence $(A^{\prime\prime},B^{\prime\prime})$ also violates the matching-condition.
-Next, assume that (i) holds, but (ii) fails.
-As $|A\smallsetminus B|\geqslant 2$, Lemma 5.6 gives that $(A^{\prime},B^{\prime})$ satisfies the matching-condition.
-The failure of (ii) yields $|B^{\prime}\smallsetminus A^{\prime}|\leqslant 1$.
-Hence applying Lemma 5.6 to $(B^{\prime},A^{\prime})$ (instead of $(A,B)$) yields that $(A^{\prime\prime},B^{\prime\prime})$ violates the matching-condition.
+Suppose next that ˜(iii) holds, but ˜(i) or ˜(ii) fails.
+By ˜(iii) and Lemma˜5.4 for $(B,A)$, no two edges in $S(A^{\prime},B^{\prime})$ share an endvertex in $B^{\prime}\smallsetminus A^{\prime}$.
+If ˜(i) fails, we have $|A\smallsetminus B|\leqslant 1$, so Lemma˜5.6 gives that $(A^{\prime},B^{\prime})$ violates the matching-condition, and hence $(A^{\prime\prime},B^{\prime\prime})$ also violates the matching-condition.
+Next, assume that ˜(i) holds, but ˜(ii) fails.
+As $|A\smallsetminus B|\geqslant 2$, Lemma˜5.6 gives that $(A^{\prime},B^{\prime})$ satisfies the matching-condition.
+The failure of ˜(ii) yields $|B^{\prime}\smallsetminus A^{\prime}|\leqslant 1$.
+Hence applying Lemma˜5.6 to $(B^{\prime},A^{\prime})$ (instead of $(A,B)$) yields that $(A^{\prime\prime},B^{\prime\prime})$ violates the matching-condition.
 ∎
-Proof of Lemma 2.3..
-(i).
-If $G$ is not quasi-5-connected, it has a 4-separation $(A,B)$ with $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$.
-Hence the right-left-shift of $(A,B)$ is a tetra-separation of $G$ by 5.10.
-(ii). Let $(A,B)$ be a tetra-separation of $G$. Let $X:=A\smallsetminus B$ and $Y:=B\smallsetminus A$. Let $F$ denote the set of edges in $S(A,B)$. We have
-$8\leqslant|G|=|X|+|Y|+4-|F|$
+Lemma 5.11.
+Let $k\in\mathbb{N}$.
+Let $G$ be a quasi-$(k+1)$-connected graph with $\geqslant k+4$ vertices.
+Then $G$ has no mixed-$k$-separation $(A,B)$ with $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$.
+Proof.
+Assume that $G$ has a mixed-$k$-separation $(A,B)$ with $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$.
+Let $X:=A\smallsetminus B$ and $Y:=B\smallsetminus A$.
+Let $F$ denote the set of edges in $S(A,B)$.
+We have
+$k+4\leqslant|G|=|X|+|Y|+k-|F|$
 and therefore
 $|F|\leqslant(\,|X|-2)+(\,|Y|-2).$
-Both $X$ and $Y$ have size $\geqslant 2$ by Lemma 4.2 since $(A,B)$ is a tetra-separation.
-Hence both terms $|X|-2$ and $|Y|-2$ are $\geqslant 0$.
+Both $X$ and $Y$ have size $\geqslant 2$ by assumption.
+Hence both terms |X|2|X|-2 and |Y|2|Y|-2 are $\geqslant 0$.
 We may therefore bipartition $F$ into classes $F_{X}$ and $F_{Y}$ of $\leqslant|X|-2$ and $\leqslant|Y|-2$ edges, respectively.
 We obtain $X^{\prime}$ from $X$ by adding for each edge in $F_{X}$ its endvertex in $A$ to $X$, and we obtain $Y^{\prime}$ similarly.
-Then $(X^{\prime},Y^{\prime})$ is a 4-separation of $G$.
+Then $(X^{\prime},Y^{\prime})$ is a $k$-separation of $G$.
 We have $|X^{\prime}\smallsetminus Y^{\prime}|\geqslant|X|-|F_{X}|\geqslant 2$, and similarly $|Y^{\prime}\smallsetminus X^{\prime}|\geqslant 2$.
-Thus $G$ is not quasi-5-connected.
+Thus $G$ is not quasi-$(k+1)$-connected.
 ∎
+Proof of Lemma˜2.3..
+˜(i).
+If $G$ is not quasi-5-connected, it has a 4-separation $(A,B)$ with $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$.
+Hence the right-left-reduction of $(A,B)$ is a tetra-separation of $G$ by ˜5.10.
+˜(ii). Let $(A,B)$ be a tetra-separation of $G$. Then $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$ by Lemma˜4.2. Hence $G$ is not quasi-5-connected by Lemma˜5.11. ∎
 6. Understanding nestedness through connectivity
-The purpose of this section is to characterise the property of total-nestedness for tetra-separations $(A,B)$ in terms of connectivity-properties of their separators $S(A,B)$; see 6.16.
+The purpose of this section is to characterise the property of total-nestedness for tetra-separations $(A,B)$ in terms of connectivity-properties of their separators $S(A,B)$; see ˜6.16.
 This translates the problem of proving total-nestedness of $(A,B)$ to a problem about constructing various families of paths that link up the vertices and edges in the separator $S(A,B)$ in different ways.
 For an element $x\in V(G)\cup E(G)$ we let $\hat{x}$ denote the singleton $\{x\}$ if $x$ is a vertex, and the set of endvertices of $x$ if $x$ is an edge.
-For a set $X\subseteq V(G)\cup E(G)$ we write ${\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{\hat{X}}}:=\bigcup_{x\in X}\hat{x}$.
+For a set $X\subseteq V(G)\cup E(G)$ we write $^{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{\hat{X}}}:=\bigcup_{x\in X}\hat{x}$.
 A bipartition of a set is balanced if its two classes have the same size.
 Lemma 6.1.
-Let $G$ be a $k$-connected graph. Then $G$ has no $-\ell$-separation for any $\ell<k$.
+Let $G$ be a $k$-connected graph. Then $G$ has no mixed-$\ell$-separation for any $\ell<k$.
 Proof.
-Assume for a contradiction that $(A,B)$ is a $-\ell$-separation of $G$ with $\ell<k$.
+Assume for a contradiction that $(A,B)$ is a mixed-$\ell$-separation of $G$ with $\ell<k$.
 Then $S(A,B)$ contains an edge $ab$ with $a\in A\smallsetminus B$ and $b\in B\smallsetminus A$.
 We claim that $A\smallsetminus(B\cup\{a\})$ or $B\smallsetminus(A\cup\{b\})$ is non-empty.
-Indeed, otherwise $V(G)=(A\cap B)\cup\{a,b\}$, so $|G|=\ell+1\leqslant k$, contradicting that $k$-connected graphs such as $G$ have $>k$ vertices by definition.
+Indeed, otherwise $V(G)=(A\cap B)\cup\{a,b\}$, so $|G|=\ell+1\leqslant k$, contradicting that $k$-connected graphs such as $G$ have >k>k vertices by definition.
 So suppose that there is a vertex $v\in A\smallsetminus(B\cup\{a\})$.
@@
   1. (i)
-    the maximum number of independent $X$$Y$ paths in $G$;
+    the maximum number of independent XXYY paths in $G$;
   2. (ii)
     the minimum order of a mixed-separation in $G$ that strongly separates $X$ and $Y$.
-Moreover, if $G$ is $k$-connected, then (i), (ii)${}\geqslant k$.
+[…]
 Proof.
 As usual, (i) $\leqslant$ (ii), so it remains to show (i) $\geqslant$ (ii).
-Since every $X$$Y$ edge forms a path in every maximum set of independent $X$$Y$ paths in $G$, we may assume without loss of generality that $G$ contains no $X$$Y$ edges.
+Since every XXYY edge forms a path in every maximum set of independent XXYY paths in $G$, we may assume without loss of generality that $G$ contains no XXYY edges.
 Let $X^{\prime}:=N_{G}(X)$ and $Y^{\prime}:=N_{G}(Y)$.
@@
 Let $G^{\prime}:=G-(X\cup Y)$.
-[…]
-The ‘moreover’-part follows from Lemma 6.1. ∎
+Now let $\mathcal{P}$ be a maximum-size set of independent XXYY paths in $G$.
+Then $\mathcal{P}$ determines a set $\mathcal{P}^{\prime}$ of $|\mathcal{P}|$ pairwise disjoint $X^{\prime}$$Y^{\prime}$ paths in $G^{\prime}$.
+Hence $G^{\prime}$ contains an $X^{\prime}$$Y^{\prime}$ separator $S^{\prime}$ of size $|\mathcal{P}|$, by Menger’s theorem applied in $G^{\prime}$ to $X^{\prime},Y^{\prime}$ and $\mathcal{P}^{\prime}$.
+Then $S^{\prime}$ is the separator of a mixed-separation of $G$ that strongly separates $X$ and $Y$, and which has order $|S^{\prime}|=|\mathcal{P}|$.
+The ‘moreover’-part follows from Lemma˜6.1. ∎
 6.1. Half-connected
-A mixed-separation $(A,B)$ of a graph $G$ is half-connected if $G[A\smallsetminus B]$ or $G[B\smallsetminus A]$ is connected [Tridecomp].
+A mixed-separation $(A,B)$ of a graph $G$ is half-connected if $G[A\smallsetminus B]$ or $G[B\smallsetminus A]$ is connected [CK23].
 Lemma 6.3.
@@
 To see that $(C,D)$ is a tetra-separation, note first that both $G[C\smallsetminus D]$ and $G[D\smallsetminus C]$ are disconnected.
-Since $G$ is 4-connected, every vertex in $S(A,B)$ has a neighbour in every component of $G-S(A,B)$.
+Since $G$ is 4-connected, every vertex in $S(A,B)$ has a neighbour in every component of GS(A,B)G-S(A,B).
 So $(C,D)$ satisfies the degree-condition, and it trivially satisfies the matching-condition.
@@
 This, however, contradicts that $(C,D)$ has order two.
-Let $(C^{\prime\prime},D^{\prime\prime})$ be the left-right-shift of $(C^{\prime},D^{\prime})$, which is a tetra-separation by 5.10.
-First, note that $\hat{x}_{1}\subseteq C^{\prime}\smallsetminus D^{\prime}\subseteq C^{\prime\prime}\smallsetminus D^{\prime\prime}$ and $\hat{x}_{2}\subseteq D^{\prime}\smallsetminus C^{\prime}\subseteq D^{\prime\prime}\smallsetminus C^{\prime\prime}$ by 5.9.
+Let $(C^{\prime\prime},D^{\prime\prime})$ be the left-right-reduction of $(C^{\prime},D^{\prime})$, which is a tetra-separation by ˜5.10.
+First, note that $\hat{x}_{1}\subseteq C^{\prime}\smallsetminus D^{\prime}\subseteq C^{\prime\prime}\smallsetminus D^{\prime\prime}$ and $\hat{x}_{2}\subseteq D^{\prime}\smallsetminus C^{\prime}\subseteq D^{\prime\prime}\smallsetminus C^{\prime\prime}$ by ˜5.9.
 So, $x_{1}$ and $x_{2}$ are in the $C^{\prime\prime}$-link and the $D^{\prime\prime}$-link, respectively.
@@
 Let $(A,B)$ and $(C,D)$ be crossing tetra-separations in a graph $G$.
-Assume that the corner for $AC$ is potter, witnessed by an edge $u_{1}u_{2}$ that dangles from the vertex $u_{2}$ in the $C$-link through the $A$-link, and by an edge $u_{3}u_{4}$ that dangles from the vertex $u_{3}$ in the $A$-link through the $C$-link; see Figure 4.2.
+Assume that the corner for $AC$ is potter, witnessed by an edge $u_{1}u_{2}$ that dangles from the vertex $u_{2}$ in the $C$-link through the $A$-link, and by an edge $u_{3}u_{4}$ that dangles from the vertex $u_{3}$ in the $A$-link through the $C$-link; see Figure˜4.2.
 Then $\{u_{2},u_{3}u_{4}\}$ is $A$-weird and $\{u_{3},u_{1}u_{2}\}$ is $C$-weird.
-Let $X,Y\subseteq V(G)\cup E(G)$. An $X$$Y$ edge means an edge $xy$ with $x\in X$ and $y\in Y$.
+Let $X,Y\subseteq V(G)\cup E(G)$. An XXYY edge means an edge $xy$ with $x\in X$ and $y\in Y$.
 Definition 6.7 (0-potter-linked).
@@
 Observe that $(C,D)$ crosses $(A,B)$ since the $C$-link and the $D$-link are non-empty.
-Let $(C^{\prime},D^{\prime})$ be the left-right-shift of $(C,D)$, which is a tetra-separation by 5.10.
+Let $(C^{\prime},D^{\prime})$ be the left-right-reduction of $(C,D)$, which is a tetra-separation by ˜5.10.
 […]
-The following definition is supported by Figure 6.1.
+The following definition is supported by Figure˜6.1.
 Definition 6.9 (1-potter-linked).
@@
 We show that $\pi_{1}$ is not 1-potter-linked to $\pi_{2}$ with respect to $(A,B)$.
-For this, we show that for $(X,Y)=(A,B)$ in Definition 6.9, all of (6.9.1)(6.9.3) fail.
-Observe that $\pi_{1}$ is $A$-weird by Example 6.6, so (6.9.1) fails.
+For this, we show that for $(X,Y)=(A,B)$ in Definition˜6.9, all of (6.9.1)(6.9.3) fail.
+Observe that $\pi_{1}$ is $A$-weird by Example˜6.6, so (6.9.1) fails.
 By assumption, the corner $AC$ is the only corner that is potter.
@@
 In particular, $|C_{1}\smallsetminus D_{1}|\geqslant 2$ and $|D_{1}\smallsetminus C_{1}|\geqslant 2$.
-See Figure 6.2 for the next definition.
+See Figure˜6.2 for the next definition.
 We define $(C_{2},D_{2}):=(C_{1}\cup a_{1},(D_{1}\cup A)\smallsetminus\{v_{1}\})$.
 We have $S(C_{2},D_{2})=S(C_{1},D_{1})\cup\{a^{\prime}_{1}v_{1},a_{1}\}$ since $v_{1}$ does not have a neighbour in $A$ besides $a_{1}^{\prime}$.
-By 4-connectivity of $G$ and Lemma 6.1, $(C_{2},D_{2})$ has order exactly 4.
-Let $(C,D)$ be the left-right-shift of $(C_{2},D_{2})$.
+By 4-connectivity of $G$ and Lemma˜6.1, $(C_{2},D_{2})$ has order exactly 4.
+Let $(C,D)$ be the left-right-reduction of $(C_{2},D_{2})$.
 Since $C_{1}\smallsetminus D_{1}\subseteq C_{2}\smallsetminus D_{2}$ we have $|C_{2}\smallsetminus D_{2}|\geqslant 2$, and similarly $|D_{2}\smallsetminus C_{2}|\geqslant 2$.
-Hence $(C,D)$ is a tetra-separation by 5.10.
+Hence $(C,D)$ is a tetra-separation by ˜5.10.
 Consider the crossing-diagram for $(A,B)$ and $(C,D)$.
@@
 Hence $e_{1}\notin S(C,D)$.
-Since the end $b_{1}$ of $e_{1}$ lies in $C_{1}\smallsetminus D_{1}\subseteq C\smallsetminus D$ by 5.9, the edge $e_{1}$ lies in the $C$-link.
+Since the end $b_{1}$ of $e_{1}$ lies in $C_{1}\smallsetminus D_{1}\subseteq C\smallsetminus D$ by ˜5.9, the edge $e_{1}$ lies in the $C$-link.
 Next, we claim that the $D$-link includes $\pi_{2}$.
@@
 Now let $f$ be an edge in $\pi_{2}$.
-The endvertex of $f$ in $B\smallsetminus A$ lies in $\hat{\pi}_{2}\cap B\subseteq D_{1}\smallsetminus C_{1}\subseteq D\smallsetminus C$ by 5.9.
-The endvertex of $f$ in $A\smallsetminus B$ must be distinct from $a_{1}$ by the matching-condition, and hence is contained in $D_{2}\smallsetminus C_{2}\subseteq D\smallsetminus C$ by 5.9.
+The endvertex of $f$ in $B\smallsetminus A$ lies in $\hat{\pi}_{2}\cap B\subseteq D_{1}\smallsetminus C_{1}\subseteq D\smallsetminus C$ by ˜5.9.
+The endvertex of $f$ in $A\smallsetminus B$ must be distinct from $a_{1}$ by the matching-condition, and hence is contained in $D_{2}\smallsetminus C_{2}\subseteq D\smallsetminus C$ by ˜5.9.
 Therefore, both ends of $f$ lie in $D\smallsetminus C$ while $f\in S(A,B)$, so $f$ lies in the $D$-link.
@@
 (4.1).
-By 4.7, the corners for $AD$ and $BC$ are not potter.
+By ˜4.7, the corners for $AD$ and $BC$ are not potter.
 Assume for contradiction that the $BD$-corner is potter.
@@
 ∎
-Figure 6.2. The situation in the second half of the proof of Lemma 6.10. $\pi_{1}$ is red and $\pi_{2}$ is blue.
+Figure 6.2. The situation in the second half of the proof of Lemma˜6.10. $\pi_{1}$ is red and $\pi_{2}$ is blue.
 Definition 6.11 (2-potter-linked).
@@
 Assume that a tetra-separation $(C,D)$ crosses $(A,B)$ so that $\pi_{1}$ and $\pi_{2}$ form opposite links, say $\pi_{1}$ is the $C$-link and $\pi_{2}$ is the $D$-link, and so that exactly two corners are potter.
-By 4.7, no two adjacent corners are potter.
+By ˜4.7, no two adjacent corners are potter.
 So we can assume without loss of generality that the $AC$-corner and the $BD$-corner are potter while the other two corners are not potter.
-Then $\pi_{1}$ is $A$-weird and $\pi_{2}$ is $B$-weird by Example 6.6.
+Then $\pi_{1}$ is $A$-weird and $\pi_{2}$ is $B$-weird by Example˜6.6.
 So, (6.11.1) and (6.11.2) fail for $(X,Y)=(A,B)$.
@@
 Say (6.11.1) and (6.11.2) fail for $(X,Y)=(A,B)$; that is to say, $\pi_{1}$ is $A$-weird and $\pi_{2}$ is $B$-weird.
-For $i\in\{1,2\}$, write $\pi_{i}=:\{v_{i},e_{i}\}$ where $v_{i}$ is a vertex and $e_{i}=a_{i}b_{i}$ is an edge with $a_{i}\in A\smallsetminus B$ and $b_{i}\in B\smallsetminus A$; see Figure 6.3.
+For $i\in\{1,2\}$, write $\pi_{i}=:\{v_{i},e_{i}\}$ where $v_{i}$ is a vertex and $e_{i}=a_{i}b_{i}$ is an edge with $a_{i}\in A\smallsetminus B$ and $b_{i}\in B\smallsetminus A$; see Figure˜6.3.
 Let $a^{\prime}_{1}$ be the unique neighbour of $v_{1}$ in $A\smallsetminus\{a_{1}\}$, and let $b_{2}^{\prime}$ be the unique neighbour of $v_{2}$ in $B\smallsetminus\{b_{1}\}$.
@@
 Let $C:=B\smallsetminus\{v_{2}\})\cup\{a_{1}\}$ and $D:=(A\smallsetminus\{v_{1}\})\cup\{b_{2}\}$.
-Since $\pi_{1}$ is $A$-weird, $\pi_{2}$ is $B$-weird, and $v_{1}v_{2}\notin E(G)$ by 6.12, we have $S(C,D)=\{a^{\prime}_{1}v_{1},a_{1},b_{2},v_{2}b^{\prime}_{2}\}$.
+Since $\pi_{1}$ is $A$-weird, $\pi_{2}$ is $B$-weird, and $v_{1}v_{2}\notin E(G)$ by ˜6.12, we have $S(C,D)=\{a^{\prime}_{1}v_{1},a_{1},b_{2},v_{2}b^{\prime}_{2}\}$.
 Note that $(C,D)$ crosses $(A,B)$ since $v_{1}$ is in the $C$-link and $v_{2}$ is in the $D$-link.
@@
 The corners $AC$ and $BD$ are potter by construction.
-By 4.7, no other corner is potter.
+By ˜4.7, no other corner is potter.
 ∎
-Figure 6.3. The situation in the second half of the proof of Lemma 6.13. $\pi_{1}$ is red and $\pi_{2}$ is blue.
+Figure 6.3. The situation in the second half of the proof of Lemma˜6.13. $\pi_{1}$ is red and $\pi_{2}$ is blue.
 Corollary 6.14.
@@
 Proof.
-By 4.7, every crossing pair of tetra-separations has at most two corners that are potter. So the corollary immediately follows from Lemma 6.8, Lemma 6.10 and Lemma 6.13. ∎
+By ˜4.7, every crossing pair of tetra-separations has at most two corners that are potter. So the corollary immediately follows from Lemma˜6.8, Lemma˜6.10 and Lemma˜6.13. ∎
 Summary.
@@
 Proof.
-We combine Lemma 6.3, Lemma 6.5 and Corollary 6.14 with the (Crossing Lemma). (4.1). ∎
+We combine Lemma˜6.3, Lemma˜6.5 and Corollary˜6.14 with the (Crossing Lemma). (4.1). ∎
 7. Thickened $K_{4,m}$’s
-This section deals with the thickened $K_{4,m}$ outcome of Theorem 1; see 7.8.
+This section deals with the thickened $K_{4,m}$ outcome of Theorem˜1; see ˜7.8.
 A sprinkled $K_{4,m}$ is a graph obtained from a $K_{4,m}$ with a side $X$ of size four by possibly adding edges between vertices in $X$. We refer to $X$ as a sprinkled side.
-Let $Z\subseteq V(G)$ and $K$ a component of $G-Z$.
+Let $Z\subseteq V(G)$ and $K$ a component of GZG-Z.
 When $Z$ and $K$ are clear from context, we write $(U^{\prime}_{K},W^{\prime}_{K}):=(V(K)\cup Z,V(G-K))$.
-Let $Z$ be a set of four vertices in a 4-connected graph $G$ such that $G-Z$ has $\geqslant 4$ components.
-For each component $K$ of $G-Z$, let $(U_{K},W_{K})$ denote the left-right-shift of $(U^{\prime}_{K},W^{\prime}_{K})$.
+Let $Z$ be a set of four vertices in a 4-connected graph $G$ such that GZG-Z has $\geqslant 4$ components.
+For each component $K$ of GZG-Z, let $(U_{K},W_{K})$ denote the left-right-reduction of $(U^{\prime}_{K},W^{\prime}_{K})$.
 Lemma 7.1.
-Let $Z$ be a set of four vertices in a 4-connected graph $G$, such that $G-Z$ has $\geqslant 4$ components.
-Let $K$ be a component of $G-Z$.
-Then $(U_{K},W_{K})$ equals the left-shift of $(U^{\prime}_{K},W^{\prime}_{K})$.
+Let $Z$ be a set of four vertices in a 4-connected graph $G$, such that GZG-Z has $\geqslant 4$ components.
+Let $K$ be a component of GZG-Z.
+Then $(U_{K},W_{K})$ equals the left-reduction of $(U^{\prime}_{K},W^{\prime}_{K})$.
 In particular, $(U_{K},W_{K})\leqslant(U^{\prime}_{K},W^{\prime}_{K})$ and $W_{K}=W^{\prime}_{K}$.
 Proof.
-Let $(A,B)$ denote the left-shift of $(U^{\prime}_{K},W^{\prime}_{K})$.
-Since every vertex in $S(U^{\prime}_{K},W^{\prime}_{K})$ has $\geqslant 2$ neighbours in $W^{\prime}_{K}\smallsetminus U^{\prime}_{K}$, and since $W^{\prime}_{K}\smallsetminus U^{\prime}_{K}\subseteq B\smallsetminus A$ by 5.2, it follows that every vertex in $S(A,B)$ has $\geqslant 2$ neighbours in $B\smallsetminus A$.
+Let $(A,B)$ denote the left-reduction of $(U^{\prime}_{K},W^{\prime}_{K})$.
+Since every vertex in $S(U^{\prime}_{K},W^{\prime}_{K})$ has $\geqslant 2$ neighbours in $W^{\prime}_{K}\smallsetminus U^{\prime}_{K}$, and since $W^{\prime}_{K}\smallsetminus U^{\prime}_{K}\subseteq B\smallsetminus A$ by ˜5.2, it follows that every vertex in $S(A,B)$ has $\geqslant 2$ neighbours in $B\smallsetminus A$.
 Hence $(A,B)=(U_{K},W_{K})$ and $W^{\prime}_{K}=B=W_{K}$.
@@
 Lemma 7.2.
-Let $Z$ be a set of four vertices in a 4-connected graph $G$ such that $G-Z$ has $\geqslant 4$ components. Let $K$ be a component of $G-Z$. Then the following assertions are equivalent:
+Let $Z$ be a set of four vertices in a 4-connected graph $G$ such that GZG-Z has $\geqslant 4$ components. Let $K$ be a component of GZG-Z. Then the following assertions are equivalent:
   1. (i)
@@
   2. (ii)
-    the left-right-shift of $(U^{\prime}_{K},W^{\prime}_{K})$ is a totally-nested tetra-separation.
+    the left-right-reduction of $(U^{\prime}_{K},W^{\prime}_{K})$ is a totally-nested tetra-separation.
 Proof.
-$\neg$ (i) $\Rightarrow$ $\neg$ (ii). If $K$ has only one vertex, then $|U_{K}\smallsetminus W_{K}|=1$, so $(U^{\prime}_{K},W^{\prime}_{K})$ fails to be a tetra-separation by 5.10.
-(i) $\Rightarrow$ (ii).
+$\neg$ ˜(i) $\Rightarrow$ $\neg$ ˜(ii). If $K$ has only one vertex, then $|U_{K}\smallsetminus W_{K}|=1$, so $(U^{\prime}_{K},W^{\prime}_{K})$ fails to be a tetra-separation by ˜5.10.
+˜(i) $\Rightarrow$ ˜(ii).
 We abbreviate $(U,W):=(U_{K},W_{K})$ and $(U^{\prime},W^{\prime}):=(U^{\prime}_{K},W^{\prime}_{K})$.
-Since $K$ has $\geqslant 2$ vertices, $(U,W)$ is a tetra-separation by 5.10.
-To show that $(U,W)$ is totally-nested, by 6.16 it suffices to show that $(U,W)$ is externally 5-connected.
-Half-connected: We have $G[U\smallsetminus W]=K$ by Lemma 7.1.
+Since $K$ has $\geqslant 2$ vertices, $(U,W)$ is a tetra-separation by ˜5.10.
+To show that $(U,W)$ is totally-nested, by ˜6.16 it suffices to show that $(U,W)$ is externally 5-connected.
+Half-connected: We have $G[U\smallsetminus W]=K$ by Lemma˜7.1.
 To check the remaining aspects of external 5-connectivity, for each $z\in Z$ we denote by $z^{*}$ the element of $S(U,W)$ that equal or incident to $z$.
-By Lemma 7.1, $z^{*}$ is either equal to $z$ or $z^{*}$ is an edge between $K$ and $z$.
+By Lemma˜7.1, $z^{*}$ is either equal to $z$ or $z^{*}$ is an edge between $K$ and $z$.
 Denote by $z^{U}$ the vertex in $\hat{z}\cap U$.
-We can use the three components of $G-Z$ besides $K$ to find paths:
+We can use the three components of GZG-Z besides $K$ to find paths:
   1. ($\dagger$)
     There are three independent $z_{1}$$z_{2}$ paths in $G$ with all internal vertices outside $K$, for every two distinct vertices $z_{1},z_{2}\in Z$.
-3-linked: follows from ($\dagger$)Section 7.
+3-linked: follows from ˜($\dagger$)Section˜7.
 0-potter-linked: Let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of $S(U,W)$.
-We find three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths through $W\smallsetminus U$ by ($\dagger$)Section 7.
+We find three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths through $W\smallsetminus U$ by ˜($\dagger$)Section˜7.
 We claim that there are two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths through $K$.
@@
 Then $\pi_{1}$ is $U$-weird since each vertex in $Z$ has $\geqslant 3$ neighbours in $W\smallsetminus U$.
-By ($\dagger$)Section 7, there exist three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[W]$.
+By ˜($\dagger$)Section˜7, there exist three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[W]$.
 Hence $\{\pi_{1},\pi_{2}\}$ is 1-potter-linked with respect to $(U,W)$.
@@
 ∎
-We write $)𝜎𝑍{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{\sigma(Z)}}$ for the star $\{\,(U_{K},W_{K}):K$ is a component of $G-Z$ with $|K|\geqslant 2\,\}$.
+[…]
 Corollary 7.3.
-Let $Z$ be a set of four vertices in a 4-connected graph $G$ such that $G-Z$ has $\geqslant 4$ components. Then $\sigma(Z)$ is a star of totally-nested tetra-separations.
+Let $Z$ be a set of four vertices in a 4-connected graph $G$ such that GZG-Z has $\geqslant 4$ components. Then $\sigma(Z)$ is a star of totally-nested tetra-separations.
 Proof.
-By Lemma 7.2, every element in $\sigma(Z)$ is a totally-nested tetra-separation.
-Let $K\neq L$ be components of $G-Z$, each of size $\geqslant 2$.
-By Lemma 7.1, $(U_{K},W_{K})\leqslant(U_{K}^{\prime},W_{K}^{\prime})$ and $(U_{L},W_{L})\leqslant(U_{L}^{\prime},W_{L}^{\prime})$.
+By Lemma˜7.2, every element in $\sigma(Z)$ is a totally-nested tetra-separation.
+Let $K\neq L$ be components of GZG-Z, each of size $\geqslant 2$.
+By Lemma˜7.1, $(U_{K},W_{K})\leqslant(U_{K}^{\prime},W_{K}^{\prime})$ and $(U_{L},W_{L})\leqslant(U_{L}^{\prime},W_{L}^{\prime})$.
 With $(U_{K}^{\prime},W_{K}^{\prime})\leqslant(W_{L}^{\prime},U_{L}^{\prime})$, it follows that $(U_{K},W_{K})\leqslant(W_{L},U_{L})$.
@@
 Let $(A,B)$ be a 4-separation in a 4-connected graph $G$ such that $|A\smallsetminus B|\geqslant 2$ and every vertex in $S(A,B)$ has $\geqslant 2$ neighbours in $B\smallsetminus A$.
-Let $(A^{\prime\prime},B^{\prime\prime})$ be the left-right-shift of $(A,B)$.
+Let $(A^{\prime\prime},B^{\prime\prime})$ be the left-right-reduction of $(A,B)$.
 Then no tetra-separation $(C,D)$ of $G$ satisfies $(A^{\prime\prime},B^{\prime\prime})<(C,D)\leqslant(A,B)$.
@@
 Suppose for a contradiction that there is a tetra-separation $(C,D)$ with $(A^{\prime\prime},B^{\prime\prime})<(C,D)\leqslant(A,B)$.
-By 5.10, $(A^{\prime\prime},B^{\prime\prime})$ is a tetra-separation.
-By Lemma 7.1, we have $B=B^{\prime\prime}$.
+By ˜5.10, $(A^{\prime\prime},B^{\prime\prime})$ is a tetra-separation.
+By Lemma˜7.1, we have $B=B^{\prime\prime}$.
 So both inclusions $B^{\prime\prime}\supseteq D\supseteq B$ are actually equalities.
@@
 In particular, $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$.
-Then by Lemma 5.4$(A,B)$ satisfies the matching-condition.
+Then by Lemma˜5.4(A,B)(A,B) satisfies the matching-condition.
 Let $v$ be a vertex in $S(A,B)$.
-By 4-connectivity of $G$ and by Lemma 6.1, $v$ has a neighbour in every component of $G-S(A,B)$.
+By 4-connectivity of $G$ and by Lemma˜6.1, $v$ has a neighbour in every component of GS(A,B)G-S(A,B).
 In particular, $v$ satisfies the degree-condition.
@@
 Lemma 7.6.
-Let $Z$ be a set of four vertices in a 4-connected graph $G$ such that $G-Z$ has $\geqslant 4$ components.
+Let $Z$ be a set of four vertices in a 4-connected graph $G$ such that GZG-Z has $\geqslant 4$ components.
 Let $(E,F)$ be a totally-nested tetra-separation of $G$.
-Then there exists a component $K$ of $G-Z$ with $|K|\geqslant 2$ such that $(E,F)\leqslant(U^{\prime}_{K},W^{\prime}_{K})$ or $(F,E)\leqslant(U^{\prime}_{K},W^{\prime}_{K})$.
+Then there exists a component $K$ of GZG-Z with $|K|\geqslant 2$ such that $(E,F)\leqslant(U^{\prime}_{K},W^{\prime}_{K})$ or $(F,E)\leqslant(U^{\prime}_{K},W^{\prime}_{K})$.
 Proof.
-We denote by $\mathcal{K}$ the set of components of $G-Z$.
+We denote by $\mathcal{K}$ the set of components of GZG-Z.
 First, we show that there is a component $K\in\mathcal{K}$ such that either $E\smallsetminus Z\subseteq V(K)$ or $F\smallsetminus Z\subseteq V(K)$.
@@
 We define $A_{i}:=Z\cup\bigcup_{K\in\mathcal{K}_{i}}V(K)$.
-Then $(A_{1},A_{2})$ is a tetra-separation by Lemma 7.5.
+Then $(A_{1},A_{2})$ is a tetra-separation by Lemma˜7.5.
 Now $E\nsubseteq A_{i}$ and $F\nsubseteq A_{i}$ for both $i\in\{1,2\}$ implies that $(A_{1},A_{2})$ crosses $(E,F)$, contradicting that $(E,F)$ is totally-nested.
 Hence we have $E\subseteq V(K)\cup Z$, say, for a component $K\in\mathcal{K}$.
-Since every vertex in $Z$ has $\geqslant 2$ neighbours outside $V(K)\cup Z$, it follows that $(U^{\prime}_{K},W^{\prime}_{K})$ equals its right-shift.
-Hence $(E,F)\leqslant(U^{\prime}_{K},W^{\prime}_{K})$ by Lemma 5.7.
-Since $(E,F)$ is a tetra-separation, we have $|E\smallsetminus F|\geqslant 2$ by Lemma 5.5.
-Therefore, $|K|\geqslant 2$.
+Since every vertex in $Z$ has $\geqslant 2$ neighbours outside $V(K)\cup Z$, it follows that $(U^{\prime}_{K},W^{\prime}_{K})$ equals its right-reduction.
+Hence $(E,F)\leqslant(U^{\prime}_{K},W^{\prime}_{K})$ by Lemma˜5.7.
+Since $(E,F)$ is a tetra-separation, we have $|E\smallsetminus F|\geqslant 2$ by Lemma˜$5.$ Therefore, $|K|\geqslant 2$.
 ∎
@@
 Assume that two tetra-separations $(A,B)$ and $(C,D)$ of $G$ cross so that the vertex-centre $Z$ has size four.
-Let $m$ denote the number of trivial components of $G-Z$.
+Let $m$ denote the number of trivial components of GZG-Z.
 Then:
@@
   3. (iii)
-    If $\sigma(Z)\neq\emptyset$, then $G-Z$ has a component with $\geqslant 2$ vertices, and the torso of $\sigma(Z)$ is a thickened $K_{4,m}$ where $Z$ is the left side.
+    If $\sigma(Z)\neq\emptyset$, then GZG-Z has a component with $\geqslant 2$ vertices, and the torso of $\sigma(Z)$ is a thickened $K_{4,m}$ where $Z$ is the left side.
   4. (iv)
-    If $\sigma(Z)=\emptyset$, then every component of $G-Z$ is trivial, and $G=\tau$ is a sprinkled $K_{4,m}$ with $Z$ as left side and $m\geqslant 4$.
+    If $\sigma(Z)=\emptyset$, then every component of GZG-Z is trivial, and $G=\tau$ is a sprinkled $K_{4,m}$ with $Z$ as left side and $m\geqslant 4$.
 Proof.
-(i).
-By Corollary 7.3, $\sigma(Z)$ is a star in $N$.
+˜(i).
+By Corollary˜7.3, $\sigma(Z)$ is a star in $N$.
 To show that $\sigma(Z)$ is splitting, assume for a contradiction that it is interlaced by a totally-nested tetra-separation $(E,F)$.
-By Lemma 7.6, there is a component $K$ of $G-Z$ with $|K|\geqslant 2$ and $(E,F)\leqslant(U^{\prime}_{K},W^{\prime}_{K})$, say.
+By Lemma˜7.6, there is a component $K$ of GZG-Z with $|K|\geqslant 2$ and $(E,F)\leqslant(U^{\prime}_{K},W^{\prime}_{K})$, say.
 Since $K$ has at least two vertices, we have $(U_{K},W_{K})\in\sigma(Z)$.
-Note that $(U_{K},W_{K})\leqslant(U^{\prime}_{K},W^{\prime}_{K})$ by Lemma 7.1, which implies $(U_{K},W_{K})\not<(F,E)$.
-By Lemma 7.4, we cannot have $(U_{K},W_{K})<(E,F)$.
+Note that $(U_{K},W_{K})\leqslant(U^{\prime}_{K},W^{\prime}_{K})$ by Lemma˜7.1, which implies $(U_{K},W_{K})\not<(F,E)$.
+By Lemma˜7.4, we cannot have $(U_{K},W_{K})<(E,F)$.
 Since $(U_{K},W_{K})\in N$, we therefore have $(E,F)\leqslant(U_{K},W_{K})$ or $(F,E)\leqslant(U_{K},W_{K})$, contradicting that $(E,F)$ interlaces $\sigma(Z)$.
-(ii).
-We denote by $\sigma^{\prime}(Z)$ the star $\{\,(U^{\prime}_{K},W^{\prime}_{K}):K$ is a component of $G-Z$ with $|K|\geqslant 2\,\}$.
-Both $(A,B)$ and $(C,D)$ interlace $\sigma^{\prime}(Z)$, hence they also interlace $\sigma(Z)$ by Lemma 7.1.
-(iii) and (iv) are straightforward. ∎
+˜(ii).
+We denote by $\sigma^{\prime}(Z)$ the star $\{\,(U^{\prime}_{K},W^{\prime}_{K}):K$ is a component of GZG-Z with $|K|\geqslant 2\,\}$.
+Both $(A,B)$ and $(C,D)$ interlace $\sigma^{\prime}(Z)$, hence they also interlace $\sigma(Z)$ by Lemma˜7.1.
+˜(iii) and ˜(iv) are straightforward. ∎
 A graph $G$ is 4-angry if it is 4-connected and every tetra-separation of $G$ is crossed by a tetra-separation.
@@
 Let $X$ be a sprinkled side of $G$.
-Let $\{P_{1},\ldots,P_{4}\}$ be a partition of the other side $V(G-X)$.
+Let $\{P_{1},\ldots,P_{4}\}$ be a partition of the other side V(GX)V(G-X).
 Let $(A,B)$ be the mixed-4-separation $(P_{1}\cup P_{2}\cup X,X\cup P_{3}\cup P_{4})$ and let $(C,D)$ be the mixed-4-separation $(P_{1}\cup P_{3}\cup X,X\cup P_{2}\cup P_{4})$.
 Then $(A,B)$ and $(C,D)$ are tetra-separations with separators equal to $X$, and they cross.
-(i) $\Rightarrow$ (ii) follows from 7.7 (iii) and (iv). ∎
+(i) $\Rightarrow$ (ii) follows from ˜7.7 ˜(iii) and ˜(iv). ∎
 8. Generalised double-wheels
-To deal with the generalised double-wheel outcome of Theorem 1, we first introduce a particular cycle-decomposition in Section 8.1 for analysing the wheel-structure obtained from two tetra-separations that cross with vertex-centre of size two.
-Section 8.2 introduces machinery that exploits the structure of this cycle-decomposition to find independent paths.
-Section 8.3 employs this machinery to show that all non-trivial bags of the cycle-decomposition give rise to totally-nested tetra-separations (8.27).
-Section 8.4 shows that these totally-nested tetra-separations form a splitting star as in Theorem 1 and deduces that its torso is a generalised double-wheel (8.37).
-Finally, Section 8.5 characterises the special case where additionally all tetra-separations of the graph are crossed (8.41).
+To deal with the generalised double-wheel outcome of Theorem˜1, we first introduce a particular cycle-decomposition in Section˜8.1 for analysing the wheel-structure obtained from two tetra-separations that cross with vertex-centre of size two.
+Section˜8.2 introduces machinery that exploits the structure of this cycle-decomposition to find independent paths.
+Section˜8.3 employs this machinery to show that all non-trivial bags of the cycle-decomposition give rise to totally-nested tetra-separations (˜8.27).
+Section˜8.4 shows that these totally-nested tetra-separations form a splitting star as in Theorem˜1 and deduces that its torso is a generalised double-wheel (˜8.37).
+Finally, Section˜8.5 characterises the special case where additionally all tetra-separations of the graph are crossed (˜8.41).
 8.1. Block-bagel
@@
 Proof.
-(i). The bag $V_{t}$ is unique since $G$ is 4-connected. It follows from the degree-condition that $G[V_{t}]$ must be 2-connected.
-(ii). The first part is immediate. For the ‘moreover’-part we use that $e$ is not dangling by the (Crossing Lemma). (4.1). ∎
+˜(i). The bag $V_{t}$ is unique since $G$ is 4-connected. It follows from the degree-condition that $G[V_{t}]$ must be 2-connected.
+˜(ii). The first part is immediate. For the ‘moreover’-part we use that $e$ is not dangling by the (Crossing Lemma). (4.1). ∎
 Corollary 8.3.
 Let $G$ be a 4-connected graph with two tetra-separations $(A,B)$ and $(C,D)$ that cross with all links of size one. Then all four corners are non-empty.∎
-The bag $V_{t}$ and node $t$ in Lemma 8.2 are said to represent the $L$-link in the $K$-block-decomposition and in the $K$-block-path.
+The bag $V_{t}$ and node $t$ in Lemma˜8.2 are said to represent the $L$-link in the $K$-block-decomposition and in the $K$-block-path.
 For each corner $K$ let $\mathcal{T}^{K}=(T^{K},\mathcal{V}^{K})$ denote the $K$-block-decomposition.
@@
   • $\cdot$
-    we join the nodes $t$ and $t^{\prime}$ by an edge $)𝑜𝐿{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{o(L)}}$ in $O$ if their bags $G[V_{t}]$ and $G[V_{t^{\prime}}]$ are 2-connected;
+[…]
   • $\cdot$
-    we otherwise identify the nodes $t$ and $t^{\prime}$ to a single node $)𝑜𝐿{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{o(L)}}$ of $O$.
+[…]
 We say that $o(L)$ represents the $L$-link in $O$.
 For each node $t$ of $O$ that is not of the form $o(L)$, we define $G_{t}:=G[V^{K}_{t}]$ where $K$ is the unique corner with $t\in T^{K}$.
-For each node $t$ of $O$ with $t=o(L)$ for some link $L$, we let $G_{t}$ be the $K_{2}$-subgraph of $G$ from Lemma 8.2 (ii) that represents the $L$-link in the $K$-block-decompositions for the two corners $K$ adjacent to $L$.
+For each node $t$ of $O$ with $t=o(L)$ for some link $L$, we let $G_{t}$ be the $K_{2}$-subgraph of $G$ from Lemma˜8.2 ˜(ii) that represents the $L$-link in the $K$-block-decompositions for the two corners $K$ adjacent to $L$.
 The pair $(O,(G_{t})_{t\in V(O)})$ is the block-bagel induced by $(A,B)$ and $(C,D)$.
 It really is a block-bagel of $G-Z^{\prime}$.
-The following setting is supported by Figure 8.1.
+The following setting is supported by Figure˜8.1.
 Setting 8.4 (Block-bagel setting).
@@
 Write $O=t(0),t(1),\ldots,t(m),t(0)$ where indices are treated as elements of $\mathbb{Z}_{m+1}$.
-For every node $t\in V(O)$, we let $U_{t}^{\prime}:=V(G_{t})\cup Z^{\prime}$ and $W_{t}^{\prime}:=Z^{\prime}\cup\bigcup_{s\in V(O-t)}V(G_{s})$, and we let $(U_{t},W_{t})$ be the right-left-shift of $(U_{t}^{\prime},W_{t}^{\prime})$.
+For every node $t\in V(O)$, we let $U_{t}^{\prime}:=V(G_{t})\cup Z^{\prime}$ and $W_{t}^{\prime}:=Z^{\prime}\cup\bigcup_{s\in V(O-t)}V(G_{s})$, and we let $(U_{t},W_{t})$ be the right-left-reduction of $(U_{t}^{\prime},W_{t}^{\prime})$.
 Note that $(U_{t}^{\prime},W_{t}^{\prime})$ is a mixed-4-separation+.
@@
 The vertex $x^{\prime}$ is red, the vertex $y^{\prime}$ is green.
-Right: The right-left-shift $(U,W)$ of $(U^{\prime},W^{\prime})$.
+Right: The right-left-reduction $(U,W)$ of $(U^{\prime},W^{\prime})$.
 Here $x$ is the red edge and $y$ is the green vertex.
@@
 Otherwise, the $K$-shrimp is a path.
-The corner $K$ contains a vertex $v$ by Corollary 8.3.
+The corner $K$ contains a vertex $v$ by Corollary˜8.3.
 Then $v$ is the adhesion-vertex of a helpful edge that belongs to the corner $K$.
@@
 Otherwise the adhesion-vertex has degree $\leqslant 3$, contradicting 4-connectivity. ∎
-The following two lemmata are supported by Figure 8.2.
+The following two lemmata are supported by Figure˜8.2.
 Lemma 8.8.
@@
 ∎
-Figure 8.2. The neighbours $w_{i}$ found by Lemma 8.8 are blue. The paths found by Lemma 8.9 are red.
+Figure 8.2. The neighbours $w_{i}$ found by Lemma˜8.8 are blue. The paths found by Lemma˜8.9 are red.
 Lemma 8.9 (Bag-Path Lemma).
@@
 Let $u\in V(G_{r})$ and $v\in V(G_{s})$.
-Then there is a $u$$v$ path in $\bigcup_{t\in V(P)}G_{t}$ that internally avoids the adhesion-vertices of the edges not in $P$.
+Then there is a uuvv path in $\bigcup_{t\in V(P)}G_{t}$ that internally avoids the adhesion-vertices of the edges not in $P$.
 Proof.
-Let $u_{i}$ denote the adhesion-vertex of $t(i)\,t(i+1)$.
-We assume without loss of generality that $P=t(1)\,t(2)\ldots t(k)$.
-For each $i\in\{1,2,\ldots,k\}$, we construct a path $P_{i}$ in $G_{t(i)}$, as follows.
-For $i=1$ we claim that $P_{i}$ can be chosen to be a $u$$u_{1}$ path with interior in $G_{t(1)}-u_{0}$.
-This is clearly possible if the bag $G_{t(1)}$ is a $K_{2}$.
-Otherwise $G_{t(1)}$ is 2-connected.
-Then we find two independent $u$$u_{1}$ paths in $G_{t(1)}$, and at least one of them internally avoids $u_{0}$.
-Similarly, if $i=k$ then we let $P_{i}$ be a $u_{k-1}$$v$ path with interior in $G_{t(k)}-u_{k}$.
-Otherwise, if $2\leqslant i\leqslant k$, then we let $P_{i}$ be a $u_{i-1}$$u_{i}$ path in $G_{t(i)}$.
-Then the concatenation $P:=P_{1}P_{2}\ldots P_{k}$ is a desired $u$$v$ path.
-∎
+[…]
 Lemma 8.11 (Two-Fan Lemma).
-Assume the  (Block-bagel setting).
-(8.4).
-Let $u$ and $v$ be the two adhesion-vertices in a bag $G_{t}$.
-Assume that either $t$ is helpful or $u$ is the adhesion-vertex of a helpful edge.
-Then there exist two independent paths $P_{1},P_{2}$ with all internal vertices in $\mathring{G}_{t}$ such that $P_{1}$ is a $u$$z_{1}^{\prime}$ path and $P_{2}$ is a $u$$v$ path.
+[…]
 Proof.
-If $u$ is the adhesion-vertex of a helpful edge, then $uz_{1}^{\prime}\in E(G)$ by Lemma 8.7, and $uv\in E(G)$ since $G_{t}$ is a $K_{2}$.
+If $u$ is the adhesion-vertex of a helpful edge, then $uz_{1}^{\prime}\in E(G)$ by Lemma˜8.7, and $uv\in E(G)$ since $G_{t}$ is a $K_{2}$.
 Then we let $P_{1}:=uz_{1}^{\prime}$ and $P_{2}:=uv$.
 Otherwise, $t$ is helpful.
-By Lemma 8.8, $z^{\prime}_{1}$ has a neighbour $w$ in $\mathring{G}_{t}$.
-Since $G_{t}$ is 2-connected, it contains two independent paths $Q_{1},Q_{2}$ such that $Q_{1}$ is a $u$$w$ path and $Q_{2}$ is a $u$$v$ path.
+By Lemma˜8.8, $z^{\prime}_{1}$ has a neighbour $w$ in $\mathring{G}_{t}$.
+Since $G_{t}$ is 2-connected, it contains two independent paths $Q_{1},Q_{2}$ such that $Q_{1}$ is a uuww path and $Q_{2}$ is a uuvv path.
 Then we obtain $P_{1}$ from $Q_{1}$ by appending $z^{\prime}_{1}$, and we let $P_{2}:=Q_{2}$.
 ∎
-The following lemma is supported by Figure 8.3.
+The following lemma is supported by Figure˜8.3.
 Lemma 8.12 (Two-Paths Lemma).
-Assume the  (Block-bagel setting).
-(8.4).
-Let $P$ be an $r$$s$ path in $O$.
-Assume that either the first edge $e$ or the second node of $P$ is helpful.
-Assume that $P$ contains at least two helpful elements distinct from $r$.
-Let $u$ be the adhesion-vertex of the edge $e$.
-Then there exist two independent $u$$z_{1}^{\prime}$ paths in $G$ with all interior vertices in $\bigcup_{t\in V(\mathring{P})}G_{t}\cup\mathring{G}_{s}$.
+[…]
 Proof.
 Without loss of generality, $P=t(0)\,t(1)\ldots t(k)$ with $r=t(0)$ and $s=t(k)$.
-By assumption, either the edge $e=t(0)\,t(1)$ or the second node $t(1)$ is helpful.
-Let $u_{i}$ denote the adhesion-vertex of $t(i)\,t(i+1)$.
+By assumption, either the edge e=t(0)t(1)e=t(0)\,t(1) or the second node $t(1)$ is helpful.
+Let $u_{i}$ denote the adhesion-vertex of t(i)t(i+1)t(i)\,t(i+1).
 In particular, $u=u_{0}$.
@@
 Note that $G_{t}$ is bad if and only if either $G_{t}$ is a $K_{2}$ or $G_{t}$ is a triangle and both neighbouring bags are 2-connected.
-The following setting is supported by Figure 8.4.
+The following setting is supported by Figure˜8.4.
 Setting 8.14 (Block-bagel bag setting).
@@
 Let $U^{\prime}:=U_{t(0)}^{\prime}$ and $W^{\prime}:=W_{t(0)}^{\prime}$.
-Let $x^{\prime}$ be the adhesion-vertex of $t(-1)\,t(0)$ and let $y^{\prime}$ be the adhesion-vertex of $t(0)\,t(1)$.
+Let $x^{\prime}$ be the adhesion-vertex of t(1)t(0)t(-1)\,t(0) and let $y^{\prime}$ be the adhesion-vertex of t(0)t(1)t(0)\,t(1).
 Observe that $S(U^{\prime},W^{\prime})=\{x^{\prime},y^{\prime},z_{1}^{\prime},z_{2}^{\prime}\}$.
-Let $(U,W)$ be the right-left-shift of $(U^{\prime},W^{\prime})$.
+Let $(U,W)$ be the right-left-reduction of $(U^{\prime},W^{\prime})$.
 Let $x,y,z_{1},z_{2}\in S(U,W)$ correspond to $x^{\prime},y^{\prime},z^{\prime}_{1},z^{\prime}_{2}$, respectively.
@@
 Proof.
-We combine the (Helpful Lemma). (8.6) with Lemma 8.7 and Lemma 8.8, and we use 5.9 for $W^{\prime}\smallsetminus U^{\prime}\subseteq W\smallsetminus U$. ∎
+We combine the (Helpful Lemma). (8.6) with Lemma˜8.7 and Lemma˜8.8, and we use ˜5.9 for $W^{\prime}\smallsetminus U^{\prime}\subseteq W\smallsetminus U$. ∎
 Lemma 8.17.
@@
 Proof.
-(i) $\Rightarrow$ (ii).
+˜(i) $\Rightarrow$ ˜(ii).
 Suppose first that $G_{t(0)}$ is good.
-By Corollary 8.3, each of the four corners contains a vertex.
+By Corollary˜8.3, each of the four corners contains a vertex.
 Since $G_{t(0)}$ is 2-connected, it is included in $A\cap C$, say, so every corner except $AC$ has its vertex contained in $W^{\prime}\smallsetminus U^{\prime}$.
@@
 If the bag $G_{t(0)}$ is not a triangle, then it has $\geqslant 4$ vertices, so $|U^{\prime}\smallsetminus W^{\prime}|\geqslant 2$.
-In this case, $(U,W)$ is a tetra-separation by 5.10.
+In this case, $(U,W)$ is a tetra-separation by ˜5.10.
 Otherwise, $G_{t(0)}$ is a triangle.
@@
 Hence $|U^{\prime}\smallsetminus W^{\prime}|=1$ and the vertex $y^{\prime}$ has exactly one neighbour in $W^{\prime}\smallsetminus U^{\prime}$.
-Again by 5.10, $(U,W)$ is a tetra-separation.
+Again by ˜5.10, $(U,W)$ is a tetra-separation.
 Suppose now that all three bags $G_{t(-1)},G_{t(0)},G_{t(1)}$ are $K_{2}$’s.
@@
 Both $x^{\prime}$ and $y^{\prime}$ have exactly one neighbour in $W^{\prime}\smallsetminus U^{\prime}$ since both $G_{t(-1)}$ and $G_{t(1)}$ are $K_{2}$’s.
-Hence $(U,W)$ is a tetra-separation by 5.10.
-$\neg$ (i) $\Rightarrow$ $\neg$ (ii).
-As (i) fails, $G_{t(0)}$ either is a $K_{2}$ neighboured by at least one 2-connected bag, or $G_{t(0)}$ is a triangle neighboured by 2-connected bags.
+Hence $(U,W)$ is a tetra-separation by ˜5.10.
+$\neg$ ˜(i) $\Rightarrow$ $\neg$ ˜(ii).
+As ˜(i) fails, $G_{t(0)}$ either is a $K_{2}$ neighboured by at least one 2-connected bag, or $G_{t(0)}$ is a triangle neighboured by 2-connected bags.
 Suppose first that $G_{t(0)}$ is a $K_{2}$ neighboured by a 2-connected bag, say $G_{t(1)}$.
 Then $y^{\prime}$ has two neighbours in $G_{t(1)}$, which are also contained in $W^{\prime}\smallsetminus U^{\prime}$.
-The two vertices $z^{\prime}_{i}$ in the centre have $\geqslant 3$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$ by Lemma 8.16.
+The two vertices $z^{\prime}_{i}$ in the centre have $\geqslant 3$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$ by Lemma˜8.16.
 So only $x^{\prime}$ might have $\leqslant 1$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$.
-But $|U^{\prime}\smallsetminus W^{\prime}|=0$, so $(U,W)$ is no tetra-separation by 5.10.
+But $|U^{\prime}\smallsetminus W^{\prime}|=0$, so $(U,W)$ is no tetra-separation by ˜5.10.
 Suppose now that $G_{t(0)}$ is a triangle with both neighbouring bags 2-connected.
 Then, as above, all vertices in $S(U^{\prime},W^{\prime})$ have $\geqslant 2$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$.
-But $|U^{\prime}\smallsetminus W^{\prime}|=1$, so $(U,W)$ is no tetra-separation by 5.10.
+But $|U^{\prime}\smallsetminus W^{\prime}|=1$, so $(U,W)$ is no tetra-separation by ˜5.10.
 ∎
 Observation 8.18.
-Assume the  (Block-bagel bag setting). (8.14). Let $t(i)\,t(i+1)$ be the edge of $O$ that has $y^{W}$ as adhesion-vertex. Then either the edge $t(i)\,t(i+1)$ or the node $t(i+1)$ is helpful.∎
+Assume the  (Block-bagel bag setting).
+(8.14).
+Let t(i)t(i+1)t(i)\,t(i+1) be the edge of $O$ that has $y^{W}$ as adhesion-vertex.
+Then either the edge t(i)t(i+1)t(i)\,t(i+1) or the node $t(i+1)$ is helpful.∎
 Lemma 8.19.
@@
 For each helpful $h\in V(O)\cup E(O)$ we construct a $z^{\prime}_{1}$$z^{\prime}_{2}$ path $P_{h}$, as follows.
-If $h$ is an edge, then its adhesion-vertex send edges to both $z^{\prime}_{i}$ by Lemma 8.7, and we let $P_{h}$ be the path defined by these two edges; see Figure 8.5.
+If $h$ is an edge, then its adhesion-vertex send edges to both $z^{\prime}_{i}$ by Lemma˜8.7, and we let $P_{h}$ be the path defined by these two edges; see Figure˜8.5.
 If $h$ is a helpful node, then we use the (Bag-Path Lemma).
@@
 At most one of them meets the bag $G_{t(0)}$.
-By Lemma 8.16, both vertices $z^{\prime}_{i}$ in the centre are contained in $W$.
+By Lemma˜8.16, both vertices $z^{\prime}_{i}$ in the centre are contained in $W$.
 Hence at least three of the paths $P_{h}$ are independent $z^{\prime}_{1}$$z^{\prime}_{2}$ paths in $G[W]-\{x^{\prime},y^{\prime}\}-z^{\prime}_{1}z^{\prime}_{2}$.
 ∎
-Figure 8.5. The red paths are constructed in Lemma 8.19.
-The following lemma is supported by Figure 8.6.
+Figure 8.5. The red paths are constructed in Lemma˜8.19.
+The following lemma is supported by Figure˜8.6.
 Lemma 8.20.
@@
 Since $t(0)$ is helpful, (Helpful Lemma). (8.6) implies that
-(6) $\geqslant 3$ elements of $O-t(0)$ are helpful.
-Let $f=t(j)\,t(j+1)$ be the edge on $O$ that has $y^{W}$ as adhesion-vertex.
-Then $f$ or $t(j+1)$ is helpful by 8.18.
+(6) $\geqslant 3$ elements of Ot(0)O-t(0) are helpful.
+Let f=t(j)t(j+1)f=t(j)\,t(j+1) be the edge on $O$ that has $y^{W}$ as adhesion-vertex.
+Then $f$ or $t(j+1)$ is helpful by ˜8.18.
 If $t(j)$ is helpful, then $j=0$, and $j=1$ otherwise.
 Let $P$ be the path in $O$ that starts with $t(j),t(j+1)$ and has two helpful elements distinct from $t(j)$ and is minimal with these properties.
-The path $P$ is well-defined since $O-t(0)$ has $\geqslant 3$ helpful elements by 6.
+The path $P$ is well-defined since Ot(0)O-t(0) has $\geqslant 3$ helpful elements by ˜6.
 Moreover, the end $s$ of $P$ is distinct from $t(0)$.
@@
 By construction, the paths $Q_{1}$ and $Q_{2}$ are contained in $G[W]$.
-Now let $e=t(i)\,t(i-1)$ be the edge on $O$ that has $x^{W}$ as adhesion-vertex.
-Then $e$ or $t(i-1)$ is helpful by 8.18.
+Now let e=t(i)t(i1)e=t(i)\,t(i-1) be the edge on $O$ that has $x^{W}$ as adhesion-vertex.
+Then $e$ or t(i1)t(i-1) is helpful by ˜8.18.
 And if $t(i)$ is helpful, then $i=0$.
@@
 If the edge $e$ is helpful, then it is the first helpful element on $O$ following $t(0)$ in the direction of $e$.
-In particular, $e$ does not lie on $P$ by 6.
+In particular, $e$ does not lie on $P$ by ˜6.
 Since the bag $G_{t(i-1)}$ is a $K_{2}$, it has empty interior.
-So $Q_{3}=x^{W}z^{\prime}_{1}$ avoids $\mathring{G}_{s}$, even if $s=t(i-1)$, so $Q_{3}$ is independent from $Q_{1}$ and $Q_{2}$.
-Otherwise $t(i-1)$ is helpful.
-Then $t(i-1)$ is the first helpful element on $O$ following $t(0)$ in the direction of $e$.
-In particular, both edges on $O$ incident to $t(i-1)$ are unhelpful.
-So $t(i-1)$ does not lie on $P$ by 6.
+So $Q_{3}=x^{W}z^{\prime}_{1}$ avoids $\mathring{G}_{s}$, even if s=t(i1)s=t(i-1), so $Q_{3}$ is independent from $Q_{1}$ and $Q_{2}$.
+Otherwise t(i1)t(i-1) is helpful.
+Then t(i1)t(i-1) is the first helpful element on $O$ following $t(0)$ in the direction of $e$.
+In particular, both edges on $O$ incident to t(i1)t(i-1) are unhelpful.
+So t(i1)t(i-1) does not lie on $P$ by ˜6.
 Hence $Q_{3}$ is independent from $Q_{1}$ and $Q_{2}$.
 ∎
-Figure 8.6. The paths constructed in Lemma 8.20. $P_{1}$ is yellow, $P_{2}$ is blue, $P_{3}$ is green.
+Figure 8.6. The paths constructed in Lemma˜8.20. $P_{1}$ is yellow, $P_{2}$ is blue, $P_{3}$ is green.
 Lemma 8.21.
-Let $G$ be a graph, let $u\in U\subseteq V(G)$ and $v\in V\subseteq V(G)$. Every $u$$v$ path in $G$ contains a $U$$V$ path.∎
+Let $G$ be a graph, let $u\in U\subseteq V(G)$ and $v\in V\subseteq V(G)$. Every uuvv path in $G$ contains a UUVV path.∎
 In the following, we characterise when $(U,W)$ is a totally-nested tetra-separation.
@@
 Without loss of generality, assume that the interior of $G_{t(0)}$ is contained in the corner $AC$.
-Let $v$ be a vertex in the corner $BD$, which exists by Corollary 8.3.
+Let $v$ be a vertex in the corner $BD$, which exists by Corollary˜8.3.
 Let $t(i)$ be a node of $O$ with $v\in G_{t(i)}$.
@@
 Otherwise $w=z^{\prime}_{1}$, say.
-Then $w$ has a neighbour in $W\smallsetminus Z^{\prime}$ by Lemma 8.19, which we already know is linked to $v$ in $G[W\smallsetminus U]$.
+Then $w$ has a neighbour in $W\smallsetminus Z^{\prime}$ by Lemma˜8.19, which we already know is linked to $v$ in $G[W\smallsetminus U]$.
 Hence $(U,W)$ is half-connected.
@@
 Since the bag $G_{t(0)}$ is 2-connected, it contains two independent $x^{\prime}$$y^{\prime}$ paths.
-By Lemma 8.21, these paths contain two independent $\hat{x}$$\hat{y}$ paths which are also included in the bag $G_{t(0)}$.
+By Lemma˜8.21, these paths contain two independent $\hat{x}$$\hat{y}$ paths which are also included in the bag $G_{t(0)}$.
 We find a third $\hat{x}$$\hat{y}$ path $G$ that avoids both the bag $G_{t(0)}$ and the centre, by the (One-Path Lemma).
@@
 Case $\pi=\{z_{1},z_{2}\}$.
-We find three independent $z^{\prime}_{1}$$z^{\prime}_{2}$ paths in $G-x-y$ by Lemma 8.19.
-These contain three independent $\hat{z}_{1}$$\hat{z}_{2}$ paths by Lemma 8.21.
+We find three independent $z^{\prime}_{1}$$z^{\prime}_{2}$ paths in GxyG-x-y by Lemma˜8.19.
+These contain three independent $\hat{z}_{1}$$\hat{z}_{2}$ paths by Lemma˜8.21.
 So we are done by (6.4.3).
@@
 (8.11).
-This path contains a $\hat{x}$$\hat{z}_{1}$ path $P_{1}$, by Lemma 8.21.
-Next, we find two independent $x^{W}$$z^{\prime}_{1}$ paths $P_{2}$ and $P_{3}$ in $G$ that internally avoid $G_{t(0)}$, by Lemma 8.20.
-Both paths $P_{2}$ and $P_{3}$ are $\hat{x}$$\hat{z}_{1}$ paths by Lemma 8.16.
+This path contains a $\hat{x}$$\hat{z}_{1}$ path $P_{1}$, by Lemma˜8.21.
+Next, we find two independent $x^{W}$$z^{\prime}_{1}$ paths $P_{2}$ and $P_{3}$ in $G$ that internally avoid $G_{t(0)}$, by Lemma˜8.20.
+Both paths $P_{2}$ and $P_{3}$ are $\hat{x}$$\hat{z}_{1}$ paths by Lemma˜8.16.
 Hence the $P_{i}$ witness (6.4.3).
@@
 We claim that $z^{\prime}_{1}=z_{1}$.
-On the one hand, $z^{\prime}_{1}$ has $\geqslant 3$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$ by Lemma 8.16, so $z^{\prime}_{1}$ remains in the separator when taking the right-shift $(U^{\prime\prime},W^{\prime\prime})$ of $(U^{\prime},W^{\prime})$.
-However, $y^{\prime}$ becomes the edge $y$ of the $K_{2}$-bag $G_{t(1)}$ in the separator $S(U^{\prime\prime},W^{\prime\prime})$ by 8.15.
+On the one hand, $z^{\prime}_{1}$ has $\geqslant 3$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$ by Lemma˜8.16, so $z^{\prime}_{1}$ remains in the separator when taking the right-reduction $(U^{\prime\prime},W^{\prime\prime})$ of $(U^{\prime},W^{\prime})$.
+However, $y^{\prime}$ becomes the edge $y$ of the $K_{2}$-bag $G_{t(1)}$ in the separator $S(U^{\prime\prime},W^{\prime\prime})$ by ˜8.15.
 Hence $z^{\prime}_{1}$ has two neighbours $y^{\prime}$ and $v$ in $U^{\prime\prime}\smallsetminus W^{\prime\prime}$.
-Thus $z^{\prime}_{1}$ also remains in the separator when taking the left-shift $(U,W)$ of $(U^{\prime\prime},W^{\prime\prime})$.
+Thus $z^{\prime}_{1}$ also remains in the separator when taking the left-reduction $(U,W)$ of $(U^{\prime\prime},W^{\prime\prime})$.
 In particular, $z^{\prime}_{1}=z_{1}$.
@@
 We will find five independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G$ for (6.7.2).
-There exist two disjoint $\{z^{\prime}_{1},z^{\prime}_{2}\}$$\mathring{G}_{t(0)}$ paths $P_{1},P_{2}$ by Lemma 8.8.
+There exist two disjoint $\{z^{\prime}_{1},z^{\prime}_{2}\}$$\mathring{G}_{t(0)}$ paths $P_{1},P_{2}$ by Lemma˜8.8.
 Since $G_{t(0)}$ is 2-connected, we can extend the paths $P_{1},P_{2}$ to two disjoint $\{z^{\prime}_{1},z^{\prime}_{2}\}$$\{x^{\prime},y^{\prime}\}$ paths with interior in $\mathring{G}_{t(0)}$.
-We find three independent $\{x_{1}^{W},x_{2}^{W}\}$$\{z_{1}^{\prime},z_{2}^{\prime}\}$ paths in $G[W]$ by Lemma 8.20.
-The five paths are independent, and they contain five $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths by Lemma 8.21.
+We find three independent $\{x_{1}^{W},x_{2}^{W}\}$$\{z_{1}^{\prime},z_{2}^{\prime}\}$ paths in $G[W]$ by Lemma˜8.20.
+The five paths are independent, and they contain five $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths by Lemma˜8.21.
 Case $\pi_{1}=\{x,z_{1}\}$ and $\pi_{2}=\{y,z_{2}\}$.
@@
 There are two independent $x^{\prime}$$y^{\prime}$ paths in $G_{t(0)}$ by 2-connectivity.
-We find three independent $z^{\prime}_{1}$$z^{\prime}_{2}$ paths in $G[W]-\{x^{\prime},y^{\prime}\}$ by Lemma 8.19.
+We find three independent $z^{\prime}_{1}$$z^{\prime}_{2}$ paths in $G[W]-\{x^{\prime},y^{\prime}\}$ by Lemma˜8.19.
 The latter three paths are internally disjoint from the former two.
-By Lemma 8.21, the five paths contain five independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths.
+By Lemma˜8.21, the five paths contain five independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths.
 All remaining cases are symmetric to one of the two cases above. ∎
@@
 Assume $\pi=\{x,z_{1}\}$.
-The vertex $z^{\prime}_{1}$ in the centre has $\geqslant 3$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$ by Lemma 8.16.
+The vertex $z^{\prime}_{1}$ in the centre has $\geqslant 3$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$ by Lemma˜8.16.
 Hence $z^{W}_{1}=z^{\prime}_{1}$.
-Since the $\geqslant 3$ neighbours also lie in $W\smallsetminus U$ by 5.9, we further get that $z_{1}$ must be the edge in $\pi$.
+Since the $\geqslant 3$ neighbours also lie in $W\smallsetminus U$ by ˜5.9, we further get that $z_{1}$ must be the edge in $\pi$.
 Since $z_{1}$ is an edge, $x$ is a vertex which has exactly one neighbour in $G[W]-z_{1}^{W}$.
 Hence $G_{t(-1)}$ is a $K_{2}$.
-But then $x$ is the edge of $G_{t(-1)}$ by 8.15, a contradiction.
+But then $x$ is the edge of $G_{t(-1)}$ by ˜8.15, a contradiction.
 Assume $\pi=\{x,y\}$.
 Say $x$ is a vertex and $y$ is an edge.
-Then $x=x^{\prime}$, so $G_{t(-1)}$ is 2-connected by 8.15.
+Then $x=x^{\prime}$, so $G_{t(-1)}$ is 2-connected by ˜8.15.
 Hence $x$ has two neighbours $u,v$ in $G_{t(-1)}$.
@@
 Proof.
-Let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of $S(U,W)$. Neither $\pi_{i}$ is $W$-weird by Lemma 8.25. In particular, $\{\pi_{1},\pi_{2}\}$ is 2-potter-linked.
+Let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of $S(U,W)$. Neither $\pi_{i}$ is $W$-weird by Lemma˜8.25. In particular, $\{\pi_{1},\pi_{2}\}$ is 2-potter-linked.
 To show that $\{\pi_{1},\pi_{2}\}$ is 1-potter-linked, it suffices to find three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[W]$.
-If one of the $\pi_{i}$ is of the form $\{x,y\}$, then we find three such paths by Lemma 8.20 and Lemma 8.21.
+If one of the $\pi_{i}$ is of the form $\{x,y\}$, then we find three such paths by Lemma˜8.20 and Lemma˜8.21.
 Otherwise $z_{1}\in\pi_{1}$ and $z_{2}\in\pi_{2}$, say.
-Then we find three desired paths by Lemma 8.19 and Lemma 8.21.
+Then we find three desired paths by Lemma˜8.19 and Lemma˜8.21.
 ∎
@@
 (8.14) with $t=t(0)$.
-Then $(U_{t},W_{t})$ is a tetra-separation by Lemma 8.17.
-It suffices to show that $(U_{t},W_{t})$ is externally 5-connected, by 6.16.
-This follows from Lemma 8.22 (half-connected), Lemma 8.23 (3-linked), Lemma 8.24 (0-potter-linked) and Lemma 8.26 (1-potter-linked and 2-potter-linked).
+Then $(U_{t},W_{t})$ is a tetra-separation by Lemma˜8.17.
+It suffices to show that $(U_{t},W_{t})$ is externally 5-connected, by ˜6.16.
+This follows from Lemma˜8.22 (half-connected), Lemma˜8.23 (3-linked), Lemma˜8.24 (0-potter-linked) and Lemma˜8.26 (1-potter-linked and 2-potter-linked).
 ∎
@@
 […]
-The following two lemmata are supported by Figure 8.7.
+The following two lemmata are supported by Figure˜8.7.
 Lemma 8.30 (Link-Moving Lemma for $K_{2}$-bags).
@@
 (8.4).
-Let $P$ and $Q$ be the two paths that are the components of $O-o(C)-o(D)$.
+Let $P$ and $Q$ be the two paths that are the components of Oo(C)o(D)O-o(C)-o(D).
 Assume $o(A)\in P$ and $o(B)\in Q$.
@@
 Define $d$ similarly.
-Let $s$ be a node in the same component of $P-o(A)$ as $c$ such that $G_{s}=K_{2}$.
+Let $s$ be a node in the same component of Po(A)P-o(A) as $c$ such that $G_{s}=K_{2}$.
 Let $P^{\prime}:=dP\mathring{s}$ and $Q^{\prime}:=\mathring{s}Pc\cup Q$.
@@
 (4.1).
-Both endvertices of $e$ lie in the corner for $AC$ by Lemma 8.2 (i).
+Both endvertices of $e$ lie in the corner for $AC$ by Lemma˜8.2 (i).
 Hence $d$ and $e$ share no ends.
@@
 Indeed, since $ud$ is not a dangling edge, $u$ is contained in the corner $AD$ or the $A$-link with respect to $(A,B),(C,D)$.
-By Lemma 8.2 (i), no endvertex of $e$ lies in the $A$-link with respect to $(A,B),(C,D)$.
+By Lemma˜8.2 (i), no endvertex of $e$ lies in the $A$-link with respect to $(A,B),(C,D)$.
 So, we have $u\in A^{\prime}\smallsetminus B^{\prime}$.
@@
 Hence, there are at least two helpful elements that belong to $(ss^{\prime})\cup T_{AD}$, where $T_{AD}$ is the $AD$-block-path.
-By Lemmas 8.7 and 8.8, we are done.
+By Lemmas˜8.7 and 8.8, we are done.
 We also have (iii) since there is a vertex in $A\cap(D\smallsetminus C)\subseteq A^{\prime}\smallsetminus B^{\prime}$ that is separated from every vertex in $B\smallsetminus A\subseteq B^{\prime}\smallsetminus A^{\prime}$ by $(A^{\prime},B^{\prime})$.
@@
 (8.4).
-Let $P$ and $Q$ be the two paths that are the components of $O-o(C)-o(D)$.
+Let $P$ and $Q$ be the two paths that are the components of Oo(C)o(D)O-o(C)-o(D).
 Assume $o(A)\in P$ and $o(B)\in Q$.
@@
 Proof.
-If the $A$-link or the $C$-link contains a vertex, then this vertex is contained in a 2-connected bag $G_{t}$ with $t\in V(T^{\prime})$, by Lemma 8.2.
+If the $A$-link or the $C$-link contains a vertex, then this vertex is contained in a 2-connected bag $G_{t}$ with $t\in V(T^{\prime})$, by Lemma˜8.2.
 Otherwise both the $A$-link and the $C$-link consist of edges.
 Then $G_{o(A)}$ and $G_{o(C)}$ are $K_{2}$-bags.
-Each of the two bags $G_{o(A)},G_{o(C)}$ has exactly one endvertex $v_{A},v_{C}$ in the corner for $AC$ by Lemma 8.2, respectively.
+Each of the two bags $G_{o(A)},G_{o(C)}$ has exactly one endvertex $v_{A},v_{C}$ in the corner for $AC$ by Lemma˜8.2, respectively.
 If $T^{\prime}$ is empty, then $v_{A}=v_{C}$, and $A\cap C=\{v_{A}\}\cup Z^{\prime}$ has size three.
@@
 Let $T$ denote the $A^{\prime}C^{\prime}$-block-path, and let $T^{\prime}:=T-o(A^{\prime})-o(C^{\prime})$.
-By Lemma 8.32, $T^{\prime}$ is non-empty.
+By Lemma˜8.32, $T^{\prime}$ is non-empty.
 Let $\mathcal{O}^{\prime}=\mathcal{O}$ be the block-bagel induced by $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$.
@@
   • $\cdot$
-[…]
+    $S(A^{\prime\prime},B^{\prime\prime})=\{e\}\cup Z^{\prime}\cup\text{($D^{\prime}$-link)}$, and
   • $\cdot$
@@
   • $\cdot$
-[…]
+    $S(C^{\prime\prime},D^{\prime\prime})=\{e\}\cup Z^{\prime}\cup\text{($B^{\prime}$-link)}$, and
   • $\cdot$
@@
 Assume for contradiction that $\sigma(\mathcal{O})$ is interlaced by a totally-nested tetra-separation $(E,F)$ of $G$.
-By Lemma 8.33, there is a 2-connected bag $G_{t}$ of $\mathcal{O}$ with $E\subseteq U_{t}^{\prime}$, say.
-By Lemma 5.7, we have $(E,F)\leqslant(U_{t}^{\prime\prime},W_{t}^{\prime\prime})$ where $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$ is the right-shift of $(U^{\prime}_{t},W^{\prime}_{t})$.
+By Lemma˜8.33, there is a 2-connected bag $G_{t}$ of $\mathcal{O}$ with $E\subseteq U_{t}^{\prime}$, say.
+By Lemma˜5.7, we have $(E,F)\leqslant(U_{t}^{\prime\prime},W_{t}^{\prime\prime})$ where $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$ is the right-reduction of $(U^{\prime}_{t},W^{\prime}_{t})$.
 Assume first that $G_{t}$ is good.
@@
 Since $(E,F)$ interlaces $\sigma(\mathcal{O})$, we have $(U_{t},W_{t})<(E,F)$ or $(U_{t},W_{t})<(F,E)$.
-We cannot have $(U_{t},W_{t})<(F,E)$ since otherwise $\emptyset\neq E\smallsetminus F\subseteq(W_{t}\smallsetminus U_{t})\cap(U_{t}^{\prime\prime}\smallsetminus W_{t}^{\prime\prime})$ contradicts that $U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t}\subseteq U_{t}\smallsetminus W_{t}$ by 5.2.
+We cannot have $(U_{t},W_{t})<(F,E)$ since otherwise $\emptyset\neq E\smallsetminus F\subseteq(W_{t}\smallsetminus U_{t})\cap(U_{t}^{\prime\prime}\smallsetminus W_{t}^{\prime\prime})$ contradicts that $U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t}\subseteq U_{t}\smallsetminus W_{t}$ by ˜5.2.
 So $(U_{t},W_{t})<(E,F)\leqslant(U_{t}^{\prime\prime},W_{t}^{\prime\prime})$.
-Since $(U_{t},W_{t})$ is the left-shift of $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, some vertex in $S(E,F)$ must have $\leqslant 1$ neighbour in $E\smallsetminus F$, contradicting the degree-condition.
+Since $(U_{t},W_{t})$ is the left-reduction of $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, some vertex in $S(E,F)$ must have $\leqslant 1$ neighbour in $E\smallsetminus F$, contradicting the degree-condition.
 Otherwise, $G_{t}$ is bad.
@@
 Since $(E,F)\leqslant(U_{t}^{\prime\prime},W_{t}^{\prime\prime})$, we have $E\smallsetminus F\subseteq U_{t}^{\prime\prime}\smallsetminus W_{t}^{\prime\prime}$.
-Hence $|E\smallsetminus F|\leqslant 1$, which by Lemma 4.2 contradicts that $(E,F)$ is a tetra-separation.
+Hence $|E\smallsetminus F|\leqslant 1$, which by Lemma˜4.2 contradicts that $(E,F)$ is a tetra-separation.
 ∎
@@
 Proof.
-(i) $\Rightarrow$ (ii): this is 8.27.
-$\neg$ (i) $\Rightarrow$ $\neg$ (ii).
+˜(i) $\Rightarrow$ ˜(ii): this is ˜8.27.
+$\neg$ ˜(i) $\Rightarrow$ $\neg$ ˜(ii).
 Assume that $G_{t}$ is bad.
-Then $(U_{t},W_{t})$ fails to be a tetra-separation, unless $G_{t}$ and both neighbouring bags are $K_{2}$’s, by 8.27.
+Then $(U_{t},W_{t})$ fails to be a tetra-separation, unless $G_{t}$ and both neighbouring bags are $K_{2}$’s, by ˜8.27.
 In this case, both vertices of $G_{t}$ lie in $U_{t}\smallsetminus W_{t}$, so $(U_{t},W_{t})$ is distinct from all tetra-separations in the tetra-star $\sigma(\mathcal{O})$.
-Therefore, $(U_{t},W_{t})$ interlaces $\sigma(\mathcal{O})$, and so $(U_{t},W_{t})$ is crossed by another tetra-separation, by Lemma 8.34.
+Therefore, $(U_{t},W_{t})$ interlaces $\sigma(\mathcal{O})$, and so $(U_{t},W_{t})$ is crossed by another tetra-separation, by Lemma˜8.34.
 ∎
@@
 Proof.
-We combine Lemmas 2.6.3 and 2.6.4 from [Tridecomp]. ∎
+We combine Lemmas 2.6.3 and 2.6.4 from [CK23]. ∎
 Key Lemma 8.37.
@@
 Let $\mathcal{O}$ be the Block-bagel induced by $(A,B)$ and $(C,D)$, which has centre $Z$.
-By Lemma 8.29, the tetra-star $\sigma(\mathcal{O})$ is a star of totally-nested tetra-separations of $G$.
-By Lemma 8.34, the tetra-star $\sigma(\mathcal{O})$ is a splitting star of $N$.
+By Lemma˜8.29, the tetra-star $\sigma(\mathcal{O})$ is a star of totally-nested tetra-separations of $G$.
+By Lemma˜8.34, the tetra-star $\sigma(\mathcal{O})$ is a splitting star of $N$.
 Since both $(A,B)$ and $(C,D)$ interlace both $\sigma$ and $\sigma(\mathcal{O})$, we obtain $\sigma=\sigma(\mathcal{O})$.
 Hence it remains to compute the torso $\tau$ of $\sigma(\mathcal{O})$.
-The torso $\tau$ is 4-connected by Lemma 8.36.
+The torso $\tau$ is 4-connected by Lemma˜8.36.
 We can obtain $\tau$ from $\mathcal{O}$ by replacing each good bag $G_{t}$ by a $K_{4}$ on the vertex set $Z\cup\{u,v\}$ where $u$ and $v$ are the adhesion-vertices of $G_{t}$, leaving only $K_{2}$-bags and triangle-bags, and contracting some edges in $K_{2}$-bags.
@@
 Let $X$ be a graph.
-We call $X$ a double-wheel with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that $X-u-v$ is a cycle.
-So $u$ and $v$ send edges to all vertices of $X-u-v$, and the edge $uv$ may or may not exist.
-The rim-length of $X$ then means the length of the cycle $X-u-v$.
-We call $X$ a double-wheel of triangles with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that $X-u-v$ has a cycle-decomposition $(O,\mathcal{G})$ with all adhesion-sets of size 1 and into triangles.
+We call $X$ a double-wheel with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that XuvX-u-v is a cycle.
+So $u$ and $v$ send edges to all vertices of XuvX-u-v, and the edge $uv$ may or may not exist.
+The rim-length of $X$ then means the length of the cycle XuvX-u-v.
+We call $X$ a double-wheel of triangles with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that XuvX-u-v has a cycle-decomposition $(O,\mathcal{G})$ with all adhesion-sets of size 1 and into triangles.
 The rim-length of $X$ means the length of the cycle $O$.
@@
 Proof.
-Let $O$ denote the cycle $G-Z$.
+Let $O$ denote the cycle GZG-Z.
 Let $e_{1},f_{1},e_{2},f_{2}$ be the first four edges on $O$.
@@
 Proof.
-Since $G$ is a double-wheel of triangles with rim-length $\geqslant 4$, the graph $G-Z$ has a cycle-decomposition $\mathcal{O}=(O,(G_{t})_{t\in O})$ with $|O|\geqslant 4$ such that all bags are triangles.
-Let $(A,B)$ and $(C,D)$ be crossing 2-separations of $O$.
-Let $\hat{A}:=Z\cup\bigcup_{t\in A}V(G_{t})$ and define $\hat{B},\hat{C},\hat{D}$ similarly.
-Then $(\hat{A},\hat{B})$ and $(\hat{C},\hat{D})$ are tetra-separations of $G$ that cross as desired.
-∎
+[…]
 Key Lemma 8.41.
@@
 (8.6), we have $h(\mathcal{O})\geqslant 4$.
-If $\mathcal{O}$ has a good bag, then $G$ has a totally-nested tetra-separation by 8.27, contradicting that $G$ is 4-angry.
+If $\mathcal{O}$ has a good bag, then $G$ has a totally-nested tetra-separation by ˜8.27, contradicting that $G$ is 4-angry.
 Hence, every bag of $\mathcal{O}$ is bad.
-(ii) $\Rightarrow$ (iii). Follows from Lemma 8.38.
-(iii) $\Rightarrow$ (i). Follows from Lemma 8.39 and Lemma 8.40. ∎
+(ii) $\Rightarrow$ (iii). Follows from Lemma˜8.38.
+(iii) $\Rightarrow$ (i). Follows from Lemma˜8.39 and Lemma˜8.40. ∎
 9. Cycles of 2-connected torsos
-One of the outcomes of Theorem 1 is a cycle of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices.
+One of the outcomes of Theorem˜1 is a cycle of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices.
 To investigate how this outcome arises, we assume that we are given two tetra-separations $(A,B)$ and $(C,D)$ of a 4-connected graph $G$ that cross with empty vertex-centre.
-In order to construct a cycle-decomposition for analysing the cyclic structure that is determined by $(A,B)$ and $(C,D)$, we first use the corner-diagram to divide $G$ into four smaller pieces (Section 9.1).
-We then apply the Tutte-decomposition to all four pieces and merge them to obtain a cycle-decomposition of $G$ (Section 9.2).
-In Section 9.3, we summarise assumptions shared by a fair share of the lemmas to come.
-Section 9.4 provides machinery that exploits the structure of the cycle-decomposition to find independent paths.
-We use this machinery in Section 9.5 to characterise which bags of the cycle-decomposition give rise to totally-nested tetra-separations (9.41).
-Section 9.6 shows that these totally-nested tetra-separations form a splitting star as in Theorem 1, and determines the structure of its torso (9.50).
-Finally, Section 9.7 characterises the special case where additionally all tetra-separations of the graph are crossed.
+In order to construct a cycle-decomposition for analysing the cyclic structure that is determined by $(A,B)$ and $(C,D)$, we first use the corner-diagram to divide $G$ into four smaller pieces (Section˜9.1).
+We then apply the Tutte-decomposition to all four pieces and merge them to obtain a cycle-decomposition of $G$ (Section˜9.2).
+In Section˜9.3, we summarise assumptions shared by a fair share of the lemmas to come.
+Section˜9.4 provides machinery that exploits the structure of the cycle-decomposition to find independent paths.
+We use this machinery in Section˜9.5 to characterise which bags of the cycle-decomposition give rise to totally-nested tetra-separations (˜9.41).
+Section˜9.6 shows that these totally-nested tetra-separations form a splitting star as in Theorem˜1, and determines the structure of its torso (˜9.50).
+Finally, Section˜9.7 characterises the special case where additionally all tetra-separations of the graph are crossed.
 9.1. Bananas
@@
 We obtain the graph $\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{G^{*}}$} from $G$ by adding edges, as follows.
-For each $L$-link $\{x,y\}$ and each adjacent corner $K=LS$ (where $L,S$ are sides), we join the unique vertices in $\hat{x}\cap S$ and $\hat{y}\cap S$ by an edge, which we call the $K$-edge for the $L$-link; see Figure 9.1.
+For each $L$-link $\{x,y\}$ and each adjacent corner $K=LS$ (where $L,S$ are sides), we join the unique vertices in $\hat{x}\cap S$ and $\hat{y}\cap S$ by an edge, which we call the $K$-edge for the $L$-link; see Figure˜9.1.
 We will only use $G^{*}$ when $(A,B)$ and $(C,D)$ are clear from context.
@@
 Proof.
-$\neg$ (i) $\Rightarrow$ $\neg$ (ii).
+$\neg$ ˜(i) $\Rightarrow$ $\neg$ ˜(ii).
 Say $K$ is the corner for $AC$.
@@
 We will show that $AC$ is potter.
-If both endvertices of $e$ lie in the corner for $AC$, then each link adjacent to $AC$ consists of a matching of size two with the endvertices of $e$ as one side of the matching; see Figure 9.2.
+If both endvertices of $e$ lie in the corner for $AC$, then each link adjacent to $AC$ consists of a matching of size two with the endvertices of $e$ as one side of the matching; see Figure˜9.2.
 As $G$ is 4-connected, the corner for $AC$ cannot contain any other vertices besides the endvertices of $e$.
@@
 (4.1).
-$\neg$ (ii) $\Rightarrow$ $\neg$ (i) holds by definition. ∎
-Figure 9.2. The situation in the proof of Lemma 9.2
+$\neg$ ˜(ii) $\Rightarrow$ $\neg$ ˜(i) holds by definition. ∎
+Figure 9.2. The situation in the proof of Lemma˜9.2
 Lemma 9.3.
@@
 […]
-Figure 9.3. The construction of $(\hat{X},\hat{Y})$ in the proof of Lemma 9.3. The extensions of $X$ to $\hat{X}$ and of $Y$ to $\hat{Y}$ are indicated by the dashed lines.
+[…]
 9.2. Tutte-bananas and Tutte-bagels
@@
 A 2-separation $(A,B)$ of $G$ is totally-nested if it is nested with every 2-separation of $G$.
-This is equivalent to saying that $(A,B)$ is half-connected and at least one of $G[A]$ and $G[B]$ is 2-connected [Tridecomp, Corollary 3.3.3].
+This is equivalent to saying that $(A,B)$ is half-connected and at least one of $G[A]$ and $G[B]$ is 2-connected [CK23, Corollary 3.3.3].
 Assume now that $G$ is 2-connected and $N$ denotes its set of totally-nested 2-separations.
@@
 Every torso of $\mathcal{T}$ is 3-connected, a cycle or a $K_{2}$.
-Tutte introduced his decomposition in [TutteCon].
-A modern proof is included in the appendix of [Tridecomp].
+Tutte introduced his decomposition in [T66].
+A modern proof is included in the appendix of [CK23].
 The Tutte-decomposition $\mathcal{T}=(T,\mathcal{V})$ has the following two properties:
@@
 Let $K$ be one of the four corners.
-Recall that the $K$-banana is 2-connected by Lemma 9.3.
+Recall that the $K$-banana is 2-connected by Lemma˜9.3.
 The $K$-Tutte-banana is the Tutte-decomposition of the $K$-banana.
@@
 Say $K$ is the $AC$-corner and $\Lambda$ is the $A$-link.
-(i).
+˜(i).
 Let $u,v$ be the two vertices in $\Lambda$.
@@
 Let $V_{s}\in\mathcal{V}$ contain the two ends of the $AC$-edge for the $C$-link, and choose $V_{s}$ to minimise the distance from $s$ to $t$ in $T$.
-Recall that $T$ is a path by Lemma 9.6.
+Recall that $T$ is a path by Lemma˜9.6.
 If $t$ is not an end of $T$, there is $V_{r}\in\mathcal{V}$ with $t\in\mathring{r}Ts$ and $V_{r}\smallsetminus V_{t}$ non-empty, so every vertex in $V_{r}\smallsetminus V_{t}$ is separated from a vertex outside the $AC$-banana by the adhesion-set of $rt$ which has size two, contradicting 4-connectivity.
@@
 Thus the torso is 3-connected.
-(ii).
+˜(ii).
 Since $\Lambda$ contains an edge $e$, the subgraph $H$ of $G^{*}$ induced by $\hat{\Lambda}$ is either a triangle or a 4-cycle, depending on whether the other element of $\Lambda$ besides $e$ is a vertex $v$ or an edge $f$, respectively.
@@
 ∎
-The node $t$ of $T$ as in Lemma 9.7 is said to represent the $L$-link in $T$, and in the $K$-Tutte-banana.
+The node $t$ of $T$ as in Lemma˜9.7 is said to represent the $L$-link in $T$, and in the $K$-Tutte-banana.
 Lemma 9.8.
@@
 Without loss of generality, $K$ is the corner $AC$.
-Both the $A$-link and the $C$-link are represented by end-nodes $a$ and $c$ of $T$, respectively, by Lemma 9.7.
+Both the $A$-link and the $C$-link are represented by end-nodes $a$ and $c$ of $T$, respectively, by Lemma˜9.7.
 Assume now that $a=c$.
@@
 We claim that both the $A$-link and the $C$-link consist of vertices only.
-Indeed, if the $A$-link contains an edge, then the $AC$-edge $e$ for the $A$-link is distinct from the $AC$-edge for the $C$-link by Lemma 9.2, so the ends of $e$ together form a 2-separator of the $AC$-banana.
-This 2-separator gives rise to a 2-separation of the $AC$-banana that cuts off the triangle-torso or 4-cycle-torso at $a$ by Lemma 9.7, and this 2-separation is totally-nested as its separator is spanned by the edge $e$.
+Indeed, if the $A$-link contains an edge, then the $AC$-edge $e$ for the $A$-link is distinct from the $AC$-edge for the $C$-link by Lemma˜9.2, so the ends of $e$ together form a 2-separator of the $AC$-banana.
+This 2-separator gives rise to a 2-separation of the $AC$-banana that cuts off the triangle-torso or 4-cycle-torso at $a$ by Lemma˜9.7, and this 2-separation is totally-nested as its separator is spanned by the edge $e$.
 Hence the 2-separation corresponds to an edge of $T$, a contradiction.
@@
 For each corner $K$ let $\mathcal{T}^{K}=(T^{K},\mathcal{V}^{K})$ denote the $K$-Tutte-banana.
-Recall that $T^{K}$ is a path by Lemma 9.6.
+Recall that $T^{K}$ is a path by Lemma˜9.6.
 We obtain $O$ from the disjoint union $\bigsqcup_{K}T^{K}$, where $K$ ranges over the four corners, by doing the following for each link $L$: for both corners $K$ and $K^{\prime}$ adjacent to $L$ we consider the nodes $t$ and $t^{\prime}$ that represent the $L$-link in $T^{K}$ and $T^{K^{\prime}}$, respectively, and
   • $\cdot$
-    we join the nodes $t$ and $t^{\prime}$ by an edge $)𝑜𝐿{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{o(L)}}$ in $O$ if the torsos associated with these nodes in their respective Tutte-bananas are 3-connected;
+[…]
   • $\cdot$
-    we otherwise identify the nodes $t$ and $t^{\prime}$ to a single node $)𝑜𝐿{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{o(L)}}$ of $O$.
+[…]
 We say that $o(L)$ represents the $L$-link in $O$.
 For each node $t$ of $O$ that is not of the form $o(L)$, we define $G_{t}:=G[V^{K}_{t}]$ where $K$ is the unique corner with $t\in T^{K}$.
-For each node of $O$ of the form $o(L)$ for some link $L$, we consider the two nodes $t\in T^{K}$ and $t^{\prime}\in T^{K^{\prime}}$ that were identified to obtain $o(L)$, and we let $G_{o(L)}:=G[V^{K}_{t}]$ which is independent of $K$ by Lemma 9.7.
+For each node of $O$ of the form $o(L)$ for some link $L$, we consider the two nodes $t\in T^{K}$ and $t^{\prime}\in T^{K^{\prime}}$ that were identified to obtain $o(L)$, and we let $G_{o(L)}:=G[V^{K}_{t}]$ which is independent of $K$ by Lemma˜9.7.
 We refer to the pair $(O,(G_{t}:t\in V(O))$ as the Tutte-bagel induced by $(A,B)$ and $(C,D)$.
@@
 Proof.
-Let $\mathcal{O}=(O,(G_{t}:t\in V(O))$ denote the Tutte-bagel induced by $(A,B)$ and $(C,D)$. First, we show that $\mathcal{O}$ is a cycle-decomposition. That $O$ is a cycle follows from Lemma 9.8.
-(H1).
+Let $\mathcal{O}=(O,(G_{t}:t\in V(O))$ denote the Tutte-bagel induced by $(A,B)$ and $(C,D)$. First, we show that $\mathcal{O}$ is a cycle-decomposition. That $O$ is a cycle follows from Lemma˜9.8.
+˜(H1).
 Recall that there are no jumping edges and no diagonal edges by the (Crossing Lemma).
@@
 Hence every vertex and edge of $G$ appears in some $G_{t}$.
-(H2). Let $v\in G$ be a vertex and consider $O_{v}:=O[\,\{t\in V(O):v\in G_{t}\}\,]$.
+˜(H2). Let $v\in G$ be a vertex and consider $O_{v}:=O[\,\{t\in V(O):v\in G_{t}\}\,]$.
 If $v$ avoids $\hat{\Lambda}$ for every link $\Lambda$, then $v$ lies in a unique banana, and so $O_{v}$ is a subpath of the decomposition-path of the corresponding Tutte-banana.
@@
 Otherwise $H_{t}$ stems from a Tutte-banana.
-Then $H_{t}$ is 3-connected, a 4-cycle or a triangle, by Lemma 9.6.
+Then $H_{t}$ is 3-connected, a 4-cycle or a triangle, by Lemma˜9.6.
 Since every adhesion-set between two neighbouring triangle-torsos of a Tutte-banana is spanned by an edge in $G$, the same is true for $\mathcal{O}$. ∎
@@
 9.3. Settings
-The following settings are supported by Figure 9.4.
+The following settings are supported by Figure˜9.4.
 Setting 9.12 (Tutte-bagel setting).
@@
 For each node $t$ of $O$ let $U^{\prime}_{t}:=V(G_{t})$ and $W^{\prime}_{t}:=\bigcup\,\{\,V(G_{s}):s\in V(O-t)\,\}$.
-Let $(U_{t},W_{t})$ denote the right-left-shift of $(U^{\prime}_{t},W^{\prime}_{t})$.
+Let $(U_{t},W_{t})$ denote the right-left-reduction of $(U^{\prime}_{t},W^{\prime}_{t})$.
 Setting 9.13 (Tutte-bagel torso setting).
@@
 Assume that the torso $H_{t(0)}$ of $\mathcal{O}$ is 3-connected; so $H_{t(0)}$ stems from the $AC$-banana.
-Let $\{x^{\prime}_{1},x^{\prime}_{2}\}$ be the adhesion set of $t(-1)t(0)$ and let $\{y^{\prime}_{1},y^{\prime}_{2}\}$ be the adhesion set of $t(0)t(1)$.
+Let $\{x^{\prime}_{1},x^{\prime}_{2}\}$ be the adhesion set of t(1)t(0)t(-1)t(0) and let $\{y^{\prime}_{1},y^{\prime}_{2}\}$ be the adhesion set of $t(0)t(1)$.
 Let $(U^{\prime},W^{\prime}):=(U^{\prime}_{t(0)},W^{\prime}_{t(0)})$ and $(U,W):=(U_{t(0)},W_{t(0)})$.
@@
 (9.13).
-The vertex-set $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at $t(0)\,t(1)$ if the torso $H_{t(1)}$ is 3-connected, and it is the adhesion-set at $t(1)\,t(2)$ otherwise.
+The vertex-set $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at t(0)t(1)t(0)\,t(1) if the torso $H_{t(1)}$ is 3-connected, and it is the adhesion-set at t(1)t(2)t(1)\,t(2) otherwise.
 In the latter case, the torso at $t(2)$ is not a 4-cycle.
@@
 Hence we may assume that the torso $H_{t(1)}$ is a triangle.
-The next torso $H_{t(2)}$ cannot be 4-cycle by Lemma 9.4, so it is 3-connected or a triangle.
+The next torso $H_{t(2)}$ cannot be 4-cycle by Lemma˜9.4, so it is 3-connected or a triangle.
 Assume first that $H_{t(2)}$ is 3-connected.
@@
 Thus $y^{W}_{1}=y^{\prime}_{1}$.
-Since $H_{t(1)}$ is a triangle, $y_{2}$ is its free edge, so $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at $t(1)\,t(2)$.
-Otherwise it is a triangle, so the adhesion-graph at $t(1)\,t(2)$ is a $K_{2}$-subgraph of $G$.
+Since $H_{t(1)}$ is a triangle, $y_{2}$ is its free edge, so $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at t(1)t(2)t(1)\,t(2).
+Otherwise it is a triangle, so the adhesion-graph at t(1)t(2)t(1)\,t(2) is a $K_{2}$-subgraph of $G$.
 Then the $y^{\prime}_{i}$ incident to the free edge of the triangle $H_{t(1)}$ corresponds to an edge $y_{i}$.
 The $y^{\prime}_{j}$ that is the tip of $H_{t(1)}$ has $y^{W}_{i}$ as a neighbour and is incident to the free edge of $H_{t(2)}$, so $y^{\prime}_{j}=y_{j}$.
-Hence $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at $t(1)\,t(2)$.
+Hence $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at t(1)t(2)t(1)\,t(2).
 ∎
@@
 Proof.
-(i).
+˜(i).
 Suppose for a contradiction that two triangle-torsos of $\mathcal{O}$ have the same tip.
@@
 Let $e$ be the free edge of $H_{r}$, and let $e^{\prime}$ be the free edge of $H_{t}$.
-Then $\{v,e,e^{\prime}\}$ is a mixed-3-separator, contradicting 4-connectivity by Lemma 6.1.
-(ii).
+Then $\{v,e,e^{\prime}\}$ is a mixed-3-separator, contradicting 4-connectivity by Lemma˜6.1.
+˜(ii).
 First, assume that there is a node $p$ on $P$ other than $r,t$ with $v\in G_{p}$ such that the torso $H_{p}$ is a 4-cycle or 3-connected.
@@
 Proof.
-Since the Tutte-banana $\mathcal{T}$ for $K$ has no 3-connected torsos, it also has no 4-cycles as torsos by Lemma 9.4.
+Since the Tutte-banana $\mathcal{T}$ for $K$ has no 3-connected torsos, it also has no 4-cycles as torsos by Lemma˜9.4.
 Hence all torsos of $\mathcal{T}$ are triangles; in particular, both links adjacent to $K$ are represented by triangle-torsos of $\mathcal{T}$.
-Again by Lemma 9.4, it follows that all adhesion-graphs of $\mathcal{T}$ are $K_{2}$-subgraphs of $G$.
+Again by Lemma˜9.4, it follows that all adhesion-graphs of $\mathcal{T}$ are $K_{2}$-subgraphs of $G$.
 ∎
@@
 We distinguish two cases.
-[…]
+Suppose first that $\mathcal{T}$ has a 3-connected torso $H_{t}$.
+We find two disjoint $a_{i}$$H_{t}$ paths $P_{i}$ for $i=1,2$ in the $AD$-banana using Lemma˜9.16.
+We even get that both $P_{i}$ are subgraphs of $G$.
+Note that the ends of the $P_{i}$ in $H_{t}$ together form an adhesion-set.
+If $H_{t}$ is a $K_{4}$, then we link the ends of the $P_{i}$ in the bag $G_{t}$ by a path that avoids any prescribed vertex $d$ in the $D$-link and that uses no torso-edges.
+Otherwise $H_{t}$ is not a $K_{4}$, so we may use Lemma˜9.17 to link the ends of the $P_{i}$ in the bag $G_{t}$ with a path that avoids all vertices in the $D$-link and that uses no torso-edges.
 Suppose second that $\mathcal{T}$ has no 3-connected torsos.
-Then all torsos of $\mathcal{T}$ are triangles, both the $A$-link and the $D$-link are represented by triangle-torsos, and $a_{1}a_{2}\in E(G)$ by Lemma 9.18.
+Then all torsos of $\mathcal{T}$ are triangles, both the $A$-link and the $D$-link are represented by triangle-torsos, and $a_{1}a_{2}\in E(G)$ by Lemma˜9.18.
 In particular, exactly one of $a_{1},a_{2}$ lies in the $A$-link, say it is $a_{1}$.
@@
 Proof.
-By Lemma 9.14, there is $i\in\{1,2\}$ such that $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at $t(i-1)t(i)$.
+By Lemma˜9.14, there is $i\in\{1,2\}$ such that $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at t(i1)t(i)t(i-1)t(i).
 Moreover, the torso $H_{t(i)}$ is not a 4-cycle.
-We consider two cases, supported by Figure 9.5.
+We consider two cases, supported by Figure˜9.5.
 Suppose first that the torso at $t(i)$ is 3-connected.
@@
 Let $k\geqslant i$ be maximum such that $t(k)$ stems from the $BC$-Tutte-banana.
-By Lemma 9.16, we find two disjoint $\{z_{1},z_{2}\}$$\{u_{1},u_{2}\}$ paths $R_{1},R_{2}$ in $G_{t(i+1)}\cup\cdots\cup G_{t(k)}$, taking the trivial paths if $k=i$.
+By Lemma˜9.16, we find two disjoint $\{z_{1},z_{2}\}$$\{u_{1},u_{2}\}$ paths $R_{1},R_{2}$ in $G_{t(i+1)}\cup\cdots\cup G_{t(k)}$, taking the trivial paths if $k=i$.
 We apply the (One-Path Lemma).
@@
 Suppose otherwise that the torso at $t(i)$ is a triangle.
-Then $i=2$ and the torso at $t(1)$ is not 3-connected, by Lemma 9.14.
-Since the torso at $t(2)$ is a triangle, the torso at $t(1)$ cannot be a 4-cycle by Lemma 9.4, so it must be a triangle as well.
+Then $i=2$ and the torso at $t(1)$ is not 3-connected, by Lemma˜9.14.
+Since the torso at $t(2)$ is a triangle, the torso at $t(1)$ cannot be a 4-cycle by Lemma˜9.4, so it must be a triangle as well.
 Then the adhesion-graph of $t(1)t(2)$ is a $K_{2}$-subgraph of $G$ with vertices $y^{W}_{1},y^{W}_{2}$.
@@
 ∎
-Figure 9.5. Two paths found by Lemma 9.20. The vertices $y^{W}_{i}$ are blue. The path $P_{2}=Q_{2}$ is red. The path $P_{1}$ is green. The green vertices in the $B$-link are $u_{1},u_{2}$.
+Figure 9.5. Two paths found by Lemma˜9.20. The vertices $y^{W}_{i}$ are blue. The path $P_{2}=Q_{2}$ is red. The path $P_{1}$ is green. The green vertices in the $B$-link are $u_{1},u_{2}$.
 Lemma 9.21 (Three-Paths Lemma).
@@
 Proof.
-Recall that not both corners $AD,BD$ are potter by 4.7.
+Recall that not both corners $AD,BD$ are potter by ˜4.7.
 Suppose first that the corner $AD$ is not potter.
@@
 Note that $|W^{\prime}\smallsetminus U^{\prime}|\geqslant 2$.
-(ii) $\Rightarrow$ (i).
+˜(ii) $\Rightarrow$ ˜(i).
 Assume that $H_{t(0)}$ is good.
-If $|H_{t(0)}|\geqslant 6$, then $U^{\prime}\smallsetminus W^{\prime}$ has size $\geqslant 2$, so $(U,W)$ is a tetra-separation by 5.10.
-If $|H_{t(0)}|=5$ or $|H_{t(0)}|=4$, then $H_{t(0)}$ being good ensures via the neighbouring torsos that at least one or at least two vertices in $S(U^{\prime},W^{\prime})$ have exactly one neighbour in $W^{\prime}\smallsetminus U^{\prime}$, respectively, so $(U,W)$ is a tetra-separation by 5.10.
-(i) $\Rightarrow$ (ii).
+If $|H_{t(0)}|\geqslant 6$, then $U^{\prime}\smallsetminus W^{\prime}$ has size $\geqslant 2$, so $(U,W)$ is a tetra-separation by ˜5.10.
+If $|H_{t(0)}|=5$ or $|H_{t(0)}|=4$, then $H_{t(0)}$ being good ensures via the neighbouring torsos that at least one or at least two vertices in $S(U^{\prime},W^{\prime})$ have exactly one neighbour in $W^{\prime}\smallsetminus U^{\prime}$, respectively, so $(U,W)$ is a tetra-separation by ˜5.10.
+˜(i) $\Rightarrow$ ˜(ii).
 Assume that $(U,W)$ fails to be a tetra-separation.
-Then by 5.10, the number of vertices in $S(U^{\prime},W^{\prime})$ with $\leqslant 1$ neighbour in $W^{\prime}\smallsetminus U^{\prime}$ fails to be $\geqslant 2-|U^{\prime}\smallsetminus W^{\prime}|$.
+Then by ˜5.10, the number of vertices in $S(U^{\prime},W^{\prime})$ with $\leqslant 1$ neighbour in $W^{\prime}\smallsetminus U^{\prime}$ fails to be $\geqslant 2-|U^{\prime}\smallsetminus W^{\prime}|$.
 Hence all sufficient conditions for $H_{t(0)}$ being good fail.
@@
 Proof.
-If $S(U,W)$ contains an edge, then $G-S(U,W)$ has exactly two components by 4-connectivity, so $(U,W)$ is half-connected.
+If $S(U,W)$ contains an edge, then GS(U,W)G-S(U,W) has exactly two components by 4-connectivity, so $(U,W)$ is half-connected.
 Otherwise $(U,W)=(U^{\prime},W^{\prime})$.
 We will show that $G[W\smallsetminus U]$ is connected.
-Let $\pi$ denote the adhesion-set of $\mathcal{O}$ that is spanned by the $BD$-edge for the $D$-link by Lemma 9.11.
+Let $\pi$ denote the adhesion-set of $\mathcal{O}$ that is spanned by the $BD$-edge for the $D$-link by Lemma˜9.11.
 We claim that every vertex $v\in W\smallsetminus U$ not in $\pi$ sends a fan of two paths to $\pi$ in $G[W\smallsetminus U]$.
 Indeed, $\pi$ together with $\{x^{\prime}_{1},x^{\prime}_{2}\}$, say, separates $v$ from $G_{t(0)}$.
-Since $G$ is 4-connected, there is a fan of four paths from $v$ to $\{x^{\prime}_{1},x^{\prime}_{2}\}\cup\pi$ by Lemma 6.2.
+Since $G$ is 4-connected, there is a fan of four paths from $v$ to $\{x^{\prime}_{1},x^{\prime}_{2}\}\cup\pi$ by Lemma˜6.2.
 The two paths ending in $\pi$ are as desired.
@@
 Proof.
-If $H_{t(1)}$ is 3-connected, we have (i).
+If $H_{t(1)}$ is 3-connected, we have ˜(i).
 Otherwise $H_{t(1)}$ is a 4-cycle or a triangle.
-Since $y_{2}$ is a vertex by assumption, $H_{t(1)}$ is a triangle and $y_{2}=y^{W}_{2}$ is its tip, by Lemma 9.14.
+Since $y_{2}$ is a vertex by assumption, $H_{t(1)}$ is a triangle and $y_{2}=y^{W}_{2}$ is its tip, by Lemma˜9.14.
 Hence $y^{W}_{1}$ is incident to the free edge of $H_{t(1)}$.
-If $H_{t(2)}$ is 3-connected, then we have (i).
-Otherwise $H_{t(2)}$ is a triangle, by Lemma 9.4.
-By Lemma 9.15, $y^{W}_{1}$ is the tip of $H_{t(2)}$.
+If $H_{t(2)}$ is 3-connected, then we have ˜(i).
+Otherwise $H_{t(2)}$ is a triangle, by Lemma˜9.4.
+By Lemma˜9.15, $y^{W}_{1}$ is the tip of $H_{t(2)}$.
 Hence $y^{W}_{2}$ is not contained in the adhesion-set between $H_{t(2)}$ and $H_{t(3)}$, and in particular $y^{W}_{2}$ is not contained in $H_{t(3)}$.
-If $H_{t(3)}$ is 3-connected, we have (i).
+If $H_{t(3)}$ is 3-connected, we have ˜(i).
 Otherwise $H_{t(3)}$ also is a triangle.
-By Lemma 9.15, $y^{W}_{1}$ is incident to the free edge of $H_{t(3)}$.
-By Lemma 9.4, the adhesion-graph between the triangle-torsos $H_{t(2)}$ and $H_{t(3)}$ is a $K_{2}$-subgraph of $G$.
+By Lemma˜9.15, $y^{W}_{1}$ is incident to the free edge of $H_{t(3)}$.
+By Lemma˜9.4, the adhesion-graph between the triangle-torsos $H_{t(2)}$ and $H_{t(3)}$ is a $K_{2}$-subgraph of $G$.
 Hence the two edges of $H_{t(3)}$ incident to $y^{W}_{1}$ exist in $G$.
-Thus we have (ii).
+Thus we have ˜(ii).
 ∎
@@
 Proof.
-Assume that (i) fails.
+Assume that ˜(i) fails.
 Then there exists a cutvertex $v$ of $G_{t}$ that separates $a_{1}$ from $b_{1}$.
@@
 As $H_{t}$ is 3-connected, $H_{t}$ is a 4-wheel with rim $a_{1}a_{2}b_{1}b_{2}$.
-Since (i) fails by assumption, both torso-edges $a_{1}a_{2}$ and $b_{1}b_{2}$ must be missing from $G$.
+Since ˜(i) fails by assumption, both torso-edges $a_{1}a_{2}$ and $b_{1}b_{2}$ must be missing from $G$.
 ∎
@@
 Case $i<j$.
-This case is supported by Figure 9.6.
+This case is supported by Figure˜9.6.
 If $H_{t(i)}$ is 3-connected, we let $F_{y}$ by a $y^{W}_{1}$$\pi_{i}$ fan of two paths in $G_{t(i)}$ avoiding the vertex in $\pi_{i-1}-y^{W}_{1}$.
-Otherwise, by Lemma 9.27, $H_{t(i)}$ is a triangle avoiding $y^{W}_{2}$ such that both edges incident to $y^{W}_{1}$ in $H_{t(i)}$ exist in $G$, and we let $F_{y}$ denote the $y^{W}_{1}$$\pi_{i}$ fan of two paths defined by these two edges.
+Otherwise, by Lemma˜9.27, $H_{t(i)}$ is a triangle avoiding $y^{W}_{2}$ such that both edges incident to $y^{W}_{1}$ in $H_{t(i)}$ exist in $G$, and we let $F_{y}$ denote the $y^{W}_{1}$$\pi_{i}$ fan of two paths defined by these two edges.
 In both cases, $F_{y}$ avoids $y_{2}$.
 Similarly, we obtain an $x^{W}_{1}$$\pi_{j-1}$ fan $F_{x}\subseteq G_{t(j)}$ avoiding $x_{2}$.
-By Lemma 9.16, we find two disjoint $\pi_{i}$$\pi_{j-1}$ paths in $G_{t(i+1)}\cup\cdots\cup G_{t(j-1)}$.
+By Lemma˜9.16, we find two disjoint $\pi_{i}$$\pi_{j-1}$ paths in $G_{t(i+1)}\cup\cdots\cup G_{t(j-1)}$.
 Call these paths $P$ and $Q$.
 Then $F_{x}\cup P\cup Q\cup F_{y}$ is a union of two independent $x^{W}_{1}$$y^{W}_{1}$ paths in $G[W]-x_{2}-y_{2}$.
-Figure 9.6. Case $i<j$ in the proof of Lemma 9.29. Here, $H_{t(i)}$ is 3-connected and $H_{t(j)}$ is a triangle. The unique vertex in $\pi_{i-1}-y^{W}_{1}$ is black.
+Figure 9.6. Case $i<j$ in the proof of Lemma˜9.29. Here, $H_{t(i)}$ is 3-connected and $H_{t(j)}$ is a triangle. The unique vertex in $\pi_{i-1}-y^{W}_{1}$ is black.
 Case $i\geqslant j$.
-This case is supported by Figure 9.7.
+This case is supported by Figure˜9.7.
 We claim that the corner $BD$ is not potter, and assume for a contradiction that it is.
 Let $\pi$ denote the unique adhesion-set of the $BD$-Tutte-banana.
-Since $BD$ is potter, neither corner $BC$ nor $AD$ is potter by 4.7.
+Since $BD$ is potter, neither corner $BC$ nor $AD$ is potter by ˜4.7.
 Hence if $y_{1}$ is an edge, it neither dangles from the $C$-link through the $B$-link nor from the $B$-link through the $C$-link, by the (Crossing Lemma).
@@
 Say this is a neighbour $t$ of $o(B)$.
-Since $y^{W}_{1}$ is the tip of the triangle $H_{o(B)}$, the vertex $y^{W}_{1}$ is an end of the free edge $f$ of $H_{t}$ by Lemma 9.15.
+Since $y^{W}_{1}$ is the tip of the triangle $H_{o(B)}$, the vertex $y^{W}_{1}$ is an end of the free edge $f$ of $H_{t}$ by Lemma˜9.15.
 Let $u$ denote the other end of $f$.
@@
 Hence $x^{W}_{1}$ lies in the two triangle-torsos from the potter corner $AD$ and in the triangle-torso $H_{t}$.
-Thus $o(D)$ and $t$ are neighbours in $T$ by Lemma 9.15.
+Thus $o(D)$ and $t$ are neighbours in $T$ by Lemma˜9.15.
 In particular, $T$ has exactly three nodes.
-Since $v$ lies in the two triangle-torsos $H_{t}$ and $H_{o(B)}$ while $x^{W}_{1}$ lies in the two triangle-torsos for the potter corner $AD$, we cannot have $v=x^{W}_{1}$ by Lemma 9.15.
+Since $v$ lies in the two triangle-torsos $H_{t}$ and $H_{o(B)}$ while $x^{W}_{1}$ lies in the two triangle-torsos for the potter corner $AD$, we cannot have $v=x^{W}_{1}$ by Lemma˜9.15.
 Hence $x^{W}_{1}=u$.
-By Lemma 9.4, the two adhesion-sets of $H_{t}$ are spanned by edges in $G$.
+By Lemma˜9.4, the two adhesion-sets of $H_{t}$ are spanned by edges in $G$.
 Hence $x^{W}_{1}y^{W}_{1}$ and $x^{W}_{1}vy^{W}_{1}$ are desired paths.
-Figure 9.7. Case $i\geqslant j$ in the proof of Lemma 9.29. Left: $H_{t}$ is a triangle.
+Figure 9.7. Case $i\geqslant j$ in the proof of Lemma˜9.29. Left: $H_{t}$ is a triangle.
 Right: $H_{t}$ is 3-connected.
@@
 Let $\{x^{W}_{1},u\}$ denote the adhesion-set of the edge $o(D)t$, and let $\{y^{W}_{1},v\}$ denote the adhesion-set of the edge $o(B)t$.
-We find two independent $y^{W}_{1}$$x^{W}_{1}$ paths in $G_{t}$, and are done, unless $H_{t}$ is a 4-wheel with rim $x^{W}_{1}uy^{W}_{1}v$ and $x^{W}_{1}u,y^{W}_{1}v\notin E(G)$, by Lemma 9.28.
+We find two independent $y^{W}_{1}$$x^{W}_{1}$ paths in $G_{t}$, and are done, unless $H_{t}$ is a 4-wheel with rim $x^{W}_{1}uy^{W}_{1}v$ and $x^{W}_{1}u,y^{W}_{1}v\notin E(G)$, by Lemma˜9.28.
 But in the latter case, $u$ and $v$ have degree three in $G$, contradicting 4-connectivity.
@@
 If $x_{2}$ or $y_{2}$ is an edge, then we are done by (6.4.1), so assume that both are vertices.
-By Lemma 9.25, there is an $x^{\prime}_{1}$$y^{\prime}_{1}$ path $P^{\prime}$ in $G_{t(0)}-\{x^{\prime}_{2},y^{\prime}_{2}\}$.
-Then $P^{\prime}$ contains a $\hat{x}_{1}$$\hat{y}_{1}$ path $P$ that avoids the vertices $x_{2}$ and $y_{2}$, by Lemma 9.26.
-We find two independent $x^{W}_{1}$$y^{W}_{1}$ paths in $G[W]-x_{2}-y_{2}$ using Lemma 9.29.
+By Lemma˜9.25, there is an $x^{\prime}_{1}$$y^{\prime}_{1}$ path $P^{\prime}$ in $G_{t(0)}-\{x^{\prime}_{2},y^{\prime}_{2}\}$.
+Then $P^{\prime}$ contains a $\hat{x}_{1}$$\hat{y}_{1}$ path $P$ that avoids the vertices $x_{2}$ and $y_{2}$, by Lemma˜9.26.
+We find two independent $x^{W}_{1}$$y^{W}_{1}$ paths in $G[W]-x_{2}-y_{2}$ using Lemma˜9.29.
 Hence $\{x_{1},y_{1}\}$ is 3-linked around $S(U,W)$ by (6.4.3).
@@
 (9.21).
-So it remains to find an $x^{\prime}_{1}$$x^{\prime}_{2}$ path in $G^{\prime}:=G_{t_{0}}-\{y^{\prime}_{1},y^{\prime}_{2}\}$ (as any such path contains a suitable $\hat{x}_{1}$$\hat{x}_{2}$ path by Lemma 9.26).
+So it remains to find an $x^{\prime}_{1}$$x^{\prime}_{2}$ path in $G^{\prime}:=G_{t_{0}}-\{y^{\prime}_{1},y^{\prime}_{2}\}$ (as any such path contains a suitable $\hat{x}_{1}$$\hat{x}_{2}$ path by Lemma˜9.26).
 Assume for a contradiction that such a path is missing.
-Then $H_{t(0)}$ is a $K_{4}$ by Lemma 9.17.
-Since both $y_{1}$ and $y_{2}$ are vertices, the torso $H_{t(1)}$ is 3-connected by Lemma 9.14.
+Then $H_{t(0)}$ is a $K_{4}$ by Lemma˜9.17.
+Since both $y_{1}$ and $y_{2}$ are vertices, the torso $H_{t(1)}$ is 3-connected by Lemma˜9.14.
 Since $H_{t(0)}$ is good, it follows that $H_{t(-1)}$ is a 4-cycle.
@@
 Suppose first that $\pi_{1}=\{x_{1},x_{2}\}$ and $\pi_{2}=\{y_{1},y_{2}\}$.
-We find two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths $P_{1},P_{2}$ in $G[W]$ using Lemmas 9.16 and 9.14.
+We find two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths $P_{1},P_{2}$ in $G[W]$ using Lemmas˜9.16 and 9.14.
 The torso $H_{t(0)}$ is 3-connected by assumption, and hence it contains three independent $\{x^{\prime}_{1},x^{\prime}_{2}\}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ paths by (Menger for independent paths).
@@
 None of these paths use torso-edges, so all three paths are subgraphs of the bag $G_{t(0)}$.
-The three $\{x^{\prime}_{1},x^{\prime}_{2}\}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ paths can be shortened to three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths $Q_{1},Q_{2},Q_{3}$ by Lemma 9.26.
+The three $\{x^{\prime}_{1},x^{\prime}_{2}\}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ paths can be shortened to three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths $Q_{1},Q_{2},Q_{3}$ by Lemma˜9.26.
 Hence $\pi_{1}$ is 0-potter-linked to $\pi_{2}$.
@@
 (9.21).
-By Lemma 9.28, we find two independent $\{x^{\prime}_{1},y^{\prime}_{1}\}$$\{x^{\prime}_{2},y^{\prime}_{2}\}$ paths in the bag $G_{t(0)}$, which shorten to two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths $P_{4},P_{5}$ in $G_{t(0)}$ by Lemma 9.26, unless $H_{t(0)}$ is a 4-wheel and both torso-edges $x^{\prime}_{1}x^{\prime}_{2},y^{\prime}_{1}y^{\prime}_{2}$ are missing in $G$.
+By Lemma˜9.28, we find two independent $\{x^{\prime}_{1},y^{\prime}_{1}\}$$\{x^{\prime}_{2},y^{\prime}_{2}\}$ paths in the bag $G_{t(0)}$, which shorten to two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths $P_{4},P_{5}$ in $G_{t(0)}$ by Lemma˜9.26, unless $H_{t(0)}$ is a 4-wheel and both torso-edges $x^{\prime}_{1}x^{\prime}_{2},y^{\prime}_{1}y^{\prime}_{2}$ are missing in $G$.
 But the latter is impossible: since $H_{t(0)}$ is good, some neighbouring torso would be a triangle or a 4-cycle, but then a vertex in the adhesion-set would have degree three in $G$, contradicting 4-connectivity.
@@
 Proof.
-(i).
+˜(i).
 Since $H_{t(0)}$ is 3-connected, the vertex $x^{\prime}_{1}$ has $\geqslant 3$ neighbours in $H_{t(0)}$.
@@
 Then $x^{\prime}_{1}y^{\prime}_{1}$ is a $(W\smallsetminus U)$$(U\smallsetminus W)$ edge, so $(U,W)$ violates the matching-condition, a contradiction.
-(ii).
+˜(ii).
 Assume for a contradiction that $\{x_{1},x_{2}\}$ is $W$-weird.
@@
 Moreover, $x^{\prime}_{2}$ has exactly one neighbour outside $H_{t(0)}$ by weirdness.
-But then $x_{2}$ should be an edge in $S(U,W)$ by the definition of right-left-shift of $(U^{\prime},W^{\prime})$, contradicting that $x_{2}$ is a vertex.
+But then $x_{2}$ should be an edge in $S(U,W)$ by the definition of right-left-reduction of $(U^{\prime},W^{\prime})$, contradicting that $x_{2}$ is a vertex.
 ∎
@@
 Assume otherwise, say $y_{1}$ is an inward edge.
-Then by Lemma 9.33 we have $y^{W}_{1}=y^{\prime}_{1}$ and some $x^{\prime}_{j}$ is a neighbour of $y^{W}_{1}$.
+Then by Lemma˜9.33 we have $y^{W}_{1}=y^{\prime}_{1}$ and some $x^{\prime}_{j}$ is a neighbour of $y^{W}_{1}$.
 Then $j\neq 1$ since otherwise $x^{\prime}_{1}y^{W}_{1}$ is a $(U\smallsetminus W)$$(W\smallsetminus U)$ edge, contradicting the matching-condition.
@@
 It follows that $x_{1}^{W}$ is the only neighbour of $x_{2}^{\prime}$ in $W^{\prime}\smallsetminus U^{\prime}$.
-This contradicts that $x_{2}$ is a vertex by definition of right-left-shift.
+This contradicts that $x_{2}$ is a vertex by definition of right-left-reduction.
 ∎
@@
 Case: $x_{1}$ is an inward edge.
-By Lemma 9.33, there is a neighbour $y^{\prime}_{i}$ of $x^{\prime}_{1}$ with $y^{\prime}_{i}=y^{W}_{i}$, and $\pi_{1}$ is not $W$-weird.
+By Lemma˜9.33, there is a neighbour $y^{\prime}_{i}$ of $x^{\prime}_{1}$ with $y^{\prime}_{i}=y^{W}_{i}$, and $\pi_{1}$ is not $W$-weird.
 So $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $W$ by (6.9.1).
 Since we assumed earlier that $\pi_{1}$ is weird for some side of $(U,W)$, but showed that $\pi_{1}$ is not $W$-weird, it follows that $\pi_{1}$ is $U$-weird.
-By Lemma 9.16, there exist two disjoint $\{x^{W}_{1},x^{W}_{2}\}$$\{y^{W}_{1},y^{W}_{2}\}$ paths in $\bigcup_{t\in O-t(0)}G_{t}$.
+By Lemma˜9.16, there exist two disjoint $\{x^{W}_{1},x^{W}_{2}\}$$\{y^{W}_{1},y^{W}_{2}\}$ paths in $\bigcup_{t\in O-t(0)}G_{t}$.
 Since $x_{1}$ is an inward edge and $y^{\prime}_{i}=y^{W}_{i}$, the edge $x^{\prime}_{1}y^{\prime}_{i}$ can also be written as $x^{W}_{1}y^{W}_{i}$.
@@
 Case: $x_{1}$ is an outward edge.
-If $\pi_{1}$ is $U$-weird, then we find an $x_{2}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ edge $e=x_{2}y_{i}^{\prime}$ with $y_{i}^{\prime}=y_{i}^{W}$ by Lemma 9.34.
+If $\pi_{1}$ is $U$-weird, then we find an $x_{2}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ edge $e=x_{2}y_{i}^{\prime}$ with $y_{i}^{\prime}=y_{i}^{W}$ by Lemma˜9.34.
 Since $x_{2}\in\pi_{1}$ and $y_{i}^{\prime}\in\hat{\pi}_{2}\cap W$, the edge $e$ is a $\pi_{1}$$(\hat{\pi}_{2}\cap W)$ edge, which witnesses that $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $U$ by (6.9.2).
@@
 For this, we assume that $\pi_{1}$ is $W$-weird by (6.9.1).
-By Lemma 9.35, no $y_{i}$ is an inward edge.
+By Lemma˜9.35, no $y_{i}$ is an inward edge.
 Since $H_{t(0)}$ is 3-connected, we find three independent $\{x^{\prime}_{1},x^{\prime}_{2}\}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ paths in $H_{t(0)}$ by (Menger for independent paths).
@@
 Proof.
-By symmetry and by 6.12, it suffices to show the following: if $\pi_{1}$ is $U$-weird and $\pi_{2}$ is $W$-weird, then there is a $\pi_{1}$$\pi_{2}$ edge in $G$.
+By symmetry and by ˜6.12, it suffices to show the following: if $\pi_{1}$ is $U$-weird and $\pi_{2}$ is $W$-weird, then there is a $\pi_{1}$$\pi_{2}$ edge in $G$.
 So we assume that $\pi_{1}$ is $U$-weird and $\pi_{2}$ is $W$-weird.
-Then no $y_{i}$ is an inward edge by Lemma 9.33.
+Then no $y_{i}$ is an inward edge by Lemma˜9.33.
 Without loss of generality, $y_{1}$ is an outward edge.
-Then by Lemma 9.35, no $x_{i}$ is an inward edge.
+Then by Lemma˜9.35, no $x_{i}$ is an inward edge.
 So we can assume that $x_{1}$ is an outward edge.
-Then by Lemma 9.34, we find an $x_{2}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ edge $x_{2}y_{i}^{\prime}$ with $y_{i}^{\prime}=y_{i}^{W}$.
+Then by Lemma˜9.34, we find an $x_{2}$$\{y^{\prime}_{1},y^{\prime}_{2}\}$ edge $x_{2}y_{i}^{\prime}$ with $y_{i}^{\prime}=y_{i}^{W}$.
 Note that $y_{i}$ is a vertex, so $y_{i}^{\prime}=y_{i}\in\pi_{2}$.
@@
 Then $y^{\prime}_{1}$ has $\leqslant 1$ neighbour in $W\smallsetminus\{x^{\prime}_{1}\}\supseteq V(G-G_{t(0)})$.
-But then, by the definition of $(U,W)$ via right-left-shift, $y_{1}$ should be an edge rather than the vertex $y^{\prime}_{1}$, a contradiction.
+But then, by the definition of $(U,W)$ via right-left-reduction, $y_{1}$ should be an edge rather than the vertex $y^{\prime}_{1}$, a contradiction.
 Assume otherwise that $x_{1}$ is an outward edge.
-By Lemma 9.38, $x^{W}_{1}\in A$.
+By Lemma˜9.38, $x^{W}_{1}\in A$.
 However, since $y^{\prime}_{1}$ is contained in $A\cap C$ and the Tutte-bagel induces a path-decomposition of $G[A]$, the vertex $y^{\prime}_{1}=y_{1}$ does not neighbour $x^{W}_{1}$, contradicting that $\{x_{1},y_{1}\}$ is $W$-weird.
@@
 Proof.
-By Lemma 9.39, neither $\pi_{1}$ nor $\pi_{2}$ is $W$-weird.
+By Lemma˜9.39, neither $\pi_{1}$ nor $\pi_{2}$ is $W$-weird.
 Hence $\{\pi_{1},\pi_{2}\}$ is 2-potter-linked.
@@
 Proof.
-If $H_{t}$ is bad, then $(U_{t},W_{t})$ is not even a tetra-separation by Lemma 9.23, let alone totally-nested.
+If $H_{t}$ is bad, then $(U_{t},W_{t})$ is not even a tetra-separation by Lemma˜9.23, let alone totally-nested.
 So we may assume now that $H_{t}$ is good.
@@
 (9.13) with $t=t(0)$.
-By Lemma 9.23, $(U,W)$ is a tetra-separation.
-It suffices to show that $(U,W)$ is externally 5-connected by 6.16.
-By Lemma 9.24, $(U,W)$ is half-connected.
-To see that $(U,W)$ is 3-linked, we combine Lemma 9.30 with Lemma 9.31.
-By Lemma 9.32, $(U,W)$ is 0-potter-linked.
-We get that $(U,W)$ is 1-potter-linked and 2-potter-linked from Lemmas 9.36, 9.40 and 9.37.
+By Lemma˜9.23, $(U,W)$ is a tetra-separation.
+It suffices to show that $(U,W)$ is externally 5-connected by ˜6.16.
+By Lemma˜9.24, $(U,W)$ is half-connected.
+To see that $(U,W)$ is 3-linked, we combine Lemma˜9.30 with Lemma˜9.31.
+By Lemma˜9.32, $(U,W)$ is 0-potter-linked.
+We get that $(U,W)$ is 1-potter-linked and 2-potter-linked from Lemmas˜9.36, 9.40 and 9.37.
 ∎
@@
 Assume not for a contradiction, and let $v$ be a neighbour of $u$ in the $D$-link.
-Recall that $u\notin D$ by Lemma 9.42.
+Recall that $u\notin D$ by Lemma˜9.42.
 So the edge $uv$ dangles from the $D$-link through the $A$-link.
@@
 By assumption, $o(A)$ and $s$ are not adjacent in $O$.
-By Lemma 9.15, there is precisely one torso $H_{r}$ in between the two triangle-torsos $H_{o(A)}$ and $H_{s}$, which also contains $u$.
-But $H_{r}$ has $u$ as tip since $uv$ is free in $H_{o(A)}$, which by Lemma 9.15 contradicts that $H_{s}$ also has $u$ as tip.
+By Lemma˜9.15, there is precisely one torso $H_{r}$ in between the two triangle-torsos $H_{o(A)}$ and $H_{s}$, which also contains $u$.
+But $H_{r}$ has $u$ as tip since $uv$ is free in $H_{o(A)}$, which by Lemma˜9.15 contradicts that $H_{s}$ also has $u$ as tip.
 ∎
-Figure 9.8. The situation in Lemma 9.44.
+Figure 9.8. The situation in Lemma˜9.44.
 Let $P$ be a path, and let $u,v$ be vertices on $P$.
 Recall that $uP\mathring{v}$ denotes the subpath of $P$ that starts at $u$ and ends just before the vertex $v$.
-The statement of the following lemma is supported by Figure 9.8.
+The statement of the following lemma is supported by Figure˜9.8.
 Lemma 9.44 (Link-Moving Lemma for vertices).
@@
 (9.12).
-Let $P$ and $Q$ be the two paths that are the two components of $O-o(C)-o(D)$.
+Let $P$ and $Q$ be the two paths that are the two components of Oo(C)o(D)O-o(C)-o(D).
 Assume that $o(A)\in P$ and $o(B)\in Q$.
@@
 Define $d$ similarly.
-Let $s$ be a node in the same component of $P-o(A)$ as $c$ such that $H_{s}$ is a 4-cycle or a triangle.
+Let $s$ be a node in the same component of Po(A)P-o(A) as $c$ such that $H_{s}$ is a 4-cycle or a triangle.
 Assume that if the $A$-link is represented by a triangle-torso, then $s$ is not adjacent to $o(A)$ in $P$.
@@
   2. (ii)
-    $S(A^{\prime},B^{\prime})=S\cup(D$-link)))), where $S$ consists of the two free edges of $H_{s}$ if $H_{s}$ is a 4-cycle and of the free edge and the tip of $H_{s}$ if $H_{s}$ is a triangle.
+    $S(A^{\prime},B^{\prime})=S\cup(D$-link)), where $S$ consists of the two free edges of $H_{s}$ if $H_{s}$ is a 4-cycle and of the free edge and the tip of $H_{s}$ if $H_{s}$ is a triangle.
   3. (iii)
@@
 Proof.
-(ii) holds by construction, and gives that the mixed-separation $(A^{\prime},B^{\prime})$ has order four. We show (i) next.
+˜(ii) holds by construction, and gives that the mixed-separation $(A^{\prime},B^{\prime})$ has order four. We show ˜(i) next.
 Matching-condition.
@@
 (4.1).
-Thus the $A$-link is represented by a triangle-torso, so $a\notin D$ by Lemma 9.42, a contradiction.
+Thus the $A$-link is represented by a triangle-torso, so $a\notin D$ by Lemma˜9.42, a contradiction.
 Degree-condition.
@@
 If $H_{r}$ is 3-connected, then $v$ has two neighbours in $G_{r}$, both of which lie in $A^{\prime}\smallsetminus B^{\prime}$ since $G_{s}\subseteq G[A\cap C]$.
-Otherwise $H_{r}$ is a triangle, and the adhesion-graph in between $H_{r}$ and $H_{s}$ is a $K_{2}$-subgraph of $G$ by Lemma 9.4.
+Otherwise $H_{r}$ is a triangle, and the adhesion-graph in between $H_{r}$ and $H_{s}$ is a $K_{2}$-subgraph of $G$ by Lemma˜9.4.
 Hence $v$ has two neighbours in $G_{r}$, both of which lie in $A^{\prime}\smallsetminus B^{\prime}$ since $r\neq o(A)$ by assumption.
@@
 Every neighbour of $v$ in $A\smallsetminus B$ also lies in $A^{\prime}\smallsetminus B^{\prime}$, except possibly when a neighbour of $v$ lies in $S$ or an edge incident to $v$ lies in $S$.
-The former exception is impossible by Lemma 9.43, while the latter is excluded by Lemma 9.42.
-This conclude the proof of (i).
-(iii).
+The former exception is impossible by Lemma˜9.43, while the latter is excluded by Lemma˜9.42.
+This conclude the proof of ˜(i).
+˜(iii).
 We have to find two vertices in $B\cup D$ that are contained in $A^{\prime}\smallsetminus B^{\prime}$ and in $B^{\prime}\smallsetminus A^{\prime}$, respectively.
 Every vertex in $B\smallsetminus A$ lies in the intersection of $B\cup D$ with $B^{\prime}\smallsetminus A^{\prime}$, and $B\smallsetminus A$ is non-empty.
-Any vertex in $D\smallsetminus B$ lies in $A^{\prime}\smallsetminus B^{\prime}$ by Lemma 9.42, so it remains to show that $D\smallsetminus B$ is non-empty.
+Any vertex in $D\smallsetminus B$ lies in $A^{\prime}\smallsetminus B^{\prime}$ by Lemma˜9.42, so it remains to show that $D\smallsetminus B$ is non-empty.
 If $D\smallsetminus B$ is empty, then the $D$-link consists of two edges.
@@
 (9.12).
-Let $P$ and $Q$ be the two paths that are the two components of $O-o(C)-o(D)$.
+Let $P$ and $Q$ be the two paths that are the two components of Oo(C)o(D)O-o(C)-o(D).
 Assume that $o(A)\in P$ and $o(B)\in Q$.
@@
   2. (ii)
-    $S(A^{\prime},B^{\prime})=\pi\cup(D$-link)))), where $\pi$ is the adhesion-set of $ac$.
+    $S(A^{\prime},B^{\prime})=\pi\cup(D$-link)), where $\pi$ is the adhesion-set of $ac$.
   3. (iii)
@@
 Proof.
-The proof is analogous to the proof of Lemma 9.44, except that everything is straightforward because the two 3-connected torsos at the ends of $e$ both have disjoint adhesion-sets and have all their vertices contained in $A\cap C$.
+The proof is analogous to the proof of Lemma˜9.44, except that everything is straightforward because the two 3-connected torsos at the ends of $e$ both have disjoint adhesion-sets and have all their vertices contained in $A\cap C$.
 ∎
@@
 Let $i:=1$ if $H_{t(1)}$ is 3-connected, and $i:=2$ otherwise.
-Similarly, let $j:=-1$ if $H_{t(-1)}$ is 3-connected, and $j:=-2$ otherwise.
-The right-shift of $(U^{\prime}_{t},W^{\prime}_{t})$ has sides
+Similarly, let j:=1j:=-1 if $H_{t(-1)}$ is 3-connected, and j:=2j:=-2 otherwise.
+The right-reduction of $(U^{\prime}_{t},W^{\prime}_{t})$ has sides
 $V(G_{t(0)})\quad\text{and}\quad\bigcup\,\{\,V(G_{t(\ell)}):i\leqslant\ell\leqslant j\text{ in the cyclic ordering of }\mathbb{Z}_{m+1}\,\}.$
 Proof.
-This is just a different perspective on Lemma 9.14. ∎
+This is just a different perspective on Lemma˜9.14. ∎
 Definition 9.47 (Tetra-star $\sigma(\mathcal{O})$).
@@
 Let $(E,F)$ be a totally-nested tetra-separation of $G$ that interlaces $\sigma$.
-Then there is a 3-connected torso $H_{t}$ of $\mathcal{O}$ such that $(E,F)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$ or $(F,E)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, where $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$ denotes the right-shift of $(U^{\prime}_{t},W^{\prime}_{t})$.
+Then there is a 3-connected torso $H_{t}$ of $\mathcal{O}$ such that $(E,F)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$ or $(F,E)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, where $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$ denotes the right-reduction of $(U^{\prime}_{t},W^{\prime}_{t})$.
 Proof.
@@
 Otherwise the torso $H_{x}$ is a triangle or 4-cycle.
-By Lemma 9.18, $H_{x}$ is a triangle and both the $A$-link and the $C$-link are represented by triangle-torsos.
+By Lemma˜9.18, $H_{x}$ is a triangle and both the $A$-link and the $C$-link are represented by triangle-torsos.
 In particular, $A\cap C=3$.
@@
 (9.44) produces a $G$-lift.
-This concludes the proof that 7 holds.
-For every admissible $(M,N)$ we use 7 to choose a $G$-lift $(\bar{M},\bar{N})$ so that for $(N,M)$ we choose $(\bar{N},\bar{M})$.
+This concludes the proof that ˜7 holds.
+For every admissible $(M,N)$ we use ˜7 to choose a $G$-lift $(\bar{M},\bar{N})$ so that for $(N,M)$ we choose $(\bar{N},\bar{M})$.
 We claim that exactly one of $E\subseteq\bar{M}$ and $E\subseteq\bar{N}$ holds, for all $(M,N)$.
@@
 We let
-$\omega$ :=${}:=\{\,(M,N):(M,N)\text{ is admissable and }\bar{N}\supseteq E\,\}$
-$\Omega$ :=${}:=\bigcap\,\{\,N\smallsetminus M:(M,N)\in\omega\,\}\cap V(T^{\prime}).$
+$\omega$ ${}:=\{\,(M,N):(M,N)\text{ is admissable and }\bar{N}\supseteq E\,\}$
+$\Omega$ ${}:=\bigcap\,\{\,N\smallsetminus M:(M,N)\in\omega\,\}\cap V(T^{\prime}).$
 We claim that $\Omega\neq\emptyset$. Assume for a contradiction that $\Omega$ is empty.
@@
 Hence $T$ is a $K_{2}$ both whose endvertices have triangle or 4-cycle torsos that represent links.
-By Lemma 9.4, both torsos are triangles.
+By Lemma˜9.4, both torsos are triangles.
 Hence the corner $AC$ is potter, so $A\cap C=2$, contradicting that $E\subseteq A\cap C$ where $E$ has size $\geqslant 4$.
@@
 Hence $\Omega$ consists of exactly one node $t^{*}$, and its torso $H_{t^{*}}$ is 3-connected.
-To show that $(E,F)\leqslant$ the right-shift of $(U^{\prime}_{t},W^{\prime}_{t})$, by Lemma 9.46 and Lemma 5.7 it remains to show $E\subseteq V(G_{t^{*}})$.
+To show that $(E,F)\leqslant$ the right-reduction of $(U^{\prime}_{t},W^{\prime}_{t})$, by Lemma˜9.46 and Lemma˜5.7 it remains to show $E\subseteq V(G_{t^{*}})$.
 Assume for a contradiction that there is a vertex $v\in E\smallsetminus V(G_{t^{*}})$.
@@
 Let $r$ be the penultimate node on the path $sTt^{*}$.
-Assume first that the torso $H_{r}$ is 3-connected.
-Then the edge $rt^{*}$ is admissible.
-Let $(M,N)\in\omega$ with $S(M,N)=\{rt^{*}\}$.
-Note that $r\in M\smallsetminus N$ and $t^{*}\in N\smallsetminus M$ since $\{t^{*}\}=\Omega$.
-Let $(\hat{M},\hat{N})$ be the $\beta$-lift of $(M,N)$.
-Observe that $v\in\hat{M}\smallsetminus\hat{N}$.
-Therefore, $v\in\bar{M}\smallsetminus\bar{N}$, contradicting that $E\subseteq\bar{N}$.
-Assume otherwise that the torso $H_{r}$ is either a triangle or a 4-cycle.
-Then $r$ is admissible or $r$ is one of $o(A)$ and $o(C)$.
-Assume first that $r$ is admissible.
-Let $(M,N)\in\omega$ with $S(M,N)=\{r\}$.
-Note that $r,s\in M$ and $t^{*}\in N\smallsetminus M$.
-Let $(\hat{M},\hat{N})$ be the $\beta$-lift of $(M,N)$.
-Then $v\in\hat{M}\smallsetminus\hat{N}$, so $v\in\bar{M}\smallsetminus\bar{N}$, contradicting that $E\subseteq\bar{N}$.
-Assume otherwise that $r=o(A)$, say.
-Then $H_{r}$ is a triangle or 4-cycle representing the $A$-link.
-Since $v$ does not lie in $H_{t^{*}}$, we in particular have that $v$ does not lie in the adhesion-set of the edge $rt^{*}$.
-But then $v$ lies in $D\smallsetminus C$, contradicting that $E\subseteq C$.
-∎
+[…]
+[…]
 Lemma 9.49.
@@
 Assume for a contradiction that $\sigma$ is interlaced by a totally-nested tetra-separation $(E,F)$ of $G$.
-By Lemma 9.48, there is a 3-connected torso $H_{t}$ of $\mathcal{O}$ such that $(E,F)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, say, where $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$ denotes the right-shift of $(U^{\prime}_{t},W^{\prime}_{t})$.
+By Lemma˜9.48, there is a 3-connected torso $H_{t}$ of $\mathcal{O}$ such that $(E,F)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, say, where $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$ denotes the right-reduction of $(U^{\prime}_{t},W^{\prime}_{t})$.
 Assume first that $H_{t}$ is good.
-Then $(U_{t},W_{t})$ is a totally-nested tetra-separation by 9.41.
+Then $(U_{t},W_{t})$ is a totally-nested tetra-separation by ˜9.41.
 In particular, $(U_{t},W_{t})\in\sigma$.
 Since $(E,F)$ interlaces $\sigma$, we have one of $(U_{t},W_{t})<(E,F)$ and $(U_{t},W_{t})<(F,E)$.
-We cannot have $(U_{t},W_{t})<(F,E)$, since otherwise $\emptyset\neq E\smallsetminus F\subseteq(W_{t}\smallsetminus U_{t})\cap(U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t})$ contradicts the fact that $U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t}\subseteq U_{t}\smallsetminus W_{t}$ by 5.2.
+We cannot have $(U_{t},W_{t})<(F,E)$, since otherwise $\emptyset\neq E\smallsetminus F\subseteq(W_{t}\smallsetminus U_{t})\cap(U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t})$ contradicts the fact that $U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t}\subseteq U_{t}\smallsetminus W_{t}$ by ˜5.2.
 Hence $(U_{t},W_{t})<(E,F)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$.
-Since $(U_{t},W_{t})$ is the left-shift of $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, some vertex in $S(E,F)$ must have $\leqslant 1$ neighbour in $E\smallsetminus F$, contradicting the degree-condition.
+Since $(U_{t},W_{t})$ is the left-reduction of $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, some vertex in $S(E,F)$ must have $\leqslant 1$ neighbour in $E\smallsetminus F$, contradicting the degree-condition.
 Assume otherwise that $G_{t}$ is bad.
@@
 Since $(E,F)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, we have $E\smallsetminus F\subseteq U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t}$.
-Therefore, $|E\smallsetminus F|\leqslant 1$, which by Lemma 5.5 contradicts that $(E,F)$ is a tetra-separation.
+Therefore, $|E\smallsetminus F|\leqslant 1$, which by Lemma˜5.5 contradicts that $(E,F)$ is a tetra-separation.
 ∎
@@
 (4.1), all links have size two.
-By Lemma 9.10, $(A,B)$ and $(C,D)$ induce a Tutte-bagel $\mathcal{O}$.
-By 9.41, the tetra-star $\sigma(\mathcal{O})$ is a star of totally-nested tetra-separations of $G$.
-By Lemma 9.49, the tetra-star $\sigma(\mathcal{O})$ is a splitting star of $N$.
+By Lemma˜9.10, $(A,B)$ and $(C,D)$ induce a Tutte-bagel $\mathcal{O}$.
+By ˜9.41, the tetra-star $\sigma(\mathcal{O})$ is a star of totally-nested tetra-separations of $G$.
+By Lemma˜9.49, the tetra-star $\sigma(\mathcal{O})$ is a splitting star of $N$.
 Since both $(A,B)$ and $(C,D)$ interlace both $\sigma$ and $\sigma(\mathcal{O})$, we obtain $\sigma=\sigma(\mathcal{O})$.
 Hence it remains to compute the torso $\tau$ of $\sigma(\mathcal{O})$.
-The torso $\tau$ is 4-connected by Lemma 8.36.
+The torso $\tau$ is 4-connected by Lemma˜8.36.
 We can obtain $\tau$ from $\mathcal{O}$ by replacing each good torso with a $K_{4}$, leaving only 3-connected torsos with five or four vertices, and contracting some free edges in torsos that are 4-cycles or triangles.
-Each torso that is a 4-cycle neighbours two 3-connected torsos by Lemma 9.4, and the 4-cycle-torso witnesses that both 3-connected torsos are good.
+Each torso that is a 4-cycle neighbours two 3-connected torsos by Lemma˜9.4, and the 4-cycle-torso witnesses that both 3-connected torsos are good.
 Hence the free edges of the 4-cycle-torsos are contained in the separators of the elements of $\sigma(\mathcal{O})$ that stem from the good neighbouring torsos, and so they are contracted.
@@
 Phrased differently, a triangle-strip is a component of
-[…]
+$O[\,\{t\in V(O):t\text{ has a triangle-torso}\}\,].$
 Triangle-strips are paths unless all torsos of $O$ are triangles. An element of $V(O)\cup E(O)$ is furious if either
@@
 The four links injectively correspond to furious edges or vertices or nodes in triangle-strips.
-Hence it suffices to show that there is no path $rst$ in $O$ with all three nodes $r,s,t$ having triangle-torsos representing links.
+Hence it suffices to show that there is no path rstrst in $O$ with all three nodes $r,s,t$ having triangle-torsos representing links.
 Assume otherwise.
-Then two adjacent corners are potter, contradicting 4.7.
+Then two adjacent corners are potter, contradicting ˜4.7.
 ∎
@@
 We may assume that $O$ traverses the $x(i)$ in the cyclic ordering of $\mathbb{Z}_{4}$.
-Let $P_{A},P_{B}$ denote the components of $O-x(0)-x(2)$, let $A:=\bigcup_{t\in V(P_{A})}V(G_{t})$ and define $B$ similarly.
+Let $P_{A},P_{B}$ denote the components of Ox(0)x(2)O-x(0)-x(2), let $A:=\bigcup_{t\in V(P_{A})}V(G_{t})$ and define $B$ similarly.
 We use $x(1),x(3)$ to define $P_{C},P_{D}$ and $C,D$ analogously.
@@
 Case: $x(0)$ is a furious vertex.
-So $H_{x(0)}$ is a 4-cycle and $S(0)$ consists of two edges, and we only have to show (ii).
-The neighbouring torso $H_{t(1)}$ must be 3-connected by Lemma 9.4, so the adhesion-set of $x(0)t(1)$ is included in $A\smallsetminus B$ and avoids $\widehat{S(2)}$.
-Hence (ii) holds.
+So $H_{x(0)}$ is a 4-cycle and $S(0)$ consists of two edges, and we only have to show ˜(ii).
+The neighbouring torso $H_{t(1)}$ must be 3-connected by Lemma˜9.4, so the adhesion-set of $x(0)t(1)$ is included in $A\smallsetminus B$ and avoids $\widehat{S(2)}$.
+Hence ˜(ii) holds.
 Case: $x(0)$ is a furious edge.
-Then $S(0)$ is the adhesion-set of the edge $x(0)$, and we only have to show (i).
+Then $S(0)$ is the adhesion-set of the edge $x(0)$, and we only have to show ˜(i).
 Note that $t(1)$ is an end of $x(0)$, so the torso $H_{t(1)}$ is 3-connected.
@@
 But then $t(2)$ is the first candidate for $x(1)$, so $x(2)$ comes after $t(2)$.
-If $H_{t(3)}$ is not a triangle, we are done, and otherwise the tip of the triangle $H_{t(3)}$ avoids $\pi$ by Lemma 9.15.
+If $H_{t(3)}$ is not a triangle, we are done, and otherwise the tip of the triangle $H_{t(3)}$ avoids $\pi$ by Lemma˜9.15.
 Hence $\pi\subseteq A\smallsetminus B$ as desired.
@@
 Hence the edge $t(1)t(2)$ is the first candidate for $x(1)$, and so the edge $t(2)t(3)$ is the first candidate for $x(2)$.
-Therefore, the entire vertex-set of $H_{t(1)}$ is included in $A\smallsetminus B$, which ensures (i) and (ii).
+Therefore, the entire vertex-set of $H_{t(1)}$ is included in $A\smallsetminus B$, which ensures ˜(i) and ˜(ii).
 Assume otherwise that $H_{t(1)}$ is a triangle.
@@
 Note that $\pi$ contains an end of the free edge in $S(0)$ and both vertices in $\pi$ are neighbours of the tip of the triangle $H_{x(0)}$.
-Hence to show both (i) and (ii), it suffices to show that $\pi$ is included in $A\smallsetminus B$.
+Hence to show both ˜(i) and ˜(ii), it suffices to show that $\pi$ is included in $A\smallsetminus B$.
 The node $t(1)$ is the first candidate for $x(1)$, so $t(2)$ is the first candidate for $x(2)$.
-By Lemma 9.4, $H_{t(2)}$ is either a triangle or 3-connected.
+By Lemma˜9.4, $H_{t(2)}$ is either a triangle or 3-connected.
 If $H_{t(2)}$ is 3-connected, then $\pi$ is included in $A\smallsetminus B$.
@@
 If $H_{t(3)}$ is 3-connected, then $\pi$ is included in $A\smallsetminus B$.
-Otherwise, if $H_{t(3)}$ is a triangle, then its tip is distinct from the tip of $H_{t(2)}$ by Lemma 9.15, and $t(3)$ is the first candidate for $x(2)$.
+Otherwise, if $H_{t(3)}$ is a triangle, then its tip is distinct from the tip of $H_{t(2)}$ by Lemma˜9.15, and $t(3)$ is the first candidate for $x(2)$.
 So $\pi$ is included in $A\smallsetminus B$.
-This completes the proof of (i) and (ii).
+This completes the proof of ˜(i) and ˜(ii).
 Hence $(A,B)$ and $(C,D)$ are tetra-separations with $S(A,B)=S(0)\cup S(2)$ and $S(C,D)=S(1)\cup S(3)$.
@@
 Proof.
-(i) $\Rightarrow$ (ii).
+˜(i) $\Rightarrow$ ˜(ii).
 Clearly, $G\neq K_{5}$ since $K_{5}$ does not have a tetra-separation.
-By Lemma 9.10, $(A,B)$ and $(C,D)$ induce a Tutte-bagel $\mathcal{O}$.
-By 9.41, the tetra-star $\sigma(\mathcal{O})$ is a star of totally-nested tetra-separations of $G$.
+By Lemma˜9.10, $(A,B)$ and $(C,D)$ induce a Tutte-bagel $\mathcal{O}$.
+By ˜9.41, the tetra-star $\sigma(\mathcal{O})$ is a star of totally-nested tetra-separations of $G$.
 Since $G$ has no totally-nested tetra-separations by assumption, all torsos of $\mathcal{O}$ are bad.
-We get $\alpha(\mathcal{O})\geqslant 4$ from Lemma 9.51.
-(ii) $\Rightarrow$ (i) is Lemma 9.52. ∎
+We get $\alpha(\mathcal{O})\geqslant 4$ from Lemma˜9.51.
+˜(ii) $\Rightarrow$ ˜(i) is Lemma˜9.52. ∎
 10. Quasi-5-connected torsos
-This section deals with the quasi-5-connected outcome of Theorem 1; see 10.5.
+This section deals with the quasi-5-connected outcome of Theorem˜1; see ˜10.5.
 A mixed-separation+ $(A,B)$ of a graph $G$ almost interlaces a star $\sigma$ of mixed-separations+ of $G$ if $(C,D)\leqslant(A,B)$ or $(C,D)\leqslant(B,A)$ for all $(C,D)\in\sigma$.
-The difference between ‘interlaces’ and ‘almost interlaces’ is ‘$<$’ versus $\leqslant$’ in the definitions.
+The difference between ‘interlaces’ and ‘almost interlaces’ is ‘<<’ versus $\leqslant$’ in the definitions.
 Lemma 10.1.
-(Analogue of [Tridecomp, Lemma 2.6.7]) Let $G$ be a 4-connected graph.
-Let $\sigma$ be a star of mixed-4-separations of $G$ satisfying the matching-condition.
-Let $(A,B)$ be a separation of the torso $\tau$ of $\sigma$.
-Then there exists a mixed-separation $(\hat{A},\hat{B})$ of $G$ such that:
+[…]
   1. (i)
-    the order of $(\hat{A},\hat{B})$ is at most the order of $(A,B)$;
+[…]
   2. (ii)
-    $(\hat{A},\hat{B})$ almost interlaces $\sigma$;
+[…]
   3. (iii)
@@
 Proof.
-The proof is the same as the proof of [Tridecomp, Lemma 2.6.7], with one exception: we replace Lemma 1.2.3 with the matching-condition. ∎
+The proof is the same as the proof of [CK23, Lemma 2.6.7], with one exception: we replace Lemma 1.2.3 with the matching-condition. ∎
 Lemma 10.2.
@@
 Let $(C,D)$ be a mixed-4-separation+ in $G$ with $(A,B)\leqslant(C,D)$.
-Let $(C^{\prime},D^{\prime})$ denote either the left-right-shift or the right-left-shift of $(C,D)$.
+Let $(C^{\prime},D^{\prime})$ denote either the left-right-reduction or the right-left-reduction of $(C,D)$.
 Then $(A,B)\leqslant(C^{\prime},D^{\prime})$.
@@
 Suppose that $|W\cap(A\smallsetminus B)|\geqslant 1$.
-Let $(A^{\prime},B^{\prime})$ be either the left-right-sift or the right-left-shift of $(A,B)$.
+Let $(A^{\prime},B^{\prime})$ be either the left-right-sift or the right-left-reduction of $(A,B)$.
 Then $(U,W)<(A^{\prime},B^{\prime})$.
 Proof.
-We have $(U,W)\leqslant(A^{\prime},B^{\prime})$ by Lemma 10.2.
-Recall that $A\smallsetminus B\subseteq A^{\prime}\smallsetminus B^{\prime}$ by 5.9.
+We have $(U,W)\leqslant(A^{\prime},B^{\prime})$ by Lemma˜10.2.
+Recall that $A\smallsetminus B\subseteq A^{\prime}\smallsetminus B^{\prime}$ by ˜5.9.
 Hence $W\cap(A^{\prime}\smallsetminus B^{\prime})$ is nonempty, so in particular the inclusion $W\supseteq B^{\prime}$ is proper.
@@
 Proof.
-By Lemma 10.1, there is a mixed-$k$-separation+ $(\hat{A},\hat{B})$ of $G$ such that
+[…]
   • $\cdot$
-    $k\leqslant 4$ by (i);
+    $k\leqslant 4$ by ˜(i);
   • $\cdot$
-    $(\hat{A},\hat{B})$ almost interlaces $\sigma$ by (ii);
+[…]
   • $\cdot$
-    $|\hat{A}\smallsetminus\hat{B}|\geqslant 2$ and $|\hat{B}\smallsetminus\hat{A}|\geqslant 2$ by (iii); and
+    $|\hat{A}\smallsetminus\hat{B}|\geqslant 2$ and $|\hat{B}\smallsetminus\hat{A}|\geqslant 2$ by ˜(iii); and
   • $\cdot$
-    for every $(U,W)\in\sigma$ we have $|W\cap(\hat{A}\smallsetminus\hat{B})|\geqslant 2$ and $|W\cap(\hat{B}\smallsetminus\hat{A})|\geqslant 2$ by (iv).
-Since $G$ is 4-connected, $(\hat{A},\hat{B})$ is a mixed-4-separation by Lemma 6.1.
-Let $(\bar{A},\bar{B})$ denote the left-right-shift of $(\hat{A},\hat{B})$.
-Then $(\bar{A},\bar{B})$ is a tetra-separation by 5.10.
-By Lemma 10.3, $(\bar{A},\bar{B})$ interlaces $\sigma$.
-∎
+    for every $(U,W)\in\sigma$ we have $|W\cap(\hat{A}\smallsetminus\hat{B})|\geqslant 2$ and $|W\cap(\hat{B}\smallsetminus\hat{A})|\geqslant 2$ by ˜(iv).
+[…]
 Key Lemma 10.5.
@@
 Proof.
-By Lemma 8.36, $\tau$ is 4-connected or a $K_{4}$, and we are done in the latter case. By the contrapositive of Lemma 10.4, every 4-separation of $\tau$ has a side of size five. ∎
+By Lemma˜8.36, $\tau$ is 4-connected or a $K_{4}$, and we are done in the latter case. By the contrapositive of Lemma˜10.4, every 4-separation of $\tau$ has a side of size five. ∎
 11. An Angry Theorem for 4-connectivity
 A key step in proving Tutte’s decomposition theorem is to characterise the graphs that are 2-angry: those that are 2-connected and all whose 2-separators are crossed – hence the name.
-Tutte proved the first angry theorem: the 2-angry graphs are precisely the 3-connected graphs and the cycles [TutteCon].
-For a modern proof, see [Tridecomp, Theorem 3.3.4].
-An angry theorem for 3-connectivity has been proved using tri-separations [Tridecomp, Theorem 1.1.5].
+Tutte proved the first angry theorem: the 2-angry graphs are precisely the 3-connected graphs and the cycles [T66].
+For a modern proof, see [CK23, Theorem 3.3.4].
+An angry theorem for 3-connectivity has been proved using tri-separations [CK23, Theorem 1.1.5].
 However, like here, the angry theorem was not used in the proof of the decomposition theorem.
-Still, angry theorems have recently found an application in group theory: the local version of the 2-angry theorem [Local2sep] was used as a key ingredient in [StallingsNilpotent], and versions for higher connectivity are expected to be needed to continue in the direction of [StallingsNilpotent].
+Still, angry theorems have recently found an application in group theory: the local version of the 2-angry theorem [C22] was used as a key ingredient in [CKKT24], and versions for higher connectivity are expected to be needed to continue in the direction of [CKKT24].
 Here we prove an angry theorem for 4-connectivity.
@@
 Assume that $G$ is 4-angry.
-If $G$ has no tetra-separation, then (i) by Lemma 2.3.
+If $G$ has no tetra-separation, then ˜(i) by Lemma˜2.3.
 So $G$ has a tetra-separation $(A,B)$, which is crossed by a tetra-separation $(C,D)$ as $G$ is 4-angry.
@@
   • $\cdot$
-    If $\ell=0$, then (iv) by 7.8.
+    If $\ell=0$, then ˜(iv) by ˜7.8.
   • $\cdot$
-    If $\ell=1$, then (iii) by 8.41.
+    If $\ell=1$, then ˜(iii) by ˜8.41.
   • $\cdot$
-    If $\ell=2$, then (ii) by 9.53 and since $G\neq K_{5}$ by $|G|\geqslant 8$.
-Backward implication. We use one of Lemma 2.3, 7.8, 8.41 and 9.53. Here Lemma 2.3 needs $|G|\geqslant 8$. ∎
+    If $\ell=2$, then ˜(ii) by ˜9.53 and since $G\neq K_{5}$ by $|G|\geqslant 8$.
+Backward implication. We use one of Lemma˜2.3, ˜7.8, ˜8.41 and ˜9.53. Here Lemma˜2.3 needs $|G|\geqslant 8$. ∎
+[…]
+Lemma 12.1.
+Let $G$ be a quasi-4-connected graph.
+Let $(A^{\Delta},B^{\Delta})$ be a $k$-separation of $G^{\Delta}$.
+Let $U$ be the set of all degree-three vertices of $G$.
+We obtain $A^{s}$ from $A^{\Delta}$ by adding every vertex $u\in U$ with $\Delta(u)\subseteq G^{\Delta}[A^{\Delta}]$.
+We obtain $B^{s}$ from $B^{\Delta}$ by adding every vertex $u\in U$ with $\Delta(u)\subseteq G^{\Delta}[B^{\Delta}]$ but $\Delta(u)\not\subseteq G^{\Delta}[A^{\Delta}]$.
+Then $(A^{s},B^{s})$ is a $k$-separation of $G^{s}$ with $A^{s}\cap V(G^{\Delta})=A^{\Delta}$ and $B^{s}\cap V(G^{\Delta})=B^{\Delta}$ and $A^{s}\cap B^{s}=A^{\Delta}\cap B^{\Delta}$, see [CK23, Lemma 2.6.4].
+Let $A:=A^{s}\cap V(G)$ and $B:=B^{s}\cap V(G)$.
+Then $(A,B)$ is a mixed-separation of $G$ of order $\leqslant k$.∎
+Lemma 12.2.
+Let $G$ be a quasi-4-connected graph on $\geqslant 7$ vertices. Then $G^{\Delta}$ is 4-connected.
+Proof.
+Assume for a contradiction that $G^{\Delta}$ is not 4-connected.
+Let $(A^{\Delta},B^{\Delta})$ be a separation of $G^{\Delta}$ of least possible order $k$, so $k\leqslant 3$.
+From $(A^{\Delta},B^{\Delta})$ we obtain a mixed-separation $(A,B)$ of $G$ of order $\leqslant k$ as in Lemma˜12.1.
+We claim that $A\smallsetminus B$ and $B\smallsetminus A$ have size $\geqslant 2$.
+It suffices to show that $|A\smallsetminus B|\geqslant 2$ (the proof for $|B\smallsetminus A|\geqslant 2$ is a subproof of that for $|A\smallsetminus B|\geqslant 2$).
+Since $(A^{\Delta},B^{\Delta})$ is proper, there is a vertex $x\in A^{\Delta}\smallsetminus B^{\Delta}$.
+We distinguish two cases.
+Assume first that $x$ is a subdividing vertex of an edge $e=uv$ of $G$.
+Consider the triangles $\Delta(u),\Delta(v)\subseteq G^{\Delta}$ that were added around $u$ and $v$, respectively.
+Both contain the vertex $x$, so both are included in $G^{\Delta}[A^{\Delta}]$ but not in $G^{\Delta}[B^{\Delta}]$.
+Hence both $u$ and $v$ lie in $A^{s}\smallsetminus B^{s}$.
+Since both $u$ and $v$ are in $G$, they also lie in $A\smallsetminus B$.
+Hence $|A\smallsetminus B|\geqslant 2$.
+Assume otherwise that $x$ is a vertex in $G^{\Delta}$ that is also a vertex in $G$.
+Then $x$ is not contained in $U$, so $x$ has degree $\geqslant 4$ in $G$.
+Now $x\in A^{s}\smallsetminus B^{s}$, and $x\in A\smallsetminus B$.
+Since $(A,B)$ has order $\leqslant 3$, at least one of the four neighbours of $x$ in $G$ also lies in $A\smallsetminus B$.
+Hence $|A\smallsetminus B|\geqslant 2$.
+Therefore, both $A\smallsetminus B$ and $B\smallsetminus A$ have size $\geqslant 2$. Hence $G$ is not quasi-4-connected by Lemma˜5.11, a contradiction. ∎
+Example 12.3.
+The number 7 in Lemma˜12.2 is optimal. To see this, let $G$ be a cycle of $C_{4}$-bags of length 3. Then $G^{\Delta}$ is a generalised 4-wheel, which is 3-connected but not quasi-4-connected.
 13. Vertex-transitive graphs
-[…]
+In this section, we prove Figure˜1.8.
+Recall that an $H$-expansion of a $|V(H)|$-regular graph $G$ is a graph obtained from $G$ by replacing the vertices $v$ of $G$ with pairwise disjoint copies $H_{v}$ of $H$ and by replacing every edge $e=uv$ of $G$ with an $H_{u}$$H_{v}$ edge $\hat{e}$ so that $\hat{e}_{1}$ and $\hat{e}_{2}$ share no ends for every two $e_{1}\neq e_{2}\in E(G)$.
+Thus, the edges $\hat{e}$ form a perfect matching of the $H$-expansion.
+Note that if $H$ is complete then the $H$-expansion of $G$ is unique up to isomorphism.
+The clique-expansion of a graph $G$ is the graph $G^{\prime}$ that is obtained from $G$ by replacing every vertex $v\in V(G)$ by a clique $C_{v}$ of size the degree of $v$, and every edge $e=uv$ by a $C_{u}$$C_{v}$ edge $\hat{e}$ such that these edges $\hat{e}$ form a perfect matching in $G^{\prime}$.
 Recall that a graph $G$ is arc-transitive if for every two ordered pairs $(u_{1},v_{1}),(u_{2},v_{2})$ of vertices forming edges $u_{1}v_{1},u_{2}v_{2}\in E(G)$, there is an automorphism $\varphi$ of $G$ such that $(\varphi(u_{1}),\varphi(v_{1}))=(u_{2},v_{2})$.
 Theorem 13.1.
-[Godsil, Theorem 3.4.2] Every finite vertex-transitive $d$-regular graph is $k$-connected for $k:=\lceil\frac{2}{3}(d+1)\rceil$.
+[GR01, Theorem 3.4.2] Every finite vertex-transitive $d$-regular graph is $k$-connected for $k:=\lceil\frac{2}{3}(d+1)\rceil$.
 Corollary 13.2.
@@
   • $\cdot$
-    cycle of $K_{4}$-bags of length $\ell$ if all bags of $\mathcal{O}$ are $K_{4}$’s and all adhesion-sets of $\mathcal{O}$ are disjoint (Figure 1.7);
+    cycle of $K_{4}$-bags of length $\ell$ if all bags of $\mathcal{O}$ are $K_{4}$’s and all adhesion-sets of $\mathcal{O}$ are disjoint (Figure˜1.9);
   • $\cdot$
-    cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\ell$ if the bags along $O$ alternate between being $K_{4}$’s and being 4-cycles, and all adhesion-sets of $\mathcal{O}$ are disjoint (Figure 1.8);
+    cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\ell$ if the bags along $O$ alternate between being $K_{4}$’s and being 4-cycles, and all adhesion-sets of $\mathcal{O}$ are disjoint (Figure˜1.10);
   • $\cdot$
-    cycle of triangle-bags of length $\ell$ if all bags are triangles, and every two non-consecutive edges on $O$ have disjoint adhesion-sets (Figure 1.9);
+    cycle of triangle-bags of length $\ell$ if all bags are triangles, and every two non-consecutive edges on $O$ have disjoint adhesion-sets (Figure˜1.12);
   • $\cdot$
-    cycle of $K_{2,2}$-bags of length $\ell$ if all bags of $\mathcal{O}$ are $K_{2,2}$’s, all torsos of $\mathcal{O}$ are $K_{4}$’s and all adhesion-sets of $\mathcal{O}$ are disjoint (Figure 1.10).
-Observation 12.4.
+    cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length $\ell$ if the torsos along $O$ alternate between being $K_{4}$’s and $C_{4}$’s, and every adhesion-set of $\mathcal{O}$ is independent and all adhesion-sets of $\mathcal{O}$ are disjoint (so the $K_{4}$-torsos have $K_{2,2}$-bags) (Figure˜1.11);
+  • $\cdot$
+    cycle of $C_{4}$-bags of length $\ell$ if all bags of $\mathcal{O}$ are $C_{4}$’s and all adhesion-sets of $\mathcal{O}$ induce disjoint $K_{2}$’s (Figure˜1.13);
+  • $\cdot$
+    cycle of $K_{2,2}$-bags of length $\ell$ if all bags of $\mathcal{O}$ are $K_{2,2}$’s, all torsos of $\mathcal{O}$ are $K_{4}$’s and all adhesion-sets of $\mathcal{O}$ are disjoint (Figure˜1.14).
+Observation 13.4.
   • $\cdot$
@@
 Proof.
-Let $Z$ be the centre of the crossing-diagram, which consists of four vertices by the (Crossing Lemma). (4.1). Note that $G-Z$ has at least four components.
+Let $Z$ be the centre of the crossing-diagram, which consists of four vertices by the (Crossing Lemma). (4.1). Note that GZG-Z has at least four components.
 […]
-By 4-connectivity of $G$, each of the $\geqslant 4$ vertices of $G-Z$ has neighbourhood equal to $Z$.
+By 4-connectivity of $G$, each of the $\geqslant 4$ vertices of GZG-Z has neighbourhood equal to $Z$.
 So, $G$ is 4-regular by vertex-transitivity.
-In particular, $G-Z$ has exactly four vertices, and $G[Z]$ is independent.
+In particular, GZG-Z has exactly four vertices, and $G[Z]$ is independent.
 So, $G=K_{4,4}$.
@@
 (8.14) with $t=t(0)$, which we assume in the following.
-By 8.27, $(U,W)$ is a totally-nested tetra-separation of $G$.
-By Lemma 12.5, the separator $S(U,W)$ consists of four edges, and either $G[U]=K_{4}$ or $G[W]=K_{4}$.
+By ˜8.27, $(U,W)$ is a totally-nested tetra-separation of $G$.
+By Lemma˜13.5, the separator $S(U,W)$ consists of four edges, and either $G[U]=K_{4}$ or $G[W]=K_{4}$.
 It follows with vertex-transitivity that $G$ is 4-regular and that the neighbourhood around each vertex induces $K_{1}\sqcup K_{3}$ in $G$.
@@
 Write $\mathcal{O}=:(O,(G_{t})_{t\in V(O)})$, and let $u,v$ be the two vertices in the centre.
-By Lemma 12.8, every bag of $\mathcal{O}$ is bad.
+By Lemma˜13.8, every bag of $\mathcal{O}$ is bad.
 In particular, every bag of $\mathcal{O}$ is a $K_{2}$ or a triangle.
@@
 Proof.
-By Lemma 12.9, every bag of the block-bagel $\mathcal{O}=(O,\mathcal{G})$ induced by $(A,B)$ and $(C,D)$ is a $K_{2}$.
+By Lemma˜13.9, every bag of the block-bagel $\mathcal{O}=(O,\mathcal{G})$ induced by $(A,B)$ and $(C,D)$ is a $K_{2}$.
 Hence every edge of $O$ is helpful, so $|O|\geqslant 4$ by the (Helpful Lemma).
@@
 Let $u,v$ be the vertices in the centre.
-By Lemma 8.7 it follows that $|O|=4$ and $uv\notin E(G)$.
+By Lemma˜8.7 it follows that $|O|=4$ and $uv\notin E(G)$.
 Then $G$ is an octahedron.
@@
 (9.13) with $t=t(0)$.
-By 9.41$(U,W)$ is a totally-nested tetra-separation.
-By Lemma 12.5, $S(U,W)$ consists of four edges, and we have either $G[U]=K_{4}$ or $G[W]=K_{4}$.
+By ˜9.41(U,W)(U,W) is a totally-nested tetra-separation.
+By Lemma˜13.5, $S(U,W)$ consists of four edges, and we have either $G[U]=K_{4}$ or $G[W]=K_{4}$.
 We claim that $G[U]=K_{4}$.
@@
 Then $G[W]=G_{s}$ for some bag $G_{s}\neq G_{t}$.
-Hence $O$ has length $\leqslant 4$, two 3-connected bags (namely $G_{s},G_{t}$), and all other bags have triangles or 4-cycles as torsos by Lemma 9.14.
-But then $\alpha(\mathcal{O})<4$ contradicts Lemma 9.51.
+Hence $O$ has length $\leqslant 4$, two 3-connected bags (namely $G_{s},G_{t}$), and all other bags have triangles or 4-cycles as torsos by Lemma˜9.14.
+But then $\alpha(\mathcal{O})<4$ contradicts Lemma˜9.51.
 Hence $G[U]=K_{4}$.
 Next, we claim that $S(U,W)$ contains no inward edge, and assume for a contradiction that $x_{1}\in S(U,W)$ is an inward edge with endvertex $u$ in the interior of $G_{t}$.
-By Lemma 9.33, some $y^{\prime}_{i}$, say $y^{\prime}_{1}$, is a neighbour of $x^{\prime}_{1}$ and $y^{W}_{1}=y^{\prime}_{1}$.
+By Lemma˜9.33, some $y^{\prime}_{i}$, say $y^{\prime}_{1}$, is a neighbour of $x^{\prime}_{1}$ and $y^{W}_{1}=y^{\prime}_{1}$.
 In particular, $y_{1}$ cannot be an outward edge.
@@
 Write $O=t(0)\,t(1)\ldots t(m)\,t(0)$ with indices in $\mathbb{Z}_{m+1}$.
-By Lemma 12.11, if a torso such as $H_{t(0)}$ is good, then its bag is a $K_{4}$ and the two neighbouring torsos $H_{t(-1)}$ and $H_{t(1)}$ are 4-cycles.
-Then, by Lemma 9.4, the torsos $H_{t(-2)}$ and $H_{t(2)}$ are 3-connected.
+By Lemma˜13.11, if a torso such as $H_{t(0)}$ is good, then its bag is a $K_{4}$ and the two neighbouring torsos $H_{t(-1)}$ and $H_{t(1)}$ are 4-cycles.
+Then, by Lemma˜9.4, the torsos $H_{t(-2)}$ and $H_{t(2)}$ are 3-connected.
 Hence, both of these torsos are good as they neighbour torsos that are 4-cycles.
 By induction, we get that $O$ has even length, that bags $G_{t(i)}$ are $K_{4}$’s for even $i$, and that torsos $H_{t(i)}$ are 4-cycles for odd $i$.
-By Lemma 9.51, there are $\geqslant 4$ bags that are $K_{4}$’s.
+By Lemma˜9.51, there are $\geqslant 4$ bags that are $K_{4}$’s.
 It follows that $G$ is a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 8$.
@@
 Write $\mathcal{O}=:(O,\mathcal{G})$.
-By Lemma 12.12, all bags of $\mathcal{O}$ are bad.
+By Lemma˜13.12, all bags of $\mathcal{O}$ are bad.
 Let $H_{t}$ be a torso of $\mathcal{O}$ on five vertices.
@@
 It follows by induction that all torsos of $\mathcal{O}$ have exactly five vertices.
-By Lemma 12.13, each bag contains a 2-matching between its two adhesion-sets, and the interior vertex of the bag sends edges to all four other vertices in the bag.
+By Lemma˜13.13, each bag contains a 2-matching between its two adhesion-sets, and the interior vertex of the bag sends edges to all four other vertices in the bag.
 By 4-regularity and since every bag looks like this, these are all the edges that there are in the bags.
 Since each bag has a unique interior vertex, the set $X$ of all these vertices has size $|O|$.
-Since all torsos are 3-connected, we have $|O|\geqslant 4$ by Lemma 9.51.
+Since all torsos are 3-connected, we have $|O|\geqslant 4$ by Lemma˜9.51.
 If $|O|\geqslant 5$, then every vertex in $X$ is contained in exactly two induced 4-cycles in $G$, while every vertex in an adhesion set is contained in exactly one induced 4-cycle in $G$, contradicting vertex-transitivity.
 Therefore, we have $|O|=4$.
-Consider the graph $H:=G-X$, which is either an 8-cycle or a disjoint union of two 4-cycles.
-If $H$ is an 8-cycle, then every vertex in $X$ is contained in two induced 4-cycles of $G$ while every vertex in $V(G-X)$ is only contained in one induced 4-cycle of $G$, contradicting vertex-transitivity.
+Consider the graph H:=GXH:=G-X, which is either an 8-cycle or a disjoint union of two 4-cycles.
+If $H$ is an 8-cycle, then every vertex in $X$ is contained in two induced 4-cycles of $G$ while every vertex in V(GX)V(G-X) is only contained in one induced 4-cycle of $G$, contradicting vertex-transitivity.
 Therefore $H$ is a disjoint union of two 4-cycles, and it follows that $G$ is the line graph of the cube.
@@
 Proof.
-If $\mathcal{O}$ has a good torso, then $G$ is a cycle alternating between $K_{4}$-bags and $C_{4}$-bags by Lemma 12.12, and it is of length $\geqslant 8$ by Lemma 9.51.
+If $\mathcal{O}$ has a good torso, then $G$ is a cycle alternating between $K_{4}$-bags and $C_{4}$-bags by Lemma˜13.12, and it is of length $\geqslant 8$ by Lemma˜9.51.
 Hence, we may assume that all torsos of $\mathcal{O}$ are bad.
 We claim that $\mathcal{O}$ has no 4-cycle as torso.
-Indeed, if $\mathcal{O}$ has a 4-cycle as a torso, then the neighbouring torsos are 3-connected by Lemma 9.4, and they are good, a contradiction.
+Indeed, if $\mathcal{O}$ has a 4-cycle as a torso, then the neighbouring torsos are 3-connected by Lemma˜9.4, and they are good, a contradiction.
 Thus,
@@
 For this, let $H_{t}$ and $H_{s}$ be neighbouring torsos with $H_{t}$ a triangle.
-By 8, the torso $H_{s}$ is either a triangle or a $K_{4}$.
+By ˜8, the torso $H_{s}$ is either a triangle or a $K_{4}$.
 Assume for a contradiction that $H_{s}=K_{4}$.
@@
 Then $v$ has exactly one neighbour in the bag $G_{r}$.
-By 8, the torso $H_{r}$ can only be a triangle or a $K_{4}$.
+By ˜8, the torso $H_{r}$ can only be a triangle or a $K_{4}$.
 As $v$ has exactly one neighbour in the bag $G_{r}$, the torso $H_{r}$ cannot be a $K_{4}$, so it must be a triangle.
-But then the edge $vw$ exists by Lemma 9.4, so $v$ has two neighbours in $G_{r}$, giving it degree five in $G$, contradicting 4-regularity.
+But then the edge $vw$ exists by Lemma˜9.4, so $v$ has two neighbours in $G_{r}$, giving it degree five in $G$, contradicting 4-regularity.
 Therefore, every triangle-torso of $\mathcal{O}$ is neighboured by two triangle-torsos.
-So if $\mathcal{O}$ has at least one triangle-torsos, then all torsos are triangles, and all adhesion-sets are spanned by edges by Lemma 9.4.
-Hence $G$ is a cycle of triangle-bags, and it has length $\geqslant 6$ by Lemma 9.51 and since $G\neq K_{5}$.
+So if $\mathcal{O}$ has at least one triangle-torsos, then all torsos are triangles, and all adhesion-sets are spanned by edges by Lemma˜9.4.
+Hence $G$ is a cycle of triangle-bags, and it has length $\geqslant 6$ by Lemma˜9.51 and since $G\neq K_{5}$.
 So we may assume that $\mathcal{O}$ has no triangle-torsos.
-Then by 8, all torsos of $\mathcal{O}$ are $K_{4}$’s.
+Then by ˜8, all torsos of $\mathcal{O}$ are $K_{4}$’s.
 Then either all adhesion-sets are spanned by edges in $G$, in which case $G$ is a cycle of $K_{4}$-bags; or none of them are, in which case $G$ is a cycle of $K_{2,2}$-bags.
-In both cases, the length is $\geqslant 4$ by Lemma 9.51.
+In both cases, the length is $\geqslant 4$ by Lemma˜9.51.
 ∎
@@
 Proof.
-We combine Lemmas 12.14 and 12.15. ∎
+We combine Lemmas˜13.14 and 13.15. ∎
 Corollary 13.17.
@@
 By the (Crossing Lemma).
-(4.1), all links have the same size 00, $1$ or $2$.
-If all links have size 00, then $G$ is a cycle of $K_{2,2}$-bags of length $4$ by Lemma 12.7.
-If all links have size $1$, then $G$ is a cycle of $K_{2,2}$-bags of length $3$ by Lemma 12.10.
-Otherwise all links have size $2$, and we apply Lemma 12.16.
+(4.1), all links have the same size 0, 11 or $2$.
+If all links have size 0, then $G$ is a cycle of $K_{2,2}$-bags of length $4$ by Lemma˜13.7.
+If all links have size $1$, then $G$ is a cycle of $K_{2,2}$-bags of length $3$ by Lemma˜13.10.
+Otherwise all links have size $2$, and we apply Lemma˜13.16.
 ∎
-Observation 12.18.
-Let $G^{\prime}$ be the expansion of a graph $G$.
+Observation 13.18.
+Let $G^{\prime}$ be the clique-expansion of a graph $G$.
 Let $v$ be a vertex of degree $d$ in $G$, and let $C_{v}$ be the clique of $G^{\prime}$ that corresponds to $v$.
 Then every vertex in $C_{v}$ has degree $d$ in $G^{\prime}$.∎
-Observation 12.19.
-Let $G$ be a graph of minimum degree at least $2$, and let $G^{\prime}$ be the expansion of $G$.
+Observation 13.19.
+Let $G$ be a graph of minimum degree at least $2$, and let $G^{\prime}$ be the clique-expansion of $G$.
 Then there is a unique partition $\mathcal{P}$ of $V(G^{\prime})$ into (the vertex-sets of) $|\mathcal{P}|=|V(G)|$ cliques.
-Moreover, the contraction-minor $G^{\prime}/\mathcal{P}$ is isomorphic to $G$.∎
-Observation 12.20.
-The expansion of a graph $G$ is isomorphic to the graph $H$ with
-$V(H)$ =${}=\{(u,v),(v,u):uv\in E(G)\}\text{ and}$
-$E(H)$ =${}=\Big{\{}\{(u,v),(v,u)\}:uv\in E(G)\Big{\}}\,\cup\,\Big{\{}\{(u,v_{1}),(u,v_{2})\}:uv_{1},uv_{2}\in E(G)\Big{\}}.$
+Moreover, the contraction-minor $G^{\prime}/\mathcal{P}$ is isomorphic to GG.∎
+Observation 13.20.
+The clique-expansion of a graph $G$ is isomorphic to the graph $H$ with
+$V(H)$ ${}=\{(u,v),(v,u):uv\in E(G)\}\text{ and}$
+$E(H)$ ${}=\Big{\{}\{(u,v),(v,u)\}:uv\in E(G)\Big{\}}\,\cup\,\Big{\{}\{(u,v_{1}),(u,v_{2})\}:uv_{1},uv_{2}\in E(G)\Big{\}}.$
 Lemma 13.21.
@@
 […]
-(i) $\Rightarrow$ (ii).
+˜(i) $\Rightarrow$ ˜(ii).
 Let $(u_{1},v_{1}),(u_{2},v_{2})$ be ordered pairs of vertices in $G$ with $u_{i}v_{i}\in E(G)$.
@@
 Moreover, we have
-[…]
+$\text{ }xy\in E(G)$
 $\Leftrightarrow$ there is a $\gamma^{-1}(x)$$\gamma^{-1}(y)$ edge in $G^{\prime}$
@@
 Similarly, $\varphi(v_{1})=v_{2}$.
-(ii) $\Rightarrow$ (i).
-We assume without loss of generality that $G^{\prime}$ takes the form described in 12.20.
+˜(ii) $\Rightarrow$ ˜(i).
+We assume without loss of generality that $G^{\prime}$ takes the form described in ˜13.20.
 Let $x_{1},x_{2}\in V(G^{\prime})$ be two vertices.
 We will find an automorphism $\varphi^{\prime}$ of $G^{\prime}$ that maps $x_{1}$ to $x_{2}$.
-We write $x_{i}=:(u_{i},v_{i})$ using 12.20, so $u_{i}v_{i}$ is an edge of $G$.
+We write $x_{i}=:(u_{i},v_{i})$ using ˜13.20, so $u_{i}v_{i}$ is an edge of $G$.
 Since $G$ is arc-transitive, there is an automorphism $\varphi$ of $G$ such that $\varphi(u_{1})=u_{2}$ and $\varphi(v_{1})=v_{2}$.
@@
 Proof.
-By Lemma 12.21, it suffices to show that $G$ is not arc-transitive.
+By Lemma˜13.21, it suffices to show that $G$ is not arc-transitive.
 Indeed, if $u$ is a vertex in $A$, then $G[N(u)]=K_{1}\sqcup K_{r-1}$ with $r-1\geqslant 2$.
@@
 Proof.
-Follows from Corollary 12.22. ∎
+Follows from Corollary˜13.22. ∎
 Corollary 13.24.
@@
 In particular, it is not arc-transitive.
-By Lemma 12.21, the $K_{4}$-expansion of a cycle of triangle-bags of length $\geqslant 6$ is not vertex-transitive.
+By Lemma˜13.21, the $K_{4}$-expansion of a cycle of triangle-bags of length $\geqslant 6$ is not vertex-transitive.
 ∎
@@
 Proof.
-This follows from Corollary 12.17 together with the following observations.
+This follows from Corollary˜13.17 together with the following observations.
 First, a cycle of $K_{4}$-bags of length $\geqslant 4$ is not 4-regular.
-Second, a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 8$ is not arc-transitive by Corollary 12.22.
-Finally, a cycle of triangle-bags of length $\geqslant 6$ is not arc-transitive by Corollary 12.24.
+Second, a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 8$ is not arc-transitive by Corollary˜13.22.
+Finally, a cycle of triangle-bags of length $\geqslant 6$ is not arc-transitive by Corollary˜13.24.
 ∎
 Lemma 13.26.
-Let $G$ be a vertex-transitive essentially 5-connected graph.
+Let $G$ be a vertex-transitive essentially 5-connected finite graph.
 Then either $G$ is quasi-5-connected or there is a partition $\mathcal{P}$ of $V(G)$ into vertex-sets of $K_{4}$’s such that the edges in $G$ with ends in different classes of $\mathcal{P}$ form a perfect matching in $G$.
 Proof.
-If $G$ does not have a tetra-separation, then $G$ is quasi-5-connected by Lemma 2.3 (i).
+If $G$ does not have a tetra-separation, then $G$ is quasi-5-connected by Lemma˜2.3 (i).
 Hence, we may assume that $G$ has a tetra-separation $(A,B)$.
@@
 Lemma 13.27.
-Let $G$ be a vertex-transitive essentially 5-connected graph. Then $G$ is one of the following:
+Let $G$ be a vertex-transitive essentially 5-connected finite graph. Then $G$ is one of the following:
   • $\cdot$
@@
 Assume that $G$ is not quasi-5-connected.
-Then by Lemma 12.26, there is a partition $\mathcal{P}$ of $V(G)$ into the vertex-sets of $K_{4}$’s such that edges between distinct partition-classes of $\mathcal{P}$ form a perfect matching in $G$.
+Then by Lemma˜13.26, there is a partition $\mathcal{P}$ of $V(G)$ into the vertex-sets of $K_{4}$’s such that edges between distinct partition-classes of $\mathcal{P}$ form a perfect matching in $G$.
 We distinguish the following cases.
-[…]
-[…]
-[…]
+Case 1: there are two parts $P_{1}\neq P_{2}$ of $\mathcal{P}$ such that there are at least three $P_{1}$$P_{2}$ edges $e_{1},e_{2},e_{3}$ in $G$.
+Let $f_{i}\notin\{e_{1},e_{2},e_{3}\}$ be the remaining edge incident with a vertex in $P_{i}$.
+By 4-connectivity of $G$, $\{f_{1},f_{2}\}$ is not a mixed-2-separator.
+It follows that $f_{1}=f_{2}$, so $G$ is a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length 4.
+Case 2: there are two parts $P_{1}\neq P_{2}$ of $\mathcal{P}$ such that there are exactly two $P_{1}$$P_{2}$ edges in $G$.
+Let $e_{1},e_{2}$ be the two edges incident with $P_{1}$ but not with $P_{2}$, and let $f_{1},f_{2}$ be the two edges incident with $P_{2}$ but not with $P_{1}$.
+We define $(A,B):=(P_{1}\cup P_{2},V(G)\smallsetminus(P_{1}\cup P_{2}))$.
+Then $S(A,B)=\{e_{1},e_{2},f_{1},f_{2}\}$, so $(A,B)$ is a tetra-separation.
+By assumption, there is a side of $(A,B)$ that induces a $K_{4}$.
+Since $|A|=8$, we must have $G[B]=K_{4}$.
+It follows that $G$ is a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $6$.
+Case 3: for every two parts $P_{1}\neq P_{2}$ of $\mathcal{P}$, there is at most one $P_{1}$$P_{2}$ edge in $G$.
+Then we define $H:=G/\mathcal{P}$, which is a 4-regular graph.
+Moreover, observe that $G$ is the $K_{4}$-expansion of $H$.
+Then by Lemma˜13.21HH is arc-transitive.
+By Corollary˜13.2HH is 4-connected.
+∎
 Corollary 13.28.
@@
 Proof.
-Follows from Lemma 12.27 together with the following observations.
-First, a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\ell\in\{4,6\}$ is not arc-transitive by Corollary 12.22.
-Second, the $K_{4}$-expansion of a graph is not arc-transitive by Corollary 12.23.
+Follows from Lemma˜13.27 together with the following observations.
+First, a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\ell\in\{4,6\}$ is not arc-transitive by Corollary˜13.22.
+Second, the $K_{4}$-expansion of a graph is not arc-transitive by Corollary˜13.23.
 ∎
-Proof of 2.
+Proposition 13.29.
+A 4-connected finite graph $G$ is vertex-transitive if and only if $G$ either is
+  1. (i)
+    a quasi-5-connected vertex-transitive graph,
+  2. (ii)
+    the $K_{4}$-expansion of a quasi-5-connected arc-transitive graph,
+  3. (iii)
+    a cycle of $K_{4}$-bags of length $\geqslant 4$ (Figure˜1.9),
+  4. (iv)
+    a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 4$ (Figure˜1.10),
+  5. (v)
+    a cycle of triangle-bags of length $\geqslant 6$ (Figure˜1.12),
+or $G$ is one of the following graphs or a $K_{4}$-expansion thereof:
+  1. (vi)
+    a cycle of $K_{2,2}$-bags of length $\geqslant 3$ (Figure˜1.14),
+  2. (vii)
+    the line-graph of the cube (Figure˜1.15).
+Proof.
 Let $G$ be a 4-connected finite graph.
-[…]
-[…]
+($\Rightarrow$).
+Assume that $G$ is vertex-transitive.
+If $G$ has two tetra-separations that cross, then we are done by Corollary˜13.17.
+Otherwise, every tetra-separation of $G$ is totally-nested.
+Then $G$ is essentially 5-connected by Corollary˜13.6.
+By Lemma˜13.27GG is either quasi-5-connected, or a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\ell\in\{4,6\}$, or the $K_{4}$-expansion of an arc-transitive 4-connected 4-regular graph.
+In the first two cases, we are done.
+So assume that $G$ is the $K_{4}$-expansion of an arc-transitive 4-connected 4-regular graph $H$.
+In particular, $H$ is vertex-transitive.
+If $H$ has two tetra-separations that cross, then we apply Corollary˜13.25 to get that $H$ is the line graph of the cube or that $H$ is a cycle of $K_{2,2}$-bags of length $\geqslant 3$.
+Here we are done.
+Otherwise, every tetra-separation of $H$ is totally-nested.
+By Corollary˜13.6HH is essentially 5-connected.
+By Corollary˜13.28HH quasi 5-connected.
+Thus, $G$ is the $K_{4}$-expansion of an arc-transitive quasi-5-connected graph, and we are done.
 ($\Leftarrow$).
 Clearly, if $G$ is one of the graphs in (i), (iii), (iv), (v), then $G$ is vertex-transitive.
-If (ii) $G$ is the $K_{4}$-expansion of a quasi-5-connected arc-transitive graph, then $G$ is vertex-transitive by Lemma 12.21.
+If (ii) $G$ is the $K_{4}$-expansion of a quasi-5-connected arc-transitive graph, then $G$ is vertex-transitive by Lemma˜13.21.
 Both (vi) a cycle of $K_{2,2}$-bags and (vii) the line graph of the cube are arc-transitive.
-If $G$ is a $K_{4}$-expansion of (vi) or (vii), then $G$ is vertex-transitive by Lemma 12.21.
+If $G$ is a $K_{4}$-expansion of (vi) or (vii), then $G$ is vertex-transitive by Lemma˜13.21.
+∎
+The line graph $L=L(G)$ of a graph $G$ is the graph on the vertex set $V(L):=E(G)$ such that two vertices in $L$ are adjacent if and only if their corresponding edges in $G$ share an endvertex.
+A graph $G$ is internally-4-connected if it is 3-connected and every 3-separation has a side that induces a claw. Note that internally-4-connected graphs are quasi-4-connected.
+Key Lemma 13.30.
+Let $G$ be a 3-regular quasi-4-connected graph. Then either $G$ is internally-4-connected and has girth $\geqslant 4$, or $G=K_{3}\square K_{2}$.
+Proof.
+Suppose that $G$ is not internally-4-connected.
+It suffices to show that either $G$ is a cycle of $C_{4}$-bags of length 3 or $G$ has girth $\geqslant 4$.
+Note that $G$ has $\geqslant 5$ vertices as $G\neq K_{4}$.
+There is a vertex $v$ in $G$ of degree 3 that has two adjacent neighbours $a$ and $b$.
+Let $c$ be the third neighbour of $v$.
+First, we show that $ac,bc\notin E(G)$.
+Assume $ac\in E(G)$.
+Then the pair $(\{v,a,b,c\},V(G)\smallsetminus\{v,a\})$ is a 2-separation with separator $\{b,c\}$ since $|V(G)|\geqslant 5$, contradicting that $G$ is 3-connected.
+It follows that $(A,B):=(\{v,a,b\},V(G)\smallsetminus\{v,a,b\})$ is a 3-edge-cut.
+Since $G$ is 3-connected and since the set $N(\{a,b\})$ separates $\{a,b\}$ from $c$, we must have $|N(\{a,b\})|\geqslant 3$.
+In particular, no two edges in $S(A,B)$ share an endvertex.
+We define $a^{\prime},b^{\prime}$ and $v^{\prime}$ such that $S(A,B)=\{aa^{\prime},bb^{\prime},vv^{\prime}\}$.
+Note that $N(\{a,b\})=\{a^{\prime},b^{\prime},v\}$ separates $\{a,b\}$ from $V(G)\smallsetminus\{v,a,b,a^{\prime},b^{\prime}\}$.
+By quasi-4-connectivity of $G$, it follows that $V(G)=\{v,a,b,v^{\prime},a^{\prime},b^{\prime}\}$.
+And by 3-regularity of $G$, it follows that $v^{\prime}a^{\prime}b^{\prime}$ is a triangle.
+So, $G=K_{3}\square K_{2}$.
+∎
+Recall that $E(v)$ denotes the set of all edges incident with a vertex $v$ in a graph $G$.
+Lemma 13.31.
+Let $G$ be a connected graph and let $L=L(G)$ be the line graph of $G$. If $(A,B)$ is an edge-cut in $G$ with $E(G[A])\neq\emptyset\neq E(G[B])$, then
+$(A^{\prime},B^{\prime}):=(E(G[A])\cup S(A,B)\>,\>E(G[B])\cup S(A,B))$
+is a genuine separation of $L$. If $(A,B)$ has order $k$, then $(A^{\prime},B^{\prime})$ has order $k$ as well.
+Proof.
+Since $E(G)=E(G[A])\,\dot{\cup}\,E(G[B])\,\dot{\cup}\,S(A,B)$, we get $A^{\prime}\cup B^{\prime}=V(L)$ and $A^{\prime}\cap B^{\prime}=S(A,B)$.
+Note that for every vertex $v\in A$ we have $E(v)\subseteq A^{\prime}$.
+Similarly, for every vertex $v\in B$ we have $E(v)\subseteq B^{\prime}$.
+It follows that there is no $(A^{\prime}\smallsetminus B^{\prime})$$(B^{\prime}\smallsetminus A^{\prime})$ edge in $L$.
+From $E(G[A])\neq\emptyset\neq E(G[B])$ we get that $(A^{\prime},B^{\prime})$ is proper.
+Hence, $(A^{\prime},B^{\prime})$ is a separation of $L$ that has the same order as $(A,B)$.
+Lemma 13.32.
+Let $G$ be a connected graph and let $L=L(G)$ be the line graph of $G$. If $(A^{\prime},B^{\prime})$ is a genuine separation of $L$, then
+$(A,B):=(\{v\in V(G)\mid E(v)\subseteq A^{\prime}\}\>,\>\{v\in V(G)\mid E(v)\subseteq B^{\prime}\})$
+is a mixed-separation of $G$ with $A\cap B=\{v\in V(G)\mid E(v)\subseteq A^{\prime}\cap B^{\prime}\}$.
+Proof.
+Note that for all $v\in V(G)$, the set $E(v)\subseteq V(L)$ induces a clique in $L$.
+Hence, $E(v)\subseteq A^{\prime}$ or $E(v)\subseteq B^{\prime}$ for all $v\in V(G)$.
+From the definition of $(A,B)$, it follows that $A\cup B=V(G)$.
+Since $(A^{\prime},B^{\prime})$ is proper, it follows that there exist vertices $v_{A},v_{B}\in V(G)$ with $E(v_{A})\subseteq A^{\prime}\nsupseteq E(v_{B})$ and $E(v_{B})\subseteq B^{\prime}\nsupseteq E(v_{A})$; that is, $v_{A}\in A\smallsetminus B$ and $v_{B}\in B\smallsetminus A$.
+Hence, $(A,B)$ is a mixed-separation.
+Moreover, a vertex $v$ is contained in $A\cap B$ if and only if $E(v)\subseteq A^{\prime}\cap B^{\prime}$.
+∎
+Lemma 13.33.
+Let $G$ be a 3-regular internally-4-connected graph and let $L=L(G)$ be the line graph of $G$. Then all of the following assertions hold:
+  1. (i)
+    $L$ is 4-regular and every vertex $v\in V(L)$ is contained in exactly two triangles $\Delta_{1}$ and $\Delta_{2}$ in $L$. Moreover, $V(\Delta_{1}\cap\Delta_{2})=\{v\}$.
+  2. (ii)
+    $L$ is 4-connected.
+  3. (iii)
+    For every triangle $\Delta\subseteq L$ there is a vertex $v$ in $G$ such that $E(v)=V(\Delta)$.
+  4. (iv)
+    For every edge $e\in E(L)$ there is a unique triangle $\Delta(e)$ in $L$ that contains $e$.
+Proof.
+(i) follows from the definition of the line graph and from $G$ having girth $\geqslant 4$.
+(ii). Suppose for a contradiction that there is a genuine $(\leqslant 3)$-separation $(A^{\prime},B^{\prime})$ of $L$. Let
+$(A,B):=(\{v\in V(G)\mid E(v)\subseteq A^{\prime}\}\>,\>\{v\in V(G)\mid E(v)\subseteq B^{\prime}\})\leavevmode\nobreak\ ,$
+which is a mixed-separation of $G$ by Lemma˜13.32.
+First, suppose $A\cap B\neq\emptyset$.
+Then $A^{\prime}\cap B^{\prime}$ induces a $K_{3}$ in $L$.
+By (i), every vertex $v$ of $L$ is contained in exactly two triangles, which only intersect in $v$.
+It follows that $(A,B)$ is a 1-separation of $G$, a contradiction.
+Hence, we may assume that $(A,B)$ is an edge-cut.
+Let $e=uv\in S(A,B)$, say with $u\in A$ and $v\in B$.
+Then $E(u)\subseteq A^{\prime}$ and $E(v)\subseteq B^{\prime}$ by definition of $(A,B)$.
+Let $\Delta_{u}$ and $\Delta_{v}$ be the triangles in $L$ induced by $E(u)$ and $E(v)$, respectively.
+From $V(\Delta_{u})\subseteq A^{\prime}$ and $V(\Delta_{v})\subseteq B^{\prime}$ and $V(\Delta_{u}\cap\Delta_{v})=\{e\}$ (by (i)) it follows that $e\in S(A^{\prime},B^{\prime})$.
+Hence, $S(A,B)\subseteq S(A^{\prime},B^{\prime})$; in particular $(A,B)$ has order $\leqslant 3$.
+Since $G$ is internally-4-connected, $(A,B)$ is an atomic edge-cut with $S(A,B)=S(A^{\prime},B^{\prime})$.
+Now $(A^{\prime},B^{\prime})$ is not proper, a contradiction.
+(iii).
+From $G$ being 3-regular and internally-4-connected, it follows that $G$ has girth $\geqslant 4$.
+If $\Delta$ is a triangle in $L$ such that there is no vertex $v$ in $G$ with $V(\Delta)=E(v)$, then $V(\Delta)$ is the set of edges of a triangle in $G$, contradicting that $G$ has girth $\geqslant 4$.
+(iv).
+Every edge $e_{1}e_{2}\in E(L)$ is contained in at least one triangle: consider the endvertex $v\in G$ shared by $e_{1}$ and $e_{2}$, then $E(v)$ induces a triangle in $L$ that contains the edge $e_{1}e_{2}$.
+Assume for a contradiction that there is an edge $e_{1}e_{2}\in E(L)$ that is contained in at least two triangles $\Delta_{1}$ and $\Delta_{2}$ of $L$.
+By (iii), there are two vertices $v_{1},v_{2}\in V(G)$ such that $E(v_{i})=V(\Delta_{i})$ for $i\in\{1,2\}$.
+Then $e_{1},e_{2}\in E(v_{1})\cap E(v_{2})$, contradicting that $G$ is simple.
+∎
+Lemma 13.34.
+Let $G$ be a finite, 3-regular, internally-4-connected, vertex-transitive graph, and let $L=L(G)$ be its line graph.
+Let $(A^{\prime},B^{\prime})$ be a mixed-4-separation of $L$ with $|A^{\prime}\smallsetminus B^{\prime}|\geqslant 2$ and $|B^{\prime}\smallsetminus A^{\prime}|\geqslant 2$.
+Then $S(A^{\prime},B^{\prime})\subseteq V(L)$.
+In particular, all tetra-separations of $L$ are genuine 4-separations.
+Proof.
+First, observe that we can assume $G\neq K_{4}$.
+By Lemma˜13.33LL is 4-connected.
+Let $(A^{\prime},B^{\prime})$ be a a mixed-4-separation of $L$ with $|A^{\prime}\smallsetminus B^{\prime}|\geqslant 2$ and $|B^{\prime}\smallsetminus A^{\prime}|\geqslant 2$.
+The left-right-reduction $(A^{\prime\prime},B^{\prime\prime})$ of $(A^{\prime},B^{\prime})$ is a tetra-separation by ˜5.10.
+Moreover, it satisfies $A^{\prime\prime}\smallsetminus B^{\prime\prime}\supseteq A^{\prime}\smallsetminus B^{\prime}$ and $B^{\prime\prime}\smallsetminus A^{\prime\prime}\supseteq B^{\prime}\smallsetminus A^{\prime}$ by ˜5.9, and every edge in $S(A^{\prime},B^{\prime})$ is in $S(A^{\prime\prime},B^{\prime\prime})$.
+Hence, it suffices to show that $S(A^{\prime\prime},B^{\prime\prime})\subseteq V(L)$.
+Assume for a contradiction that $e_{1}\in S(A^{\prime\prime},B^{\prime\prime})$ is an edge in $L$.
+By Lemma˜13.33 (iv), there is a unique triangle $\Delta_{1}$ in $L$ that contains $e_{1}$.
+Let $v$ be the third vertex in $\Delta_{1}$ besides the endvertices of $e_{1}$.
+Since $\Delta_{1}$ is a triangle with the edge $e_{1}$ contained in $S(A^{\prime\prime},B^{\prime\prime})$, we get that $v$ is contained in the separator $S(A^{\prime\prime},B^{\prime\prime})$ as well.
+By Lemma˜13.33 (i), there is a triangle $\Delta_{2}\neq\Delta_{1}$ in $L$ such that $\Delta_{1}$ and $\Delta_{2}$ meet exactly in the vertex $v$.
+Let $e_{2}$ denote the edge in $\Delta_{2}$ whose endvertices are distinct from $v$.
+Since $L$ is 4-regular, the degree-condition for $v\in S(A^{\prime\prime},B^{\prime\prime})$ yields $e_{2}\in S(A^{\prime\prime},B^{\prime\prime})$.
+So we know three elements of $S(A^{\prime\prime},B^{\prime\prime})$: they are $e_{1},v$ and $e_{2}$.
+The fourth element $u$ of $S(A^{\prime\prime},B^{\prime\prime})$ must be a vertex, as otherwise we could argue as for $e$ to find a third triangle contributing both an edge and a vertex to $S(A^{\prime\prime},B^{\prime\prime})$, contradicting that $\Delta_{1}$ and $\Delta_{2}$ already use all four edges incident to $v$.
+By Lemma˜13.33 (i), there are two triangles $\Delta,\Delta^{\prime}$ in $L$ that meet exactly in the vertex $u$.
+By 4-regularity of $L$ and the degree-condition for $u\in S(A^{\prime},B^{\prime})$, we have $\Delta-u\subseteq G[A^{\prime}\smallsetminus B^{\prime}]$ and $\Delta^{\prime}-u\subseteq G[B^{\prime}\smallsetminus A^{\prime}]$, say.
+Now consider the mixed-separation $(A,B)$ of $G$ that is induced by $(A^{\prime\prime},B^{\prime\prime})$ as in Lemma˜13.32.
+The separator of $(A,B)$ consists of two vertices $x_{1},x_{2}$ that correspond to the triangles $\Delta_{1},\Delta_{2}$, respectively, and of one edge $f$ that corresponds to the vertex $u$ (by Lemma˜13.33 (iii)).
+In particular, $(A,B)$ is a mixed-3-separation of $G$.
+Also note that $x_{1}x_{2}\in E(G)$ as the triangles $\Delta_{1},\Delta_{2}$ share the vertex $u$.
+Since $G\neq K_{4}$, there is a proper side of $(A,B)$, say $B\smallsetminus A$, that contains at least two vertices.
+Let $x_{3}$ be the endvertex of $f$ in $B\smallsetminus A$.
+Note that $(A\cup f,B)$ is a genuine 3-separation with separator $\{x_{1},x_{2},x_{3}\}$.
+Since $G$ is internally-4-connected, it follows that $x_{1}x_{2}\notin E(G)$, a contradiction.
+∎
+Let $G$ be a graph. A tetra-cut of $G$ is a 4-edge-cut of $G$ that satisfies the matching-condition. A tetra-cut of $G$ is totally-nested if it is nested with every tetra-cut of $G$.
+Observation 13.35.
+If $G$ is a 2-connected graph and $(A,B)$ is a non-atomic edge-cut of $G$ then $E(G[A])\neq\emptyset\neq E(G[B])$.∎
+Lemma 13.36.
+Let $G$ be a finite, 3-regular, internally-4-connected, vertex-transitive graph and let $L=L(G)$ be its line graph.
+Let the map $(A,B)\mapsto(A^{\prime},B^{\prime})$ send non-atomic edge-cuts $(A,B)$ of $G$ to genuine separations $(A^{\prime},B^{\prime})$ of $L$ as in the statement of Lemma˜13.31.
+Then this map restricts to a bijection from the set of tetra-cuts of $G$ to the set of tetra-separations of $L$.
+Proof.
+First note that $E(G[A])\neq\emptyset\neq E(G[B])$ for every nonatomic edge-cut $(A,B)$ of $G$ by ˜13.35.
+Hence, the prerequisites of Lemma˜13.31 are fulfilled.
+Since $G$ has no isolated vertices, the map $(A,B)\mapsto(A^{\prime},B^{\prime})$ is injective.
+We claim that every tetra-cut $(A,B)$ is mapped to a tetra-separation $(A^{\prime},B^{\prime})$.
+We get from Lemma˜13.31 that $(A^{\prime},B^{\prime})$ is a 4-separation of $L$.
+So it remains to check the degree-condition for the vertices in the separator of $(A^{\prime},B^{\prime})$.
+For this, let $v$ be an arbitrary vertex in $S(A^{\prime},B^{\prime})$.
+By Lemma˜13.33 (i), the vertex $v$ is contained in exactly two triangles $\Delta_{1},\Delta_{2}$, and the two triangles meet precisely in the vertex $v$.
+Each triangle $\Delta_{i}\subseteq L$ corresponds to an endvertex $x_{i}$ of $v$ viewed as an edge of $G$.
+Since $(A,B)$ is a tetra-cut, we have $x_{1}\in A\smallsetminus B$ and $x_{2}\in B\smallsetminus A$, say.
+Each $\Delta_{i}$ determines a 3-star in $G$ with centre $x_{i}$, and the edges in $S(A,B)$ form a matching from which each 3-star already contains $v$ (viewed as an edge of $G$).
+Hence $\Delta_{1}\subseteq L[A^{\prime}]$ and $v$ is the only vertex of $\Delta_{1}$ contained in $S(A^{\prime},B^{\prime})$.
+So $v$ has $\geqslant 2$ neighbours in $L[A^{\prime}\smallsetminus B^{\prime}]$.
+Similarly, $v$ has $\geqslant 2$ neighbours in $L[B^{\prime}\smallsetminus A^{\prime}]$.
+Therefore, $(A^{\prime},B^{\prime})$ is a tetra-separation of $L$.
+To conclude the proof, it remains to show that every tetra-separation of $L$ equals $(A^{\prime},B^{\prime})$ for some tetra-cut $(A,B)$.
+For this, let $(C,D)$ be an arbitrary tetra-separation of $L$.
+By Lemma˜13.34, $(C,D)$ is a genuine 4-separation.
+Let $(A,B)$ be the mixed-separation in $G$ as defined in Lemma˜13.32 applied to $(C,D)$.
+Then we claim that $(A,B)$ is a tetra-cut and $(A^{\prime},B^{\prime})=(C,D)$.
+By the degree-condition and Lemma˜13.33 (i), for every $v\in S(C,D)$ there are two triangles $\Delta_{1}(v),\Delta_{2}(v)$ with $\Delta_{1}(v)\smallsetminus\{v\}\subseteq C\smallsetminus D$ and $\Delta_{2}(v)\smallsetminus\{v\}\subseteq D\smallsetminus C$.
+That means, for two distinct $v,w\in S(C,D)$ we have $\Delta_{i}(v)\neq\Delta_{i}(w)$ for $i\in\{1,2\}$.
+With Lemma˜13.33 it follows that $(A,B)$ is a tetra-cut.
+Moreover, if $v\in C\smallsetminus D$ then $v\in E(G[A])$, and it follows that $v\in A^{\prime}\smallsetminus B^{\prime}$.
+Analogously, $v\in D\smallsetminus C$ implies $v\in B^{\prime}\smallsetminus A^{\prime}$.
+On the other hand, if $v\in C\cap D$ then $v\in S(A,B)$, and it follows that $v\in A^{\prime}\cap B^{\prime}$.
+Hence, $(A^{\prime},B^{\prime})=(C,D)$.
+∎
+Lemma 13.37.
+Let $G$ be a finite, 3-regular, internally-4-connected, vertex-transitive graph and let $L=L(G)$ be its line graph.
+Let $(A,B)\mapsto(A^{\prime},B^{\prime})$ be the bijection from the set of tetra-cuts $(A,B)$ of $G$ to the set of tetra-separations $(A^{\prime},B^{\prime})$ of $L$.
+Then $(A,B)\leqslant(C,D)$ if and only if $(A^{\prime},B^{\prime})\leqslant(C^{\prime},D^{\prime})$.
+Proof.
+If $(A,B)\leqslant(C,D)$ then $A\subseteq C$.
+Hence, $E(G[A])\subseteq E(G[C])$ and $S(A,B)\subseteq E(G[C])\cup S(C,D)$.
+It follows that $A^{\prime}\subseteq C^{\prime}$.
+Analogously, $B^{\prime}\supseteq D^{\prime}$.
+On the other hand, assume $(A^{\prime},B^{\prime})\leqslant(C^{\prime},D^{\prime})$.
+Let $v$ be a vertex in $A$.
+Recall that $E(v)$ denotes the edges incident with $v$ in $G$.
+Then $E(v)\subseteq A^{\prime}\subseteq C^{\prime}$.
+From $E(v)\subseteq C^{\prime}$ it follows that $v\in C$.
+Hence $A\subseteq C$ and, analogously, $B\supseteq D$.
+∎
+Lemma 13.38.
+Let $G$ be a finite 3-regular internally-4-connected vertex-transitive graph and let $(A_{1},A_{2})$ be a totally-nested tetra-cut in $G$. Then $G[A_{1}]=C_{4}$ or $G[A_{2}]=C_{4}$.
+Proof.
+Let $M$ be the union of the $\operatorname{Aut}(G)$-orbits of $(A_{1},A_{2})$ and $(A_{2},A_{1})$.
+Let $(C,D)$ be $\leqslant$-minimal in $M$; that is, $(C^{\prime},D^{\prime})\not<(C,D)$ for every $(C^{\prime},D^{\prime})\in M$.
+We claim that every vertex in $C$ is incident to an edge in $S(C,D)$.
+Assume for a contradiction that $v\in C$ is not.
+Let $u\in C\smallsetminus D$ be an endvertex of one of the four edges in $S(C,D)$ and let $\varphi\in\operatorname{Aut}(G)$ such that $\varphi(u)=v$.
+Then $(\varphi(C),\varphi(D))<(C,D)$ or $(\varphi(D),\varphi(C))<(C,D)$, contradicting the $\leqslant$-minimality of $(C,D)$.
+Finally, since $G$ is internally-4-connected, every vertex has degree $\geqslant 3$, so $G[C]$ is a $C_{4}$. The tetra-cut $(A_{1},A_{2})$ shares the above properties of $(C,D)$ since $(C,D)\in M$. ∎
+Lemma 13.39.
+Let $G$ be an internally-4-connected graph. Then every $(\leqslant 3)$-edge-cut of $G$ is an atomic 3-edge-cut.
+Proof.
+Since $G$ is 3-connected, it does not contain a $(\leqslant 2)$-edge-cut.
+Assume that $G$ has a 3-edge-cut $(A,B)$ with separator $S(A,B)=\{e_{1},e_{2},e_{3}\}$.
+Let $a_{i}$ be the endvertex of $e_{i}$ in $A$, and let $b_{i}$ be the endvertex of $e_{i}$ in $B$.
+First, assume that $(A,B)$ satisfies the matching-condition.
+Since $\{a_{1},a_{2},b_{3}\}$ is a 3-separator separating $a_{3}$ from $\{b_{1},b_{2}\}$, we must have $a_{1}a_{2}\notin E(G)$ by $G$ being internally-4-connected.
+Since $\{b_{1},a_{2},a_{3}\}$ is a 3-separator separating $a_{1}$ from $\{b_{2},b_{3}\}$, we must have $a_{1}a_{2}\in E(G)$ by $G$ being internally-4-connected.
+This is a contradiction.
+Second, assume that $(A,B)$ does not satisfy the matching-condition.
+If $(A,B)$ is not atomic, we have without loss of generality $a_{1}=a_{2}\neq a_{3}$ and $b_{1}\neq b_{3}$.
+Then, however, $\{a_{1},b_{3}\}$ is a 2-separator separating $a_{3}$ from $b_{1}$.
+Hence, $(A,B)$ is atomic.
+∎
+Lemma 13.40.
+Let $G$ be an internally-4-connected graph. Let $(A,B)$ and $(C,D)$ be two crossing tetra-cuts of $G$. Then they cross with all links of size 2 and empty centre.
+Proof.
+Recall that $L(X,Y)$ denotes the corner-separator for the corner $XY$. Note that, since $(A,B)$ and $(C,D)$ are edge-cuts, their corner-separators are edge-cuts as well.
+First assume that the centre is nonempty.
+Say, there is a diagonal edge from the $AD$-corner to the $BC$-corner.
+By Lemma˜4.3, either $|L(A,C)|\leqslant 3$ or $|L(B,D)|\leqslant 3$, say the former.
+By Lemma˜13.39, $L(A,C)$ is an atomic 3-edge-cut, contradicting that $(A,B)$ and $(C,D)$ satisfy the matching-condition.
+Similarly, we show that every link has size 2.
+Otherwise, there would exist a corner-separator of order $\leqslant 3$, say $|L(A,C)|\leqslant 3$.
+By Lemma˜13.39, $L(A,C)$ is an atomic 3-edge-cut, contradicting that $(A,B)$ and $(C,D)$ satisfy the matching-condition.
+∎
+Lemma 13.41.
+Let $G$ be a 3-regular internally-4-connected graph and let $L=L(G)$ be its line graph.
+Let $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$ be two crossing tetra-separations of $L$.
+Then they cross with all links of size 2 and empty centre.
+Proof.
+We combine Lemma˜13.36 with Lemma˜13.40. ∎
+Lemma 13.42.
+Let $G$ be a 3-regular internally-4-connected vertex-transitive finite graph and let $L=L(G)$ be its line graph.
+Let $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$ be two tetra-separations of $L$ that cross with all links of size 2.
+Let $\mathcal{O}=(O,\mathcal{G})$ be the Tutte-bagel of $L$ induced by $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$.
+Then every torso of $\mathcal{O}$ is 3-connected and has $\geqslant 5$ vertices.
+Proof.
+We may assume the (Tutte-bagel setting). (9.12) for $L$, but with $(A^{\prime},B^{\prime})$ in place of $(A,B)$ and $(C^{\prime},D^{\prime})$ in place of $(C,D)$.
+Assume that there is a torso $H_{s}$ in $\mathcal{O}$ that is a triangle or 4-cycle.
+Say, $H_{s}$ stems from the $A^{\prime}C^{\prime}$-Tutte-path.
+Since $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$ are genuine separations by Lemma˜13.34, no link contains an edge.
+It follows that all conditions for the (Link-Moving Lemma for vertices).
+(9.44) are fulfilled.
+By the (Link-Moving Lemma for vertices).
+(9.44(i), we get a tetra-separation of $L$ that contains at least one edge in its separator.
+This contradicts Lemma˜13.34.
+Therefore, every torso $H_{s}$ of $\mathcal{O}$ is 3-connected.
+Assume for a contradiction that there is a torso $H_{s}$ of $\mathcal{O}$ with $\leqslant 4$ vertices.
+Then $H_{s}=K_{4}$.
+By Lemma˜13.33 (iv), every edge $e\in E(G_{s})$ is contained in a unique triangle.
+If some adhesion set of $s$ induces an edge $e$, then $e$ is contained in two triangles.
+Otherwise, the edges in $G_{s}$ are not contained in any triangle.
+This is a contradiction.
+∎
+Let $G$ be a graph. A cycle-partition of $G$ is a pair $\mathcal{O}=(O,(V_{t})_{t\in V(O)})$ consisting of a cycle $O$ and vertex sets $V_{t}$ such that
+  1. (i)
+    $\{\,V_{t}\mid t\in V(O)\,\}$ is a partition of $V(G)$; and
+  2. (ii)
+    every edge $e\in E(G)$ either has its ends in a unique $V_{t}$ or there is an edge $tt^{\prime}\in E(O)$ such that $e$ has its ends in $V_{t}$ and $V_{t^{\prime}}$.
+We call the sets $V_{t}$ the bags of $\mathcal{O}$. For every $tt^{\prime}\in E(O)$, we call $E_{G}(V_{t},V_{t^{\prime}})$ the adhesion-set of $tt^{\prime}$.
+Lemma 13.43.
+Let $G$ be a 3-regular internally-4-connected vertex-transitive finite graph.
+Assume that $G$ has two crossing tetra-cuts.
+Then $G$ admits a cycle-partition $\mathcal{O}=(O,(V_{t})_{t\in O})$ with $|O|\geqslant 4$ such that all of the following are satisfied:
+  1. (i)
+    All adhesion-sets of $\mathcal{O}$ are 2-matchings;
+  2. (ii)
+    Every bag of $\mathcal{O}$ induces a $C_{4}$ or a $K_{2}$;
+  3. (iii)
+    Let $t\in V(O)$ and let $t_{1},t_{2}$ be its neighbours on $O$.
+Let $U_{i}$ be the set of endvertices of edges in the adhesion-set for $tt_{i}$ in $V_{t}$.
+If $V_{t}$ induces a 4-cycle, then this 4-cycle alternates between $U_{1}$ and $U_{2}$.
+Proof.
+Let $L=L(G)$ be the line graph of $G$.
+Let $(A,B)\mapsto(A^{\prime},B^{\prime})$ be the bijection from Lemma˜13.36 between tetra-cuts $(A,B)$ in $G$ and tetra-separations $(A^{\prime},B^{\prime})$ in $L$.
+By Lemma˜13.37(A,B)(A,B) is a totally-nested tetra-cut of $G$ if and only if $(A^{\prime},B^{\prime})$ is a totally-nested tetra-separation of $L$.
+Let $(A,B)$ and $(C,D)$ be two crossing tetra-cuts in $G$.
+By Lemma˜13.37 and Lemma˜13.41, the tetra-separations $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$ cross with links of size 2 and empty centre.
+Let $\mathcal{O}^{\prime}=(O^{\prime},(G^{\prime}_{t})_{t\in V(O^{\prime})})$ be the Tutte-bagel of $L$ induced by $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$.
+We denote by $H^{\prime}_{t}$ the torso of $t\in V(O^{\prime})$ in $\mathcal{O}^{\prime}$.
+By Lemma˜13.42, every torso of $\mathcal{O}^{\prime}$ is 3-connected on $\geqslant 5$ vertices.
+Recall that $E(v)$ denotes the set of edges incident with the vertex $v$ in $G$.
+We define $\mathcal{O}=(O,(V_{t})_{t\in V(O)})$ with $O:=O^{\prime}$ and $V_{t}:=\{v\in V(G)\mid E(v)\subseteq V_{t}\}$.
+We show that $\mathcal{O}$ is a cycle-partition of $G$.
+First, note that for every $v\in V(G)$, the set $E(v)$ induces a triangle in the graph $L$, which in turn is contained in a unique torso $H^{\prime}_{t}$ of $\mathcal{O}^{\prime}$.
+Hence, $\{\,V_{t}\mid t\in V(O)\,\}$ is a partition of $V(G)$.
+Second, let $e$ be an edge in $G$.
+Then $e$ is a vertex in $L$ which is contained in two triangles $\Delta_{1}(e)$ and $\Delta_{2}(e)$ in $L$.
+If $\Delta_{1}(e)$ and $\Delta_{2}(e)$ are contained in the same torso of $\mathcal{O}^{\prime}$, then both endvertices of $e$ are contained in the same bag of $\mathcal{O}$.
+Otherwise, since every torso of $\mathcal{O}^{\prime}$ is 3-connected the adhesion-sets of $\mathcal{O}^{\prime}$ are pairwise disjoint by Lemma˜9.4, so $\Delta_{1}(e)$ and $\Delta_{2}(e)$ are contained in two torsos of $\mathcal{O}^{\prime}$, whose corresponding nodes $t_{1},t_{2}\in V(O^{\prime})$ are neighbours on $O^{\prime}$.
+Then, the endvertices of $e$ are contained in $V_{t_{1}}$ and $V_{t_{2}}$, where $t_{1}t_{2}\in E(O)$.
+This shows that $\mathcal{O}$ is a cycle-partition of $G$.
+It remains to show that (i)(iii) are satisfied for the cycle-partition $\mathcal{O}$ of $G$.
+(i).
+Let $tt^{\prime}\in E(O)$.
+We show that the adhesion-set of $tt^{\prime}$ is a 2-matching.
+First, note that the adhesion-set of $tt^{\prime}$ in $\mathcal{O}$ is equal to the adhesion-set of $tt^{\prime}$ in $\mathcal{O}^{\prime}$.
+Since the latter consists of two vertices of $L$, the former consists of two edges of $G$.
+Since every torso of $\mathcal{O}^{\prime}$ is 3-connected and $L$ is 4-regular, it follows that the adhesion-set of $tt^{\prime}$ in $\mathcal{O}^{\prime}$ is independent.
+Hence, the adhesion-set of $tt^{\prime}$ in $\mathcal{O}$ is a matching.
+[…]
+Now assume that $H^{\prime}_{t}$ is bad.
+Then $V(H^{\prime}_{t})$ consists of exactly five vertices.
+Let $v$ be the vertex in the interior of $H^{\prime}_{t}$.
+By 4-regularity of $L$, the vertex $v$ is adjacent to every vertex in $H^{\prime}_{t}-v$.
+Since every torso of $\mathcal{O}^{\prime}$ is 3-connected, every vertex $u$ in an adhesion-set of $H^{\prime}_{t}$ has degree 2 in $G^{\prime}_{t}$.
+It follows that the torso $H^{\prime}_{t}$ is a 4-wheel, that $G^{\prime}_{t}$ is the union of two triangles that meet exactly in $v$, and that $V_{t}$ induces a $K_{2}$ in $G$.
+(iii).
+We denote the vertices in $U_{i}$ by $u_{i}$ and $u_{i}^{\prime}$.
+Assume for contradiction that $V_{t}$ induces a 4-cycle $Z$ that does not alternate between $U_{1}$ and $U_{2}$.
+As shown in the proof of (ii), the tetra-cut $(V_{t},V(G)\smallsetminus V_{t})$ is totally-nested.
+Then without loss of generality $Z=u_{1}u_{1}^{\prime}u_{2}^{\prime}u_{2}u_{1}$.
+Then, however, $(V_{t_{1}}\cup U_{1},V(G)\smallsetminus(V_{t_{1}}\cup U_{1}))$ is a tetra-cut that crosses $(V_{t},V(G)\smallsetminus V_{t})$, a contradiction.
+∎
+Key Lemma 13.44.
+Let $G$ be a 3-regular internally-4-connected vertex-transitive graph.
+Assume that $G$ has two crossing tetra-cuts.
+Then $G$ is a cycle of $C_{4}$-bags of length $\geqslant 4$, or a cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length $\geqslant 8$.
+Proof.
+By Lemma˜13.43GG admits a cycle-partition $\mathcal{O}=(O,(V_{t})_{t\in O})$ with $|O|\geqslant 4$ such that properties (i)(iii) from Lemma˜13.43 are fulfilled.
+Let $O=:t_{1}t_{2}\ldots t_{\ell}t_{1}$.
+For $t=t_{i}\in V(O)$ we denote by $U_{t,1}$ the set of endvertices of the edges in the adhesion-set for $tt_{i-1}$ that lie in $V_{t}$; and by $U_{t,2}$ we denote the set of endvertices of the edges in the adhesion-set for $tt_{i+1}$ that lie in $V_{t}$.
+We show that either every bag $V_{t}$ of $\mathcal{O}$ induces a $K_{2}$ or every bag $V_{t}$ of $\mathcal{O}$ induces a $C_{4}$ alternating between $U_{t,1}$ and $U_{t,2}$.
+In the former case, it follows that $G$ is a cycle of $C_{4}$-bags of length $\geqslant 4$, while in the latter case it follows that $G$ is a cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length $\geqslant 8$.
+Assume for a contradiction that there is a bag of $\mathcal{O}$ that induces a $K_{2}$, and that there is a bag of $\mathcal{O}$ that induces a $C_{4}$.
+Then, without loss of generality, one of the following cases occurs.
+The bags for $t_{1}$, $t_{2}$ and $t_{3}$ induce a $K_{2}$ while the bag for $t_{4}$ induces a $C_{4}$.
+Then every vertex in $V_{t_{2}}$ is contained in two $C_{4}$’s of $G$ while every vertex in $V_{t_{4}}$ is contained in exactly one $C_{4}$ of $G$, contradicting vertex-transitivity.
+The bags for $t_{2}$ and $t_{3}$ induce a $K_{2}$ while the bags for $t_{1}$ and $t_{4}$ induce a $C_{4}$.
+Let $u\in V_{t_{2}}$.
+Then $u$ is contained in exactly two 5-cycles $Z,Z^{\prime}$.
+The intersection $Z\cap Z^{\prime}$ is a path with three edges, and $u$ is an internal vertex of this path.
+The neighbour of $u$ in $V_{t_{1}}$ also lies on $Z$ and $Z^{\prime}$, but it is not an internal vertex of $Z\cap Z^{\prime}$, contradicting vertex-transitivity.
+The bag for $t_{2}$ induces a $K_{2}$ while the bags for $t_{1}$ and $t_{3}$ induce a $C_{4}$.
+Then every vertex in $V_{t_{2}}$ is contained in exactly four 5-cycles of $G$, while every vertex in $V_{t_{1}}\cap N(V_{t_{2}})$ is contained in at most three 5-cycles.
+This contradicts vertex-transitivity.
+∎
+Recall that a graph $G$ is 2-quasi-5-connected if $G$ is quasi-4-connected and every 4-separation $(A,B)$ of $G$ satisfies $|A\smallsetminus B|\leqslant 2$ or $|B\smallsetminus A|\leqslant 2$.
+Lemma 13.45.
+Let $G$ be a 3-regular quasi-4-connected graph. Assume that $G$ has no tetra-cut. Then $G$ is 2-quasi-5-connected.
+Proof.
+If $G$ is not 2-quasi-5-connected, it has a 4-separation $(A,B)$ with $|A\smallsetminus B|\geqslant 3$ and $|B\smallsetminus A|\geqslant 3$.
+Since $G$ is quasi-4-connected, it follows that $(A,B)$ is tight.
+Let $S(A,B)=\{v_{1},v_{2},v_{3},v_{4}\}$.
+First we show that if $v_{i}$ only has one neighbour $u_{i}$ in $A\smallsetminus B$ and $v_{j}\neq v_{i}$ only has one neighbour $u_{j}$ in $A\smallsetminus B$, then $u_{i}\neq u_{j}$.
+Indeed, otherwise $(A^{\prime},B^{\prime}):=((A\cup\{u_{i}\})\smallsetminus\{v_{i},v_{j}\},B\cup\{u_{i}\})$ is a 3-separation of $G$ with $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$, contradicting that $G$ is quasi-4-connected.
+By the above argument, it follows that the left-right-reduction of $(A,B)$ satisfies the matching-condition.
+Since $G$ is 3-regular and $(A,B)$ is tight, every vertex $v_{i}\in S(A,B)$ either has exactly one neighbour in $A\smallsetminus B$, or $v_{i}$ has exactly one neighbour in $B\smallsetminus A$ and no neighbour in $A\cap B$.
+Therefore, the left-right-reduction of $(A,B)$ is a 4-edge-cut and hence a tetra-cut of $G$.
+∎
+Example 13.46.
+The toroidal hex-grid on a torus shows that 2-quasi-5-connected in Lemma˜13.45 is best possible.
+Lemma 13.47.
+Let $G$ be a 3-regular internally-4-connected vertex-transitive graph such that every tetra-cut of $G$ is totally-nested.
+Then either $G$ is 2-quasi-5-connected or there is a canonical partition $\mathcal{P}=\{\,A_{i}:i\in I\,\}$ of $V(G)$ such that
+  1. (i)
+    each $A_{i}$ induces a $C_{4}$;
+  2. (ii)
+    the edges in $G$ with ends in different classes of $\mathcal{P}$ form a perfect matching in $G$.
+Proof.
+If $G$ has no tetra-cut, then $G$ is 2-quasi-5-connected by Lemma˜13.45.
+So we may assume that $G$ has at least one tetra-cut.
+Let $\sigma:=\{\,(A_{i},B_{i}):i\in I\,\}$ be the set of all tetra-cuts of $G$ whose side $A_{i}$ induces a 4-cycle.
+Since all tetra-cuts of $G$ are totally-nested, $\sigma$ is a star of tetra-cuts.
+Hence the sets $A_{i}$ are pairwise disjoint.
+Since $G$ is vertex-transitive, we have $\bigcup_{i\in I}A_{i}=V(G)$.
+Therefore, $\mathcal{P}:=\{\,A_{i}:i\in I\,\}$ is a canonical partition of $G$ into vertex-sets that induce 4-cycles.
+Let $M$ denote the set of all edges in $G$ with ends in different classes of $\mathcal{P}$.
+Every vertex $v$ of $G$ lies in exactly one $A_{i}$, and since $G$ is 3-regular, it follows that $v$ is incident to exactly one edge in $M$.
+Hence $M$ is a perfect matching.
+∎
+Key Lemma 13.48.
+Let $G$ be a 3-regular internally-4-connected vertex-transitive graph. Assume that every tetra-cut in $G$ is totally-nested. Then $G$ is one of the following:
+  • $\cdot$
+    a cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length 6;
+  • $\cdot$
+    one of the three $C_{4}$-expansions of the line graph of the cube shown in Figure˜1.15;
+  • $\cdot$
+    a vertex-transitive $C_{4}$-expansion of a cycle of $K_{2,2}$-bags of length $\geqslant 3$;
+  • $\cdot$
+    a $C_{4}$-expansion of a quasi-5-connected 4-regular arc-transitive graph;
+  • $\cdot$
+    2-quasi-5-connected.
+Proof.
+Assume that $G$ is not 2-quasi-5-connected.
+Then $V(G)$ has a canonical partition $\mathcal{P}$ into vertex-sets that induce 4-cycles by Lemma˜13.47.
+Fix $U\in\mathcal{P}$ and let $\{\,U_{i}:i\in I\,\}$ be the set of classes $U_{i}\in\mathcal{P}$ that contain neighbours of $U$.
+Let $M$ denote the set of edges of $G$ with ends in different classes of $\mathcal{P}$, so $M$ is a perfect matching by Lemma˜13.47.
+Let $F\subseteq M$ denote the set of edges leaving $U$.
+Case: all four edges in $F$ end in the same $U_{i}$.
+Then $F=M$ and so $V(G)=U\cup U_{i}$.
+Both $U$ and $U_{i}$ induce 4-cycles, so either $G=C_{4}\square K_{2}$ which is a cycle of $C_{4}$-bags of length 4, or $G$ is the life belt.
+However, in both cases $G$ has two crossing tetra-cuts, so neither case occurs.
+Case: exactly three edges in $F$ end in the same $U_{i}$.
+This case is impossible: Then the vertex in $U$ and the vertex in $U_{i}$ that are not incident with these three edges form a 2-separator of $G$, contradicting that $G$ is internally-4-connected.
+Case: there is a $U_{i}$ in which exactly two edges from $F$ end.
+Then four edges from $M$ leave $U\cup U_{i}$.
+Hence $U\cup U_{i}$ is a side of a tetra-cut of $G$.
+Since this tetra-cut is totally-nested by assumption, it has a side that induces a 4-cycle by Lemma˜13.38.
+This side can only be $V(G)\smallsetminus(U\cup U_{i})$, which then must be a $U_{j}$ for some $j\in I\smallsetminus\{i\}$.
+Now $U$ and $U_{i}$ each send exactly two edges to $U_{j}$.
+If $G$ is a cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length 6, then we are done.
+Otherwise the 4-cycle $O$ induced by $U$, say, first traverses the two neighbours $u_{i},u_{i}^{\prime}$ of $U_{i}$ in $U$, and then traverses the two neighbours $u_{j},u_{j}^{\prime}$ of $U_{j}$ in $U$.
+But then (U,V(GU))(U,V(G-U)) and $(U_{i}\cup\{u_{i},u_{i}^{\prime}\},U_{j}\cup\{u_{j},u_{j}^{\prime}\})$ are two crossing tetra-cuts in $G$, contradicting our assumption.
+Case: $|I|=4$ and exactly one edge from $F$ ends in each $U_{i}$.
+Then $G/\mathcal{P}$ is 4-regular and 4-connected.
+Moreover, $G/\mathcal{P}$ is arc-transitive, as $\mathcal{P}$ is unique and every vertex of $G$ is incident with a unique edge in $F$.
+Then by Corollaries˜13.6 and 13.25, the graph $G/\mathcal{P}$ is one of the following:
+  • $\cdot$
+    the line-graph of the cube,
+  • $\cdot$
+    a cycle of $K_{2,2}$-bags of length $\geqslant 3$, or
+  • $\cdot$
+    essentially-5-connected.
+Assume first that $G/\mathcal{P}$ is the line-graph of the cube, so $G$ is some $C_{4}$-expansion thereof.
+Since $G$ is vertex-transitive, there are only three possible $C_{4}$-expansions: these are precisely the three depicted in Figure˜1.15.
+Assume finally that $G/\mathcal{P}$ is essentially-5-connected.
+Then $G/\mathcal{P}$ is quasi-5-connected by Corollary˜13.28.
+Hence $G$ is the $C_{4}$-expansion of a quasi-5-connected 4-regular arc-transitive graph.
+∎
+Proposition 13.49.
+A 3-regular quasi-4-connected finite graph $G$ is vertex-transitive if and only if $G$ is either
+  1. (i)
+    2-quasi-5-connected and vertex-transitive,
+  2. (ii)
+    a vertex-transitive $C_{4}$-expansion of a quasi-5-connected 4-regular arc-transitive graph,
+  3. (iii)
+    a cycle of $C_{4}$-bags of length $\geqslant 4$ or $K_{3}\square K_{2}$,
+  4. (iv)
+    a cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length $\geqslant 6$,
+  5. (v)
+    one of the three $C_{4}$-expansions of the line-graph of the cube shown in Figure˜1.15, or
+  6. (vi)
+    a vertex-transitive $C_{4}$-expansion of a cycle of $K_{2,2}$-bags of length $\geqslant 3$.
+Proof.
+($\Rightarrow$).
+By ˜13.30, we may assume that $G$ is internally 4-connected and has girth $\geqslant 4$.
+If $G$ has two crossing tetra-cuts, then we are done by ˜13.44.
+Otherwise every tetra-cut of $G$ is totally-nested.
+Then we are done by ˜13.48.
+($\Leftarrow$) is straightforward. ∎
+Proof of Figure˜1.8.
+Proposition 13.50.
+[CK23, Corollary 2]. Every vertex-transitive finite connected $G$ is either
+  • $\cdot$
+    3-connected and every non-trivial strong tri-separation of $G$ is a 3-edge-cut with a side that induces a triangle,
+  • $\cdot$
+    a cycle, $K_{2}$ or $K_{1}$.
+Proposition 13.51.
+Every vertex-transitive finite connected graph $G$ is either
+  • $\cdot$
+    internally 4-connected and vertex-transitive,
+  • $\cdot$
+    the $K_{3}$-expansion of a quasi-4-connected 4-regular arc-transitive graph,
+  • $\cdot$
+    a cycle, $K_{3}\square K_{2}$, $K_{2}$ or $K_{1}$.
+Proof.
+By Proposition˜13.50, we may assume that $G$ is 3-connected and every non-trivial strong tri-separation of $G$ is a 3-edge-cut with a side that induces a triangle.
+If $G$ has no non-trivial strong tri-separation, then $G$ is internally-4-connected by [CK23, 1.2.8].
+Hence we may assume that $G$ has at least one non-trivial strong tri-separation, and by assumption it has a side that induces a triangle.
+Let $\{\,(A_{i},B_{i}):i\in I\,\}=\sigma$ be the set of all non-trivial strong tri-separations of $G$ with $G[A_{i}]=K_{3}$.
+Since each $S(A_{i},B_{i})$ is a 3-edge-cut, the tri-separations $(A_{i},B_{i})$ are totally-nested by [CK23, 1.3.10], so $\sigma$ is a star of tri-separations.
+Hence the $A_{i}$ are pairwise disjoint, and every vertex of $G$ is contained in some $A_{i}$ by vertex-transitivity, so $\mathcal{P}:=\{\,A_{i}:i\in I\,\}$ is a partition of $V(G)$.
+Since each $S(A_{i},B_{i})$ is a 3-matching, the set $M$ of cross-edges of $\mathcal{P}$ is a perfect matching.
+If two edges in $M$ join $A_{i}$ to $A_{j}$ for some $i\neq j\in I$, then the two vertices in $A_{i}\cup A_{j}$ that are not incident to these edges form a 2-separator of $G$ contradicting 3-connectivity, unless $G=K_{3}\square K_{2}$.
+Hence we may assume that at most one edge in $M$ joins any $A_{i}$ to any $A_{j}$, so $G/\mathcal{P}$ is 3-regular.
+Since $G$ is vertex-transitive and the $K_{3}$-expansion of $G/\mathcal{P}$, it follows that $G/\mathcal{P}$ is arc-transitive by Lemma˜13.21.
+In particular, $G/\mathcal{P}$ is vertex-transitive, and since $G/\mathcal{P}$ is 3-regular we get that $G/\mathcal{P}$ is 3-connected by Proposition˜13.50.
+However, $G/\mathcal{P}$ has no non-atomic 3-edge-cut, as this would yield a non-trivial strong tri-separation of $G$ that is a 3-edge-cut but whose sides have size $\geqslant 9$, contradicting our assumption.
+Hence applying Proposition˜13.50 to $G/\mathcal{P}$ in combination with [CK23, 1.2.8] yields that $G/\mathcal{P}$ is quasi-4-connected.
 ∎
 Acknowledgement.
-We thank Joseph Devine for exciting discussions. We thank Agelos Georgakopoulos for telling us about Theorem 12.1 of Godsil and Royle.
-Appendix A Circular saws
+We thank a referee for valuable comments on the introduction.
+We thank another referee for bringing [LPPS22] to our attention.
+We thank Joseph Devine for exciting discussions.
+We thank Agelos Georgakopoulos for telling us about Theorem˜13.1 of Godsil and Royle.
+14. Appendix A: Circular saws
 Here we verify that the circular saws from the introduction have the desired properties.
-Lemma A.1.
+Lemma 14.1.
 Let $G$ be a $k$-regular circular saw on $\mathbb{Z}_{n}\times\mathbb{Z}_{2}$ with $n\gg k$.
@@
 Let $X$ denote the neighbourhood of $\{a,a^{\prime}\}$, and let $Y$ denote the neighbourhood of $\{b,b^{\prime}\}$.
-There are $2(k-1)$ disjoint $X$$Y$ paths in $G$.
+There are 2(k1)2(k-1) disjoint XXYY paths in $G$.
 Proof.
-For each vertex $x\in X$ we construct two $x$$Y$ paths $P^{+}_{x},P^{-}_{x}$ inductively, as follows. See Figure A.1 for an example.
+For each vertex $x\in X$ we construct two xxYY paths $P^{+}_{x},P^{-}_{x}$ inductively, as follows. See Figure˜14.1 for an example.
 In the first step of the construction of $P^{+}_{x}$, we let $P^{+}_{x}$ start and end with $x$.
@@
 Let $P_{x}$ denote the other path.
-Then $(P_{x}:x\in X)$ is a family of disjoint $X$$Y$ paths.
-There are $|X|=2(k-1)$ of these paths.
+Then $(P_{x}:x\in X)$ is a family of disjoint XXYY paths.
+There are |X|=2(k1)|X|=2(k-1) of these paths.
 ∎
-Figure A.1. Paths constructed in the proof of Lemma A.1. The edges $aa^{\prime}$ and $bb^{\prime}$ are coloured blue, the paths are coloured red.
-Lemma A.2.
+Figure 14.1. Paths constructed in the proof of Lemma˜14.1. The edges $aa^{\prime}$ and $bb^{\prime}$ are coloured blue, the paths are coloured red.
+Lemma 14.2.
 Let $G$ be a $k$-regular circular saw on $\mathbb{Z}_{n}\times\mathbb{Z}_{2}$ with $n\gg k\geqslant 3$.
@@
 Proof.
-We show first that $G$ is $k$-connected, and assume for a contradiction that some two vertices $a,b$ of $G$ are separated by a set $S$ of $k-1$ vertices.
+We show first that $G$ is $k$-connected, and assume for a contradiction that some two vertices $a,b$ of $G$ are separated by a set $S$ of k1k-1 vertices.
 Let $a^{\prime}$ be a neighbour of $a$ outside $S$, and let $b^{\prime}$ be a neighbour of $b$ outside $S$.
-Note that $12(k-1)\geqslant k+$ 1 since $k\geqslant 3$.
-Hence by Lemma A.1, there are $k+1$ disjoint $N(\{a,a^{\prime}\})$$N(\{b,b^{\prime}\})$ paths in $G$, contradicting that $|S|\leqslant k-1$.
+Note that $2(k-1)\geqslant k+1$ since $k\geqslant 3$.
+Hence by Lemma˜14.1, there are $k+1$ disjoint $N(\{a,a^{\prime}\})$$N(\{b,b^{\prime}\})$ paths in $G$, contradicting that $|S|\leqslant k-1$.
 Hence $G$ is $k$-connected.
@@
 Then there are vertices $a^{\prime}\in N(a)\smallsetminus S(A,B)$ and $b^{\prime}\in N(b)\smallsetminus S(A,B)$.
-Now Lemma A.1 yields $k+1$ disjoint $N(\{a,a^{\prime}\})$$N(\{b,b^{\prime}\})$ paths in $G$, contradicting that $(A,B)$ has order $k$.
+Now Lemma˜14.1 yields $k+1$ disjoint $N(\{a,a^{\prime}\})$$N(\{b,b^{\prime}\})$ paths in $G$, contradicting that $(A,B)$ has order $k$.
 ∎
-Appendix B Comparison of strict tri-separations and tri-separations
-The strict tri-separation-decomposition (2) is different from the tri-separation decomposition [Tridecomp, Theorem 1] as can be seen in Figures B.1 and B.2.
-However, the strict tri-separation-decomposition can be viewed as a refinement of the tri-separation decomposition via Lemma B.1.
-Figure B.1. The totally-nested strict tri-separations
-Figure B.2. The totally-nested non-trivial tri-separations
+15. Appendix B: Comparison of strict tri-separations and tri-separations
+The strict tri-separation-decomposition (˜2) is different from the tri-separation decomposition [CK23, Theorem 1] as can be seen in Figures˜15.1 and 15.2.
+However, the strict tri-separation-decomposition can be viewed as a refinement of the tri-separation decomposition via Lemma˜15.1.
+Figure 15.1. The totally-nested strict tri-separations
+Figure 15.2. The totally-nested non-trivial tri-separations
 A tri-separation $(A,B)$ is negligible if $|A\smallsetminus B|\leqslant 1$ or $|B\smallsetminus A|\leqslant 1$.
-Lemma B.1.
+Lemma 15.1.
 Let $G$ be a 3-connected graph.
 Let $(A,B)$ be a totally-nested non-trivial tri-separation of $G$.
-Then the left-right-shift $(A^{\prime},B^{\prime})$ of $(A,B)$ is a totally-nested strict tri-separation, or $(A,B)$ is negligible.
+Then the left-right-reduction $(A^{\prime},B^{\prime})$ of $(A,B)$ is a totally-nested strict tri-separation, or $(A,B)$ is negligible.
 Proof.
 Assume that $(A,B)$ is not negligible.
-Then $(A^{\prime},B^{\prime})$ is a strict tri-separation by 5.10 (which generalises to strict tri-separations via an apex-argument as in the proof of Corollary 3.2).
+Then $(A^{\prime},B^{\prime})$ is a strict tri-separation by ˜5.10 (which generalises to strict tri-separations via an apex-argument as in the proof of Corollary˜3.2).
 Assume for a contradiction that there is a strict tri-separation $(C,D)$ of $G$ that crosses $(A^{\prime},B^{\prime})$ with $C$-link $\gamma^{\prime}$ and $D$-link $\delta^{\prime}$.
@@
 For this, we consider the crossing-diagram of $(C,D)$ and $(A^{\prime},B^{\prime})$.
-By [Tridecomp, Lemma 1.3.10], there are two cases.
+By [CK23, Lemma 1.3.10], there are two cases.
 Case: the centre consists of one vertex and all links have size one.
@@
 But all corners are non-empty, so $(A^{\prime},B^{\prime})$ is not half-connected.
-However, $(A,B)$ is half-connected because it is totally-nested, by [Tridecomp, Lemma 2.3.4], contradicting that $(A^{\prime},B^{\prime})=(A,B)$.
+However, $(A,B)$ is half-connected because it is totally-nested, by [CK23, Lemma 2.3.4], contradicting that $(A^{\prime},B^{\prime})=(A,B)$.
 ∎
-[…]

### v3

@@
 This decomposition is similar yet different from the tri-separation decomposition [CK23].
-As an application of the decomposition for 4-connectivity, we obtain a new theorem characterising all quasi-4-connected vertex-transitive finite graphs as essentially quasi-5-connected or on a short explicit list of graphs.
+As an application of the decomposition for 4-connectivity, we obtain a new theorem [KP26] characterising all vertex-transitive finite connected graphs as essentially quasi-5-connected or on a short explicit list of graphs.
 Key words and phrases:
-4-connected, decomposition, Tutte, canonical, double wheel, 4-separation, 4-separator, tetra-separation, tetra-separator
+4-connected, decomposition, Tutte, canonical, double wheel, quasi-5-connected, 4-separation, 4-separator, tetra-separation, tetra-separator
 2020 Mathematics Subject Classification:
@@
 Indeed, different 2-separators can cross in that they cut through each other, as happens in a cycle.
-Then if we cut at one of the 2-separators, we loose the other.
+Then if we cut at one of the 2-separators, we lose the other.
 Tutte realised that the solution is to cut the graph at all 2-separators that are totally-nested in that they are not crossed by any 2-separators, and indeed no conflicts arise when cutting in this way.
@@
 For example, if $G$ is $k$-connected and the neighbourhoods of its vertices are precisely the $k$-separators of $G$, then the $k$-separators around neighbouring vertices cross, but $G$ fails to be $(k+1)$-connected.
-Such $G$ exist for all $k\geqslant 3$: take a circular saw as in Figure˜1.1, for instance.
+Such $G$ exist for all $k\geqslant 3$: take a circular saw as in fig.˜1.1, for instance.
 We believe that such graphs should not have to be decomposed any further at level $k$.
 Figure $1.$ A circular saw has vertex-set $\mathbb{Z}_{n}\times\mathbb{Z}_{2}$ with $n\gg k$ and every vertex $(v,0)$ sends edges to the $k$ vertices $(v,1),\ldots,(v+k-1,1)$.
-It is $k$-connected, and when $k\geqslant 3$ its neighbourhoods of vertices are precisely its $k$-separators (Lemma˜14.2).
+It is $k$-connected, and when $k\geqslant 3$ its neighbourhoods of vertices are precisely its $k$-separators (lemma˜13.2).
 Challenge 2.
-We consider the $k$-connected graph $G$ depicted in Figure˜1.2 if $k$ is odd and in Figure˜1.3 if $k$ is even.
+We consider the $k$-connected graph $G$ depicted in fig.˜1.2 if $k$ is odd and in fig.˜1.3 if $k$ is even.
 Every $k$-separator of $G$ is formed from a red set, a non-adjacent blue set, and if $k$ is even also the apex-vertex in the centre.
@@
 A tetra-separation of $G$ is totally-nested if it is nested with every tetra-separation of $G$.
-Figure˜1.4 shows the totally-nested tetra-separations of an example-graph similar to the one shown in Figure˜1.3.
+fig.˜1.4 shows the totally-nested tetra-separations of an example-graph similar to the one shown in fig.˜1.3.
 Figure 1.4. The totally-nested tetra-separations of a graph. Their separators are coloured.
@@
 In order to state our decomposition-result for 4-connected graphs, we need the following definitions to describe the ‘basic’ parts of the decomposition – those that are not quasi-5-connected.
-Graph-decompositions generalise tree-decompositions in that the decomposition-tree may take the form of an arbitrary graph, such as a cycle, in which case we speak of a cycle-decomposition [DJKK22].
+Graph-decompositions generalise tree-decompositions in that the decomposition-tree may take the form of an arbitrary graph, such as a cycle, in which case we speak of a cycle-decomposition [RRPJ26].
 Let $X$ be a graph.
-We call $X$ a generalised double-wheel with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that XuvX-u-v has a cycle-decomposition of adhesion 1 into $K_{2}$’s and triangles (Figure˜1.5).
-We call $X$ a cycle of $X_{1}$-torsos, …, $X_{k}$-torsos, $Y_{1}$-bags, …, $Y_{\ell}$-bags where the $X_{i}$ and $Y_{j}$ are graphs, if $X$ is 4-connected and has a cycle-decomposition with all adhesion-sets of size 2 such that for every bag either its torso is isomorphic to some $X_{i}$ or the bag itself is isomorphic to some $Y_{j}$ (Figure˜1.6).
+We call $X$ a generalised double-wheel with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that XuvX-u-v has a cycle-decomposition of adhesion 1 into $K_{2}$’s and triangles (fig.˜1.5).
+We call $X$ a cycle of $X_{1}$-torsos, …, $X_{k}$-torsos, $Y_{1}$-bags, …, $Y_{\ell}$-bags where the $X_{i}$ and $Y_{j}$ are graphs, if $X$ is 4-connected and has a cycle-decomposition with all adhesion-sets of size 2 such that for every bag either its torso is isomorphic to some $X_{i}$ or the bag itself is isomorphic to some $Y_{j}$ (fig.˜1.6).
 Recall that the only 3-connected graphs on $\leqslant 5$ vertices are $K_{4}$, the $4$-wheel, $K_{5}$ minus an edge, and $K_{5}$.
@@
 This seemingly subtle difference has a surprisingly big impact: the verbatim extension of the tri-separation approach to 4-connectivity fails ˜2.
-For 3-connected graphs, the strict tri-separation decomposition can be viewed as a refinement of the tri-separation decomposition, but the details are a bit technical; see Section˜15.
-Algorithmic aspects. It is straightforward to come up with a polynomial-time algorithm that produces the decompositions in Theorem˜1 and ˜2 for all 4- and 3-connected graphs.
-Korhonen [K24] recently showed that there is an algorithm that, given an $n$-vertex $m$-edge graph $G$ and an integer $k\geqslant 1$, returns a $k$-lean tree-decomposition of $G$ in time $k^{\mathcal{O}(k^{2})}n+\mathcal{O}(m)$.
-We believe that every 5-lean tree-decomposition of every 4-connected graph $G$ can be made into a tetra-separation decomposition of $G$ as in Theorem˜1 in linear time:
+For 3-connected graphs, the strict tri-separation decomposition can be viewed as a refinement of the tri-separation decomposition, but the details are a bit technical; see section˜14.
+Algorithmic aspects. It is straightforward to come up with a polynomial-time algorithm that produces the decompositions in Theorem˜1 and ˜2 for all 4- and 3-connected finite graphs.
+Korhonen [K25] recently showed that there is an algorithm that, given an $n$-vertex $m$-edge finite graph $G$ and an integer $k\geqslant 1$, returns a $k$-lean tree-decomposition of $G$ in time $k^{\mathcal{O}(k^{2})}n+\mathcal{O}(m)$.
+We believe that every 5-lean tree-decomposition of every 4-connected finite graph $G$ can be made into a tetra-separation decomposition of $G$ as in Theorem˜1 in linear time:
 Conjecture 1.4.
-There is an algorithm that, given a 4-connected $n$-vertex $m$-edge graph $G$, returns the tetra-separation decomposition of $G$ from Theorem˜1 in linear time $\mathcal{O}(n+m)$.
-A positive answer would also provide a linear-time algorithm for the decomposition of 3-connected graphs in ˜2 by following the short proof of ˜2.
+There is an algorithm that, given a 4-connected $n$-vertex $m$-edge finite graph $G$, returns the tetra-separation decomposition of $G$ from Theorem˜1 in linear time $\mathcal{O}(n+m)$.
+A positive answer would also provide a linear-time algorithm for the decomposition of 3-connected finite graphs in ˜2 by following the short proof of ˜2.
 Total-nestedness makes parallel computing a feasible option for computing the tetra-separation decomposition even more efficiently.
@@
 Let $H$ be an arbitrary graph.
-The following definition of $H^{\Delta}$ is supported by Figure˜1.8.
+The following definition of $H^{\Delta}$ is supported by fig.˜1.8.
 Let $U$ be the set of all vertices in $H$ of degree three.
@@
     Finally, we decompose each 4-connected $\tau^{\Delta}$ with Theorem˜1.
-The operation $\tau\mapsto\tau^{\Delta}$ harmonises well with the other decomposition steps. Indeed, we have followed all steps to prove Figure˜1.8 below.
-Vertex-transitive graphs.
-Recall that a graph $G$ is arc-transitive if for every two ordered pairs $(u_{i},v_{i})$ with $u_{i}v_{i}\in E(G)$ for $i=1,2$ there is an automorphism $\varphi$ of $G$ such that $\varphi(u_{1})=u_{2}$ and $\varphi(v_{1})=v_{2}$.
-Let $H$ be an $r$-vertex graph and let $G$ be an $r$-regular graph.
-An $H$-expansion of $G$ is a graph obtained from $G$ by replacing the vertices $v$ of $G$ with pairwise disjoint copies $H_{v}$ of $H$ and by replacing every edge $e=uv$ of $G$ with an $H_{u}$$H_{v}$ edge $\hat{e}$ so that $\hat{e}_{1}$ and $\hat{e}_{2}$ share no ends for every two $e_{1}\neq e_{2}\in E(G)$.
-Thus, the edges $\hat{e}$ form a perfect matching of the $H$-expansion.
-It was mostly222[CK23, Corollary 2] is slightly weaker than the result we mention here.
-The slight strengthening happens in Proposition 13.51 in this paper.
-shown in [CK23] that a finite connected graph $G$ is vertex-transitive if and only if $G$ is either
-  • $\cdot$
-    quasi-4-connected and vertex-transitive,
-  • $\cdot$
-    the $K_{3}$-expansion of a quasi-4-connected 4-regular arc-transitive graph,
-  • $\cdot$
-    a cycle, $K_{2}$ or $K_{1}$.
-[…]
-A graph $G$ is 2-quasi-5-connected if $G$ is quasi-4-connected and every 4-separation $(A,B)$ of $G$ satisfies $|A\smallsetminus B|\leqslant 2$ or $|B\smallsetminus A|\leqslant 2$.
-For other definitions related to Figure˜1.8, see Section˜13.
-{restatable}
-mainresultmaintransitive A quasi-4-connected finite graph $G$ is vertex-transitive if and only if $G$ either is
-  1. (i)
-    a $(\geqslant 4)$-regular quasi-5-connected vertex-transitive graph,
-  2. (ii)
-    a 3-regular 2-quasi-5-connected vertex-transitive graph,
-  3. (iii)
-    a $K_{4}$-expansion of a quasi-5-connected 4-regular arc-transitive graph, or a vertex-transitive $C_{4}$-expansion of a quasi-5-connected 4-regular arc-transitive graph,
-  4. (iv)
-    a cycle of $K_{4}$-bags of length $\geqslant 4$ (Figure˜1.9),
-  5. (v)
-    a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 4$ (Figure˜1.10),
-  6. (vi)
-    a cycle of triangle-bags of length $\geqslant 6$ (Figure˜1.12),
-  7. (vii)
-    a cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length $\geqslant 6$ (Figure˜1.11),
-  8. (viii)
-    a cycle of $C_{4}$-bags of length $\geqslant 4$ (Figure˜1.13) or $K_{3}\square K_{2}$,
-  9. (ix)
-    a cycle of $K_{2,2}$-bags of length $\geqslant 3$, the $K_{4}$-expansion thereof, or a vertex-transitive $C_{4}$-expansion thereof (Figure˜1.14), or
-  10. (x)
-    the line-graph of the cube, its $K_{4}$-expansion, or one of the three $C_{4}$-expansions of it depicted in Figure˜1.15.
-Figure 1.9.
-A cycle of $K_{4}$-bags
-Figure 1.10.
-A cycle alternating between $K_{4}$-bags and $C_{4}$-bags
-Figure 1.11.
-A cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos
-Figure 1.12.
-Cycles of triangle-bags
-Figure 1.13.
-Cycles of $C_{4}$-bags
-Figure 1.14. A cycle of $K_{2,2}$-bags, its $K_{4}$-expansion, and two vertex-transitive $C_{4}$-expansions
-Figure 1.15. The line-graph of the cube, its $K_{4}$-expansion, and its vertex-transitive $C_{4}$-expansions
+The operation $\tau\mapsto\tau^{\Delta}$ harmonises well with the other decomposition steps.
+Indeed, we have followed all steps to classify all vertex-transitive finite graphs of low connectivity [KP26].
 When canonicity and an explicit description matter.
@@
 Examples of these applications include:
+  1. $\cdot$
+[…]
   2. $\cdot$
@@
     Finite groups.
-The proof of Figure˜1.8 crucially relies on canonicity and total-nestedness.
-In a different direction, Stallings’ theorem applies tree-decomposition methods to Cayley graphs to detect product structure in groups.
+Stallings’ theorem applies tree-decomposition methods to Cayley graphs to detect product structure in groups.
 However, Stallings’ approach only works for infinite groups.
@@
 We say that $(A,B)$ interlaces $\tau$ if, roughly speaking, both $A\smallsetminus B$ and $B\smallsetminus A$ intersect $\tau$.
-Here may think of $(A,B)$ as ‘cutting through’ $\tau$.
+Here we may think of $(A,B)$ as ‘cutting through’ $\tau$.
 Whether $\tau$ is interlaced or not already yields a first outcome:
@@
 Assume now that $|Z|=0$.
-As our first step, we will use $(A,B)$ and $(C,D)$ to cut $G$ into four 2-connected pieces, arranged in the shape of a 4-cycle, and apply the Tutte-decomposition to each piece; see Figure˜1.16.
+As our first step, we will use $(A,B)$ and $(C,D)$ to cut $G$ into four 2-connected pieces, arranged in the shape of a 4-cycle, and apply the Tutte-decomposition to each piece; see fig.˜1.9.
 Since $G$ is 4-connected, the Tutte-decompositions will be path-decompositions.
@@
     3 independent paths between the two elements of $\pi_{1}$ in $G-\pi_{2}$ on the other hand.
-A major difference to tri-separations is that the connectivity-properties required of the separators are much more complex, as tetra-separations can cross in significantly more complicated ways than tri-separations; see Figure˜1.17 for an example.
+A major difference to tri-separations is that the connectivity-properties required of the separators are much more complex, as tetra-separations can cross in significantly more complicated ways than tri-separations; see fig.˜1.10 for an example.
 Figure 1.10. Three tetra-separations cross in a way that tri-separations cannot.
@@
 To this end, we develop a number of lemmata for finding such paths systematically, and then combine them to prove ˜3.
-Finding the paths takes a lot of work, especially when $O$ is short; see Figure˜1.18 for example.
+Finding the paths takes a lot of work, especially when $O$ is short; see fig.˜1.11 for example.
 Figure 1.11.
@@
 Proving ˜5, however, is a little tricky.
-We will assume for a contradicting that ˜5 fails, so some totally-nested tetra-separation $(E,F)$ will ‘cut through’ $\mathcal{O}$.
-Since $(E,F)$ is nested with both $(A,B)$ and $(C,D)$, we have $E\subseteq A\cap C$, say; see Figure˜1.19.
+We will assume for a contradiction that ˜5 fails, so some totally-nested tetra-separation $(E,F)$ will ‘cut through’ $\mathcal{O}$.
+Since $(E,F)$ is nested with both $(A,B)$ and $(C,D)$, we have $E\subseteq A\cap C$, say; see fig.˜1.12.
 We then start adjusting both $(A,B)$ and $(C,D)$: by moving bags of $\mathcal{O}$ from $A\smallsetminus B$ to $B\smallsetminus A$ and from $C\smallsetminus D$ to $D\smallsetminus C$, one bag at a time, while using the total-nestedness of $(E,F)$ to maintain the inclusion $E\subseteq A\cap C$.
@@
 […]
+Mixed-separators (without additional conditions on their vertices) have been used before, see for example Mader’s theorem [M78, R24] or [BH67, JKS24, HHSS20].
 […]
@@
 2. Essential terminology
-This section introduces all terminology used by Theorem˜1. For graph-theoretic terminology we follow [D24a].
+This section introduces all terminology used by Theorem˜1. For graph-theoretic terminology we follow [R24].
 2.1. Tetra-separations
@@
 Example 2.4.
-The lower-bound $|G|\geqslant 8$ in Lemma˜2.3 (ii) is tight: the graph depicted in Figure˜2.1 is quasi-5-connected, has $7$ vertices, but has a tetra-separation.
+The lower-bound $|G|\geqslant 8$ in lemma˜2.3 (ii) is tight: the graph depicted in fig.˜2.1 is quasi-5-connected, has $7$ vertices, but has a tetra-separation.
 Figure 2.1. A small quasi-5-connected graph with a tetra-separation (red)
@@
 2.2. Stars, mixed-tree-decompositions and torsos
-The following definition is supported by Figure˜2.2.
+The following definition is supported by fig.˜2.2.
 Let $G$ be a graph.
@@
 by making each of the vertex-sets $\lambda(\alpha_{\mathcal{T}}(s,t))$ into a clique.
-From this expanded-torso we obtain the compressed-torso, or torso for short, by contracting all edges contained in adhesion-sets of the edges $st\in E(T)$; see Figure˜2.3.
+From this expanded-torso we obtain the compressed-torso, or torso for short, by contracting all edges contained in adhesion-sets of the edges $st\in E(T)$; see fig.˜2.3.
 More generally, let $\sigma=\{\,(A_{i},B_{i}):i\in I\,\}$ be a star of mixed-separations+.
@@
 The two notions of torsos are compatible: the torso of a node $t$ of $T$ equals the torso of the star $\sigma_{t}$.
-Figure 2.3. The torsos of the mixed-tree-decomposition depicted in Figure˜$2.$
+Figure 2.3. The torsos of the mixed-tree-decomposition depicted in fig.˜$2.$
 2.3. Graph-decompositions
@@
 The outcome in theorem˜3.1 is not in general unique.
-For example, consider the graph $G$ depicted in Figure˜3.1 on the left.
+For example, consider the graph $G$ depicted in fig.˜3.1 on the left.
 Every tetra-separation of $G$ is crossed, hence $\sigma=\emptyset$ is a splitting star of the set $N(G)$ of all totally-nested tetra-separations of $G$.
-On the left of Figure˜3.1, we see two tetra-separations that interlace $\sigma$ and cross with empty vertex-centre.
+On the left of fig.˜3.1, we see two tetra-separations that interlace $\sigma$ and cross with empty vertex-centre.
 From their perspective, $G$ is a cycle of $K_{4}$-torsos and $4$-wheel torsos, see theorem˜3.1 ˜(ii).
-On the right of Figure˜3.1, we see two tetra-separations that interlace $\sigma$ and cross with a vertex-centre of size 2.
+On the right of fig.˜3.1, we see two tetra-separations that interlace $\sigma$ and cross with a vertex-centre of size 2.
 From their perspective, $G$ is a generalised wheel, see theorem˜3.1 ˜(iii).
-For another example, consider the graph $G$ depicted in Figure˜3.2 on the left.
+For another example, consider the graph $G$ depicted in fig.˜3.2 on the left.
 Again, $\sigma=\emptyset$ is a splitting star of $N(G)$.
-On the left of Figure˜3.1, we see two tetra-separations that interlace $\sigma$ and cross with empty vertex-centre.
+On the left of fig.˜3.2, we see two tetra-separations that interlace $\sigma$ and cross with empty vertex-centre.
 From their perspective, $G$ is a cycle of $K_{4}$-torsos, see theorem˜3.1 ˜(ii).
-On the right of Figure˜3.1, we see two tetra-separations that interlace $\sigma$ and cross with a vertex-centre of size 4.
+On the right of fig.˜3.2, we see two tetra-separations that interlace $\sigma$ and cross with a vertex-centre of size 4.
 From their perspective, $G$ is a $K_{4,4}$, see theorem˜3.1 ˜(iv).
@@
 In this section, we characterise the ways in which tetra-separations can cross. The main result here is the (Crossing Lemma). (4.1).
-We require the following terminology for corner diagrams from [CK23, §1.3], which is supported by Figure˜4.1.
+We require the following terminology for corner diagrams from [CK23, §1.3], which is supported by fig.˜4.1.
 Let $(A,B)$ and $(C,D)$ be mixed-separations+ of a graph $G$.
@@
 The left blue edge dangles from the $A$-link through the $C$-link.
-The following definition is supported by Figure˜4.2.
+The following definition is supported by fig.˜4.2.
 The corner for $AC$ is potter if it is empty and its adjacent links have size two and there exists a path $u_{1}u_{2}u_{3}u_{4}$ in $G$ such that $u_{1}u_{2}$ dangles from the vertex $u_{2}$ in the $C$-link through the $A$-link, and $u_{3}u_{4}$ dangles from the vertex $u_{3}$ in the $A$-link through the $C$-link.
@@
 Moreover, there are no diagonal edges and no jumping edges. If there is an edge that dangles from one link through another, then we are in case (iii) and the corner in between the two links is potter.
-A mixed-separation $(A,B)$ is an edge-cut if $S(A\cap B)\subseteq E(G)$. An edge-cut $(A,B)$ is atomic if $|A|=1$ or $|B|=1$.
+A mixed-separation $(A,B)$ is an edge-cut if $S(A,B)\subseteq E(G)$. An edge-cut $(A,B)$ is atomic if $|A|=1$ or $|B|=1$.
 Lemma 4.2.
@@
 [CK23, Lemma 1.3.4] For two mixed-separations $(A,B)$ and $(C,D)$ of $G$, we have
-$|L(A,C)|+|L(B,D)|\leqslant|S(A,B)|+|S(C,D)|\leavevmode\nobreak\ .$
+$|L(A,C)|+|L(B,D)|\leqslant|S(A,B)|+|S(C,D)|\penalty 10000\ .$
 Moreover, if we have equality, there are no jumping edges, and every diagonal edge has its endvertices in the corners for $AC$ and $BD$.
@@
 Let $(A,B)$ and $(C,D)$ be crossing tetra-separations in a graph $G$.
-Assume that the corner for $AC$ is potter, witnessed by an edge $u_{1}u_{2}$ that dangles from the vertex $u_{2}$ in the $C$-link through the $A$-link, and by an edge $u_{3}u_{4}$ that dangles from the vertex $u_{3}$ in the $A$-link through the $C$-link; see Figure˜4.2.
+Assume that the corner for $AC$ is potter, witnessed by an edge $u_{1}u_{2}$ that dangles from the vertex $u_{2}$ in the $C$-link through the $A$-link, and by an edge $u_{3}u_{4}$ that dangles from the vertex $u_{3}$ in the $A$-link through the $C$-link; see fig.˜4.2.
 Then $\{u_{2},u_{3}u_{4}\}$ is $A$-weird and $\{u_{3},u_{1}u_{2}\}$ is $C$-weird.
@@
 Let $(C^{\prime},D^{\prime})$ be the left-right-reduction of $(C,D)$, which is a tetra-separation by ˜5.10.
-[…]
-The following definition is supported by Figure˜6.1.
+We claim that $(C^{\prime},D^{\prime})$ crosses $(A,B)$ so that $\pi_{1}$ and $\pi_{2}$ form opposite links and no corner is potter.
+Since $\hat{\pi}_{1}\subseteq C\smallsetminus D\subseteq C^{\prime}\smallsetminus D^{\prime}$ and $\hat{\pi}_{2}\subseteq D\smallsetminus C\subseteq D^{\prime}\smallsetminus C^{\prime}$ by ˜5.9, $\pi_{1}$ forms the $C^{\prime}$-link while $\pi_{2}$ forms the $D^{\prime}$-link.
+In particular, $(C^{\prime},D^{\prime})$ crosses $(A,B)$.
+Assume for contradiction that some corner is potter, say the corner $AC^{\prime}$.
+Then $\pi_{1}$ is $A$-weird and hence contains an edge $e$.
+Let $a$ be the endvertex of $e$ in $A\smallsetminus B$.
+So, $a\in\hat{\pi}_{1}\subseteq C^{\prime}\smallsetminus D^{\prime}$.
+Therefore, $a$ is contained in the corner $AC^{\prime}$.
+In particular, the corner $AC^{\prime}$ is not empty, and so it is not potter, a contradiction.
+∎
+The following definition is supported by fig.˜6.1.
 Definition 6.9 (1-potter-linked).
@@
 Note that this is not symmetric in $\pi_{1},\pi_{2}$.
-We say that $\{\pi_{1},\pi_{2}\}$ is 1-potter-linked with respect to $(A,B)$ if for both $X\in\{A,B\}$$\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $X$ and $\pi_{2}$ is 1-potter-linked to $\pi_{1}$ around $X$ with respect to $(A,B)$.
+We say that $\{\pi_{1},\pi_{2}\}$ is 1-potter-linked with respect to $(A,B)$ if for both $X\in\{A,B\}$, $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $X$ and $\pi_{2}$ is 1-potter-linked to $\pi_{1}$ around $X$ with respect to $(A,B)$.
 A tetra-separation $(A,B)$ is 1-potter-linked if $\{\pi_{1},\pi_{2}\}$ is 1-potter-linked for every balanced bipartition $\{\pi_{1},\pi_{2}\}$ of $S(A,B)$.
@@
 In particular, $|C_{1}\smallsetminus D_{1}|\geqslant 2$ and $|D_{1}\smallsetminus C_{1}|\geqslant 2$.
-See Figure˜6.2 for the next definition.
+See fig.˜6.2 for the next definition.
 We define $(C_{2},D_{2}):=(C_{1}\cup a_{1},(D_{1}\cup A)\smallsetminus\{v_{1}\})$.
@@
 Say (6.11.1) and (6.11.2) fail for $(X,Y)=(A,B)$; that is to say, $\pi_{1}$ is $A$-weird and $\pi_{2}$ is $B$-weird.
-For $i\in\{1,2\}$, write $\pi_{i}=:\{v_{i},e_{i}\}$ where $v_{i}$ is a vertex and $e_{i}=a_{i}b_{i}$ is an edge with $a_{i}\in A\smallsetminus B$ and $b_{i}\in B\smallsetminus A$; see Figure˜6.3.
+For $i\in\{1,2\}$, write $\pi_{i}=:\{v_{i},e_{i}\}$ where $v_{i}$ is a vertex and $e_{i}=a_{i}b_{i}$ is an edge with $a_{i}\in A\smallsetminus B$ and $b_{i}\in B\smallsetminus A$; see fig.˜6.3.
 Let $a^{\prime}_{1}$ be the unique neighbour of $v_{1}$ in $A\smallsetminus\{a_{1}\}$, and let $b_{2}^{\prime}$ be the unique neighbour of $v_{2}$ in $B\smallsetminus\{b_{1}\}$.
@@
 Proof.
-We combine Lemma˜6.3, Lemma˜6.5 and Corollary˜6.14 with the (Crossing Lemma). (4.1). ∎
 7. Thickened $K_{4,m}$’s
@@
 In particular, $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$.
-Then by Lemma˜5.4(A,B)(A,B) satisfies the matching-condition.
+Then by lemma˜5.4, $(A,B)$ satisfies the matching-condition.
 Let $v$ be a vertex in $S(A,B)$.
@@
 It really is a block-bagel of $G-Z^{\prime}$.
-The following setting is supported by Figure˜8.1.
+The following setting is supported by fig.˜8.1.
 Setting 8.4 (Block-bagel setting).
@@
 Otherwise the adhesion-vertex has degree $\leqslant 3$, contradicting 4-connectivity. ∎
-The following two lemmata are supported by Figure˜8.2.
+The following two lemmata are supported by fig.˜8.2.
 Lemma 8.8.
@@
 ∎
-The following lemma is supported by Figure˜8.3.
+The following lemma is supported by fig.˜8.3.
 Lemma 8.12 (Two-Paths Lemma).
@@
 Note that $G_{t}$ is bad if and only if either $G_{t}$ is a $K_{2}$ or $G_{t}$ is a triangle and both neighbouring bags are 2-connected.
-The following setting is supported by Figure˜8.4.
+The following setting is supported by fig.˜8.4.
 Setting 8.14 (Block-bagel bag setting).
@@
 For each helpful $h\in V(O)\cup E(O)$ we construct a $z^{\prime}_{1}$$z^{\prime}_{2}$ path $P_{h}$, as follows.
-If $h$ is an edge, then its adhesion-vertex send edges to both $z^{\prime}_{i}$ by Lemma˜8.7, and we let $P_{h}$ be the path defined by these two edges; see Figure˜8.5.
+If $h$ is an edge, then its adhesion-vertex send edges to both $z^{\prime}_{i}$ by lemma˜8.7, and we let $P_{h}$ be the path defined by these two edges; see fig.˜8.5.
 If $h$ is a helpful node, then we use the (Bag-Path Lemma).
@@
 Figure 8.5. The red paths are constructed in lemma˜8.19.
-The following lemma is supported by Figure˜8.6.
+The following lemma is supported by fig.˜8.6.
 Lemma 8.20.
@@
 Proof.
-[…]
-The following two lemmata are supported by Figure˜8.7.
+By ˜8.27, $\sigma(\mathcal{O})$ is a set of totally-nested tetra-separations.
+It remains to show that for all good nodes $s\neq t$, we have $(U_{s},W_{s})\leqslant(W_{t},U_{t})$.
+We have $U_{s}\subseteq V(G_{s})$ and $U_{t}\subseteq V(G_{t})$.
+Moreover, by ˜8.15, for all nodes $r\in V(O-s)$ that have a 2-connected bag, we have $V(G_{r})\subseteq W_{s}$.
+Therefore, $U_{t}\subseteq V(G_{t})\subseteq W_{s}$.
+Similarly, $U_{s}\subseteq V(G_{s})\subseteq W_{t}$.
+It follows that $(U_{s},W_{s})\leqslant(W_{t},U_{t})$.
+∎
+The following two lemmata are supported by fig.˜8.7.
 Lemma 8.30 (Link-Moving Lemma for $K_{2}$-bags).
@@
 Since $B\smallsetminus A\subseteq B^{\prime}\smallsetminus A^{\prime}$, every vertex in $S(A^{\prime},B^{\prime})$ has at least two neighbours in $B^{\prime}\smallsetminus A^{\prime}$.
-If $d$ is a vertex, then we claim that every neighbour $u$ of $d$ in $A\smallsetminus B$ also also lies in $A^{\prime}\smallsetminus B^{\prime}$.
+If $d$ is a vertex, then we claim that every neighbour $u$ of $d$ in $A\smallsetminus B$ also lies in $A^{\prime}\smallsetminus B^{\prime}$.
 Indeed, since $ud$ is not a dangling edge, $u$ is contained in the corner $AD$ or the $A$-link with respect to $(A,B),(C,D)$.
@@
 We obtain the graph $\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{G^{*}}$} from $G$ by adding edges, as follows.
-For each $L$-link $\{x,y\}$ and each adjacent corner $K=LS$ (where $L,S$ are sides), we join the unique vertices in $\hat{x}\cap S$ and $\hat{y}\cap S$ by an edge, which we call the $K$-edge for the $L$-link; see Figure˜9.1.
+For each $L$-link $\{x,y\}$ and each adjacent corner $K=LS$ (where $L,S$ are sides), we join the unique vertices in $\hat{x}\cap S$ and $\hat{y}\cap S$ by an edge, which we call the $K$-edge for the $L$-link; see fig.˜9.1.
 We will only use $G^{*}$ when $(A,B)$ and $(C,D)$ are clear from context.
@@
 We will show that $AC$ is potter.
-If both endvertices of $e$ lie in the corner for $AC$, then each link adjacent to $AC$ consists of a matching of size two with the endvertices of $e$ as one side of the matching; see Figure˜9.2.
+If both endvertices of $e$ lie in the corner for $AC$, then each link adjacent to $AC$ consists of a matching of size two with the endvertices of $e$ as one side of the matching; see fig.˜9.2.
 As $G$ is 4-connected, the corner for $AC$ cannot contain any other vertices besides the endvertices of $e$.
@@
 9.3. Settings
-The following settings are supported by Figure˜9.4.
+The following settings are supported by fig.˜9.4.
 Setting 9.12 (Tutte-bagel setting).
@@
 Moreover, the torso $H_{t(i)}$ is not a 4-cycle.
-We consider two cases, supported by Figure˜9.5.
+We consider two cases, supported by fig.˜9.5.
 Suppose first that the torso at $t(i)$ is 3-connected.
@@
 Case $i<j$.
-This case is supported by Figure˜9.6.
+This case is supported by fig.˜9.6.
 If $H_{t(i)}$ is 3-connected, we let $F_{y}$ by a $y^{W}_{1}$$\pi_{i}$ fan of two paths in $G_{t(i)}$ avoiding the vertex in $\pi_{i-1}-y^{W}_{1}$.
@@
 Case $i\geqslant j$.
-This case is supported by Figure˜9.7.
+This case is supported by fig.˜9.7.
 We claim that the corner $BD$ is not potter, and assume for a contradiction that it is.
@@
 Recall that $uP\mathring{v}$ denotes the subpath of $P$ that starts at $u$ and ends just before the vertex $v$.
-The statement of the following lemma is supported by Figure˜9.8.
+The statement of the following lemma is supported by fig.˜9.8.
 Lemma 9.44 (Link-Moving Lemma for vertices).
@@
 The number 7 in lemma˜12.2 is optimal. To see this, let $G$ be a cycle of $C_{4}$-bags of length 3. Then $G^{\Delta}$ is a generalised 4-wheel, which is 3-connected but not quasi-4-connected.
-13. Vertex-transitive graphs
-In this section, we prove Figure˜1.8.
-Recall that an $H$-expansion of a $|V(H)|$-regular graph $G$ is a graph obtained from $G$ by replacing the vertices $v$ of $G$ with pairwise disjoint copies $H_{v}$ of $H$ and by replacing every edge $e=uv$ of $G$ with an $H_{u}$$H_{v}$ edge $\hat{e}$ so that $\hat{e}_{1}$ and $\hat{e}_{2}$ share no ends for every two $e_{1}\neq e_{2}\in E(G)$.
-Thus, the edges $\hat{e}$ form a perfect matching of the $H$-expansion.
-Note that if $H$ is complete then the $H$-expansion of $G$ is unique up to isomorphism.
-The clique-expansion of a graph $G$ is the graph $G^{\prime}$ that is obtained from $G$ by replacing every vertex $v\in V(G)$ by a clique $C_{v}$ of size the degree of $v$, and every edge $e=uv$ by a $C_{u}$$C_{v}$ edge $\hat{e}$ such that these edges $\hat{e}$ form a perfect matching in $G^{\prime}$.
-Recall that a graph $G$ is arc-transitive if for every two ordered pairs $(u_{1},v_{1}),(u_{2},v_{2})$ of vertices forming edges $u_{1}v_{1},u_{2}v_{2}\in E(G)$, there is an automorphism $\varphi$ of $G$ such that $(\varphi(u_{1}),\varphi(v_{1}))=(u_{2},v_{2})$.
-Theorem 13.1.
-[GR01, Theorem 3.4.2] Every finite vertex-transitive $d$-regular graph is $k$-connected for $k:=\lceil\frac{2}{3}(d+1)\rceil$.
-Corollary 13.2.
-Every finite $d$-regular graph with $d\geqslant 4$ is 4-connected.∎
-A graph $G$ is essentially 5-connected if $G$ is 4-connected and every tetra-separation $(A_{1},A_{2})$ of $G$ has only edges in its separator and at least one side $A_{i}$ with $G[A_{i}]=K_{4}$.
-Definition 13.3.
-Let $G$ be a graph with a cycle-decomposition $\mathcal{O}=(O,(G_{t})_{t\in V(O)})$. Let $\ell$ denote the length of the cycle $O$. We say that $G$ is a
-  • $\cdot$
-    cycle of $K_{4}$-bags of length $\ell$ if all bags of $\mathcal{O}$ are $K_{4}$’s and all adhesion-sets of $\mathcal{O}$ are disjoint (Figure˜1.9);
-  • $\cdot$
-    cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\ell$ if the bags along $O$ alternate between being $K_{4}$’s and being 4-cycles, and all adhesion-sets of $\mathcal{O}$ are disjoint (Figure˜1.10);
-  • $\cdot$
-    cycle of triangle-bags of length $\ell$ if all bags are triangles, and every two non-consecutive edges on $O$ have disjoint adhesion-sets (Figure˜1.12);
-  • $\cdot$
-    cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length $\ell$ if the torsos along $O$ alternate between being $K_{4}$’s and $C_{4}$’s, and every adhesion-set of $\mathcal{O}$ is independent and all adhesion-sets of $\mathcal{O}$ are disjoint (so the $K_{4}$-torsos have $K_{2,2}$-bags) (Figure˜1.11);
-  • $\cdot$
-    cycle of $C_{4}$-bags of length $\ell$ if all bags of $\mathcal{O}$ are $C_{4}$’s and all adhesion-sets of $\mathcal{O}$ induce disjoint $K_{2}$’s (Figure˜1.13);
-  • $\cdot$
-    cycle of $K_{2,2}$-bags of length $\ell$ if all bags of $\mathcal{O}$ are $K_{2,2}$’s, all torsos of $\mathcal{O}$ are $K_{4}$’s and all adhesion-sets of $\mathcal{O}$ are disjoint (Figure˜1.14).
-Observation 13.4.
-  • $\cdot$
-    Every cycle of $K_{4}$-bags of length $\ell$ is isomorphic to $K_{2}\boxtimes C_{\ell}$.
-  • $\cdot$
-    Every cycle of $K_{4}$-bags of length 3 is isomorphic to $K_{6}$.
-  • $\cdot$
-    Every cycle of $K_{2,2}$-bags of length 3 is isomorphic to the octahedron.
-  • $\cdot$
-    Every cycle of $K_{2,2}$-bags of length 4 is isomorphic to $K_{4,4}$.
-  • $\cdot$
-    Every cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\ell\in\{4,6\}$ is essentially 5-connected.
-  • $\cdot$
-    Every cycle of triangle-bags of length $\ell$ is isomorphic to $C_{\ell}^{2}$ (the graph obtained from the $\ell$-cycle $C_{\ell}$ by joining every vertex $v$ to all vertices of distance at most 2 from $v$ on $C_{\ell}$).
-  • $\cdot$
-    Every cycle of triangle-bags of length 5 is isomorphic to $K_{5}$. ∎
-Lemma 13.5.
-Let $G$ be a vertex-transitive 4-connected finite graph, and let $(A_{1},A_{2})$ be a totally-nested tetra-separation in $G$.
-Then $S(A_{1},A_{2})$ consists of four edges, and $(A_{1},A_{2})$ has a side $A_{i}$ with $G[A_{i}]=K_{4}$.
-Proof.
-Let $M$ be the union of the $\operatorname{Aut}(G)$-orbits of $(A_{1},A_{2})$ and $(A_{2},A_{1})$.
-Let $(C,D)$ be $\leqslant$-minimal in $M$; that is, $(C^{\prime},D^{\prime})\not<(C,D)$ for every $(C^{\prime},D^{\prime})\in M$.
-We claim that $S(C,D)$ contains no vertex.
-Assume otherwise that $u\in S(C,D)$ is a vertex.
-Then let $v$ be an arbitrary vertex in $C\smallsetminus D$ and let $\varphi\in\operatorname{Aut}(G)$ such that $\varphi(u)=v$.
-Then $(\varphi(C),\varphi(D))<(C,D)$ or $(\varphi(D),\varphi(C))<(C,D)$, contradicting the $\leqslant$-minimality of $(C,D)$.
-Hence $S(C,D)$ consists of four edges.
-We claim that every vertex in $C$ is incident to an edge in $S(C,D)$.
-Assume for a contradiction that $v\in C$ is not.
-Let $u\in C\smallsetminus D$ be an endvertex of one of the four edges in $S(C,D)$ and let $\varphi\in\operatorname{Aut}(G)$ such that $\varphi(u)=v$.
-Then $(\varphi(C),\varphi(D))<(C,D)$ or $(\varphi(D),\varphi(C))<(C,D)$, contradicting the $\leqslant$-minimality of $(C,D)$.
-Finally, since $G$ is 4-connected, every vertex has degree $\geqslant 4$, so $G[C]$ is a $K_{4}$. The tetra-separation $(A_{1},A_{2})$ shares the above properties of $(C,D)$ since $(C,D)\in M$. ∎
-Corollary 13.6.
-Let $G$ be a vertex-transitive 4-connected finite graph such that every tetra-separation is totally-nested. Then $G$ is essentially 5-connected.∎
-Lemma 13.7.
-Let $G$ be a vertex-transitive 4-connected finite graph that has two tetra-separations $(A,B),(C,D)$ that cross with all links empty. Then $G=K_{4,4}$, which is a cycle of $K_{2,2}$-bags of length 4.
-Proof.
-Let $Z$ be the centre of the crossing-diagram, which consists of four vertices by the (Crossing Lemma). (4.1). Note that GZG-Z has at least four components.
-[…]
-By 4-connectivity of $G$, each of the $\geqslant 4$ vertices of GZG-Z has neighbourhood equal to $Z$.
-So, $G$ is 4-regular by vertex-transitivity.
-In particular, GZG-Z has exactly four vertices, and $G[Z]$ is independent.
-So, $G=K_{4,4}$.
-∎
-Lemma 13.8.
-Let $G$ be a vertex-transitive 4-connected finite graph that has two tetra-separations $(A,B),(C,D)$ that cross with all links of size one.
-Let $\mathcal{O}$ be the block-bagel induced by $(A,B)$ and $(C,D)$.
-Then every bag of $\mathcal{O}$ is bad.
-Proof.
-Assume for a contradiction that $\mathcal{O}=:(O,\mathcal{G})$ has a good bag $G_{t}$.
-Then we have the (Block-bagel bag setting).
-(8.14) with $t=t(0)$, which we assume in the following.
-By ˜8.27, $(U,W)$ is a totally-nested tetra-separation of $G$.
-By Lemma˜13.5, the separator $S(U,W)$ consists of four edges, and either $G[U]=K_{4}$ or $G[W]=K_{4}$.
-It follows with vertex-transitivity that $G$ is 4-regular and that the neighbourhood around each vertex induces $K_{1}\sqcup K_{3}$ in $G$.
-By the (Helpful Lemma).
-(8.6), $O$ has at least four helpful edges or vertices.
-By 4-regularity, there are exactly four helpful edges or nodes of $O$.
-For each helpful edge, we colour its adhesion-vertex red.
-For each helpful node $t$, there exists a unique interior vertex of $G_{t}$ that is a neighbour of $z^{\prime}_{1}$ (by 4-regularity), and we colour this vertex red.
-Then exactly four vertices have been coloured red.
-By the above, the red vertices induced a $K_{1}\sqcup K_{3}$ in $G-Z^{\prime}$, where $Z^{\prime}$ is the centre of the crossing-diagram.
-None of the red vertices in the $K_{3}$ can be an interior vertex of a bag, so all three are adhesion-vertices.
-But then $O$ would have to be a triangle, contradicting that $O$ has four helpful edges or nodes.
-∎
-Lemma 13.9.
-Let $G$ be a vertex-transitive 4-connected finite graph that has two tetra-separations $(A,B),(C,D)$ that cross with all links of size one.
-Let $\mathcal{O}$ be the Block-bagel induced by $(A,B)$ and $(C,D)$.
-Then every bag of $\mathcal{O}$ is a $K_{2}$.
-Proof.
-Write $\mathcal{O}=:(O,(G_{t})_{t\in V(O)})$, and let $u,v$ be the two vertices in the centre.
-By Lemma˜13.8, every bag of $\mathcal{O}$ is bad.
-In particular, every bag of $\mathcal{O}$ is a $K_{2}$ or a triangle.
-Assume for a contradiction that $\mathcal{O}$ has a bag that is a triangle.
-Since every bag of $\mathcal{O}$ is bad, it follows that all bags of $\mathcal{O}$ are triangles.
-Let $G_{t}$ be an arbitrary bag of $\mathcal{O}$, and let $x$ be the vertex in $G_{t}$ that is not an adhesion-vertex.
-By 4-connectivity of $G$, we have $xu,xv\in E(G)$ and $x$ has degree 4.
-Hence $G$ is 4-regular, and since $G_{t}$ was arbitrary, we get $|O|\leqslant 4$.
-As every node of $O$ is helpful, we get $|O|\geqslant 4$ from the (Helpful Lemma).
-(8.6), so $O$ has length 4.
-Hence each of $u,v$ has four neighbours, one in the interior of each bag, so by 4-regularity the edge $uv$ cannot exist in $G$.
-Therefore, $u$ is not contained in a triangle.
-But all bags are triangles, contradicting vertex-transitivity.
-∎
-Lemma 13.10.
-Let $G$ be a vertex-transitive 4-connected finite graph that has two tetra-separations $(A,B),(C,D)$ that cross with all links of size one.
-Then $G$ is an octahedron, which is a cycle of $K_{2,2}$-bags of length 3.
-Proof.
-By Lemma˜13.9, every bag of the block-bagel $\mathcal{O}=(O,\mathcal{G})$ induced by $(A,B)$ and $(C,D)$ is a $K_{2}$.
-Hence every edge of $O$ is helpful, so $|O|\geqslant 4$ by the (Helpful Lemma).
-(8.6).
-Since each vertex of $G$ contained in $\bigcup_{t\in V(O)}G_{t}$ has degree $\leqslant 4$ in $G$, the graph $G$ is 4-regular.
-Let $u,v$ be the vertices in the centre.
-By Lemma˜8.7 it follows that $|O|=4$ and $uv\notin E(G)$.
-Then $G$ is an octahedron.
-∎
-Lemma 13.11.
-Let $G$ be a vertex-transitive 4-connected finite graph that has two tetra-separations $(A,B),(C,D)$ that cross with all links of size two.
-Let $\mathcal{O}$ be the Tutte-bagel induced by $(A,B)$ and $(C,D)$.
-Let $H_{t}$ be a good torso of $\mathcal{O}$.
-Then the bag $G_{t}$ is a $K_{4}$ and its two neighbouring torsos are 4-cycles.
-Proof.
-We may assume the (Tutte-bagel torso setting).
-(9.13) with $t=t(0)$.
-By ˜9.41(U,W)(U,W) is a totally-nested tetra-separation.
-By Lemma˜13.5, $S(U,W)$ consists of four edges, and we have either $G[U]=K_{4}$ or $G[W]=K_{4}$.
-We claim that $G[U]=K_{4}$.
-Assume otherwise, so $G[W]=K_{4}$.
-Then $G[W]=G_{s}$ for some bag $G_{s}\neq G_{t}$.
-Hence $O$ has length $\leqslant 4$, two 3-connected bags (namely $G_{s},G_{t}$), and all other bags have triangles or 4-cycles as torsos by Lemma˜9.14.
-But then $\alpha(\mathcal{O})<4$ contradicts Lemma˜9.51.
-Hence $G[U]=K_{4}$.
-Next, we claim that $S(U,W)$ contains no inward edge, and assume for a contradiction that $x_{1}\in S(U,W)$ is an inward edge with endvertex $u$ in the interior of $G_{t}$.
-By Lemma˜9.33, some $y^{\prime}_{i}$, say $y^{\prime}_{1}$, is a neighbour of $x^{\prime}_{1}$ and $y^{W}_{1}=y^{\prime}_{1}$.
-In particular, $y_{1}$ cannot be an outward edge.
-Since $S(U,W)$ consists of four edges and $y_{1}$ is one of them, $y_{1}$ must be an inward edge.
-Let $u$ be the endvertex of $y_{1}$ in the interior of $G_{t}$.
-Then $u\neq v$ by the matching-condition for $(U,W)$.
-Let $w$ be a neighbour of $x^{\prime}_{1}$ in $G$ but outside of $G_{t}$.
-Then the vertex-set $\{u,v,w\}$ is independent in $G$ and consists entirely of neighbours of $x^{\prime}_{1}$.
-Since $G[U]=K_{4}$ and $S(U,W)$ consists of four edges, $G$ is 4-regular and every vertex lies in a $K_{4}$ by vertex-transitivity.
-But $\{u,v,w\}$ witnesses that $x^{\prime}_{1}$ does not lie in a $K_{4}$, a contradiction.
-Therefore, $S(U,W)$ contains no inward edge.
-So all edges in $S(U,W)$ are outward edges. Hence $G_{t}=K_{4}$ and both neighbouring torsos are 4-cycles. ∎
-Lemma 13.12.
-Let $G$ be a vertex-transitive 4-connected finite graph that has two tetra-separations $(A,B),(C,D)$ that cross with all links of size two.
-Let $\mathcal{O}$ be the Tutte-bagel induced by $(A,B)$ and $(C,D)$.
-If some torso of $\mathcal{O}$ is good, then $G$ is a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 8$.
-Proof.
-Write $\mathcal{O}=:(O,(G_{t})_{t\in V(O)})$.
-Write $O=t(0)\,t(1)\ldots t(m)\,t(0)$ with indices in $\mathbb{Z}_{m+1}$.
-By Lemma˜13.11, if a torso such as $H_{t(0)}$ is good, then its bag is a $K_{4}$ and the two neighbouring torsos $H_{t(-1)}$ and $H_{t(1)}$ are 4-cycles.
-Then, by Lemma˜9.4, the torsos $H_{t(-2)}$ and $H_{t(2)}$ are 3-connected.
-Hence, both of these torsos are good as they neighbour torsos that are 4-cycles.
-By induction, we get that $O$ has even length, that bags $G_{t(i)}$ are $K_{4}$’s for even $i$, and that torsos $H_{t(i)}$ are 4-cycles for odd $i$.
-By Lemma˜9.51, there are $\geqslant 4$ bags that are $K_{4}$’s.
-It follows that $G$ is a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 8$.
-∎
-Lemma 13.13.
-Let $\mathcal{O}$ be a Tutte-bagel of a 4-connected graph.
-Let $G_{t}$ be a bag of $\mathcal{O}$ on five vertices.
-Let $\pi,\pi^{\prime}$ be the two adhesion-sets at $G_{t}$, and let $v$ be the interior vertex of $G_{t}$.
-Then the bag $G_{t}$ contains a full matching between $\pi$ and $\pi^{\prime}$, and $v$ sends edges to all four vertices in $\pi\cup\pi^{\prime}$.
-Proof.
-Since $G$ is 4-connected, $v$ sends edges to all four vertices in $\pi\cup\pi^{\prime}$.
-Since the torso $H_{t}$ is 3-connected with $|H_{t}|=5$, it is either a $4$-wheel, a $K_{5}^{-}$ or a $K_{5}$.
-In each case, $H_{t}$ contains a full matching between $\pi$ and $\pi^{\prime}$, and both edges in the matching also exist in the bag $G_{t}$.
-∎
-Lemma 13.14.
-Let $G$ be a vertex-transitive 4-connected finite graph that has two tetra-separations $(A,B),(C,D)$ that cross with all links of size two.
-Let $\mathcal{O}$ be the Tutte-bagel induced by $(A,B)$ and $(C,D)$.
-Assume that $\mathcal{O}$ has a torso on five vertices.
-Then $G$ is the line graph of the cube.
-Proof.
-Write $\mathcal{O}=:(O,\mathcal{G})$.
-By Lemma˜13.12, all bags of $\mathcal{O}$ are bad.
-Let $H_{t}$ be a torso of $\mathcal{O}$ on five vertices.
-Since the unique interior vertex $x$ of $H_{t}$ has degree 4, the graph $G$ is 4-regular.
-Since $H_{t}$ is bad, it is neighboured by a 3-connected torso $H_{s}$.
-Since $H_{s}$ is bad as well, it has $\leqslant 5$ vertices.
-That is, the torso of $H_{s}$ either is a $K_{4}$ or has exactly five vertices.
-We claim that $H_{s}$ has exactly five vertices.
-Otherwise $H_{s}=K_{4}$.
-Let $\{u,v\}$ denote the adhesion-set in between the torsos $H_{t}$ and $H_{s}$.
-Then $uv\notin E(G)$ by 4-regularity of $G$.
-Let $\{u^{\prime},v^{\prime}\}$ denote the other adhesion-set at $H_{s}$.
-Since $v$ is contained in at least two triangles in $G$, we must have $u^{\prime}v^{\prime}\in E(G)$.
-The vertex $v$ is contained in exactly two triangles: one from $H_{s}$ and the other is $vu^{\prime}v^{\prime}$.
-The vertex $v^{\prime}$, however, is contained in at least two triangles (namely $vu^{\prime}v^{\prime}$ and $uu^{\prime}v^{\prime}$), two of which intersect in the edge $u^{\prime}v^{\prime}$.
-This contradicts the vertex-transitivity of $G$.
-Hence $H_{s}$ has exactly five vertices.
-It follows by induction that all torsos of $\mathcal{O}$ have exactly five vertices.
-By Lemma˜13.13, each bag contains a 2-matching between its two adhesion-sets, and the interior vertex of the bag sends edges to all four other vertices in the bag.
-By 4-regularity and since every bag looks like this, these are all the edges that there are in the bags.
-Since each bag has a unique interior vertex, the set $X$ of all these vertices has size $|O|$.
-Since all torsos are 3-connected, we have $|O|\geqslant 4$ by Lemma˜9.51.
-If $|O|\geqslant 5$, then every vertex in $X$ is contained in exactly two induced 4-cycles in $G$, while every vertex in an adhesion set is contained in exactly one induced 4-cycle in $G$, contradicting vertex-transitivity.
-Therefore, we have $|O|=4$.
-Consider the graph H:=GXH:=G-X, which is either an 8-cycle or a disjoint union of two 4-cycles.
-If $H$ is an 8-cycle, then every vertex in $X$ is contained in two induced 4-cycles of $G$ while every vertex in V(GX)V(G-X) is only contained in one induced 4-cycle of $G$, contradicting vertex-transitivity.
-Therefore $H$ is a disjoint union of two 4-cycles, and it follows that $G$ is the line graph of the cube.
-∎
-Lemma 13.15.
-Let $G$ be a vertex-transitive 4-connected finite graph that has two tetra-separations $(A,B),(C,D)$ that cross with all links of size two.
-Let $\mathcal{O}$ be the Tutte-bagel induced by $(A,B)$ and $(C,D)$.
-Assume that every bag of $\mathcal{O}$ has $\leqslant 4$ vertices.
-Then $G$ is one of the following graphs:
-  • $\cdot$
-    a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 8$;
-  • $\cdot$
-    a cycle of triangle-bags of length $\geqslant 6$;
-  • $\cdot$
-    a cycle of $K_{2,2}$-bags of length $\geqslant 4$;
-  • $\cdot$
-    a cycle of $K_{4}$-bags of length $\geqslant 4$.
-Proof.
-If $\mathcal{O}$ has a good torso, then $G$ is a cycle alternating between $K_{4}$-bags and $C_{4}$-bags by Lemma˜13.12, and it is of length $\geqslant 8$ by Lemma˜9.51.
-Hence, we may assume that all torsos of $\mathcal{O}$ are bad.
-We claim that $\mathcal{O}$ has no 4-cycle as torso.
-Indeed, if $\mathcal{O}$ has a 4-cycle as a torso, then the neighbouring torsos are 3-connected by Lemma˜9.4, and they are good, a contradiction.
-Thus,
-(8) all torsos are either triangles or $K_{4}$’s.
-We claim that if some torso of $\mathcal{O}$ is a triangle, then also its neighbouring torsos are triangles.
-For this, let $H_{t}$ and $H_{s}$ be neighbouring torsos with $H_{t}$ a triangle.
-By ˜8, the torso $H_{s}$ is either a triangle or a $K_{4}$.
-Assume for a contradiction that $H_{s}=K_{4}$.
-Let $\{u,v\}$ be the adhesion-set in between $H_{t}$ and $H_{s}$, so that $v$ is the tip of the triangle-torso $H_{t}$, say.
-Since every vertex has degree $\geqslant 4$, the edge $uv$ must exist in $G$, and $G$ must be 4-regular.
-Let $w$ be the vertex of the triangle $H_{t}$ besides $u$ and $v$.
-Let $G_{r}$ be the bag neighbouring $G_{t}$ besides $G_{s}$.
-Then $v$ has exactly one neighbour in the bag $G_{r}$.
-By ˜8, the torso $H_{r}$ can only be a triangle or a $K_{4}$.
-As $v$ has exactly one neighbour in the bag $G_{r}$, the torso $H_{r}$ cannot be a $K_{4}$, so it must be a triangle.
-But then the edge $vw$ exists by Lemma˜9.4, so $v$ has two neighbours in $G_{r}$, giving it degree five in $G$, contradicting 4-regularity.
-Therefore, every triangle-torso of $\mathcal{O}$ is neighboured by two triangle-torsos.
-So if $\mathcal{O}$ has at least one triangle-torsos, then all torsos are triangles, and all adhesion-sets are spanned by edges by Lemma˜9.4.
-Hence $G$ is a cycle of triangle-bags, and it has length $\geqslant 6$ by Lemma˜9.51 and since $G\neq K_{5}$.
-So we may assume that $\mathcal{O}$ has no triangle-torsos.
-Then by ˜8, all torsos of $\mathcal{O}$ are $K_{4}$’s.
-Then either all adhesion-sets are spanned by edges in $G$, in which case $G$ is a cycle of $K_{4}$-bags; or none of them are, in which case $G$ is a cycle of $K_{2,2}$-bags.
-In both cases, the length is $\geqslant 4$ by Lemma˜9.51.
-∎
-Lemma 13.16.
-Let $G$ be a vertex-transitive 4-connected finite graph that has two tetra-separations $(A,B),(C,D)$ that cross with all links of size two. Then $G$ is one of the following:
-  • $\cdot$
-    a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 8$;
-  • $\cdot$
-    a cycle of triangle-bags of length $\geqslant 6$;
-  • $\cdot$
-    a cycle of $K_{2,2}$-bags of length $\geqslant 4$;
-  • $\cdot$
-    a cycle of $K_{4}$-bags of length $\geqslant 4$;
-  • $\cdot$
-    line graph of the cube.
-Proof.
-We combine Lemmas˜13.14 and 13.15. ∎
-Corollary 13.17.
-Let $G$ be a vertex-transitive 4-connected finite graph that has two tetra-separations $(A,B),(C,D)$ that cross. Then $G$ is one of the following:
-  • $\cdot$
-    a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 8$;
-  • $\cdot$
-    a cycle of triangle-bags of length $\geqslant 6$;
-  • $\cdot$
-    a cycle of $K_{2,2}$-bags of length $\geqslant 3$;
-  • $\cdot$
-    a cycle of $K_{4}$-bags of length $\geqslant 4$;
-  • $\cdot$
-    line graph of the cube.
-Proof.
-By the (Crossing Lemma).
-(4.1), all links have the same size 0, 11 or $2$.
-If all links have size 0, then $G$ is a cycle of $K_{2,2}$-bags of length $4$ by Lemma˜13.7.
-If all links have size $1$, then $G$ is a cycle of $K_{2,2}$-bags of length $3$ by Lemma˜13.10.
-Otherwise all links have size $2$, and we apply Lemma˜13.16.
-∎
-Observation 13.18.
-Let $G^{\prime}$ be the clique-expansion of a graph $G$.
-Let $v$ be a vertex of degree $d$ in $G$, and let $C_{v}$ be the clique of $G^{\prime}$ that corresponds to $v$.
-Then every vertex in $C_{v}$ has degree $d$ in $G^{\prime}$.∎
-Observation 13.19.
-Let $G$ be a graph of minimum degree at least $2$, and let $G^{\prime}$ be the clique-expansion of $G$.
-Then there is a unique partition $\mathcal{P}$ of $V(G^{\prime})$ into (the vertex-sets of) $|\mathcal{P}|=|V(G)|$ cliques.
-Moreover, the contraction-minor $G^{\prime}/\mathcal{P}$ is isomorphic to GG.∎
-Observation 13.20.
-The clique-expansion of a graph $G$ is isomorphic to the graph $H$ with
-$V(H)$ ${}=\{(u,v),(v,u):uv\in E(G)\}\text{ and}$
-$E(H)$ ${}=\Big{\{}\{(u,v),(v,u)\}:uv\in E(G)\Big{\}}\,\cup\,\Big{\{}\{(u,v_{1}),(u,v_{2})\}:uv_{1},uv_{2}\in E(G)\Big{\}}.$
-Lemma 13.21.
-Let $G$ be an $r$-regular graph with $r\geqslant 3$. Then the following assertions are equivalent:
-  1. (i)
-    the $K_{r}$-expansion of $G$ is vertex-transitive;
-  2. (ii)
-    $G$ is arc-transitive.
-Proof.
-[…]
-˜(i) $\Rightarrow$ ˜(ii).
-Let $(u_{1},v_{1}),(u_{2},v_{2})$ be ordered pairs of vertices in $G$ with $u_{i}v_{i}\in E(G)$.
-We will find an automorphism $\varphi$ of $G$ such that $\varphi(u_{1})=u_{2}$ and $\varphi(v_{1})=v_{2}$.
-For this, we abbreviate $x_{i}:=(u_{i},v_{i})$.
-Let $P_{i}\in\mathcal{P}$ be the unique partition-class of $\mathcal{P}$ that contains $x_{i}$.
-Since $G^{\prime}$ is vertex-transitive by assumption, there is an automorphism $\varphi^{\prime}$ of $G^{\prime}$ that maps $x_{1}$ to $x_{2}$.
-Note that $\varphi$ then maps $P_{1}$ to $P_{2}$.
-We define $\varphi:=\gamma\circ\varphi^{\prime}\circ\gamma^{-1}$ by abuse of notation, using that $\operatorname{Aut}(G^{\prime})$ acts on $\mathcal{P}$.
-Clearly, $\varphi$ is a bijection $V(G)\to V(G)$.
-Moreover, we have
-$\text{ }xy\in E(G)$
-$\Leftrightarrow$ there is a $\gamma^{-1}(x)$$\gamma^{-1}(y)$ edge in $G^{\prime}$
-$\Leftrightarrow$ there is a $\varphi^{\prime}(\gamma^{-1}(x))$$\varphi^{\prime}(\gamma^{-1}(y))$ edge in $G^{\prime}$
-$\Leftrightarrow$ $\varphi(x)\varphi(y)\in E(G)$.
-So, $\varphi$ is an automorphism of $G$. Moreover, we have
-$\varphi(u_{1})=\gamma(\varphi^{\prime}(\gamma^{-1}(u_{1})))=\gamma(\varphi^{\prime}(P_{1}))=\gamma(P_{2})=u_{2}.$
-Similarly, $\varphi(v_{1})=v_{2}$.
-˜(ii) $\Rightarrow$ ˜(i).
-We assume without loss of generality that $G^{\prime}$ takes the form described in ˜13.20.
-Let $x_{1},x_{2}\in V(G^{\prime})$ be two vertices.
-We will find an automorphism $\varphi^{\prime}$ of $G^{\prime}$ that maps $x_{1}$ to $x_{2}$.
-We write $x_{i}=:(u_{i},v_{i})$ using ˜13.20, so $u_{i}v_{i}$ is an edge of $G$.
-Since $G$ is arc-transitive, there is an automorphism $\varphi$ of $G$ such that $\varphi(u_{1})=u_{2}$ and $\varphi(v_{1})=v_{2}$.
-Now $\varphi$ defines an automorphism $\varphi^{\prime}$ of $G^{\prime}$ that sends every vertex $x=(u,v)$ to the vertex $\varphi^{\prime}(x):=(\varphi(u),\varphi(v))$.
-The automorphism $\varphi^{\prime}$ satisfies
-$\varphi^{\prime}(x_{1})=(\varphi(u_{1}),\varphi(v_{1}))=(u_{2},v_{2})=x_{2}.\qed$
-Corollary 13.22.
-Let $G$ be an $r$-regular graph with $r\geqslant 3$.
-Let $(A,B)$ be a mixed-$r$-separation of $G$ such that $S(A,B)$ consists only of edges and $S(A,B)$ is a matching and $G[A]=K_{r}$.
-Then $G$ is not arc-transitive.
-Moreover, the $K_{r}$-expansion of $G$ is not vertex-transitive.
-Proof.
-By Lemma˜13.21, it suffices to show that $G$ is not arc-transitive.
-Indeed, if $u$ is a vertex in $A$, then $G[N(u)]=K_{1}\sqcup K_{r-1}$ with $r-1\geqslant 2$.
-Let $v_{1}$ be the isolated vertex in $G[N(u)]$ and let $v_{2}\neq v_{1}$ be another vertex in $N(u)$.
-Then there is no automorphism of $G$ that maps $(u,v_{1})$ to $(u,v_{2})$.
-∎
-Corollary 13.23.
-Let $r\geqslant 3$.
-The $K_{r}$-expansion of an $r$-regular graph $G$ is not arc-transitive.
-In particular, the $K_{r}$-expansion of the $K_{r}$-expansion of an $r$-regular graph $G$ is not vertex-transitive.
-Proof.
-Follows from Corollary˜13.22. ∎
-Corollary 13.24.
-A cycle of triangle-bags of length $\geqslant 6$ is not arc-transitive. In particular, the $K_{4}$-expansion of a cycle of triangle-bags of length $\geqslant 6$ is not vertex-transitive.
-Proof.
-Let $G$ be a cycle of triangle-bags of length $\geqslant 6$, and let $\mathcal{O}$ be the corresponding cycle-decomposition.
-Note that every free edge of $\mathcal{O}$ is contained in exactly one triangle of $G$, while every other edge of $G$ is contained in exactly two triangles.
-So, $G$ is not edge-transitive.
-In particular, it is not arc-transitive.
-By Lemma˜13.21, the $K_{4}$-expansion of a cycle of triangle-bags of length $\geqslant 6$ is not vertex-transitive.
-∎
-Corollary 13.25.
-Let $G$ be a finite graph that is arc-transitive, 4-connected, 4-regular and has two tetra-separations that cross.
-Then $G$ is the line graph of the cube or a cycle of $K_{2,2}$-bags of length $\geqslant 3$.
-Proof.
-This follows from Corollary˜13.17 together with the following observations.
-First, a cycle of $K_{4}$-bags of length $\geqslant 4$ is not 4-regular.
-Second, a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 8$ is not arc-transitive by Corollary˜13.22.
-Finally, a cycle of triangle-bags of length $\geqslant 6$ is not arc-transitive by Corollary˜13.24.
-∎
-Lemma 13.26.
-Let $G$ be a vertex-transitive essentially 5-connected finite graph.
-Then either $G$ is quasi-5-connected or there is a partition $\mathcal{P}$ of $V(G)$ into vertex-sets of $K_{4}$’s such that the edges in $G$ with ends in different classes of $\mathcal{P}$ form a perfect matching in $G$.
-Proof.
-If $G$ does not have a tetra-separation, then $G$ is quasi-5-connected by Lemma˜2.3 (i).
-Hence, we may assume that $G$ has a tetra-separation $(A,B)$.
-Since $G$ is essentially 5-connected, the separator of $(A,B)$ consists of four edges and there is a side, say $A$, that induces a $K_{4}$ in $G$.
-Let $u\in A$ and observe that $G[N(u)]=K_{1}\sqcup K_{3}$; in particular, $u$ is contained in exactly one $K_{4}$.
-By vertex-transitivity, so is every vertex of $G$.
-It follows that $G$ admits a partition of its vertex-set into the vertex-sets of $K_{4}$’s.
-Moreover, $G$ is 4-regular since $u$ has degree 4 in $G$.
-Then it follows that the edges between distinct partition-classes of $\mathcal{P}$ form a perfect matching in $G$.
-∎
-Lemma 13.27.
-Let $G$ be a vertex-transitive essentially 5-connected finite graph. Then $G$ is one of the following:
-  • $\cdot$
-    quasi-5-connected;
-  • $\cdot$
-    a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\ell\in\{4,6\}$;
-  • $\cdot$
-    the $K_{4}$-expansion of an arc-transitive 4-connected 4-regular graph.
-Proof.
-Assume that $G$ is not quasi-5-connected.
-Then by Lemma˜13.26, there is a partition $\mathcal{P}$ of $V(G)$ into the vertex-sets of $K_{4}$’s such that edges between distinct partition-classes of $\mathcal{P}$ form a perfect matching in $G$.
-We distinguish the following cases.
-Case 1: there are two parts $P_{1}\neq P_{2}$ of $\mathcal{P}$ such that there are at least three $P_{1}$$P_{2}$ edges $e_{1},e_{2},e_{3}$ in $G$.
-Let $f_{i}\notin\{e_{1},e_{2},e_{3}\}$ be the remaining edge incident with a vertex in $P_{i}$.
-By 4-connectivity of $G$, $\{f_{1},f_{2}\}$ is not a mixed-2-separator.
-It follows that $f_{1}=f_{2}$, so $G$ is a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length 4.
-Case 2: there are two parts $P_{1}\neq P_{2}$ of $\mathcal{P}$ such that there are exactly two $P_{1}$$P_{2}$ edges in $G$.
-Let $e_{1},e_{2}$ be the two edges incident with $P_{1}$ but not with $P_{2}$, and let $f_{1},f_{2}$ be the two edges incident with $P_{2}$ but not with $P_{1}$.
-We define $(A,B):=(P_{1}\cup P_{2},V(G)\smallsetminus(P_{1}\cup P_{2}))$.
-Then $S(A,B)=\{e_{1},e_{2},f_{1},f_{2}\}$, so $(A,B)$ is a tetra-separation.
-By assumption, there is a side of $(A,B)$ that induces a $K_{4}$.
-Since $|A|=8$, we must have $G[B]=K_{4}$.
-It follows that $G$ is a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $6$.
-Case 3: for every two parts $P_{1}\neq P_{2}$ of $\mathcal{P}$, there is at most one $P_{1}$$P_{2}$ edge in $G$.
-Then we define $H:=G/\mathcal{P}$, which is a 4-regular graph.
-Moreover, observe that $G$ is the $K_{4}$-expansion of $H$.
-Then by Lemma˜13.21HH is arc-transitive.
-By Corollary˜13.2HH is 4-connected.
-∎
-Corollary 13.28.
-Let $G$ be an arc-transitive essentially 5-connected graph. Then $G$ is quasi-5-connected.
-Proof.
-Follows from Lemma˜13.27 together with the following observations.
-First, a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\ell\in\{4,6\}$ is not arc-transitive by Corollary˜13.22.
-Second, the $K_{4}$-expansion of a graph is not arc-transitive by Corollary˜13.23.
-∎
-Proposition 13.29.
-A 4-connected finite graph $G$ is vertex-transitive if and only if $G$ either is
-  1. (i)
-    a quasi-5-connected vertex-transitive graph,
-  2. (ii)
-    the $K_{4}$-expansion of a quasi-5-connected arc-transitive graph,
-  3. (iii)
-    a cycle of $K_{4}$-bags of length $\geqslant 4$ (Figure˜1.9),
-  4. (iv)
-    a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 4$ (Figure˜1.10),
-  5. (v)
-    a cycle of triangle-bags of length $\geqslant 6$ (Figure˜1.12),
-or $G$ is one of the following graphs or a $K_{4}$-expansion thereof:
-  1. (vi)
-    a cycle of $K_{2,2}$-bags of length $\geqslant 3$ (Figure˜1.14),
-  2. (vii)
-    the line-graph of the cube (Figure˜1.15).
-Proof.
-Let $G$ be a 4-connected finite graph.
-($\Rightarrow$).
-Assume that $G$ is vertex-transitive.
-If $G$ has two tetra-separations that cross, then we are done by Corollary˜13.17.
-Otherwise, every tetra-separation of $G$ is totally-nested.
-Then $G$ is essentially 5-connected by Corollary˜13.6.
-By Lemma˜13.27GG is either quasi-5-connected, or a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\ell\in\{4,6\}$, or the $K_{4}$-expansion of an arc-transitive 4-connected 4-regular graph.
-In the first two cases, we are done.
-So assume that $G$ is the $K_{4}$-expansion of an arc-transitive 4-connected 4-regular graph $H$.
-In particular, $H$ is vertex-transitive.
-If $H$ has two tetra-separations that cross, then we apply Corollary˜13.25 to get that $H$ is the line graph of the cube or that $H$ is a cycle of $K_{2,2}$-bags of length $\geqslant 3$.
-Here we are done.
-Otherwise, every tetra-separation of $H$ is totally-nested.
-By Corollary˜13.6HH is essentially 5-connected.
-By Corollary˜13.28HH quasi 5-connected.
-Thus, $G$ is the $K_{4}$-expansion of an arc-transitive quasi-5-connected graph, and we are done.
-($\Leftarrow$).
-Clearly, if $G$ is one of the graphs in (i), (iii), (iv), (v), then $G$ is vertex-transitive.
-If (ii) $G$ is the $K_{4}$-expansion of a quasi-5-connected arc-transitive graph, then $G$ is vertex-transitive by Lemma˜13.21.
-Both (vi) a cycle of $K_{2,2}$-bags and (vii) the line graph of the cube are arc-transitive.
-If $G$ is a $K_{4}$-expansion of (vi) or (vii), then $G$ is vertex-transitive by Lemma˜13.21.
-∎
-The line graph $L=L(G)$ of a graph $G$ is the graph on the vertex set $V(L):=E(G)$ such that two vertices in $L$ are adjacent if and only if their corresponding edges in $G$ share an endvertex.
-A graph $G$ is internally-4-connected if it is 3-connected and every 3-separation has a side that induces a claw. Note that internally-4-connected graphs are quasi-4-connected.
-Key Lemma 13.30.
-Let $G$ be a 3-regular quasi-4-connected graph. Then either $G$ is internally-4-connected and has girth $\geqslant 4$, or $G=K_{3}\square K_{2}$.
-Proof.
-Suppose that $G$ is not internally-4-connected.
-It suffices to show that either $G$ is a cycle of $C_{4}$-bags of length 3 or $G$ has girth $\geqslant 4$.
-Note that $G$ has $\geqslant 5$ vertices as $G\neq K_{4}$.
-There is a vertex $v$ in $G$ of degree 3 that has two adjacent neighbours $a$ and $b$.
-Let $c$ be the third neighbour of $v$.
-First, we show that $ac,bc\notin E(G)$.
-Assume $ac\in E(G)$.
-Then the pair $(\{v,a,b,c\},V(G)\smallsetminus\{v,a\})$ is a 2-separation with separator $\{b,c\}$ since $|V(G)|\geqslant 5$, contradicting that $G$ is 3-connected.
-It follows that $(A,B):=(\{v,a,b\},V(G)\smallsetminus\{v,a,b\})$ is a 3-edge-cut.
-Since $G$ is 3-connected and since the set $N(\{a,b\})$ separates $\{a,b\}$ from $c$, we must have $|N(\{a,b\})|\geqslant 3$.
-In particular, no two edges in $S(A,B)$ share an endvertex.
-We define $a^{\prime},b^{\prime}$ and $v^{\prime}$ such that $S(A,B)=\{aa^{\prime},bb^{\prime},vv^{\prime}\}$.
-Note that $N(\{a,b\})=\{a^{\prime},b^{\prime},v\}$ separates $\{a,b\}$ from $V(G)\smallsetminus\{v,a,b,a^{\prime},b^{\prime}\}$.
-By quasi-4-connectivity of $G$, it follows that $V(G)=\{v,a,b,v^{\prime},a^{\prime},b^{\prime}\}$.
-And by 3-regularity of $G$, it follows that $v^{\prime}a^{\prime}b^{\prime}$ is a triangle.
-So, $G=K_{3}\square K_{2}$.
-∎
-Recall that $E(v)$ denotes the set of all edges incident with a vertex $v$ in a graph $G$.
-Lemma 13.31.
-Let $G$ be a connected graph and let $L=L(G)$ be the line graph of $G$. If $(A,B)$ is an edge-cut in $G$ with $E(G[A])\neq\emptyset\neq E(G[B])$, then
-$(A^{\prime},B^{\prime}):=(E(G[A])\cup S(A,B)\>,\>E(G[B])\cup S(A,B))$
-is a genuine separation of $L$. If $(A,B)$ has order $k$, then $(A^{\prime},B^{\prime})$ has order $k$ as well.
-Proof.
-Since $E(G)=E(G[A])\,\dot{\cup}\,E(G[B])\,\dot{\cup}\,S(A,B)$, we get $A^{\prime}\cup B^{\prime}=V(L)$ and $A^{\prime}\cap B^{\prime}=S(A,B)$.
-Note that for every vertex $v\in A$ we have $E(v)\subseteq A^{\prime}$.
-Similarly, for every vertex $v\in B$ we have $E(v)\subseteq B^{\prime}$.
-It follows that there is no $(A^{\prime}\smallsetminus B^{\prime})$$(B^{\prime}\smallsetminus A^{\prime})$ edge in $L$.
-From $E(G[A])\neq\emptyset\neq E(G[B])$ we get that $(A^{\prime},B^{\prime})$ is proper.
-Hence, $(A^{\prime},B^{\prime})$ is a separation of $L$ that has the same order as $(A,B)$.
-Lemma 13.32.
-Let $G$ be a connected graph and let $L=L(G)$ be the line graph of $G$. If $(A^{\prime},B^{\prime})$ is a genuine separation of $L$, then
-$(A,B):=(\{v\in V(G)\mid E(v)\subseteq A^{\prime}\}\>,\>\{v\in V(G)\mid E(v)\subseteq B^{\prime}\})$
-is a mixed-separation of $G$ with $A\cap B=\{v\in V(G)\mid E(v)\subseteq A^{\prime}\cap B^{\prime}\}$.
-Proof.
-Note that for all $v\in V(G)$, the set $E(v)\subseteq V(L)$ induces a clique in $L$.
-Hence, $E(v)\subseteq A^{\prime}$ or $E(v)\subseteq B^{\prime}$ for all $v\in V(G)$.
-From the definition of $(A,B)$, it follows that $A\cup B=V(G)$.
-Since $(A^{\prime},B^{\prime})$ is proper, it follows that there exist vertices $v_{A},v_{B}\in V(G)$ with $E(v_{A})\subseteq A^{\prime}\nsupseteq E(v_{B})$ and $E(v_{B})\subseteq B^{\prime}\nsupseteq E(v_{A})$; that is, $v_{A}\in A\smallsetminus B$ and $v_{B}\in B\smallsetminus A$.
-Hence, $(A,B)$ is a mixed-separation.
-Moreover, a vertex $v$ is contained in $A\cap B$ if and only if $E(v)\subseteq A^{\prime}\cap B^{\prime}$.
-∎
-Lemma 13.33.
-Let $G$ be a 3-regular internally-4-connected graph and let $L=L(G)$ be the line graph of $G$. Then all of the following assertions hold:
-  1. (i)
-    $L$ is 4-regular and every vertex $v\in V(L)$ is contained in exactly two triangles $\Delta_{1}$ and $\Delta_{2}$ in $L$. Moreover, $V(\Delta_{1}\cap\Delta_{2})=\{v\}$.
-  2. (ii)
-    $L$ is 4-connected.
-  3. (iii)
-    For every triangle $\Delta\subseteq L$ there is a vertex $v$ in $G$ such that $E(v)=V(\Delta)$.
-  4. (iv)
-    For every edge $e\in E(L)$ there is a unique triangle $\Delta(e)$ in $L$ that contains $e$.
-Proof.
-(i) follows from the definition of the line graph and from $G$ having girth $\geqslant 4$.
-(ii). Suppose for a contradiction that there is a genuine $(\leqslant 3)$-separation $(A^{\prime},B^{\prime})$ of $L$. Let
-$(A,B):=(\{v\in V(G)\mid E(v)\subseteq A^{\prime}\}\>,\>\{v\in V(G)\mid E(v)\subseteq B^{\prime}\})\leavevmode\nobreak\ ,$
-which is a mixed-separation of $G$ by Lemma˜13.32.
-First, suppose $A\cap B\neq\emptyset$.
-Then $A^{\prime}\cap B^{\prime}$ induces a $K_{3}$ in $L$.
-By (i), every vertex $v$ of $L$ is contained in exactly two triangles, which only intersect in $v$.
-It follows that $(A,B)$ is a 1-separation of $G$, a contradiction.
-Hence, we may assume that $(A,B)$ is an edge-cut.
-Let $e=uv\in S(A,B)$, say with $u\in A$ and $v\in B$.
-Then $E(u)\subseteq A^{\prime}$ and $E(v)\subseteq B^{\prime}$ by definition of $(A,B)$.
-Let $\Delta_{u}$ and $\Delta_{v}$ be the triangles in $L$ induced by $E(u)$ and $E(v)$, respectively.
-From $V(\Delta_{u})\subseteq A^{\prime}$ and $V(\Delta_{v})\subseteq B^{\prime}$ and $V(\Delta_{u}\cap\Delta_{v})=\{e\}$ (by (i)) it follows that $e\in S(A^{\prime},B^{\prime})$.
-Hence, $S(A,B)\subseteq S(A^{\prime},B^{\prime})$; in particular $(A,B)$ has order $\leqslant 3$.
-Since $G$ is internally-4-connected, $(A,B)$ is an atomic edge-cut with $S(A,B)=S(A^{\prime},B^{\prime})$.
-Now $(A^{\prime},B^{\prime})$ is not proper, a contradiction.
-(iii).
-From $G$ being 3-regular and internally-4-connected, it follows that $G$ has girth $\geqslant 4$.
-If $\Delta$ is a triangle in $L$ such that there is no vertex $v$ in $G$ with $V(\Delta)=E(v)$, then $V(\Delta)$ is the set of edges of a triangle in $G$, contradicting that $G$ has girth $\geqslant 4$.
-(iv).
-Every edge $e_{1}e_{2}\in E(L)$ is contained in at least one triangle: consider the endvertex $v\in G$ shared by $e_{1}$ and $e_{2}$, then $E(v)$ induces a triangle in $L$ that contains the edge $e_{1}e_{2}$.
-Assume for a contradiction that there is an edge $e_{1}e_{2}\in E(L)$ that is contained in at least two triangles $\Delta_{1}$ and $\Delta_{2}$ of $L$.
-By (iii), there are two vertices $v_{1},v_{2}\in V(G)$ such that $E(v_{i})=V(\Delta_{i})$ for $i\in\{1,2\}$.
-Then $e_{1},e_{2}\in E(v_{1})\cap E(v_{2})$, contradicting that $G$ is simple.
-∎
-Lemma 13.34.
-Let $G$ be a finite, 3-regular, internally-4-connected, vertex-transitive graph, and let $L=L(G)$ be its line graph.
-Let $(A^{\prime},B^{\prime})$ be a mixed-4-separation of $L$ with $|A^{\prime}\smallsetminus B^{\prime}|\geqslant 2$ and $|B^{\prime}\smallsetminus A^{\prime}|\geqslant 2$.
-Then $S(A^{\prime},B^{\prime})\subseteq V(L)$.
-In particular, all tetra-separations of $L$ are genuine 4-separations.
-Proof.
-First, observe that we can assume $G\neq K_{4}$.
-By Lemma˜13.33LL is 4-connected.
-Let $(A^{\prime},B^{\prime})$ be a a mixed-4-separation of $L$ with $|A^{\prime}\smallsetminus B^{\prime}|\geqslant 2$ and $|B^{\prime}\smallsetminus A^{\prime}|\geqslant 2$.
-The left-right-reduction $(A^{\prime\prime},B^{\prime\prime})$ of $(A^{\prime},B^{\prime})$ is a tetra-separation by ˜5.10.
-Moreover, it satisfies $A^{\prime\prime}\smallsetminus B^{\prime\prime}\supseteq A^{\prime}\smallsetminus B^{\prime}$ and $B^{\prime\prime}\smallsetminus A^{\prime\prime}\supseteq B^{\prime}\smallsetminus A^{\prime}$ by ˜5.9, and every edge in $S(A^{\prime},B^{\prime})$ is in $S(A^{\prime\prime},B^{\prime\prime})$.
-Hence, it suffices to show that $S(A^{\prime\prime},B^{\prime\prime})\subseteq V(L)$.
-Assume for a contradiction that $e_{1}\in S(A^{\prime\prime},B^{\prime\prime})$ is an edge in $L$.
-By Lemma˜13.33 (iv), there is a unique triangle $\Delta_{1}$ in $L$ that contains $e_{1}$.
-Let $v$ be the third vertex in $\Delta_{1}$ besides the endvertices of $e_{1}$.
-Since $\Delta_{1}$ is a triangle with the edge $e_{1}$ contained in $S(A^{\prime\prime},B^{\prime\prime})$, we get that $v$ is contained in the separator $S(A^{\prime\prime},B^{\prime\prime})$ as well.
-By Lemma˜13.33 (i), there is a triangle $\Delta_{2}\neq\Delta_{1}$ in $L$ such that $\Delta_{1}$ and $\Delta_{2}$ meet exactly in the vertex $v$.
-Let $e_{2}$ denote the edge in $\Delta_{2}$ whose endvertices are distinct from $v$.
-Since $L$ is 4-regular, the degree-condition for $v\in S(A^{\prime\prime},B^{\prime\prime})$ yields $e_{2}\in S(A^{\prime\prime},B^{\prime\prime})$.
-So we know three elements of $S(A^{\prime\prime},B^{\prime\prime})$: they are $e_{1},v$ and $e_{2}$.
-The fourth element $u$ of $S(A^{\prime\prime},B^{\prime\prime})$ must be a vertex, as otherwise we could argue as for $e$ to find a third triangle contributing both an edge and a vertex to $S(A^{\prime\prime},B^{\prime\prime})$, contradicting that $\Delta_{1}$ and $\Delta_{2}$ already use all four edges incident to $v$.
-By Lemma˜13.33 (i), there are two triangles $\Delta,\Delta^{\prime}$ in $L$ that meet exactly in the vertex $u$.
-By 4-regularity of $L$ and the degree-condition for $u\in S(A^{\prime},B^{\prime})$, we have $\Delta-u\subseteq G[A^{\prime}\smallsetminus B^{\prime}]$ and $\Delta^{\prime}-u\subseteq G[B^{\prime}\smallsetminus A^{\prime}]$, say.
-Now consider the mixed-separation $(A,B)$ of $G$ that is induced by $(A^{\prime\prime},B^{\prime\prime})$ as in Lemma˜13.32.
-The separator of $(A,B)$ consists of two vertices $x_{1},x_{2}$ that correspond to the triangles $\Delta_{1},\Delta_{2}$, respectively, and of one edge $f$ that corresponds to the vertex $u$ (by Lemma˜13.33 (iii)).
-In particular, $(A,B)$ is a mixed-3-separation of $G$.
-Also note that $x_{1}x_{2}\in E(G)$ as the triangles $\Delta_{1},\Delta_{2}$ share the vertex $u$.
-Since $G\neq K_{4}$, there is a proper side of $(A,B)$, say $B\smallsetminus A$, that contains at least two vertices.
-Let $x_{3}$ be the endvertex of $f$ in $B\smallsetminus A$.
-Note that $(A\cup f,B)$ is a genuine 3-separation with separator $\{x_{1},x_{2},x_{3}\}$.
-Since $G$ is internally-4-connected, it follows that $x_{1}x_{2}\notin E(G)$, a contradiction.
-∎
-Let $G$ be a graph. A tetra-cut of $G$ is a 4-edge-cut of $G$ that satisfies the matching-condition. A tetra-cut of $G$ is totally-nested if it is nested with every tetra-cut of $G$.
-Observation 13.35.
-If $G$ is a 2-connected graph and $(A,B)$ is a non-atomic edge-cut of $G$ then $E(G[A])\neq\emptyset\neq E(G[B])$.∎
-Lemma 13.36.
-Let $G$ be a finite, 3-regular, internally-4-connected, vertex-transitive graph and let $L=L(G)$ be its line graph.
-Let the map $(A,B)\mapsto(A^{\prime},B^{\prime})$ send non-atomic edge-cuts $(A,B)$ of $G$ to genuine separations $(A^{\prime},B^{\prime})$ of $L$ as in the statement of Lemma˜13.31.
-Then this map restricts to a bijection from the set of tetra-cuts of $G$ to the set of tetra-separations of $L$.
-Proof.
-First note that $E(G[A])\neq\emptyset\neq E(G[B])$ for every nonatomic edge-cut $(A,B)$ of $G$ by ˜13.35.
-Hence, the prerequisites of Lemma˜13.31 are fulfilled.
-Since $G$ has no isolated vertices, the map $(A,B)\mapsto(A^{\prime},B^{\prime})$ is injective.
-We claim that every tetra-cut $(A,B)$ is mapped to a tetra-separation $(A^{\prime},B^{\prime})$.
-We get from Lemma˜13.31 that $(A^{\prime},B^{\prime})$ is a 4-separation of $L$.
-So it remains to check the degree-condition for the vertices in the separator of $(A^{\prime},B^{\prime})$.
-For this, let $v$ be an arbitrary vertex in $S(A^{\prime},B^{\prime})$.
-By Lemma˜13.33 (i), the vertex $v$ is contained in exactly two triangles $\Delta_{1},\Delta_{2}$, and the two triangles meet precisely in the vertex $v$.
-Each triangle $\Delta_{i}\subseteq L$ corresponds to an endvertex $x_{i}$ of $v$ viewed as an edge of $G$.
-Since $(A,B)$ is a tetra-cut, we have $x_{1}\in A\smallsetminus B$ and $x_{2}\in B\smallsetminus A$, say.
-Each $\Delta_{i}$ determines a 3-star in $G$ with centre $x_{i}$, and the edges in $S(A,B)$ form a matching from which each 3-star already contains $v$ (viewed as an edge of $G$).
-Hence $\Delta_{1}\subseteq L[A^{\prime}]$ and $v$ is the only vertex of $\Delta_{1}$ contained in $S(A^{\prime},B^{\prime})$.
-So $v$ has $\geqslant 2$ neighbours in $L[A^{\prime}\smallsetminus B^{\prime}]$.
-Similarly, $v$ has $\geqslant 2$ neighbours in $L[B^{\prime}\smallsetminus A^{\prime}]$.
-Therefore, $(A^{\prime},B^{\prime})$ is a tetra-separation of $L$.
-To conclude the proof, it remains to show that every tetra-separation of $L$ equals $(A^{\prime},B^{\prime})$ for some tetra-cut $(A,B)$.
-For this, let $(C,D)$ be an arbitrary tetra-separation of $L$.
-By Lemma˜13.34, $(C,D)$ is a genuine 4-separation.
-Let $(A,B)$ be the mixed-separation in $G$ as defined in Lemma˜13.32 applied to $(C,D)$.
-Then we claim that $(A,B)$ is a tetra-cut and $(A^{\prime},B^{\prime})=(C,D)$.
-By the degree-condition and Lemma˜13.33 (i), for every $v\in S(C,D)$ there are two triangles $\Delta_{1}(v),\Delta_{2}(v)$ with $\Delta_{1}(v)\smallsetminus\{v\}\subseteq C\smallsetminus D$ and $\Delta_{2}(v)\smallsetminus\{v\}\subseteq D\smallsetminus C$.
-That means, for two distinct $v,w\in S(C,D)$ we have $\Delta_{i}(v)\neq\Delta_{i}(w)$ for $i\in\{1,2\}$.
-With Lemma˜13.33 it follows that $(A,B)$ is a tetra-cut.
-Moreover, if $v\in C\smallsetminus D$ then $v\in E(G[A])$, and it follows that $v\in A^{\prime}\smallsetminus B^{\prime}$.
-Analogously, $v\in D\smallsetminus C$ implies $v\in B^{\prime}\smallsetminus A^{\prime}$.
-On the other hand, if $v\in C\cap D$ then $v\in S(A,B)$, and it follows that $v\in A^{\prime}\cap B^{\prime}$.
-Hence, $(A^{\prime},B^{\prime})=(C,D)$.
-∎
-Lemma 13.37.
-Let $G$ be a finite, 3-regular, internally-4-connected, vertex-transitive graph and let $L=L(G)$ be its line graph.
-Let $(A,B)\mapsto(A^{\prime},B^{\prime})$ be the bijection from the set of tetra-cuts $(A,B)$ of $G$ to the set of tetra-separations $(A^{\prime},B^{\prime})$ of $L$.
-Then $(A,B)\leqslant(C,D)$ if and only if $(A^{\prime},B^{\prime})\leqslant(C^{\prime},D^{\prime})$.
-Proof.
-If $(A,B)\leqslant(C,D)$ then $A\subseteq C$.
-Hence, $E(G[A])\subseteq E(G[C])$ and $S(A,B)\subseteq E(G[C])\cup S(C,D)$.
-It follows that $A^{\prime}\subseteq C^{\prime}$.
-Analogously, $B^{\prime}\supseteq D^{\prime}$.
-On the other hand, assume $(A^{\prime},B^{\prime})\leqslant(C^{\prime},D^{\prime})$.
-Let $v$ be a vertex in $A$.
-Recall that $E(v)$ denotes the edges incident with $v$ in $G$.
-Then $E(v)\subseteq A^{\prime}\subseteq C^{\prime}$.
-From $E(v)\subseteq C^{\prime}$ it follows that $v\in C$.
-Hence $A\subseteq C$ and, analogously, $B\supseteq D$.
-∎
-Lemma 13.38.
-Let $G$ be a finite 3-regular internally-4-connected vertex-transitive graph and let $(A_{1},A_{2})$ be a totally-nested tetra-cut in $G$. Then $G[A_{1}]=C_{4}$ or $G[A_{2}]=C_{4}$.
-Proof.
-Let $M$ be the union of the $\operatorname{Aut}(G)$-orbits of $(A_{1},A_{2})$ and $(A_{2},A_{1})$.
-Let $(C,D)$ be $\leqslant$-minimal in $M$; that is, $(C^{\prime},D^{\prime})\not<(C,D)$ for every $(C^{\prime},D^{\prime})\in M$.
-We claim that every vertex in $C$ is incident to an edge in $S(C,D)$.
-Assume for a contradiction that $v\in C$ is not.
-Let $u\in C\smallsetminus D$ be an endvertex of one of the four edges in $S(C,D)$ and let $\varphi\in\operatorname{Aut}(G)$ such that $\varphi(u)=v$.
-Then $(\varphi(C),\varphi(D))<(C,D)$ or $(\varphi(D),\varphi(C))<(C,D)$, contradicting the $\leqslant$-minimality of $(C,D)$.
-Finally, since $G$ is internally-4-connected, every vertex has degree $\geqslant 3$, so $G[C]$ is a $C_{4}$. The tetra-cut $(A_{1},A_{2})$ shares the above properties of $(C,D)$ since $(C,D)\in M$. ∎
-Lemma 13.39.
-Let $G$ be an internally-4-connected graph. Then every $(\leqslant 3)$-edge-cut of $G$ is an atomic 3-edge-cut.
-Proof.
-Since $G$ is 3-connected, it does not contain a $(\leqslant 2)$-edge-cut.
-Assume that $G$ has a 3-edge-cut $(A,B)$ with separator $S(A,B)=\{e_{1},e_{2},e_{3}\}$.
-Let $a_{i}$ be the endvertex of $e_{i}$ in $A$, and let $b_{i}$ be the endvertex of $e_{i}$ in $B$.
-First, assume that $(A,B)$ satisfies the matching-condition.
-Since $\{a_{1},a_{2},b_{3}\}$ is a 3-separator separating $a_{3}$ from $\{b_{1},b_{2}\}$, we must have $a_{1}a_{2}\notin E(G)$ by $G$ being internally-4-connected.
-Since $\{b_{1},a_{2},a_{3}\}$ is a 3-separator separating $a_{1}$ from $\{b_{2},b_{3}\}$, we must have $a_{1}a_{2}\in E(G)$ by $G$ being internally-4-connected.
-This is a contradiction.
-Second, assume that $(A,B)$ does not satisfy the matching-condition.
-If $(A,B)$ is not atomic, we have without loss of generality $a_{1}=a_{2}\neq a_{3}$ and $b_{1}\neq b_{3}$.
-Then, however, $\{a_{1},b_{3}\}$ is a 2-separator separating $a_{3}$ from $b_{1}$.
-Hence, $(A,B)$ is atomic.
-∎
-Lemma 13.40.
-Let $G$ be an internally-4-connected graph. Let $(A,B)$ and $(C,D)$ be two crossing tetra-cuts of $G$. Then they cross with all links of size 2 and empty centre.
-Proof.
-Recall that $L(X,Y)$ denotes the corner-separator for the corner $XY$. Note that, since $(A,B)$ and $(C,D)$ are edge-cuts, their corner-separators are edge-cuts as well.
-First assume that the centre is nonempty.
-Say, there is a diagonal edge from the $AD$-corner to the $BC$-corner.
-By Lemma˜4.3, either $|L(A,C)|\leqslant 3$ or $|L(B,D)|\leqslant 3$, say the former.
-By Lemma˜13.39, $L(A,C)$ is an atomic 3-edge-cut, contradicting that $(A,B)$ and $(C,D)$ satisfy the matching-condition.
-Similarly, we show that every link has size 2.
-Otherwise, there would exist a corner-separator of order $\leqslant 3$, say $|L(A,C)|\leqslant 3$.
-By Lemma˜13.39, $L(A,C)$ is an atomic 3-edge-cut, contradicting that $(A,B)$ and $(C,D)$ satisfy the matching-condition.
-∎
-Lemma 13.41.
-Let $G$ be a 3-regular internally-4-connected graph and let $L=L(G)$ be its line graph.
-Let $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$ be two crossing tetra-separations of $L$.
-Then they cross with all links of size 2 and empty centre.
-Proof.
-We combine Lemma˜13.36 with Lemma˜13.40. ∎
-Lemma 13.42.
-Let $G$ be a 3-regular internally-4-connected vertex-transitive finite graph and let $L=L(G)$ be its line graph.
-Let $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$ be two tetra-separations of $L$ that cross with all links of size 2.
-Let $\mathcal{O}=(O,\mathcal{G})$ be the Tutte-bagel of $L$ induced by $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$.
-Then every torso of $\mathcal{O}$ is 3-connected and has $\geqslant 5$ vertices.
-Proof.
-We may assume the (Tutte-bagel setting). (9.12) for $L$, but with $(A^{\prime},B^{\prime})$ in place of $(A,B)$ and $(C^{\prime},D^{\prime})$ in place of $(C,D)$.
-Assume that there is a torso $H_{s}$ in $\mathcal{O}$ that is a triangle or 4-cycle.
-Say, $H_{s}$ stems from the $A^{\prime}C^{\prime}$-Tutte-path.
-Since $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$ are genuine separations by Lemma˜13.34, no link contains an edge.
-It follows that all conditions for the (Link-Moving Lemma for vertices).
-(9.44) are fulfilled.
-By the (Link-Moving Lemma for vertices).
-(9.44(i), we get a tetra-separation of $L$ that contains at least one edge in its separator.
-This contradicts Lemma˜13.34.
-Therefore, every torso $H_{s}$ of $\mathcal{O}$ is 3-connected.
-Assume for a contradiction that there is a torso $H_{s}$ of $\mathcal{O}$ with $\leqslant 4$ vertices.
-Then $H_{s}=K_{4}$.
-By Lemma˜13.33 (iv), every edge $e\in E(G_{s})$ is contained in a unique triangle.
-If some adhesion set of $s$ induces an edge $e$, then $e$ is contained in two triangles.
-Otherwise, the edges in $G_{s}$ are not contained in any triangle.
-This is a contradiction.
-∎
-Let $G$ be a graph. A cycle-partition of $G$ is a pair $\mathcal{O}=(O,(V_{t})_{t\in V(O)})$ consisting of a cycle $O$ and vertex sets $V_{t}$ such that
-  1. (i)
-    $\{\,V_{t}\mid t\in V(O)\,\}$ is a partition of $V(G)$; and
-  2. (ii)
-    every edge $e\in E(G)$ either has its ends in a unique $V_{t}$ or there is an edge $tt^{\prime}\in E(O)$ such that $e$ has its ends in $V_{t}$ and $V_{t^{\prime}}$.
-We call the sets $V_{t}$ the bags of $\mathcal{O}$. For every $tt^{\prime}\in E(O)$, we call $E_{G}(V_{t},V_{t^{\prime}})$ the adhesion-set of $tt^{\prime}$.
-Lemma 13.43.
-Let $G$ be a 3-regular internally-4-connected vertex-transitive finite graph.
-Assume that $G$ has two crossing tetra-cuts.
-Then $G$ admits a cycle-partition $\mathcal{O}=(O,(V_{t})_{t\in O})$ with $|O|\geqslant 4$ such that all of the following are satisfied:
-  1. (i)
-    All adhesion-sets of $\mathcal{O}$ are 2-matchings;
-  2. (ii)
-    Every bag of $\mathcal{O}$ induces a $C_{4}$ or a $K_{2}$;
-  3. (iii)
-    Let $t\in V(O)$ and let $t_{1},t_{2}$ be its neighbours on $O$.
-Let $U_{i}$ be the set of endvertices of edges in the adhesion-set for $tt_{i}$ in $V_{t}$.
-If $V_{t}$ induces a 4-cycle, then this 4-cycle alternates between $U_{1}$ and $U_{2}$.
-Proof.
-Let $L=L(G)$ be the line graph of $G$.
-Let $(A,B)\mapsto(A^{\prime},B^{\prime})$ be the bijection from Lemma˜13.36 between tetra-cuts $(A,B)$ in $G$ and tetra-separations $(A^{\prime},B^{\prime})$ in $L$.
-By Lemma˜13.37(A,B)(A,B) is a totally-nested tetra-cut of $G$ if and only if $(A^{\prime},B^{\prime})$ is a totally-nested tetra-separation of $L$.
-Let $(A,B)$ and $(C,D)$ be two crossing tetra-cuts in $G$.
-By Lemma˜13.37 and Lemma˜13.41, the tetra-separations $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$ cross with links of size 2 and empty centre.
-Let $\mathcal{O}^{\prime}=(O^{\prime},(G^{\prime}_{t})_{t\in V(O^{\prime})})$ be the Tutte-bagel of $L$ induced by $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$.
-We denote by $H^{\prime}_{t}$ the torso of $t\in V(O^{\prime})$ in $\mathcal{O}^{\prime}$.
-By Lemma˜13.42, every torso of $\mathcal{O}^{\prime}$ is 3-connected on $\geqslant 5$ vertices.
-Recall that $E(v)$ denotes the set of edges incident with the vertex $v$ in $G$.
-We define $\mathcal{O}=(O,(V_{t})_{t\in V(O)})$ with $O:=O^{\prime}$ and $V_{t}:=\{v\in V(G)\mid E(v)\subseteq V_{t}\}$.
-We show that $\mathcal{O}$ is a cycle-partition of $G$.
-First, note that for every $v\in V(G)$, the set $E(v)$ induces a triangle in the graph $L$, which in turn is contained in a unique torso $H^{\prime}_{t}$ of $\mathcal{O}^{\prime}$.
-Hence, $\{\,V_{t}\mid t\in V(O)\,\}$ is a partition of $V(G)$.
-Second, let $e$ be an edge in $G$.
-Then $e$ is a vertex in $L$ which is contained in two triangles $\Delta_{1}(e)$ and $\Delta_{2}(e)$ in $L$.
-If $\Delta_{1}(e)$ and $\Delta_{2}(e)$ are contained in the same torso of $\mathcal{O}^{\prime}$, then both endvertices of $e$ are contained in the same bag of $\mathcal{O}$.
-Otherwise, since every torso of $\mathcal{O}^{\prime}$ is 3-connected the adhesion-sets of $\mathcal{O}^{\prime}$ are pairwise disjoint by Lemma˜9.4, so $\Delta_{1}(e)$ and $\Delta_{2}(e)$ are contained in two torsos of $\mathcal{O}^{\prime}$, whose corresponding nodes $t_{1},t_{2}\in V(O^{\prime})$ are neighbours on $O^{\prime}$.
-Then, the endvertices of $e$ are contained in $V_{t_{1}}$ and $V_{t_{2}}$, where $t_{1}t_{2}\in E(O)$.
-This shows that $\mathcal{O}$ is a cycle-partition of $G$.
-It remains to show that (i)(iii) are satisfied for the cycle-partition $\mathcal{O}$ of $G$.
-(i).
-Let $tt^{\prime}\in E(O)$.
-We show that the adhesion-set of $tt^{\prime}$ is a 2-matching.
-First, note that the adhesion-set of $tt^{\prime}$ in $\mathcal{O}$ is equal to the adhesion-set of $tt^{\prime}$ in $\mathcal{O}^{\prime}$.
-Since the latter consists of two vertices of $L$, the former consists of two edges of $G$.
-Since every torso of $\mathcal{O}^{\prime}$ is 3-connected and $L$ is 4-regular, it follows that the adhesion-set of $tt^{\prime}$ in $\mathcal{O}^{\prime}$ is independent.
-Hence, the adhesion-set of $tt^{\prime}$ in $\mathcal{O}$ is a matching.
-[…]
-Now assume that $H^{\prime}_{t}$ is bad.
-Then $V(H^{\prime}_{t})$ consists of exactly five vertices.
-Let $v$ be the vertex in the interior of $H^{\prime}_{t}$.
-By 4-regularity of $L$, the vertex $v$ is adjacent to every vertex in $H^{\prime}_{t}-v$.
-Since every torso of $\mathcal{O}^{\prime}$ is 3-connected, every vertex $u$ in an adhesion-set of $H^{\prime}_{t}$ has degree 2 in $G^{\prime}_{t}$.
-It follows that the torso $H^{\prime}_{t}$ is a 4-wheel, that $G^{\prime}_{t}$ is the union of two triangles that meet exactly in $v$, and that $V_{t}$ induces a $K_{2}$ in $G$.
-(iii).
-We denote the vertices in $U_{i}$ by $u_{i}$ and $u_{i}^{\prime}$.
-Assume for contradiction that $V_{t}$ induces a 4-cycle $Z$ that does not alternate between $U_{1}$ and $U_{2}$.
-As shown in the proof of (ii), the tetra-cut $(V_{t},V(G)\smallsetminus V_{t})$ is totally-nested.
-Then without loss of generality $Z=u_{1}u_{1}^{\prime}u_{2}^{\prime}u_{2}u_{1}$.
-Then, however, $(V_{t_{1}}\cup U_{1},V(G)\smallsetminus(V_{t_{1}}\cup U_{1}))$ is a tetra-cut that crosses $(V_{t},V(G)\smallsetminus V_{t})$, a contradiction.
-∎
-Key Lemma 13.44.
-Let $G$ be a 3-regular internally-4-connected vertex-transitive graph.
-Assume that $G$ has two crossing tetra-cuts.
-Then $G$ is a cycle of $C_{4}$-bags of length $\geqslant 4$, or a cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length $\geqslant 8$.
-Proof.
-By Lemma˜13.43GG admits a cycle-partition $\mathcal{O}=(O,(V_{t})_{t\in O})$ with $|O|\geqslant 4$ such that properties (i)(iii) from Lemma˜13.43 are fulfilled.
-Let $O=:t_{1}t_{2}\ldots t_{\ell}t_{1}$.
-For $t=t_{i}\in V(O)$ we denote by $U_{t,1}$ the set of endvertices of the edges in the adhesion-set for $tt_{i-1}$ that lie in $V_{t}$; and by $U_{t,2}$ we denote the set of endvertices of the edges in the adhesion-set for $tt_{i+1}$ that lie in $V_{t}$.
-We show that either every bag $V_{t}$ of $\mathcal{O}$ induces a $K_{2}$ or every bag $V_{t}$ of $\mathcal{O}$ induces a $C_{4}$ alternating between $U_{t,1}$ and $U_{t,2}$.
-In the former case, it follows that $G$ is a cycle of $C_{4}$-bags of length $\geqslant 4$, while in the latter case it follows that $G$ is a cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length $\geqslant 8$.
-Assume for a contradiction that there is a bag of $\mathcal{O}$ that induces a $K_{2}$, and that there is a bag of $\mathcal{O}$ that induces a $C_{4}$.
-Then, without loss of generality, one of the following cases occurs.
-The bags for $t_{1}$, $t_{2}$ and $t_{3}$ induce a $K_{2}$ while the bag for $t_{4}$ induces a $C_{4}$.
-Then every vertex in $V_{t_{2}}$ is contained in two $C_{4}$’s of $G$ while every vertex in $V_{t_{4}}$ is contained in exactly one $C_{4}$ of $G$, contradicting vertex-transitivity.
-The bags for $t_{2}$ and $t_{3}$ induce a $K_{2}$ while the bags for $t_{1}$ and $t_{4}$ induce a $C_{4}$.
-Let $u\in V_{t_{2}}$.
-Then $u$ is contained in exactly two 5-cycles $Z,Z^{\prime}$.
-The intersection $Z\cap Z^{\prime}$ is a path with three edges, and $u$ is an internal vertex of this path.
-The neighbour of $u$ in $V_{t_{1}}$ also lies on $Z$ and $Z^{\prime}$, but it is not an internal vertex of $Z\cap Z^{\prime}$, contradicting vertex-transitivity.
-The bag for $t_{2}$ induces a $K_{2}$ while the bags for $t_{1}$ and $t_{3}$ induce a $C_{4}$.
-Then every vertex in $V_{t_{2}}$ is contained in exactly four 5-cycles of $G$, while every vertex in $V_{t_{1}}\cap N(V_{t_{2}})$ is contained in at most three 5-cycles.
-This contradicts vertex-transitivity.
-∎
-Recall that a graph $G$ is 2-quasi-5-connected if $G$ is quasi-4-connected and every 4-separation $(A,B)$ of $G$ satisfies $|A\smallsetminus B|\leqslant 2$ or $|B\smallsetminus A|\leqslant 2$.
-Lemma 13.45.
-Let $G$ be a 3-regular quasi-4-connected graph. Assume that $G$ has no tetra-cut. Then $G$ is 2-quasi-5-connected.
-Proof.
-If $G$ is not 2-quasi-5-connected, it has a 4-separation $(A,B)$ with $|A\smallsetminus B|\geqslant 3$ and $|B\smallsetminus A|\geqslant 3$.
-Since $G$ is quasi-4-connected, it follows that $(A,B)$ is tight.
-Let $S(A,B)=\{v_{1},v_{2},v_{3},v_{4}\}$.
-First we show that if $v_{i}$ only has one neighbour $u_{i}$ in $A\smallsetminus B$ and $v_{j}\neq v_{i}$ only has one neighbour $u_{j}$ in $A\smallsetminus B$, then $u_{i}\neq u_{j}$.
-Indeed, otherwise $(A^{\prime},B^{\prime}):=((A\cup\{u_{i}\})\smallsetminus\{v_{i},v_{j}\},B\cup\{u_{i}\})$ is a 3-separation of $G$ with $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$, contradicting that $G$ is quasi-4-connected.
-By the above argument, it follows that the left-right-reduction of $(A,B)$ satisfies the matching-condition.
-Since $G$ is 3-regular and $(A,B)$ is tight, every vertex $v_{i}\in S(A,B)$ either has exactly one neighbour in $A\smallsetminus B$, or $v_{i}$ has exactly one neighbour in $B\smallsetminus A$ and no neighbour in $A\cap B$.
-Therefore, the left-right-reduction of $(A,B)$ is a 4-edge-cut and hence a tetra-cut of $G$.
-∎
-Example 13.46.
-The toroidal hex-grid on a torus shows that 2-quasi-5-connected in Lemma˜13.45 is best possible.
-Lemma 13.47.
-Let $G$ be a 3-regular internally-4-connected vertex-transitive graph such that every tetra-cut of $G$ is totally-nested.
-Then either $G$ is 2-quasi-5-connected or there is a canonical partition $\mathcal{P}=\{\,A_{i}:i\in I\,\}$ of $V(G)$ such that
-  1. (i)
-    each $A_{i}$ induces a $C_{4}$;
-  2. (ii)
-    the edges in $G$ with ends in different classes of $\mathcal{P}$ form a perfect matching in $G$.
-Proof.
-If $G$ has no tetra-cut, then $G$ is 2-quasi-5-connected by Lemma˜13.45.
-So we may assume that $G$ has at least one tetra-cut.
-Let $\sigma:=\{\,(A_{i},B_{i}):i\in I\,\}$ be the set of all tetra-cuts of $G$ whose side $A_{i}$ induces a 4-cycle.
-Since all tetra-cuts of $G$ are totally-nested, $\sigma$ is a star of tetra-cuts.
-Hence the sets $A_{i}$ are pairwise disjoint.
-Since $G$ is vertex-transitive, we have $\bigcup_{i\in I}A_{i}=V(G)$.
-Therefore, $\mathcal{P}:=\{\,A_{i}:i\in I\,\}$ is a canonical partition of $G$ into vertex-sets that induce 4-cycles.
-Let $M$ denote the set of all edges in $G$ with ends in different classes of $\mathcal{P}$.
-Every vertex $v$ of $G$ lies in exactly one $A_{i}$, and since $G$ is 3-regular, it follows that $v$ is incident to exactly one edge in $M$.
-Hence $M$ is a perfect matching.
-∎
-Key Lemma 13.48.
-Let $G$ be a 3-regular internally-4-connected vertex-transitive graph. Assume that every tetra-cut in $G$ is totally-nested. Then $G$ is one of the following:
-  • $\cdot$
-    a cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length 6;
-  • $\cdot$
-    one of the three $C_{4}$-expansions of the line graph of the cube shown in Figure˜1.15;
-  • $\cdot$
-    a vertex-transitive $C_{4}$-expansion of a cycle of $K_{2,2}$-bags of length $\geqslant 3$;
-  • $\cdot$
-    a $C_{4}$-expansion of a quasi-5-connected 4-regular arc-transitive graph;
-  • $\cdot$
-    2-quasi-5-connected.
-Proof.
-Assume that $G$ is not 2-quasi-5-connected.
-Then $V(G)$ has a canonical partition $\mathcal{P}$ into vertex-sets that induce 4-cycles by Lemma˜13.47.
-Fix $U\in\mathcal{P}$ and let $\{\,U_{i}:i\in I\,\}$ be the set of classes $U_{i}\in\mathcal{P}$ that contain neighbours of $U$.
-Let $M$ denote the set of edges of $G$ with ends in different classes of $\mathcal{P}$, so $M$ is a perfect matching by Lemma˜13.47.
-Let $F\subseteq M$ denote the set of edges leaving $U$.
-Case: all four edges in $F$ end in the same $U_{i}$.
-Then $F=M$ and so $V(G)=U\cup U_{i}$.
-Both $U$ and $U_{i}$ induce 4-cycles, so either $G=C_{4}\square K_{2}$ which is a cycle of $C_{4}$-bags of length 4, or $G$ is the life belt.
-However, in both cases $G$ has two crossing tetra-cuts, so neither case occurs.
-Case: exactly three edges in $F$ end in the same $U_{i}$.
-This case is impossible: Then the vertex in $U$ and the vertex in $U_{i}$ that are not incident with these three edges form a 2-separator of $G$, contradicting that $G$ is internally-4-connected.
-Case: there is a $U_{i}$ in which exactly two edges from $F$ end.
-Then four edges from $M$ leave $U\cup U_{i}$.
-Hence $U\cup U_{i}$ is a side of a tetra-cut of $G$.
-Since this tetra-cut is totally-nested by assumption, it has a side that induces a 4-cycle by Lemma˜13.38.
-This side can only be $V(G)\smallsetminus(U\cup U_{i})$, which then must be a $U_{j}$ for some $j\in I\smallsetminus\{i\}$.
-Now $U$ and $U_{i}$ each send exactly two edges to $U_{j}$.
-If $G$ is a cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length 6, then we are done.
-Otherwise the 4-cycle $O$ induced by $U$, say, first traverses the two neighbours $u_{i},u_{i}^{\prime}$ of $U_{i}$ in $U$, and then traverses the two neighbours $u_{j},u_{j}^{\prime}$ of $U_{j}$ in $U$.
-But then (U,V(GU))(U,V(G-U)) and $(U_{i}\cup\{u_{i},u_{i}^{\prime}\},U_{j}\cup\{u_{j},u_{j}^{\prime}\})$ are two crossing tetra-cuts in $G$, contradicting our assumption.
-Case: $|I|=4$ and exactly one edge from $F$ ends in each $U_{i}$.
-Then $G/\mathcal{P}$ is 4-regular and 4-connected.
-Moreover, $G/\mathcal{P}$ is arc-transitive, as $\mathcal{P}$ is unique and every vertex of $G$ is incident with a unique edge in $F$.
-Then by Corollaries˜13.6 and 13.25, the graph $G/\mathcal{P}$ is one of the following:
-  • $\cdot$
-    the line-graph of the cube,
-  • $\cdot$
-    a cycle of $K_{2,2}$-bags of length $\geqslant 3$, or
-  • $\cdot$
-    essentially-5-connected.
-Assume first that $G/\mathcal{P}$ is the line-graph of the cube, so $G$ is some $C_{4}$-expansion thereof.
-Since $G$ is vertex-transitive, there are only three possible $C_{4}$-expansions: these are precisely the three depicted in Figure˜1.15.
-Assume finally that $G/\mathcal{P}$ is essentially-5-connected.
-Then $G/\mathcal{P}$ is quasi-5-connected by Corollary˜13.28.
-Hence $G$ is the $C_{4}$-expansion of a quasi-5-connected 4-regular arc-transitive graph.
-∎
-Proposition 13.49.
-A 3-regular quasi-4-connected finite graph $G$ is vertex-transitive if and only if $G$ is either
-  1. (i)
-    2-quasi-5-connected and vertex-transitive,
-  2. (ii)
-    a vertex-transitive $C_{4}$-expansion of a quasi-5-connected 4-regular arc-transitive graph,
-  3. (iii)
-    a cycle of $C_{4}$-bags of length $\geqslant 4$ or $K_{3}\square K_{2}$,
-  4. (iv)
-    a cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length $\geqslant 6$,
-  5. (v)
-    one of the three $C_{4}$-expansions of the line-graph of the cube shown in Figure˜1.15, or
-  6. (vi)
-    a vertex-transitive $C_{4}$-expansion of a cycle of $K_{2,2}$-bags of length $\geqslant 3$.
-Proof.
-($\Rightarrow$).
-By ˜13.30, we may assume that $G$ is internally 4-connected and has girth $\geqslant 4$.
-If $G$ has two crossing tetra-cuts, then we are done by ˜13.44.
-Otherwise every tetra-cut of $G$ is totally-nested.
-Then we are done by ˜13.48.
-($\Leftarrow$) is straightforward. ∎
-Proof of Figure˜1.8.
-Proposition 13.50.
-[CK23, Corollary 2]. Every vertex-transitive finite connected $G$ is either
-  • $\cdot$
-    3-connected and every non-trivial strong tri-separation of $G$ is a 3-edge-cut with a side that induces a triangle,
-  • $\cdot$
-    a cycle, $K_{2}$ or $K_{1}$.
-Proposition 13.51.
-Every vertex-transitive finite connected graph $G$ is either
-  • $\cdot$
-    internally 4-connected and vertex-transitive,
-  • $\cdot$
-    the $K_{3}$-expansion of a quasi-4-connected 4-regular arc-transitive graph,
-  • $\cdot$
-    a cycle, $K_{3}\square K_{2}$, $K_{2}$ or $K_{1}$.
-Proof.
-By Proposition˜13.50, we may assume that $G$ is 3-connected and every non-trivial strong tri-separation of $G$ is a 3-edge-cut with a side that induces a triangle.
-If $G$ has no non-trivial strong tri-separation, then $G$ is internally-4-connected by [CK23, 1.2.8].
-Hence we may assume that $G$ has at least one non-trivial strong tri-separation, and by assumption it has a side that induces a triangle.
-Let $\{\,(A_{i},B_{i}):i\in I\,\}=\sigma$ be the set of all non-trivial strong tri-separations of $G$ with $G[A_{i}]=K_{3}$.
-Since each $S(A_{i},B_{i})$ is a 3-edge-cut, the tri-separations $(A_{i},B_{i})$ are totally-nested by [CK23, 1.3.10], so $\sigma$ is a star of tri-separations.
-Hence the $A_{i}$ are pairwise disjoint, and every vertex of $G$ is contained in some $A_{i}$ by vertex-transitivity, so $\mathcal{P}:=\{\,A_{i}:i\in I\,\}$ is a partition of $V(G)$.
-Since each $S(A_{i},B_{i})$ is a 3-matching, the set $M$ of cross-edges of $\mathcal{P}$ is a perfect matching.
-If two edges in $M$ join $A_{i}$ to $A_{j}$ for some $i\neq j\in I$, then the two vertices in $A_{i}\cup A_{j}$ that are not incident to these edges form a 2-separator of $G$ contradicting 3-connectivity, unless $G=K_{3}\square K_{2}$.
-Hence we may assume that at most one edge in $M$ joins any $A_{i}$ to any $A_{j}$, so $G/\mathcal{P}$ is 3-regular.
-Since $G$ is vertex-transitive and the $K_{3}$-expansion of $G/\mathcal{P}$, it follows that $G/\mathcal{P}$ is arc-transitive by Lemma˜13.21.
-In particular, $G/\mathcal{P}$ is vertex-transitive, and since $G/\mathcal{P}$ is 3-regular we get that $G/\mathcal{P}$ is 3-connected by Proposition˜13.50.
-However, $G/\mathcal{P}$ has no non-atomic 3-edge-cut, as this would yield a non-trivial strong tri-separation of $G$ that is a 3-edge-cut but whose sides have size $\geqslant 9$, contradicting our assumption.
-Hence applying Proposition˜13.50 to $G/\mathcal{P}$ in combination with [CK23, 1.2.8] yields that $G/\mathcal{P}$ is quasi-4-connected.
-∎
-Acknowledgement.
-We thank a referee for valuable comments on the introduction.
-We thank another referee for bringing [LPPS22] to our attention.
-We thank Joseph Devine for exciting discussions.
-We thank Agelos Georgakopoulos for telling us about Theorem˜13.1 of Godsil and Royle.
 13. Appendix A: Circular saws
@@
 Proof.
-For each vertex $x\in X$ we construct two xxYY paths $P^{+}_{x},P^{-}_{x}$ inductively, as follows. See Figure˜14.1 for an example.
+For each vertex $x\in X$ we construct two xxYY paths $P^{+}_{x},P^{-}_{x}$ inductively, as follows. See fig.˜13.1 for an example.
 In the first step of the construction of $P^{+}_{x}$, we let $P^{+}_{x}$ start and end with $x$.
@@
 ∎
-Figure 14.1. Paths constructed in the proof of Lemma˜14.1. The edges $aa^{\prime}$ and $bb^{\prime}$ are coloured blue, the paths are coloured red.
+Figure 13.1. Paths constructed in the proof of lemma˜13.1. The edges $aa^{\prime}$ and $bb^{\prime}$ are coloured blue, the paths are coloured red.
 Lemma 13.2.
@@
 Note that $2(k-1)\geqslant k+1$ since $k\geqslant 3$.
-Hence by Lemma˜14.1, there are $k+1$ disjoint $N(\{a,a^{\prime}\})$$N(\{b,b^{\prime}\})$ paths in $G$, contradicting that $|S|\leqslant k-1$.
+Hence by lemma˜13.1, there are $k+1$ disjoint $N(\{a,a^{\prime}\})$$N(\{b,b^{\prime}\})$ paths in $G$, contradicting that $|S|\leqslant k-1$.
 Hence $G$ is $k$-connected.
@@
 Then there are vertices $a^{\prime}\in N(a)\smallsetminus S(A,B)$ and $b^{\prime}\in N(b)\smallsetminus S(A,B)$.
-Now Lemma˜14.1 yields $k+1$ disjoint $N(\{a,a^{\prime}\})$$N(\{b,b^{\prime}\})$ paths in $G$, contradicting that $(A,B)$ has order $k$.
+Now lemma˜13.1 yields $k+1$ disjoint $N(\{a,a^{\prime}\})$$N(\{b,b^{\prime}\})$ paths in $G$, contradicting that $(A,B)$ has order $k$.
 ∎
 14. Appendix B: Comparison of strict tri-separations and tri-separations
-The strict tri-separation-decomposition (˜2) is different from the tri-separation decomposition [CK23, Theorem 1] as can be seen in Figures˜15.1 and 15.2.
-However, the strict tri-separation-decomposition can be viewed as a refinement of the tri-separation decomposition via Lemma˜15.1.
+The strict tri-separation-decomposition (˜2) is different from the tri-separation decomposition [CK23, Theorem 1] as can be seen in figs.˜14.1 and 14.2.
+However, the strict tri-separation-decomposition can be viewed as a refinement of the tri-separation decomposition via lemma˜14.1.
 Figure 14.1. The totally-nested strict tri-separations
@@
 ∎
+Acknowledgement.
+We thank a referee for valuable comments on the introduction.
+We thank another referee for bringing [LPPS22] to our attention.
+We thank Joseph Devine for exciting discussions.
+We thank Agelos Georgakopoulos for telling us about a theorem of Godsil and Royle that we use in [KP26].